### v1

# Existence of genus 2 minimal surfaces in 3-spheres. I

Abstract.
We prove that every 3-sphere of positive Ricci curvature contains some embedded minimal surface of genus 2.
We also establish a theorem for more general 3-manifolds that relates the existence of genus $g$ minimal surfaces to topological properties regarding the set of all embedded singular surfaces of genus $\leq g$.
1. Introduction
Enumerative problems in differential geometry have a long history. For example, we have the following well-known conjectures:
  • In 1905, Poincaré conjectured that every Riemannian 2-sphere has at least 3 simple closed geodesics.
  • In 1982, S.-T. Yau conjectured every Riemannian 3-sphere has at least 4 embedded minimal spheres [Yau82].
  • In 1989, B. White conjectured every Riemannian 3-sphere has at least 5 embedded minimal tori [Whi89].
Among these three conjectures, only the first one was fully solved: Building on the work of Lyusternik-Schnirelmann [LS47], M. Grayson gave a proof [Gra89] using curve shortening flow. Yau’s conjecture was solved in the case of bumpy metric and positive Ricci curvature by Wang-Zhou [WZ23]: They developed a multiplicity one thereom for Simon-Smith min-max theory, using some regularity results from [WZ25] and from Sarnataro-Stryker’s work  [SS23] (see also [Smi82, Whi91, HK19]). Recently, the first two authors proved B. White’s conjecture in the case where the ambient metric has positive Ricci curvature [CL24]. See also X. Li and Z. Wang’s recent work on the existence of 4 minimal tori in 3-spheres with positive Ricci curvature, and 9 minimal tori in generic metrics with positive Ricci curvature [LW24].
Following this line of inquiry, in this paper we prove the following theorem.
Theorem 1.1.
Every Riemannian 333-sphere of positive Ricci curvature contains some embedded minimal surface of genus 222.
In fact, we expect the existence of multiple minimal surfaces of genus 2, but this would be addressed in a separate paper. More generally, in view of the three conjectures above, it is natural to pose the following question.
Question 1.2.
In a given Riemannian 333-manifold, how can one construct minimal surfaces of a specific genus $g$?
In recent years, min-max theory has been really successful in constructing minimal hypersurfaces [MN14, MN16, MN17, IMN18, LMN18, Zho20, Li22, Li23, Son23]. However, it is usually difficult to pin down the topology of the min-max minimal hypersurface obtained. For example, in the Simon-Smith min-max theory, the minimal surfaces produced only have genus upper bound, but not lower bound in general. One strategy to resolve this issue is to impose symmetry on the ambient manifold and the surfaces of interest: This is known as equivariant min-max theory [Ket16, CFS20, BNS21, FKS24]. In this paper, we will directly tackle this issue without imposing symmetry. Instead, we establish a general theorem that reduces the construction of a genus $g$ minimal surface to a topological problem of understanding the structure of “the space of all surfaces” in the given 3-manifold. This theorem not only enables us to prove Theorem 1.1 but also serves as a tool to tackle Question 1.2.
Here is our setting: In a closed Riemannian 3-manifold $M$, we let $\mathcal{S}(M)$ be the set of all surfaces embedded in $M$ that possibly have finitely many singularities. For some technical reasons, we will focus on surfaces that are orientable, and separate $M$ into two regions. Note that the notion of genus extends naturally to singular surfaces, so we may let $\mathcal{S}_{\leq g}(M)\subset\mathcal{S}(M)$ be the set of all singular surfaces of genus $\leq g$: Readers may refer to §2 for the precise definitions.
The following is our general theorem, which essentially states that if there is a family $\Phi$ of singular surfaces with genus $\leq g$ that detects some non-trivial relative structure of the pair $(\mathcal{S}_{\leq g}(M),\mathcal{S}_{\leq g-1}(M))$, then $\Phi$ can be used to produce a minimal surface of genus $g$: See Figure 1 for a schematic illustration.
Figure 1. This schematic shows the set $\mathcal{S}_{\leq g}(M)$ of all surfaces with genus at most $g$, possibly with singularities. The gray cap below represents $\mathcal{S}_{\leq g-1}(M)$, while the blue part represents the image of a map $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}_{\leq g}(M)$, detecting some non-trivial relative structure of the pair $(\mathcal{S}_{\leq g}(M),\mathcal{S}_{\leq g-1}(M))$.
Theorem 1.3.
Let $M$ be a closed orientable Riemannian 333-manifold with positive Ricci curvature, and $g$ be a positive integer. Suppose there exists a Simon-Smith family
$\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}_{\leq g}(M)$
that cannot be deformed, via pinch-off processes, into a map into $\mathcal{S}_{\leq g-1}(M)$. Then $M$ contains an orientable, embedded minimal surface of genus $g$ with area at most $\max_{x\in X}\operatorname{area}(\Phi(x))$.
The definitions of “Simon-Smith family” and “deformation via pinch-off processes” are given in Definition 2.7 and Definition 2.20 respectively. Note that we will not equip $\mathcal{S}_{\leq g}(M)$ with a topology, but instead, the notion of Simon-Smith family includes mild conditions on the map $\Phi$ to ensure the applicability of the Simon-Smith min-max theory. A pinch-off process refers to a deformation process for elements in $\mathcal{S}(M)$ that consists of any combination of the following three procedures: (1) isotopy, (2) neck-pinch surgery, and (3) shrinking connected components to points (see Figure 2). Such processes resemble the evolution under mean curvature flow with good singularities—roughly speaking, those without fattening. In particular, a pinch-off process does not increase genus. This methodology has been introduced in [CL24], and readers may refer to [CL24, §1.2.2 - §1.2.4, §1.2.6].
Figure 2. This is an example of pinch-off process. It has one neck-pinch surgery, and one connected component shrunk to a point.
The above theorem builds upon numerous important and foundational results in Simon-Smith min-max theory  [Smi82, CD03, DP10, Ket19, Zho20, MN21, WZ23]. We also note that the proof of the theorem requires repetitively running min-max procedure, so the area of the resulting genus $g$ minimal surface may be less than the Simon-Smith width associated with the initial family $\Phi$.
With Theorem 1.3, we can prove Theorem 1.1 by showing that the pair $(\mathcal{S}_{\leq 2}(S^{3}),\mathcal{S}_{\leq 1}(S^{3}))$ has some non-trivial relative structure, in the sense that there exists a map $\Psi$ into $\mathcal{S}_{\leq 2}(S^{3})$ that cannot be deformed via pinch-off processes to a map into $\mathcal{S}_{\leq 1}(S^{3})$ (this will be proved in §3.3). In fact, we will construct a map $\Psi$ with a 777-dimensional domain. Intuitively, this is compatible with the fact that the genus 222 Lawson surface in the unit 333-sphere has Morse index 777, as proven by Kapouleas-Wiygul [KW20].
Remark 1.4.
With respect to the assumption of positive Ricci curvature, a primary purpose is to guarantee the Frankel property, that every embedded minimal surface is connected. Without this, heuristically, it is possible that while the pair $(\mathcal{S}_{\leq g}(M),\mathcal{S}_{\leq g-1}(M))$ has some non-trivial relative structure, the “genus $g$” minimal surface produced is however just the union of multiple minimal surfaces of lower genus whose total genus sums to $g$. Furthermore, positive Ricci curvature enables the application of the multiplicity one theorem by Wang-Zhou [WZ22], which will play a crucial role in our argument.
In addition, we assume that the ambient 3-manifold is orientable, as most existing results in min-max theory are proved under this assumption. Removing it would require adapting the min-max theory to non-orientable 3-manifolds, which is plausible but involves additional technical work.
For additional works on the construction of geodesics or minimal surfaces with controlled topological type, we refer the reader to [Str84, GJ86, Zho16, BP23, HK23, Ko23, Ko23a, GP24].
We will now describe the main techniques involved, and sketch the proofs of the above theorems. Throughout, we assume the ambient manifold $M$ has positive Ricci curvature.
1.1. Obtaining an optimal family via repetitive min-max and pinch-off processes
A key technique underlying Theorem 1.1 and Thoerem 1.3 is the construction of an “optimal family $\Phi^{\prime}$ from a given Simon-Smith family $\Phi$. We begin with a heuristic description of what we mean by an “optimal family”, and then proceed to a more detailed explanation.
1.1.1. Inspiration: Mean curvature flow
In $M$, let $\Phi$ be a family of embedded surfaces, possibly with singularities. Recall that mean curvature flow is a gradient flow for the area functional. For each member $\Phi(x)$ of the family, let $\{\Phi^{t}(x)\}_{t\in[0,\infty)}$ denote the evolution under the mean curvature flow (formally ignoring issues related to long-time existence due to the non-uniqueness issue). We then consider the family $\Phi^{t}$, take the limit as $t\to\infty$, and define the limiting family
$\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}=\lim_{t\to\infty}\Phi^{t}\,.$
Note that the process of obtaining $\Phi^{\prime}$ from $\Phi$ resembles a homotopy between families of surfaces.
It is expected that the family $\Phi^{\prime}$ may contain some minimal surfaces $\Gamma_{1},\dots,\Gamma_{n}$, and admits a decomposition
$\Phi^{\prime}=\Phi^{\prime}_{1}\cup\dots\cup\Phi^{\prime}_{n}\,,$
where each $\Phi^{\prime}_{i}$, resembles the “unstable manifold” associated with the minimal surface $\Gamma_{i}$, consisting of ancient mean curvature flow lines emanating from $\Gamma_{i}$: See Figure 3. (Though, as we will later see, in practice we can only obtain genus control but not area control for each part $\Phi^{\prime}_{i}$.)
However, the non-uniqueness issue in the mean curvature flow prevents us to obtain $\Phi^{\prime}$ in this way. Fortunately, we only need $\Phi^{\prime}$ to admit the decomposition property above. Admittedly, this discussion is somewhat vague and inaccurate. Let us now give a more concrete explanation of the essential properties of $\Phi^{\prime}$.
Figure 3. The figure on the left is the original family $\Phi$, while the figure on the right is $\Phi^{\prime}$, with the three critical points representing the minimal surfaces $\Gamma_{1},\Gamma_{2}$ and $\Gamma_{3}$ detected, and the gradient flow lines representing ancient mean curvature flow originating from the minimal surfaces (as $t\to-\infty$). For example, the red region represents the “unstable manifold” associated with the critical point at the left mountain peak.
1.1.2. Optimal family
Let $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ be a family of singular surfaces of genus $\leq g$. Assume that $M$ has positive Ricci curvature, and that all embedded minimal surfaces are non-degenerate and have linearly independent area over $\mathbb{Z}$. Then we can “deform” $\Phi$ to obtain an “optimal” family $\Phi^{\prime}$ with the following property. There exist:
  • orientable embedded minimal surfaces $\Gamma_{1},\dots,\Gamma_{n}$ of genus $\leq g$,
  • subcomplexes $D_{0},D_{1},\dots,D_{n}\subset X$ that cover $X$,
such that:
  1. (1)
    $\Phi^{\prime}$ is constructed from $\Phi$ via pinch-off processes, which do not increase genus.
  2. (2)
    For each $i=1,\dots,n$, the genus of each member of $\Phi^{\prime}|_{D_{i}}$ is at most that of $\Gamma_{i}$, while each member of $\Phi^{\prime}|_{D_{0}}$ has genus 0.
  3. (3)
    For each $i=1,\dots,n$, the family $\Phi^{\prime}|_{D_{i}}$ can be constructed from some other family $\Xi_{i}$ via pinch-off processes such that each member of $\Xi_{i}$ is close to $\Gamma_{i}$, and has genus the same as $\Gamma_{i}$; in our construction, $\Xi_{i}$ is not necessarily a subfamily of $\Phi^{\prime}$.
Readers may refer to Figure 4 for a schematic illustration for the optimal family $\Phi^{\prime}$, and Theorem 3.1 for the precise statements. Heuristically, each $\Phi^{\prime}|_{D_{i}}$ still resembles the “unstable manifold” associated to $\Gamma_{i}$. However, unlike actual unstable manifolds, the members of $\Phi^{\prime}|_{D_{i}}$ only admit genus control, not area control.
Figure 4. The three critical points are $\Gamma_{1},\Gamma_{2},$ and $\Gamma_{3}$, from left to right. The entire red part is $\Phi^{\prime}|_{D_{1}}$. The gray part at the bottom is $\Phi|_{D_{0}}$ where the genus is zero.
In practice, to construct the optimal family $\Phi^{\prime}$, we iteratively apply the Simon-Smith min-max theorem to $\Phi$ and the deformation from $\Phi$ to $\Phi^{\prime}$ will be given by “pinch-off processes,” as illustrated by Figure 2.
Remark 1.5.
The feature that we get exactly one “unstable manifold” $\Phi^{\prime}|_{D_{i}}$ for each critical point $\Gamma_{i}$ would be especially useful for finding multiple minimal surfaces of designated genus: Counting the number of these pieces $\Phi^{\prime}|_{D_{i}}$ would give information about the number of distinct minimal surfaces of that genus. This was crucially used by the first two authors to find 5 minimal tori in [CL24]. But in this paper, since we just need to find one minimal surface of genus $g$ in the main theorems, this feature will not be utilized to its full strength.
1.2. Sketch of proof for Theorem 1.3
Suppose that $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}_{\leq g}(M)$ is a Simon-Smith family that cannot be deformed (via pinch-off processes) to a map into $\mathcal{S}_{\leq g-1}(M)$. For simplicity, we will write $\mathcal{S}_{\leq g}=\mathcal{S}_{\leq g}(M)$. Suppose by contradiction $M$ does not have any embedded minimal surfaces of genus $g$. As a result, for the optimal family $\Phi^{\prime}$ associated with the given family $\Phi$, all the minimal surfaces $\Gamma_{1},\dots,\Gamma_{n}$ mentioned in §1.1.2 must have genus $\leq g-1$. Hence, by item (2) in §1.1.2, the family $\Phi^{\prime}$ maps into $\mathcal{S}_{\leq g-1}$. So we have deformed $\Phi$ to become a map into $\mathcal{S}_{\leq g-1}$. This contradicts the assumption on $\Phi$.
1.3. Sketch of proof for Theorem 1.1 via Theorem 1.3
Let $(S^{3},\mathbf{g}_{0})$ be a Riemannian 333-sphere with positive Ricci curvature. First, in §1.3.1 we will define a 7-parameter family $\Psi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}_{\leq 2}$ of singular surfaces of genus 2$\leq$ 2≤ 2 in $S^{3}$. Then in §1.3.2, we show that $\Psi$ cannot be deformed to become a map into $\mathcal{S}_{\leq 1}$. This means the pair $(\mathcal{S}_{\leq 2},\mathcal{S}_{\leq 1})$ has some non-trivial relative structure, and by Theorem 1.3, we obtain a genus 2 minimal surface.
1.3.1. A 777-parameter family $\Psi$
We construct in the unit 3-sphere $S^{3}\subset\mathbb{R}^{4}$ a 7-parameter family of singular surfaces with genus 2$\leq$ 2≤ 2 as follows.
Consider an ${\mathbb{RP}}^{5}$-family $\Phi_{5}$ of surfaces in $S^{3}$: For each $[a_{0}\mathrel{\mathop{\ordinarycolon}}a_{1}\mathrel{\mathop{\ordinarycolon}}\dots\mathrel{\mathop{\ordinarycolon}}a_{5}]\in{\mathbb{RP}}^{5}$, $\Phi_{5}$ maps it to the following zero set in $S^{3}$:
$\{a_{0}+a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3}+a_{4}x_{4}+a_{5}x_{1}x_{2}=0\}\,.$
This family is a 5-sweepout in the Almgren-Pitts min-max sense, but it is not a family allowed in the Simon-Smith min-max setting, as it contains, for example, the set $\{x_{1}x_{2}=0\}$, which is a union of two equators intersecting perpendicularly along the great circle $C\mathrel{\mathop{\ordinarycolon}}=\{x_{1}=x_{2}=0\}$. Thus, we need to desingularize it, by replacing the intersection region with some handles (like Scherk tower), which would produce singular surfaces of genus 1 or 2. Below, we will first explain how to desingularize the union of equators $x_{1}x_{2}=0$: In fact, we will introduce a $\mathbb{D}$-family of ways to desingularize it ($\mathbb{D}$ denotes the closed unit 2-disc). After that, we describe how to desingularize the whole family $\Phi_{5}$, which contains other surfaces with singular curves: Namely, surfaces given by $(x_{1}+a_{1})(x_{2}+a_{2})=0$. Note, after desingularization, the surfaces are still allowed to have finitely many singularities.
We introduce a coordinate system $(x_{1},x_{2},\alpha)$ for $S^{3}$ (on the complement of a great circle $\{x_{3}=x_{4}=0\}$), define by the relation
$(x_{1},x_{2},x_{3},x_{4})=(x_{1},x_{2},\sqrt{1-x_{1}^{2}-x_{2}^{2}}\cos\alpha,\sqrt{1-x_{1}^{2}-x_{2}^{2}}\sin\alpha)\,,$
where $(x_{1},x_{2})$ lies in the open unit disc in $\mathbb{R}^{2}$ and $\alpha\in[0,2\pi)$. Note that $(0,0,\alpha)$ parametrizes the great circle $C=\{x_{1}=x_{2}=0\}$. Here are two examples of surfaces in $S^{3}$ that desingularize $\{x_{1}x_{2}=0\}$, which, geometrically, are obtained by slightly opening up the original surface $\{x_{1}x_{2}=0\}$ along the circle $C$ (see Figure 5):
  • The zero set $\Sigma_{1}\mathrel{\mathop{\ordinarycolon}}=\{x_{1}x_{2}+\epsilon\cos 2\alpha=0\}$.
  • The zero set $\Sigma_{2}\mathrel{\mathop{\ordinarycolon}}=\{x_{1}x_{2}+\epsilon\cos 3\alpha=0\}$.
Figure 5. This show the genus one surface $\Sigma_{1}$ and the genus two surface $\Sigma_{2}$. Note the vertical axis represents the great circle $C$.
Here $\epsilon>0$ is a fixed small number. In fact, the first example is a smooth genus one surface, while the second is a smooth genus 2 surface. We note that this way of desingularizing a pair of intersecting spheres (or planes) by adding some small term to the equation was previously used by the first author in [Chu23].
Now, we shift the handles of the above genus one surface $\Sigma_{1}$ along the circle $C$ (see Figure 6):
  • For each $\theta\in[0,2\pi)$, consider the zero set $\{x_{1}x_{2}+\epsilon\cos(\theta+2\alpha)=0\}$.
This is an $S^{1}$-family of smooth tori ($\theta=0$ and $\theta=2\pi$ give the same surface), and we denote this family by $\mathcal{T}$. A crucial feature is that, this family $\mathcal{T}$ is actually homologically non-trivial in the space of all unknotted tori in $S^{3}$ [JM10, Theorem 4].
Figure 6. This shows an $S^{1}$-family $\mathcal{T}$ of smooth tori, for the part near the great circle $C$. From left to right, we have $\theta=0,\pi/2,\pi,3\pi/2$. Note the handles are being shifted down.
Then, we interpolate between $\Sigma_{2}$ and this $S^{1}$-family $\mathcal{T}$ as follows:
  • For each $(r,\theta)$ in the closed unit 2-disc $\mathbb{D}$ (under polar coordinate), we consider the zero set
    (1.1) $\{x_{1}x_{2}+\epsilon\left(r\cos(\theta+2\alpha)+(1-r)\cos 3\alpha\right)=0\}\,.$
Namely, the center of $\mathbb{D}$ represents $\Sigma_{2}$, while the boundary $\partial\mathbb{D}$ represents the $S^{1}$-family $\mathcal{T}$ of tori. This yields the desired $\mathbb{D}$-family of desingularizations of the union of equators, $\{x_{1}x_{2}=0\}$.
Finally, we desingularize the whole family $\Phi_{5}$. In fact, we will consider a $\mathbb{D}$-family of desingularization of $\Phi_{5}$, i.e., we will define an $({\mathbb{RP}}^{5}\times\mathbb{D})$-family $\Psi$ of singular surfaces. Namely, for each $a=[a_{0}\mathrel{\mathop{\ordinarycolon}}a_{1}\mathrel{\mathop{\ordinarycolon}}\dots\mathrel{\mathop{\ordinarycolon}}a_{5}]\in{\mathbb{RP}}^{5}$ and $(r,\theta)\in\mathbb{D}$, we let $\Psi$ map them to the zero set (in $S^{3}$)
[…]
$+a_{5}\left[x_{1}x_{2}+\epsilon(r\cos(\theta+2\alpha)+(1-r)\cos 3\alpha)\right]=0\}\,.$
Up to some minor modification, we can guarantee that each singular surface $\Psi(a,(r,\theta))$ has genus 2$\leq$ 2≤ 2, so $\Psi$ maps into $\mathcal{S}_{\leq 2}$. Moreover, it is easy to check that for each $z\in\mathbb{D}$, the subfamily $\Psi|_{{\mathbb{RP}}^{5}\times\{z\}}$ is a 5-sweepout in the Almgren-Pitts min-max sense.
1.3.2. $\Psi$ cannot be deformed into $\mathcal{S}_{\leq 1}$
As explained at the beginning of §1.3, to prove Theorem 1.1 via Theorem 1.3, it suffices to show that $\Psi$ cannot be deformed (via pinch-off processes) to a map into $\mathcal{S}_{\leq 1}$.
Suppose the otherwise, that it can be deformed to be a map $\Psi^{\prime}\mathrel{\mathop{\ordinarycolon}}{\mathbb{RP}}^{5}\times\mathbb{D}\to\mathcal{S}_{\leq 1}$ and denote
$Y\mathrel{\mathop{\ordinarycolon}}={\mathbb{RP}}^{5}\times\mathbb{D}=\mathrm{dmn}(\Psi)=\mathrm{dmn}(\Psi^{\prime}).$
We can split $Y$ into a disjoint union, $Z_{0}\cup Z_{1}$, where $Z_{k}$ is the set of parameters corresponding to genus $k$ singular surfaces under the optimal family $\Psi^{\prime}$, for $k=0,1$. Note that:
  • For each $z\in\mathbb{D}$, the subfamily $\Psi^{\prime}|_{{\mathbb{RP}}^{5}\times\{z\}}$ of $\Psi^{\prime}$ is a 5-sweepout, as it is constructed from the 555-sweepout $\Psi|_{{\mathbb{RP}}^{5}\times\{z\}}$. In fact, more generally, for each 5-cycle $\Gamma$ in ${\mathbb{RP}}^{5}\times\mathbb{D}$ homologous to ${\mathbb{RP}}^{5}\times\{z\}$, $\Psi^{\prime}|_{\Gamma}$ is a 5-sweepout. Hence, $Z_{0}$ cannot contain any such $\Gamma$, because no family of genus 0 surfaces in $S^{3}$ can be a 5-sweepout (see Proposition 2.28).
  • Let $D$ be the subset
[…]
    Note $\Psi|_{D}$ is exactly the $\mathbb{D}$-family of singular surfaces with genus 1 or 2 that we previously constructed in §1.3.1 using (1.1). Then, the intersection number of $[D]\in H_{2}(Y,\partial Y)$ and $[{\mathbb{RP}}^{5}\times\{z\}]\in H_{5}(Y)$ is 1: See Figure 7. Here, we use $\mathbb{Z}_{2}$-coefficients for all homology and cohomology groups.
From these two facts, and that $Y=Z_{0}\cup Z_{1}$, one can deduce using elementary algebraic topology that there exists some 2-chain $D^{\prime}\subset Z_{1}$, with boundary $\partial D^{\prime}\subset Z_{1}\cap\partial Y$, such that
$[D^{\prime}]=[D]\in H_{2}(Y,\partial Y),$
which in particular implies $[\partial D^{\prime}]=[\partial D]$ in $H_{1}(\partial Y).$ For simplicity, let us here just assume $D^{\prime}=D$.
Figure 7. The blue disc is $D$, and the red circle is any ${\mathbb{RP}}^{5}\times\{z\}$ (this picture is not accurate as the intersection is actually transverse).
Note, by definition, the family $\Psi^{\prime}|_{D}$ consists only of genus one singular surfaces, as $D\subset Z_{1}$. Let us consider the following family $H$: Recall that by assumption the family $\Psi^{\prime}$ is obtained from $\Psi$ via the pinch-off process, so there is a “homotopy” $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times D\to\mathcal{S}_{\leq 2}$ that bridges the subfamily $\Psi|_{\partial D}$ to $\Psi^{\prime}|_{\partial D}$: See Figure 8.
Figure 8. The blue disc on top is $\Psi|_{D}$, the blue disc below is $\Psi^{\prime}|_{D}$, and the red lines represent a cylinder, which is the bridge $H$ joining the boundaries of $\Psi|_{D}$ and $\Psi^{\prime}|_{D}$. Note that the boundary of the blue disc on top is equal to the $S^{1}$-family $\mathcal{T}$ of smooth tori.
Now, we observe that:
  • $\Psi^{\prime}|_{\partial D}$ is a family of genus one singular surfaces.
  • $\Psi|_{\partial D}=\mathcal{T}$ is a family of smooth tori.
  • The homotopy $H$ that bridges $\Psi|_{\partial D}$ to $\Psi^{\prime}|_{\partial D}$ does not increase genus.
Hence, $H|_{\partial D}$ also consists only of genus one singular surfaces. Consequently, if we glue the families $\Psi^{\prime}|_{D}$ and $H$ together along the common boundary part $\Psi^{\prime}|_{\partial D}$, and denote the new family by $\Psi^{\prime}|_{D}+H$, then:
  • The boundary of $\Psi^{\prime}|_{D}+H$ is $\Psi|_{\partial D}=\mathcal{T}$.
  • All members of $\Psi^{\prime}|_{D}+H$ has genus one.
In conclusion, we have shown that $\mathcal{T}$ is, in some sense, “homologically trivial” in the space of all genus one singular surfaces.
However, we will show that, the family $\mathcal{T}$ is actually “homologically non-trivial”, in the sense that it cannot be the boundary of any 2-parameter family of genus one singular surfaces. This is closely related to the fact that the loop $\mathcal{T}$ is homologically non-trivial in the space of all embedded smooth tori in $S^{3}$. This leads to a contradiction.
1.4. Further directions
As mentioned earlier, we expect to construct multiple genus 2 minimal surfaces in 3-spheres with positive Ricci curvature, by investigating the relative structure of the pair $(\mathcal{S}_{\leq 2},\mathcal{S}_{\leq 1})$ more thoroughly.
Theorem 1.3 and the existence theorem for optimal family (Theorem 3.1), which hold all the min-max machinery within, transform the problem of constructing minimal surfaces of specific genus into a topological problem. Namely, the task becomes finding a suitable family $\Phi$ that detects some non-trivial structure of $(\mathcal{S}_{\leq g},\mathcal{S}_{\leq g-1})$, so that it cannot be deformed (via pinch-off processes) to a map into $\mathcal{S}_{\leq g-1}$.
Beyond metrics with positive Ricci curvature, one can also study other classes of metric. The case of bumpy metrics would be a rather different setting, where one can utilize Morse inequality type arguments [LW24], and hence may expect more minimal surfaces compared to the non-bumpy case. In contrast, the case of metrics that are non-bumpy and do not have positive Ricci curvature likely require new techniques, due to the presence of degenerate minimal surfaces. For example, it remains unknown whether every 333-sphere equipped with a metric of non-negative Ricci curvature must contain at least 222 minimal spheres, or at least 111 minimal torus.
1.5. Organization
In Part I, after introducing the necessary terminologies in §2, we prove the two main theorems in §3. Then in Part II, we provide the proof of various results stated in Part I: See the table of contents in page 1.
Acknowledgment
We would like to thank Chi Cheuk Tsang for the helpful discussions, and Boyu Zhang for his contribution to Lemma 7.12. We also want to thank Ryan Budney, André Neves, Daniel Stern, Shmuel Weinberger, and Xin Zhou for their interests.
This material is based upon work supported by the National Science Foundation under Grant No. DMS-1928930, while the first author was in residence at the Simons Laufer Mathematical Sciences Institute (formerly MSRI) in Berkeley, California, during the Fall semester of 2024. The first and the second author were partially supported by the AMS-Simons travel grant.
Part I Main arguments
2. Preliminaries
2.1. Notations
All simplicial complexes and cubical complexes are assumed to be finite. All chains, homology groups, and cohomology groups are assumed to have $\mathbb{Z}_{2}$-coefficients, unless specified otherwise.
  • $\mathbf{I}_{n}(M;\mathbb{Z}_{2})$: the set of integral $n$-dimensional currents in $M$ with $\mathbb{Z}_{2}$-coefficients.
  • $\mathcal{Z}_{n}(M;\mathbb{Z}_{2})\subset\mathbf{I}_{n}(M;\mathbb{Z}_{2})$: the subset that consists of elements $T$ such that $T=\partial Q$ for some $Q\in\mathbf{I}_{n+1}(M;\mathbb{Z}_{2})$ (such $T$ are also called flat $k$-cycles).
  • $\mathcal{Z}_{n}(M;\nu;\mathbb{Z}_{2})$ with $\nu=\mathcal{F},\mathbf{F},\mathbf{M}$: the set $\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$ equipped with the three topologies given corresponding respectively to the flat norm $\mathcal{F}$, the $\mathbf{F}$-metric, and the mass norm $\mathbf{M}$ (see [Pit81] and the survey paper [MN20]). For the flat norm, there are two definitions that, by the isoperimetric inequality, would induce the same topology:
    $\mathcal{F}(T)\mathrel{\mathop{\ordinarycolon}}=\inf\{\mathbf{M}(P)+\mathbf{M}(Q)\mathrel{\mathop{\ordinarycolon}}T=P+\partial Q\}$
    and
    $\mathcal{F}(T)\mathrel{\mathop{\ordinarycolon}}=\inf\{\mathbf{M}(Q)\mathrel{\mathop{\ordinarycolon}}T=\partial Q\}.$
    In this paper, we will use the second definition.
  • $\mathcal{V}_{n}(M)$: the closure, in the varifold weak topology, of the space of $n$-dimensional rectifiable varifolds in $M$.
  • $\mathcal{C}(M)$: the space of Caccioppoli sets in $M$, equipped with the metric induced by the Hausdorff measure of the symmetric difference.
  • $\partial^{*}\Omega$: the reduced boundary of $\Omega\in\mathcal{C}(M)$.
  • $\nu_{\Omega}$: the inward pointing normal of $\Omega\in\mathcal{C}(M)$.
  • $\mathinner{\!\left\lVert V\right\rVert}$: the Radon measure induced on $M$ by $V\in\mathcal{V}_{n}(M)$.
  • $|T|$: the varifold in $\mathcal{V}_{n}(M)$ induced by a current $T\in\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$, or a countable $n$-rectifiable set $T$. In the same spirit, given a map $\Phi$ into $\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$, the associated map into $\mathcal{V}_{n}(M)$ is denoted by $|\Phi|$.
  • $\operatorname{spt}(\cdot)$: the support of a current or a measure.
[…]
  • $[\mathcal{W}]\mathrel{\mathop{\ordinarycolon}}=\{[\Sigma_{i}]\}\subset\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$ for a set $\mathcal{W}$ of varifolds $\{V_{i}\}\subset\mathcal{V}_{n}(M)$, each associated with a countably $n$-rectifiable set $\Sigma_{i}$.
  • $\mathbf{B}^{\nu}_{\varepsilon}(\cdot)$: the open $\varepsilon$-neighborhood of an element or a subset of the space $\mathcal{Z}_{n}(M;\nu;\mathbb{Z}_{2})$.
  • $\mathbf{B}^{\mathbf{F}}_{\varepsilon}(\cdot)$: the open $\varepsilon$-neighborhood of an element or a subset of $\mathcal{V}_{n}(M)$ under the $\mathbf{F}$-metric.
  • $\Gamma^{\infty}(M)$: the set of smooth Riemannian metrics on $M$.
  • $B_{r}(p)$: the open $r$-neighborhood of a point $p$.
  • $B_{r}(\{p_{1},\dots,p_{n}\})\mathrel{\mathop{\ordinarycolon}}=\cup_{i}B_{r}(p_{i})$.
  • $\mathfrak{g}(S)$: the genus of a surface $S$.
  • $b_{1}(\cdot)$: the first Betti number of a space.
  • $\mathrm{in}(S)$: the inside, open region of a surface $S$.
  • $\mathrm{out}(S)$: the outside, open region of a surface $S$.
For an $m$-dimensional cube $I^{m}=\mathbb{R}^{m}\cap\{x\mathrel{\mathop{\ordinarycolon}}0\leq x_{i}\leq 1,i=1,2,\dots,m\}$, we can give it cubical complex structures as follows.
  • $I(1,j)$: the cubical complex on $I\mathrel{\mathop{\ordinarycolon}}=[0,1]$ whose 111-cells and 0-cells are respectively
    $[0,1/3^{j}],[1/3^{j},2/3^{j}],\dots,[1-1/3^{j},1]\;\;\textrm{ and }\;\;[0],[1/3^{j}],[2/3^{j}],\dots,[1].$
  • $I(m,j)$: the cubical complex structure
    $I(m,j)=I(1,j)\otimes\dots\otimes I(1,j)\;\;(m\textrm{ times})$
    on $I^{m}$. $\alpha=\alpha_{1}\otimes\cdots\otimes\alpha_{m}$ is a $q$-cell of $I(m,j)$ if and only if each $\alpha_{i}$ is a cell in $I(1,j)$ and there are exactly $q$ 111-cells. A cell $\beta$ is a face of a cell $\alpha$ if and only if $\beta\subset\alpha$ as sets.
[…]
$X(j^{\prime})\mathrel{\mathop{\ordinarycolon}}=\{\sigma\in I(m,j^{\prime})\mathrel{\mathop{\ordinarycolon}}\sigma\cap|X|\neq\emptyset\}$
of $I(m,j^{\prime})$. For the sake of convenience, we will denote the refined cubical subcomplex by $X$ unless there is ambiguity.
2.2. Surface with singularities
Let $M$ be a closed orientable 3-manifold, equipped with a Riemannian metric (though all notions in §2.2, except those involving the $\mathbf{F}$-metric, will not depend on the metric of $M$). In this paper we will consider surfaces with finite area and finitely many singularities, which possibly include some isolated points too:
Definition 2.1 (Punctate surface).
Let $M$ be a closed orientable 3-manifold. A closed subset $S\subset M$ is a punctate surface in $M$ provided that:
  1. (1)
    The 2-dimensional Hausdorff measure $\mathcal{H}^{2}(S)$ is finite.
  2. (2)
    There exists a finite set $P\subset S$ such that $S\setminus P$ is a smooth, orientable, embedded surface.
  3. (3)
[…]
We denote by $\mathcal{S}(M)$ the set of all punctate surfaces in $M$.
[…]
Definition 2.2 (Punctate set).
Let $S\in\mathcal{S}(M)$. For any finite set $P$ such that $S\setminus P$ is smooth embedded surface, we call $P$ a punctate set for $S$.
Figure 9. This picture shows an element $S\in\mathcal{S}(M)$: It is a closed set, which contains a (black) smooth surface part and also the red points. The smallest possible punctate set for $S$ is given by the red points. Note $S$ has genus 111.
Remark 2.3.
As shown in [CL24, §2], every element $S\in\mathcal{S}(M)$ can be associated uniquely to an integral 2-current $[S]\in\mathbf{I}_{2}(M;\mathbb{Z}_{2})$ with $\partial[S]=0$. In other words, $[S]\in\mathcal{Z}_{2}(M;\mathbb{Z}_{2})$.
Hence, we can borrow the $\mathbf{F}$-metric on $\mathcal{Z}_{2}(M;\mathbb{Z}_{2})$ and impose it on $\mathcal{S}(M)$.
Definition 2.4.
We define $\mathbf{F}\mathrel{\mathop{\ordinarycolon}}\mathcal{S}(M)\times\mathcal{S}(M)\to\mathbb{R}$ by
$\mathbf{F}(S_{1},S_{2})\mathrel{\mathop{\ordinarycolon}}=\mathbf{F}([S_{1}],[S_{2}])=\mathcal{F}([S_{1}],[S_{2}])+\mathbf{F}(|S_{1}|,|S_{2}|).$
One can check that $\mathbf{F}$ is a pseudometric on $\mathcal{S}(M)$ (two members of $\mathcal{S}(M)$ differing just by finitely many points would have zero $\mathbf{F}$-distance).
Now, for any $S\in\mathcal{S}(M)$, let $S_{\operatorname{sing}}$ be the set of non-smooth points of $S$. Then by Sard’s theorem, there exists a set $E\subset(0,\infty)$ of full measure such that for every $r\in E$, $S\setminus B_{r}(S_{\operatorname{sing}})$ is a compact smooth surface with boundary: Recall that $B_{r}(P)$ denotes the union of the open $r$-balls centered at the points in $P$. Moreover, by [CM12, Chapter 1 §2.1 Lemma 1.5], for any $r_{1}>r_{2}>0$ in $E$ we have
$\mathfrak{g}(S\setminus B_{r_{1}}(S_{\operatorname{sing}}))\leq\mathfrak{g}(S\setminus B_{r_{2}}(S_{\operatorname{sing}}))\,.$
Recall $\mathfrak{g}$ denotes the genus. We remark that:
  • The genus of a smooth surface with boundary is defined as the genus of the closed surface obtained by closing up each boundary component using a disc.
  • The genus of a smooth disconnected surface is defined as the sum of the genus of each of its connected components.
  • The genus of a non-orientable surface $\Sigma$ is the number of cross-caps needed to be attached to a two-sphere in order to obtain a surface homeomorphic to $\Sigma$.
The conclusion is, the limit
${}\lim_{r\in E,r\to 0}\mathfrak{g}(S\setminus B_{r}(S_{\operatorname{sing}}))\in\mathbb{N}\cup\{\infty\}$
always exists. Based on this, we may introduce the following definition.
Definition 2.5 (Genus of a punctate surface).
Let $S\in\mathcal{S}(M)$, and let $E$ be the set of $r>0$ such that $S\setminus B_{r}(S_{\operatorname{sing}})$ is a smooth surface with boundary. Then we define the genus of $S$ by
$\mathfrak{g}(S)\mathrel{\mathop{\ordinarycolon}}=\lim_{r\in E,r\to 0}\mathfrak{g}(S\setminus B_{r}(S_{\operatorname{sing}}))\,.$
Moreover, we introduce the following notations:
  • The set of $S\in\mathcal{S}(M)$ with genus $g$ is denoted $\mathcal{S}_{g}(M)$, or simply $\mathcal{S}_{g}$.
  • The set of $S\in\mathcal{S}(M)$ with genus $\leq g$ is denoted $\mathcal{S}_{\leq g}(M)$, or simply $\mathcal{S}_{\leq g}$.
See Figure 9 for an example of a genus one element in $\mathcal{S}(M)$. Moreover, it is easy to check that given any $S\in\mathcal{S}(M)$, for any finite set $P$ that contains $S_{\operatorname{sing}}$ (i.e. $P$ is any punctate set for $S$), if we let $E$ be the set of $r>0$ such that $S\setminus B_{r}(P)$ is a smooth surface with boundary, then
$\lim_{r\in E,r\to 0}\mathfrak{g}(S\setminus B_{r}(P))=\mathfrak{g}(S)\,.$
Let us state a lemma relating genus and the $\mathbf{F}$-pseudometric.
Lemma 2.6.
Let $(M,\mathbf{g})$ be a closed Riemannian 333-manifold, and $\Sigma$ be an orientable smooth embedded surface in $M$. Then there exists some $r>0$ such that for any $S\in\mathcal{S}(M)$ with $\mathbf{F}(S,\Sigma)<r$, we have $\mathfrak{g}(S)\geq\mathfrak{g}(\Sigma)$.
The proof of this lemma is a topological argument. We include the proof in Appendix B.
2.3. Simon-Smith family
Definition 2.7 (Simon-Smith family).
Let $X$ be a cubical or simplicial complex. A map $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ is called a Simon-Smith family with parameter space $X$ if the following hold.
  1. (1)
    The map $x\mapsto\mathcal{H}^{2}(\Phi(x))$ is continuous.
  2. (2)
    (Closedness) The family $\Phi$ is “closed” in the sense that
    $\{(x,p)\in X\times M\mathrel{\mathop{\ordinarycolon}}p\in\Phi(x)\}$
    is a closed subset of $X\times M$.
  3. (3)
    ($C^{\infty}$-continuity away from singularities) For each $x\in X$, we can choose a finite set $P_{\Phi}(x)\subset\Phi(x)$ such that:
    • $\Phi(x)\backslash P_{\Phi}(x)$ is a smooth embedded orientable surface, i.e. $P_{\Phi}(x)$ is a punctate set for $\Phi(x)$.
    • For any $x_{0}\in X$ and within any open set $U\subset\subset M\backslash P_{\Phi}(x_{0})$, $\Phi(x)\to\Phi(x_{0})$ smoothly whenever $x\to x_{0}$.
    • $N(P_{\Phi})\mathrel{\mathop{\ordinarycolon}}=\sup_{x\in X}|P_{\Phi}(x)|<\infty.$
Example 2.8.
Sometimes we need to allow the set $P_{\Phi}(x)$ contain more points than just the non-smooth points of $\Phi(x)$. Consider a family $\Phi\mathrel{\mathop{\ordinarycolon}}[0,1]\to\mathcal{S}(M)$, where $\Phi(t)$ is a smooth torus for $t<1$, $\Phi(1)$ is a smooth sphere, and as $t\to 1$, the handle of the torus becomes smaller and smaller and eventually vanishes at some smooth point $p$ on the sphere $\Phi(1)$. In this case, $P_{\Phi}(1)$ should contain the point $p$.
Remark 2.9.
In previous literature, Simon-Smith families are often allowed to contain certain subsets of measure zero. In this paper, to simplify the notations, we choose to not do that. But we note that all results can modified in obvious ways to allow Simon-Smith families to contain closed subsets of measure zero. See [CL24] for such a modification.
Definition 2.10.
We say that a Simon-Smith family $\Phi$ is of genus $\leq g$ if each member of $\Phi$ has genus $\leq g$.
As shown in [CL24, §2], every Simon-Smith family is $\mathbf{F}$-continuous.
Definition 2.11 (Simon-Smith isotopy).
Two Simon-Smith families $\Phi,\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ are said to be (Simon-Smith) isotopic to each other if there exists a continuous map
$\varphi\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\operatorname{Diff}^{\infty}(M)$
such that:
  1. (1)
    $\varphi(0,x)=\operatorname{Id}$ for all $x\in X$.
  2. (2)
    $\varphi(1,x)(\Phi(x))=\Phi^{\prime}(x)$ for all $x\in X$.
Note that $\textrm{Diff}^{\infty}(M)$ denotes the group of diffeomorphisms on $M$ equipped with the $C^{\infty}$ topology. Moreover, the set of all families isotopic to a Simon-Smith family $\Phi$ is called the (Simon-Smith) isotopy class associated to $\Phi$, and is denoted by $\Lambda(\Phi)$.
Remark 2.12.
  • For any choice of sets $P_{\Phi}(x)$, where $x\in X$, that satisfies the conditions in Definition 2.7 for $\Phi$, and for any isotopy $\varphi$, the sets $P_{\Phi^{\prime}}(x)\mathrel{\mathop{\ordinarycolon}}=\varphi(1,x)\circ P_{\Phi}(x)$ would satisfy Definition 2.7 for the family $\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}=\varphi(1,\cdot)\circ\Phi\in\Lambda(\Phi)$. Consequently, we may assume that $\sup_{x}|P_{\Phi}(x)|<\infty$ is a constant within the isotopy class $\Lambda(\Phi)$.
  • If $\Phi$ is of genus $\leq g$, then so is every $\Phi^{\prime}\in\Lambda(\Phi)$.
2.4. Min-max theorems
In this section we will introduce the min-max theorems. We assume $(M,\mathbf{g})$ is a closed orientable Riemannian 3-manifold.
Definition 2.13 (Simon-Smith width).
Given a isotopy class $\Lambda$ of Simon-Smith families, its width is defined by
$\mathbf{L}(\Lambda)\mathrel{\mathop{\ordinarycolon}}=\inf_{\Phi\in\Lambda}\sup_{x\in X}\mathcal{H}^{2}(\Phi(x))\,.$
Definition 2.14 (Minimizing sequence and min-max sequence).
A sequence $\{\Phi_{i}\}$ in $\Lambda$ is said to be minimizing if
$\lim_{i\to\infty}\sup_{x\in X}\mathcal{H}^{2}(\Phi_{i}(x))=\mathbf{L}(\Lambda)\,.$
If $\{\Phi_{i}\}$ is a minimizing sequence in $\Lambda$ and $\{x_{i}\}\subset X$ satisfies
$\lim_{i\to\infty}\mathcal{H}^{2}(\Phi_{i}(x_{i}))=\mathbf{L}(\Lambda)\,,$
then $\{\Phi_{i}(x_{i})\}$ is called a min-max sequence. For a minimizing sequence $\{\Phi_{i}\}$, we define its critical set $\mathbf{C}(\{\Phi_{i}\})$ to be the set of all subsequential varifold-limit of its min-max sequences:
$\mathbf{C}(\{\Phi_{i}\})\mathrel{\mathop{\ordinarycolon}}=\{V=\lim_{j}|\Phi_{i_{j}}(x_{j})|\mathrel{\mathop{\ordinarycolon}}x_{j}\in X,\|V\|(M)=\mathbf{L}(\Lambda)\}\,.$
And $\{\Phi_{i}\}$ is called pulled-tight if every varifold in $\mathbf{C}(\{\Phi_{i}\})$ is stationary.
Definition 2.15.
Let $(M,{\mathbf{g}})$ be a closed orientable Riemannian 333-manifold and $L>0$.
  1. (1)
    Let $\mathcal{W}_{L}(M,{\mathbf{g}})$ be the set of all varifolds $W\in\mathcal{V}_{2}(M)$, with $\|W\|(M)=L$, of the form
    (2.1) $W=m_{1}|\Gamma_{1}|+\cdots+m_{l}|\Gamma_{l}|\,,$
    where $m_{1},\dots,m_{l}$ are positive integers, and $\Gamma_{1},\dots,\Gamma_{l}$ are disjoint smooth, connected, embedded minimal surfaces in $(M,{\mathbf{g}})$, with $m_{i}$ even if $\Gamma_{i}$ is non-orientable (see [Ket19, Remark 1.4]).
  2. (2)
    Let $\mathcal{W}_{L,\leq g}(M,{\mathbf{g}})$ consist of all elements $W\in\mathcal{W}_{L}(M,{\mathbf{g}})$ (of the form (2.1)) that satisfy
    (2.2) $\sum_{j\in I_{O}}m_{j}\mathfrak{g}(\Gamma_{j})+\frac{1}{2}\sum_{j\in I_{N}}m_{j}(\mathfrak{g}(\Gamma_{j})-1)\leq g\,,$
    where $\Gamma_{j}$ is orientable if $j\in I_{O}$ and non-orientable if $j\in I_{N}$.
For simplicity, when the ambient manifold $(M,{\mathbf{g}})$ is clear, we may omit $(M,{\mathbf{g}})$ in the notations.
Theorem 2.16 (Simon-Smith min-max theorem).
Let $(M,\mathbf{g})$ be a closed orientable Riemannian 333-manifold. Suppose we have a Simon-Smith family $\Phi$ of genus $\leq g$ such that $L\mathrel{\mathop{\ordinarycolon}}=\mathbf{L}(\Lambda(\Phi))>0$. Let $r>0$. Then there exists a minimizing sequence $\{\Phi_{i}\}$ in $\Lambda(\Phi)$ such that:
  1. (1)
    The sequence $\{\Phi_{i}\}$ is pulled-tight and
    $\mathbf{C}(\{\Phi_{i}\})\cap\mathcal{W}_{L,\leq g}\neq\emptyset\,.$
  2. (2)
    There exists some
    $W=m_{1}|\Gamma_{1}|+\cdots+m_{l}|\Gamma_{l}|\in\mathbf{C}(\{\Phi_{i}\})\cap\mathcal{W}_{L,\leq g},$
    under the notation in Definition 2.15 (1), such that:
    1. (a)
      If $\Gamma_{j}$ is unstable and two-sided, then $m_{j}=1$.
    2. (b)
      If $\Gamma_{j}$ is one-sided, then its connected double cover is stable and $m_{j}$ is even.
  3. (3)
    Furthermore, there exists $\eta>0$ such that for all sufficiently large $i$,
    $\mathcal{H}^{2}(\Phi_{i}(x))\geq L-\eta\implies|\Phi_{i}(x)|\in\mathbf{B}^{\mathbf{F}}_{r}(\mathcal{W}_{L,\leq g})\,.$
The above theorem builds upon the foundational results  [Smi82, CD03, DP10, Ket19, MN21, WZ23], and was proven in [CL24]: See Theorem 2.16 therein (we can guarantee that $m_{i}$ is even if $\Gamma_{i}$ is one-sided [Ket19, Remark 1.4]). Note that since $M$ is orientable, a surface is two-sided if and only if it is orientable.
We will also need the theorem below which roughly says if we assume all min-max minimal surfaces are of multiplicity one then there exists some family where in the “large-area region” we can guarantee closeness to the minimal surfaces in the $\mathbf{F}$-norm for currents (instead of just the varifold norm).
Theorem 2.17 ($\mathbf{F}$-closeness).
Let $(M,{\mathbf{g}})$ be a closed orientable Riemannian 333-manifold, and $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ be a Simon-Smith family of genus $\leq g$ such that $L\mathrel{\mathop{\ordinarycolon}}=\mathbf{L}(\Lambda(\Phi))>0$. Suppose that $\mathcal{W}_{L,\leq g}$ consists of only finitely many varifolds, each induced by some multiplicity one embedded minimal surface. Let $r>0$. Then there exists some $\eta>0$ and $\Phi^{\prime}\in\Lambda(\Phi)$ such that
$\mathcal{H}^{2}(\Phi^{\prime}(x))\geq L-\eta\implies[\Phi^{\prime}(x)]\in\mathbf{B}^{\mathbf{F}}_{r}([\mathcal{W}_{L,\leq g}])\,.$
This theorem was proven in [CL24, §3].
2.5. Pinch-off process and interpolation
In this section, we state an important interpolation result, Theorem 2.21, which in some sense significantly improves Theorem 2.17. Namely, Theorem 2.21 says, given a family $\Phi$ that produces a multiplicity one, non-degenerate min-max minimal surface $\Gamma$, we can construct a “better” family $\Phi^{\prime}\in\Lambda(\Phi)$ such that on the large-area region, each element has the same genus as $\Gamma$. From a topological point of view, this is much stronger than Theorem 2.17, in which we could only guarantee that, for some family $\Phi^{\prime\prime}\in\Lambda(\Phi)$, on the large-area region, each element is close to $\Gamma$ in the $\mathbf{F}$-norm for currents, while allowing the possibility it has larger genus than $\Gamma$. In fact, to obtain the desired $\Phi^{\prime}$ from $\Phi^{\prime\prime}$, we need to perform a so-called pinch-off process continuously across the whole region of large area in $\Phi^{\prime\prime}$. This is in fact a genus non-increasing process that ensures at the end, on the large-area region, all elements have the same genus as $\Gamma$. This step is highly non-trivial, and this key feature of Theorem 2.21 will be crucially used in the proof of Theorem 1.3.
Let us introduce the pinch-off process, which first appeared in [CL24]. Essentially, it refers to any deformation process (i.e. a path in $\mathcal{S}(M)$) with finitely many spacetime singular points, which correspond to neck-pinch surgeries and shrinking some components into points: See Figure 2. Note that this definition is purely differential topological.
Definition 2.18.
[…]
If $(t,p)$ is not equal to any $(t_{i},p_{i})$, then:
  1. (1)
    There exists a closed neighborhood $J\subset[a,b]$ around $t$ and a ball $U\subset M$ around $p$ such that $\{\Gamma(t)\}_{t\in J}$ deforms by isotopy within $U$.
If $(t,p)$ is equal to some $(t_{i},p_{i})$, then there exists a closed neighborhood $J\subset[a,b]$ around $t$ and a ball $U\subset M$ around $p$ such that the family $\{\Gamma(t^{\prime})\cap U\}_{t^{\prime}\in J}$ of surfaces is of one of the following types: (Note it is possible that $t=b$.)
  1. (2)
    Surgery : The initial surface deforms by isotopy for $t^{\prime}<t$, but then at time $t$ becomes, up to diffeomorphism, a double cone with $p$ as the cone point (as shown in the second picture in Figure 2). Afterwards, it either remains a double cone or splits into two smooth discs: In either case it is allowed to deform by isotopy.
  2. (3)
    Shrink: For each $t^{\prime}\in J$, $\Gamma(t^{\prime})$ does not intersect $\partial U$. Moreover, for the family $\{\Gamma(t^{\prime})\cap U\}_{t^{\prime}\in J}$, the initial surface deforms by isotopy in $U$ when $t^{\prime}<t$, but then becomes just the point $p$ at time $t$, and this point moves smoothly after time $t.$
Note, all pinch-off processes are Simon-Smith families. Note also that, in (3), since $\Gamma(t^{\prime})$ does not intersect $\partial U$, the part we are shrinking must be some connected components. The reason for allowing the double cone to remain after time $t$ in case (2) will be addressed in Remark 6.3. Now, the lemma below follows directly from the definition of pinch-off process.
Lemma 2.19.
Along a pinch-off process, genus is non-increasing.
Definition 2.20.
Let $\Phi,\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ be Simon-Smith families. Suppose we have a Simon-Smith family $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}(M)$ such that:
  • $H(0,\cdot)=\Phi$ and $H(1,\cdot)=\Phi^{\prime}$.
  • For each $x\in X$, $t\mapsto H(t,x)$ is a pinch-off process.
Then we say $\Phi$ can be deformed via pinch-off processes to $\Phi^{\prime}$.
We now state the crucial interpolation theorem described at the beginning of §2.5.
Theorem 2.21 (Interpolation).
In a closed orientable Riemannian 333-manifold $(M,{\mathbf{g}})$, let $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ be a Simon-Smith family of genus $\leq g$ with $X$ a cubical complex. Assume that $L\mathrel{\mathop{\ordinarycolon}}=\mathbf{L}(\Lambda(\Phi))>0$, and $\mathcal{W}_{L,\leq g}=\{|\Gamma|\}$ where $\Gamma$ is a nondegenerate embedded closed minimal surface. Then for every $r>0$, there exists a Simon-Smith family $\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}_{\leq g}$ with the following properties:
  1. (1)
    There exists an $\eta>0$ such that for all $x\in X$ with $\mathcal{H}^{2}(\Phi^{\prime}(x))\geq L-\eta$,
    $\Phi^{\prime}(x)\in\mathcal{S}_{\mathfrak{g}(\Gamma)}\cap\mathbf{B}^{\mathbf{F}}_{r}(\Gamma).$
  2. (2)
    $\Phi$ can be deformed via pinch-off processes to $\Phi^{\prime}$.
We postpone the proof of Theorem 2.21 to Section 5, which details the pinch-off constructions, and Appendix D.
As a special case, assuming $L=0$ in the previous theorem, we can perform a pinch-off process to deform the whole family to surfaces of genus 0.
Proposition 2.22.
In a closed orientable Riemannian 333-manifold $(M,{\mathbf{g}})$, suppose that $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ is a Simon-Smith family of genus $\leq g$ with $X$ a cubical complex and $\mathbf{L}(\Lambda(\Phi))=0$. Then there exists a Simon-Smith family $\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}_{0}$, and $\Phi$ can be deformed via pinch-off processes to $\Phi^{\prime}$.
The above proposition is a direct corollary of Proposition 5.2; see Section 5.4.
We also need the following proposition on interpolations, which will be proved in §5.
Proposition 2.23.
Given a closed orientable Riemannian 333-manifold $(M,{\mathbf{g}})$, and integers $0\leq g_{0}<g_{1}$, suppose that $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ is a Simon-Smith family of genus $\leq g_{1}$ with $X$ a cubical complex. Let $K$ be an arbitrary compact subset in $X$, and assume $\Phi|_{K}$ maps into $\mathcal{S}_{\leq g_{0}}$.
Then there exists a subcomplex $Z\subset X$, whose interior contains $K$, of the following property: There exists a Simon-Smith family $\Psi^{\prime}\mathrel{\mathop{\ordinarycolon}}Z\to\mathcal{S}_{\leq g_{0}}$, and $\Phi|_{Z}$ can be deformed via pinch-off processes to $\Psi^{\prime}$.
2.6. Almgren-Pitts min-max theory
Let $(M,\mathbf{g})$ be any closed Riemannian manifold.
Definition 2.24.
Let $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{Z}_{2}(M;\mathbb{Z}_{2})$ be an $\mathbf{F}$-continuous map, with $X$ being a finite simplicial complex. The (Almgren-Pitts) homotopy class $\Pi(\Phi)$ associated to $\Phi$ is the set of all $\mathbf{F}$-continuous maps $\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{Z}_{2}(M;\mathbb{Z}_{2})$ homotopic to $\Phi$ in the flat topology.
Note that every finite cubical complex is homeomorphic to a finite simplicial complex and vice versa (see [BP02, §4]). So we could actually allow $X$ to be a finite cubical complex as well.
Definition 2.25.
The width of a Almgren-Pitts homotopy class $\Pi$ of maps with parameter space $X$ is defined by
$\mathbf{L}(\Pi)\mathrel{\mathop{\ordinarycolon}}=\inf_{\Phi\in\Pi}\sup_{x\in X}\mathbf{M}(\Phi(x)).$
Now let us also introduce the $p$-widths. By the Almgren isomorphism theorem [Alm62] (see also [LMN18, §2.5]), when equipped with the flat topology, $\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$ is weakly homotopic equivalent to $\mathbb{R}\mathbb{P}^{\infty}$. Thus we can denote its cohomology ring by $\mathbb{Z}_{2}[\bar{\lambda}]$, where $\bar{\lambda}\in H^{1}(\mathcal{Z}_{n}(M;\mathbb{Z}_{2}),\mathbb{Z}_{2})$ is the generator.
Definition 2.26 ($p$-sweepouts).
Let $\mathcal{P}_{p}$ be the set of all $\mathbf{F}$-continuous maps $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{Z}_{2}(M;\mathbb{Z}_{2})$, where $X$ is a finite simplicial complex, such that the pullback $\Phi^{*}(\bar{\lambda}^{p})\neq 0$. Elements of $\mathcal{P}_{p}$ are called $p$-sweepouts.
Definition 2.27 (Almgren-Pitts $p$-width).
[…]
$\omega_{p}(M,g)\mathrel{\mathop{\ordinarycolon}}=\inf_{\Phi\in\mathcal{P}_{p}}\sup_{x\in\mathrm{dmn}(\Phi)}\mathbf{M}(\Phi(x)).$
As in Simon-Smith min-max theory, we can define the notion of minimizing sequences, critical sets, and min-max sequences for the $p$-widths and for an Almgren-Pitts homotopy class.
For our purpose, the result we need from Almgren-Pitts min-max theory is that the 555-width of the unit 3-sphere $(S^{3},\bar{g})$ is greater than $4\pi$. In fact, building on Marques-Neves’ resolution of the Willmore conjecture [MN14], C. Nurser [Nur16] proved that
$\omega_{5}(S^{3},\bar{g})=2\pi^{2}\,.$
As a side note, not too much is known regarding the $p$-widths of the unit 3-sphere: We know the first four widths are equal to $4\pi$, and by [MN14, Nur16, CL23, Mar23],
$\omega_{5}(S^{3},\bar{g})=\dots=\omega_{8}(S^{3},\bar{g})=2\pi^{2},\;\;\omega_{9}(S^{3},\bar{g}),\dots,\omega_{13}(S^{3},\bar{g})\in(2\pi^{2},8\pi).$
It is unknown whether $\omega_{14}(S^{3},\bar{g})\leq 8\pi$ is true.
We will use the result $\omega_{5}(S^{3},\bar{g})>4\pi$ in the following proposition.
Proposition 2.28.
No Simon-Smith family of genus 0 in $S^{3}$ can be a 555-sweepout.
Note that this statement does not involve the metric of $S^{3}$. This proposition was proven in [CL24] already, but the proof is quite short so we include it here:
Proof.
Let $\Phi$ be a Simon-Smith family of genus 0 in $S^{3}$, and let us equip $S^{3}$ with the unit round metric. Using the Simon-Smith min-max theorem and the multiplicity one theorem, the Simon-Smith width $\mathbf{L}(\Lambda(\Phi))$ is equal to the multiplicity one equator, which has area $4\pi$. Noting $\Lambda(\Phi)\subset\Pi(\Phi)$, as every Simon-Smith isotopy is an Almgren-Pitts homotopy, we know that the Almgren-Pitts width must be at most the Simon-Smith width, i.e.
$\mathbf{L}(\Pi(\Phi))\leq\mathbf{L}(\Lambda(\Phi))=4\pi.$
Recalling $\omega_{5}(S^{3},\bar{g})>4\pi$, we know $\Phi$ cannot be a 5-sweepout.∎
3. Main proofs
In this section, we prove Theorem 1.1 and 1.3. We will take the following steps:
  1. (1)
    We prove Theorem 1.3.
  2. (2)
    We construct a suitable Simon-Smith family $\Psi$ of genus 2$\leq$ 2≤ 2.
  3. (3)
    Finally, we prove Theorem 1.1. We will show that the family $\Psi$ constructed in Step 2 satisfies the assumptions in Theorem 1.3, and then apply Theorem 1.3.
3.1. Proof of Theorem 1.3
First, let us recall the statement of Theorem 1.3, which we will prove in this section (while leaving the proof of some important ingredient to later sections): Let $(M,\bf g)$ be a closed orientable Riemannian 333-manifold with positive Ricci curvature, and $g$ be a positive integer. Suppose there exists some Simon-Smith family
$\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}_{\leq g}$
that cannot be “deformed via pinch-off processes” to become a map into $\mathcal{S}_{\leq g-1}$. Then $(M,\bf g)$ contains some orientable embedded minimal surface of genus $g$ and area at most $\max\mathcal{H}^{2}\circ\Phi$.
Let us prove Theorem 1.3 by contradiction: Suppose that no embedded genus $g$ minimal surface has area at most $\max\mathcal{H}^{2}\circ\Phi$. This implies, by Choi-Schoen compactness [CS85], that for some $\delta_{0}>0$, no orientable embedded genus $g$ minimal surface has area at most $\max\mathcal{H}^{2}\circ\Phi+\delta_{0}$. Thus, using [MN21, Proposition 8.5] and Choi-Schoen compactness [CS85], we can perturb the metric $\mathbf{g}$ to ${\mathbf{g}}^{\prime}$ such that $(M,{\mathbf{g}}^{\prime})$ has the following properties:
  1. (1)
    It has positive Ricci curvature.
  2. (2)
    All embedded minimal surfaces are non-degenerate.
  3. (3)
    The areas of the embedded minimal surfaces are linearly independent over $\mathbb{Z}$.
  4. (4)
    There does not exist an orientable embedded minimal surface of genus $g$ and has area at most $\max\mathcal{H}^{2}_{\mathbf{g}^{\prime}}\circ\Phi+\delta_{0}/2$.
We claim that the following theorem would immediately show that $\Phi$ can be deformed via pinch-off processes to become a map into $\mathcal{S}_{\leq g-1}$, thereby contradicting the assumption of Theorem 1.3.
Theorem 3.1 (Obtaining an optimal family via repetitive min-max).
Let $(M,\mathbf{g})$ be a closed orientable Riemannian 333-manifold of positive Ricci curvature, and $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ be a Simon-Smith family of genus $\leq g$ where $g\geq 1$. Suppose we know $L\mathrel{\mathop{\ordinarycolon}}=\mathbf{L}(\Lambda(\Psi))>0$. Let $r>0$. Suppose also that:
  • There are only finitely many orientable embedded minimal surfaces of genus $\leq g$.
  • Every orientable embedded minimal surface of genus $\leq g-1$ is non-degenerate.
  • Let $\mathcal{A}$ be the set of all orientable embedded minimal surface of genus $\leq g-1$, and non-orientable embedded minimal surface of genus $\leq g+1$. Then for every orientable embedded minimal surface $\Gamma$ of genus $g$, numbers in the set
    $\{\operatorname{area}(\Gamma^{\prime})\mathrel{\mathop{\ordinarycolon}}\Gamma^{\prime}\in\mathcal{A}\}\cup\{\operatorname{area}(\Gamma)\}$
    are linearly independent over $\mathbb{Z}$.
Then there exist
  • a Simon-Smith family $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}_{\leq g}$,
  • distinct, orientable, multiplicity one, embedded minimal surfaces $\Gamma_{1},\Gamma_{2},\dots,\Gamma_{l}$ of genus $\leq g$ and with area at most $L$,
  • subcomplexes $D_{0},D_{1},\dots,D_{l}$ of $X$ (after refinement) that cover $X$,
such that:
  1. (1)
    $H(0,\cdot)=\Phi$, and $t\mapsto H(t,x)$ is a pinch-off process for every $x\in X$.
  2. (2)
    The family $\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}=H(1,\cdot)$ is such that $\Phi^{\prime}|_{D_{i}}$ is of genus $\leq\mathfrak{g}(\Gamma_{i})$ for every $i=1,\dots,l$, and $\Phi^{\prime}|_{D_{0}}$ is of genus 0.
  3. (3)
    For each $i=1,..,l$, there is some smooth function $f_{i}\mathrel{\mathop{\ordinarycolon}}D_{i}\to[0,1]$ such that for every $x\in D_{i}$,
    $H(f_{i}(x),x)\in\mathcal{S}_{\mathfrak{g}(\Gamma_{i})}\cap\mathbf{B}^{\mathbf{F}}_{r}(\Gamma_{i}).$
We refer readers to §1.1 for a heuristic description of this theorem, and also Figure 4 for a schematic diagram for the optimal family. We postpone its proof to §6. Note that, the three bullet points in the assumption of this theorem are much weaker than the defining properties of the perturbed metric ${\mathbf{g}}^{\prime}$ (so this theorem is stronger than what we need), but we choose to present the theorem in this form because of possible future usage. Moreover, in [CL24, Section 3], we essentially already proved a version of this theorem for the case $g=1$ in $S^{3}$.
Let us return to proving Theorem 1.3. We claim that by assuming Theorem 3.1 and applying it to $(M,\mathbf{g}^{\prime})$, $g$, $\Phi$, with $r=1$, we can immediately obtain a contradiction and conclude the proof of Theorem 1.3. Indeed, using the defining properties (1) - (3) of $(M,\mathbf{g}^{\prime})$, the assumptions of Theorem 3.1 are satisfied. So we can apply Theorem 3.1 and obtain some Simon-Smith family $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}_{\leq g}$ that satisfies the properties in Theorem 3.1. However, recall that by property (4) of $(M,\mathbf{g}^{\prime})$, there does not exist an orientable embedded minimal surface of genus $g$ that has area at most $\max\mathcal{H}^{2}_{\mathbf{g}^{\prime}}\circ\Phi$. As a result, all those minimal surfaces $\Gamma_{1},\dots,\Gamma_{l}$ produced in Theorem 3.1 must be of genus $\leq g-1$. Hence, by item (4) of the theorem, it follows that $\Xi|_{D_{i}}$ maps into $\mathcal{S}_{\leq g-1}$ for each $i=1,\dots,l$. Moreover, $\Phi^{\prime}|_{D_{0}}$ has genus 0, so we know actually the whole family $\Phi^{\prime}$ maps into $\mathcal{S}_{\leq g-1}$. This contradicts the assumption of Theorem 1.3 that $\Phi$ cannot be deformed via pinch-off processes to become a map into $\mathcal{S}_{\leq g-1}$. So we have proven Theorem 1.3.
3.2. A 7-parameter family $\Psi$
In this section, we construct a Simon-Smith family $\Psi$ that we will later apply Theorem 1.3 to.
[…]
$x=(x_{1},x_{2},\sqrt{1-x_{1}^{2}-x_{2}^{2}}\cos\alpha,\sqrt{1-x_{1}^{2}-x_{2}^{2}}\sin\alpha)$
[…]
Now, let $\mathbb{D}\subset\mathbb{R}^{2}$ be the closed unit 2-disc. We begin to define a Simon-Smith family $\Psi\mathrel{\mathop{\ordinarycolon}}{\mathbb{RP}}^{5}\times\mathbb{D}\to\mathcal{S}(M)$ of genus 2$\leq$ 2≤ 2 as follows. For each $a=[a_{0}\mathrel{\mathop{\ordinarycolon}}a_{1}\mathrel{\mathop{\ordinarycolon}}\cdots\mathrel{\mathop{\ordinarycolon}}a_{5}]\in{\mathbb{RP}}^{5}$ and $(r,\theta)\in\mathbb{D}$ (under polar coordinate), $\Psi$ sends them to the zero set of the equation
(3.1) $a_{0}+a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3}+a_{4}x_{4}$
$+a_{5}\left[x_{1}x_{2}+\rho(a,x)\left(r\cos(\theta+2\alpha)+(1-r)\cos 3\alpha\right)\right]=0$
[…]
Theorem 3.2.
There exists a suitable function $\rho\mathrel{\mathop{\ordinarycolon}}{\mathbb{RP}}^{5}\times S^{3}\to[0,1]$ such that (3.1) gives a well-defined Simon-Smith family
$\Psi\mathrel{\mathop{\ordinarycolon}}{\mathbb{RP}}^{5}\times\mathbb{D}\to\mathcal{S}(S^{3})$
of genus 2$\leq$ 2≤ 2 such that the following properties hold.
  1. (1)
    $\Psi|_{{\mathbb{RP}}^{4}\times\mathbb{D}}$ is of genus 0, where ${\mathbb{RP}}^{4}\subset{\mathbb{RP}}^{5}$ is the subset defined by $a_{5}=0$.
  2. (2)
    For any $z\in\mathbb{D}$, $\Psi|_{{\mathbb{RP}}^{5}\times\{z\}}$ is a 555-sweepout in the Almgren-Pitts sense.
  3. (3)
    $\Psi|_{{\mathbb{RP}}^{5}\times\partial\mathbb{D}}$ is of genus 1$\leq$ 1≤ 1.
  4. (4)
    If $\gamma\subset{\mathbb{RP}}^{5}\times\partial\mathbb{D}$ is any 111-cycle homologous to the loop
[…]
    within ${\mathbb{RP}}^{5}\times\partial\mathbb{D}$ such that each member of $\Psi|_{\gamma}$ has genus 111, then $\Psi|_{\gamma}$ is “homologically non-trivial” in the following sense: There does not exist a Simon-Smith family $\Phi\mathrel{\mathop{\ordinarycolon}}W\to\mathcal{S}_{1}$, where $W$ is some simplicial 222-complex, such that:
    • $\partial W\cong\gamma$.
    • $\Phi|_{\partial W}=\Psi|_{\gamma}$.
    • There is a consistent choice, for every $x\in W$, of inside and outside regions for $\Phi(x)$.
We postpone the proof of this theorem to §7. Readers may refer to §1.3.1 for a rough description of the family $\Psi$. Note that, every element of $\mathcal{S}(S^{3})$ is by definition separating (Definition 2.7), but there need not be (in this case, there actually would not be) a consistent way of choosing inside and outside regions for $\Psi(x)$, for every $x\in{\mathbb{RP}}^{5}\times\mathbb{D}$. Namely, it is possible that for some loop in ${\mathbb{RP}}^{5}\times\mathbb{D}$, when we travel around it once, the two sides of the associated surface are interchanged.
3.3. Proof of Theorem 1.1
In this section, we view $\Psi$ as a family in any given 3-sphere $(S^{3},\mathbf{g}_{0})$ of positive Ricci curvature. We would like to apply Theorem 1.3 to the family $\Psi$ and construct a genus 222 embedded minimal surface. Hence, it suffices to show that the family $\Psi\subset\mathcal{S}_{\leq 2}$ cannot be deformed via pinch-off processes to become a map into $\mathcal{S}_{\leq 1}$ (recall Definition 2.20).
[…]
In fact, we claim that if we let $\lambda$ be the only non-trivial element in $H^{1}(Y)\cong H^{1}({\mathbb{RP}}^{5})$, then $PD([O_{1}\times\mathbb{D}])=\lambda^{5}$. Indeed, since $H^{5}(Y)=\mathbb{Z}_{2}$, our claim follows from the observation that both $\lambda^{5}$ and $PD([O_{1}\times\mathbb{D}])$ pairs non-trivially with the 5-cycle ${\mathbb{RP}}^{5}\times O_{2}$.
Summarizing, we have justified that:
Lemma 3.3.
We have the following facts.
  1. (1)
    $\lambda^{5}([{\mathbb{RP}}^{5}\times O_{2}])=1.$
  2. (2)
    The cap product $[Y]\frown\lambda^{5}\in H_{2}(Y,\partial Y)$ is given by $[O_{1}\times\mathbb{D}]$.
Readers may refer to [Hat02, p.240] for this relative version of cup product.
Now, we begin to prove by contradiction that the family $\Psi$ cannot be deformed via pinch-off processes to become a map into $\mathcal{S}_{\leq 1}$. Suppose, instead, we have a Simon-Smith family $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times Y\to\mathcal{S}_{\leq 2}$ such that:
  • $H(0,\cdot)=\Psi$.
  • $\Psi^{\prime}\mathrel{\mathop{\ordinarycolon}}=H(1,\cdot)$ maps into $\mathcal{S}_{\leq 1}$.
  • For each $y\in Y$, $t\mapsto H(t,y)$ is a pinch-off process.
For $k=0,1$, let $Z_{k}\mathrel{\mathop{\ordinarycolon}}=\{y\in Y\mathrel{\mathop{\ordinarycolon}}\mathfrak{g}(\Psi^{\prime}(y))=k\}$. Note $Z_{1}\subset X$ must be open while $Z_{0}$ must be compact, by the definition of Simon-Smith family. We will prove the following proposition in a moment.
Proposition 3.4.
There exists some simplicial 222-chain $D\subset Z_{1}$ such that:
  • $\partial D\subset{\mathbb{RP}}^{5}\times\partial\mathbb{D}=\partial Y$.
  • As elements of $H_{1}({\mathbb{RP}}^{5}\times\partial\mathbb{D})$, $[\partial D]=[O_{1}\times\partial\mathbb{D}].$
We claim that one can deduce a contradiction using this proposition.
Lemma 3.5.
Proposition 3.4 implies a contradiction.
Proof.
By the second bullet point in Proposition 3.4 and item (4) of Theorem 3.2, $\Psi|_{\partial D}$ is “homologically non-trivial” in the sense that there does not exist a Simon-Smith family $\Phi\mathrel{\mathop{\ordinarycolon}}W\to\mathcal{S}_{1}$, where $W$ is some simplicial 222-subcomplex such that:
  • $\partial W\cong\partial D$.
  • $\Phi|_{\partial W}=\Psi|_{\partial D}$.
  • There is a consistent way to choose, for each $x\in W$, inside and outside regions for $\Phi(x)$.
On the other hand, if we consider the 2-chain
$E\mathrel{\mathop{\ordinarycolon}}=\{1\}\times D+[0,1]\times\partial D$
in $[0,1]\times Y$, then:
  • $\partial E=\{0\}\times\partial D.$
  • $H|_{\partial E}=\Psi|_{\partial D}$ (we are treating $\{0\}\times\partial D$ and $\partial D$ as the same).
  • By the definition of $D^{\prime}$ we know $H|_{\{1\}\times D^{\prime}}$ is of genus 1. Then, by Theorem 3.2 (3) ($\Psi|_{{\mathbb{RP}}^{5}\times\partial\mathbb{D}}$ is of genus 1), and that pinch-off process is genus non-increasing, we know $H|_{[0,1]\times\partial D}$ is also of genus one. So $H|_{E}$ would map into $\mathcal{S}_{1}$.
Furthermore, we claim that:
Lemma 3.6.
There is a consistent way to choose, for each $x\in E$, inside and outside regions for $H(x)$.
Proof.
Note, each member of $H|_{E}$ is obtained from some member of $\Psi|_{D}$ via pinch-off processes. As a result, it suffices to check that there is a consistent way to choose, for each $x\in D$, inside and outside regions for $\Psi(x)$.
Now, since $\mathfrak{g}(H(x))=1$ for every $x\in E$ as we already checked, we know $\mathfrak{g}(\Psi(x))=1$ or 2 for every $x\in D$ as pinch-off process is genus non-increasing. Hence, $D\subset Y$ actually lies within $({\mathbb{RP}}^{5}\backslash{\mathbb{RP}}^{4})\times\mathbb{D}$, where ${\mathbb{RP}}^{4}$ is defined by $a_{5}=0$ (this is because when $a_{5}=0$, the equation (3.1) could only give genus 0 elements). As a result, we can assume on $D$ that $a_{5}=1$ in the defining equations for $\Psi|_{D}$. Therefore, there in fact exists an $D$-family of functions
$\{f_{x}\mathrel{\mathop{\ordinarycolon}}{\mathbb{RP}}^{5}\times\mathbb{D}\to\mathbb{R}\}_{x\in D}$
such that $f_{x}(a,(r,\theta))=0$ defines the surface $\Psi|_{D}(x)$. Crucially, such a continuous family of $f_{x}$ is well-define not just up to a sign. Hence, we can declare that the inside region of $\Psi(x)$ is wherever $f_{x}<0$ and the outside region of $\Psi(x)$ is wherever $f_{x}>0$. This shows there is a consistent way to choose, for each $x\in D$, inside and outside regions for $\Psi(x)$. ∎
In conclusion, we have:
  • $\partial E\cong\partial D.$
  • $H|_{\partial E}=\Psi|_{\partial D}$.
  • $H|_{E}$ would map into $\mathcal{S}_{1}$.
  • There is a consistent way to choose, for each $x\in E$, inside and outside regions for $H(x)$.
This contradicts the first paragraph in the proof of Lemma 3.5, regarding that $\Psi|_{\partial D}$ is “homologically non-trivial”. This finishes the proof that, if Proposition 3.4 is true then contradiction arises. ∎
Hence, to complete our proof by contradiction (that $\Psi$ cannot be deformed via pinch-off processes to become a map into $\mathcal{S}_{\leq 1}$), we just need to prove Proposition 3.4.
Proof of Proposition 3.4..
From Lemma 3.3
$\lambda^{5}([{\mathbb{RP}}^{5}\times O_{2}])=1,\quad[Y]\frown\lambda^{5}=[O_{1}\times\mathbb{D}].$
To prove Proposition 3.4, we claim that it suffices to prove the following lemma.
Lemma 3.7.
There exists some
$\theta\in H_{2}(Z_{1},Z_{1}\cap\partial Y)$
such that the pushforward of $\theta$ under the inclusion
$(Z_{1},Z_{1}\cap\partial Y)\hookrightarrow(Y,\partial Y)$
is equal to $[O_{1}\times\mathbb{D}].$
Indeed, once we have this lemma, we write $\theta=[D]$ for some 2-subcomplex $D\subset Z_{1}$ where $\partial D\subset Z_{1}\cap\partial Y$, and $[D]=[O_{1}\times\mathbb{D}]$ as elements of $H_{2}(Y,\partial Y)$. As a result, as elements of $H_{1}({\mathbb{RP}}^{5}\times\partial\mathbb{D})$, $[\partial D]=[O_{1}\times\partial\mathbb{D}].$ This proves Proposition 3.4.
So now we just need to prove Lemma 3.7. First, let $\widetilde{Z}_{0}$ be the subcomplex of $Y$ one obtains by applying Proposition 2.23 to the family $\Psi^{\prime}$ and the compact set $Z_{0}$ (compactness of $Z_{0}$ follows immediately from the definition of Simon-Smith family). This means, $Z_{0}\subset\mathrm{int}(\widetilde{Z}_{0})$ (note $\mathrm{int}(\cdot)$ denotes the interior region), and $\Psi^{\prime}|_{\widetilde{Z}_{0}}$ can be deformed via pinch-off processes to become some map into $\mathcal{S}_{0}$.
Lemma 3.8.
The pullback $\lambda^{5}|_{\widetilde{Z}_{0}}\in H^{5}(\widetilde{Z}_{0})$ is trivial.
Proof.
Suppose the otherwise. Then by universal coefficient theorem, there exists some 5-cycle $\Gamma\subset\widetilde{Z}_{0}$ such that $[\Gamma]\in H_{5}(Y)$ satisfies $\lambda^{5}([\Gamma])=1$. Then, noting $H_{5}(Y)=\mathbb{Z}_{2}$ and $\lambda^{5}([{\mathbb{RP}}^{5}\times O_{2}])=1$, we have $[\Gamma]=[{\mathbb{RP}}^{5}\times O_{2}]$. Therefore, $\lambda^{5}([\Gamma])=1$ and $\Psi|_{\Gamma}$ is also a 555-sweepout. And since $\Psi^{\prime}|_{\Gamma}$ is obtained from $\Psi|_{\Gamma}$ via pinch-off processes, $\Psi^{\prime}|_{\Gamma}$ is also a 5-sweepout. Thus, by how we defined $\widetilde{Z}_{0}$ (as a product of Proposition 2.23), we know $\Psi^{\prime}|_{\Gamma}$ can be deformed via pinch-off processes to become some Simon-Smith family of genus 0, which also must be a 5-sweepout. This contradicts Proposition 2.28. ∎
We are now ready to prove Lemma 3.7. Let us collect the relevant facts:
  • $[Y]\frown\lambda^{5}=[O_{1}\times\mathbb{D}]$
  • $Y=Z_{0}\cup Z_{1},$ and so $Y=\mathrm{int}(\widetilde{Z}_{0})\cup Z_{1}$.
  • $\lambda^{5}|_{\mathrm{int}(\widetilde{Z}_{0})}=0$ by the above lemma.
With these facts, we claim that Lemma 3.7 follows immediately from the following elementary algebraic topological lemma. Readers may refer to [Hat02, p.240] for the following relative version of cup product.
Lemma 3.9.
In this lemma, we fix an arbitrary commutative ring for the coefficients of homology and cohomology. Let $X$ be a (not necessarily finite) simplicial complex, and $A,B\subset X$ be open subsets. Let $p\geq q\geq 0$. Let $C$ be a union of $p$-subcomplexes in $X$ with $\partial C\subset A\cup B$ (so that $[C]\in H_{p}(X,A\cup B)$), and $\omega\in H^{q}(X,A)$. Suppose that:
  • $C$ is a union of two open subsets $W_{1},W_{2}\subset C$.
  • $W_{2}\cap A=\emptyset.$
  • The pullback of $\omega$ under the inclusion $(W_{1},W_{1}\cap A)\hookrightarrow(X,A)$ is zero.
Then there exists some $\theta\in H_{p-q}(W_{2},W_{2}\cap B)$ such that the pushforward of $\theta$ under the inclusion $(W_{2},W_{2}\cap B)\hookrightarrow(X,B)$ is equal to
$[C]\frown\omega\in H_{p-q}(X,B).$
(Note that this lemma is more general than we need here: We only use the case $A=\emptyset$. But we choose to state this general version for possible future use.) We postpone the proof of Lemma 3.9 to Appendix E.3: It just uses standard algebraic topology. Using this lemma, we can easily finish the proof of Lemma 3.7. Namely, we define the following:
  • $X\mathrel{\mathop{\ordinarycolon}}=Y$,
  • $A\mathrel{\mathop{\ordinarycolon}}=\emptyset$ and $B\mathrel{\mathop{\ordinarycolon}}=\partial Y$ (note that even though here $\partial Y$ is not an open set, it is so up to homotopy, so this is not an issue),
  • $C=Y$,
  • $p\mathrel{\mathop{\ordinarycolon}}=7$, $q\mathrel{\mathop{\ordinarycolon}}=5$,
  • $\omega\mathrel{\mathop{\ordinarycolon}}=\lambda^{5}$
  • $W_{1}\mathrel{\mathop{\ordinarycolon}}=\textrm{int}(\widetilde{Z}_{0})$ and $W_{2}\mathrel{\mathop{\ordinarycolon}}=Z_{1}$, which are both open.
Note by the paragraph just before Lemma 3.9, we know that the assumptions of the lemma would be satisfied. Hence, Lemma 3.9 produces some class
$\theta\in H_{2}(Z_{1},Z_{1}\cap\partial Y)$
such that the pushforward of $\theta$ under the inclusion
$(Z_{1},Z_{1}\cap\partial Y)\hookrightarrow(Y,\partial Y)$
is equal to $[Y]\frown\lambda^{5}=[O_{1}\times\mathbb{D}].$ Now $\theta$ satisfies the conditions in Lemma 3.7. This finishes the proof of Lemma 3.7. So we have proven Proposition 3.4. ∎
Hence, as pointed out by Lemma 3.5, Proposition 3.4 gives a contradiction. This means the family $\Psi$ in fact cannot be deformed via pinch-off processes to become a map into $\mathcal{S}_{\leq 1}$, and then Theorem 1.3 immediately tells that $(M,\mathbf{g}_{0})$ has an orientable embedded minimal surface of genus 2. This finishes the proof of Theorem 1.1.
Part II Technical ingredients
4. Existence of homologically independent simple short loops
Let $(M,\mathbf{g})$ be an orientable closed Riemannian 333-manifold. We use $\Gamma$ and $\Sigma$ denote smooth embedded orientable closed surfaces in $M$, and $S$ denote a punctate surface in $\mathcal{S}(M)$. The main goal of these sections are to locate homologically independent simple short loops in a smooth surface and small balls in a punctate surface.
Recall that the systole of a surface $\Gamma$, denoted by $\operatorname{sys}\pi_{1}(\Gamma)$, is the length of the least length of a noncontractible loop in $\Gamma$. In [Gro96], Gromov introduced the notion of $\mathbb{Z}_{2}$ homological systole of $\Gamma$, denoted by $\operatorname{sys}H_{1}(\Gamma;\mathbb{Z}_{2})$, defined as the least length of a loop in $\Gamma$, which represents a non-trivial class in $H_{1}(\Gamma;\mathbb{Z}_{2})$. The remarkable systolic inequality by Loewner and Gromov states that if the area of $\Gamma$ is small, then the $\mathbb{Z}_{2}$ homological systole is also small; see Theorem 4.4. By the standard minimization process in the calculus of variations, the systole can be realized by a simple loops—specifically, simple closed geodesic.
In general, if the genus $\mathfrak{g}(\Gamma)>1$ and the area $\mathcal{H}^{2}(\Gamma)$ is small, one expects to find $\mathfrak{g}(\Gamma)$ homologically independent simple short loops in $\Gamma$. This is closely related to Bers’ constant for hyperbolic surfaces [Ber74, Ber85]. A natural idea is to perform a neck-pinch surgery along the detected loop in $M$ realizing the systole and then employ an inductive argument; see Proposition 4.1 in §4.1
In §4.2, we show that in $(M,\mathbf{g})$, for $g_{1}\geq g_{0}\in\mathbb{N}$, a smooth embedded orientable closed surface $\Gamma$ of genus $g_{1}$ must contain $g_{1}-g_{0}$ disjoint simple short loops which are homologically independent, provided that it is close to a smooth embedded orientable closed surface $\Sigma$ of genus $g_{0}$ with respect to the $\mathbf{F}$-metric.
If $\mathfrak{g}(\Gamma)>\mathfrak{g}(\Sigma)$, and $\Gamma$ lies in a tubular neighborhood $N_{\eta}(\Sigma)$ of $\Sigma$, then the map
$f\mathrel{\mathop{\ordinarycolon}}\Gamma\hookrightarrow N_{\eta}(\Sigma)\to\Sigma$
induces a non-injective map
$f_{*}\mathrel{\mathop{\ordinarycolon}}\pi_{1}(\Gamma)\to\pi_{1}(\Sigma)\,.$
It follows from Gabai’s resolution of the so called Simple Loop Conjecture [Gab85, Theorem 2.1] that there exists a non-contractible simple closed curve $\alpha\subset\Gamma$ such that $f|_{\alpha}$ is homotopically trivial. In particular, together with the Dehn’s lemma, $\alpha$ bounds a compressing disk in $M$. However, this topological result doesn’t give us any control on the length of the loop, and our result can be regarded as a quantitative and homological generalization of this simple loop conjecture.
Heuristically, if $\Gamma$ and $\Sigma$ are sufficiently close with respect to the $\mathbf{F}$-metric, a large portion $\Gamma^{\prime}$ of $\Gamma$ is homeomorphic to $\Sigma$ removing finitely many disks. Hence, $\Gamma\setminus\Gamma^{\prime}$ has small area, and one may appeal to Proposition 4.1 to detect multiple simple short loops which lie outside $\Gamma^{\prime}$, provided that its boundary also has small length. The technical details will be presented in the proof of Proposition 4.7, where we also need to deal with the case where some thin parts, such as filigrees, of $\Gamma$ is not inside a tubular neighborhood of $\Sigma$.
Next, in § 4.3, we extend this result to a genus $g_{1}$ punctate surface. However, as illustrated in Appendix C, one cannot expect these loops to have small lengths; instead, they are contained in 333-balls of small radius, i.e., small balls.
4.1. Simple short loops in surfaces with small area
If the surface $\Gamma$ is an embedded surface, then in general, the simple short loop realizing the systole might be knotted in $M$, and in this case, there is no compressing disk to perform the surgery in the ambient manifold. Therefore, in this subsection, $\Gamma$ is always regarded as an intrinsic surface so that we can carry out the inductive argument.
Proposition 4.1 (Existence of homologically independent simple short loops I).
For any $g\in\mathbb{N}$ and $\varepsilon>0$, there exists $\delta=\delta(g,\varepsilon)\in(0,\varepsilon)$ with the following property.
For any smooth orientable closed surface $\Gamma$ containing a finite set of disjoint closed disks $\{\overline{D_{i}}\}^{n}_{i=1}$ in $\Gamma$ with piecewise smooth boundary satisfying;
  1. (i)
    $\mathfrak{g}(\Gamma)\leq g$,
  2. (ii)
    $\mathcal{H}^{2}(\Gamma)<\delta$,
  3. (iii)
    $\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D_{i})<\delta$,
there exist $\mathfrak{g}(\Gamma)$ disjoint simple loops $\{c_{j}\}^{\mathfrak{g}(\Gamma)}_{j=1}$ in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$ such that:
  1. (1)
    Each $c_{j}$ has length $\mathcal{H}^{1}(c_{j})\leq\varepsilon$.
  2. (2)
    $\{[c_{j}]\}^{\mathfrak{g}(\Gamma)}_{j=1}$ is linearly independent in $H_{1}(\Gamma;\mathbb{Z}_{2})$.
In particular, $\Gamma\setminus\bigcup^{\mathfrak{g}(\Gamma)}_{j=1}c_{j}$ is a genus 0 surface with boundary.
Remark 4.2.
In this proposition, as well as Proposition 4.7, we specify a finite set of disks that serves as forbidden regions for the loops. This technical assumption allows to apply the propositions to punctate surfaces by removing neighborhoods of their punctate sets.
Remark 4.3.
By Corollary A.2, in $(M,\mathbf{g})$, small $\mathcal{H}^{2}(\Gamma)$ is equivalent to the assumption that $\mathbf{F}(\Gamma,\emptyset)$ is small. Hence, if we regard $\emptyset$ as a degenerate genus 0 surface, then the condition $\mathcal{H}^{2}(\Gamma)<\delta$ can be replaced by $\mathbf{F}(\Gamma,\emptyset)<\delta$ and the proposition is a special case of Proposition 4.7 with $\Sigma=\emptyset$.
Note that without the forbidden regions condition, this follows immediately from a quantitative estimate on the lengths of pants decomposition [BPS12, Proposition 6.3]. In the following, we prove it by applying the systolic inequality inductively and a technical result that allows loops to bypass the forbidden regions.
Theorem 4.4 (Loewner torus inequality [Pu52, Section 3], Gromov’s homological systolic inequality [Gro96, Section 2.C]).
Given a smooth orientable closed surface $\Gamma$ of positive genus, there exists a universal constant $C>0$ such that
$\operatorname{sys}H_{1}(\Gamma;\mathbb{Z}_{2})\mathrel{\mathop{\ordinarycolon}}=\inf\left\{\mathcal{H}^{1}(c)\mathrel{\mathop{\ordinarycolon}}[c]\neq 0\in H_{1}(\Gamma;\mathbb{Z}_{2})\right\}\leq C\sqrt{\mathcal{H}^{2}(\Gamma)}\,.$
In particular, there exists a simple closed geodesic $c\in\Gamma$ such that $[c]\neq 0\in H_{1}(\Gamma;\mathbb{Z}_{2})$ and $\mathcal{H}^{1}(c)\leq C\sqrt{\mathcal{H}^{2}(\Gamma)}$.
Remark 4.5.
In [Gro96], Gromov proved an optimal estimate that there exists a universal constant $\widetilde{C}>0$ such that for any $\Gamma$ of genus $g\geq 2$,
$\operatorname{sys}H_{1}(\Gamma;\mathbb{Z}_{2})\leq\widetilde{C}\frac{\log g}{\sqrt{g}}\sqrt{\mathcal{H}^{2}(\Gamma)}\,.$
This optimal result is not needed in our paper.
Lemma 4.6 (Bypassing the forbidden regions).
Let $\alpha$ be a positive real number. Suppose that in a smooth orientable closed surface $\Gamma$, there exists a simple loop $c\subset\Gamma$ and a finite set of disjoint closed disks $\{\overline{D_{i}}\}^{n}_{i=1}$ in $\Gamma$ with piecewise smooth boundary satisfying
$\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D_{i})<\alpha\,.$
Then there exists a finite union of disjoint simple loops $\widetilde{c}$ in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$ such that
$[\widetilde{c}]=[c]\in H_{1}(\Gamma;\mathbb{Z}_{2})\,,\quad\mathcal{H}^{1}(\widetilde{c})<\mathcal{H}^{1}(c)+\alpha\,.$
Proof.
Since each disk has piecewise smooth boundary, by Sard’s theorem, for each disk $D_{i}$, we can choose a larger disk $D^{\prime}_{i}\supset\overline{D_{i}}$ such that $\{\overline{D^{\prime}_{i}}\}^{n}_{i=1}$ are also pairwise disjoint, $c$ intersects $\partial D^{\prime}_{i}$ transversally, and
$\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D^{\prime}_{i})<\alpha\,.$
Since each $D^{\prime}_{i}$ is contractible, $c$ separates $D^{\prime}_{i}$ into two sets $E^{\prime}_{i,1}$ and $E^{\prime}_{i,2}$ such that
$\partial E^{\prime}_{i,1}\cap D^{\prime}_{i}=\partial E^{\prime}_{i,2}\cap D^{\prime}_{i}=c\cap D^{\prime}_{i}\,.$
We define (see figure 10)
$\widetilde{c}=\left(c\setminus\bigcup^{n}_{i=1}(\partial E^{\prime}_{i,1}\cap D^{\prime}_{i})\right)\cup\bigcup^{n}_{i=1}(\partial D^{\prime}_{i}\cap\partial E^{\prime}_{i,1})\,.$
Figure 10. Constructing $\widetilde{c}$ from $c$
Clearly, by the transversality, $\widetilde{c}$ is a union of disjoint simple loops in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$ satisfying
$[\widetilde{c}]=[c]+[\partial D^{\prime}_{i}]=[c]\in H_{1}(\Gamma;\mathbb{Z}_{2})\,,$
and
$\mathcal{H}^{1}(\widetilde{c})\leq\mathcal{H}^{1}(c)+\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D^{\prime}_{i}\cap\partial E^{\prime}_{i,1})\leq\mathcal{H}^{1}(c)+\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D^{\prime}_{i})<\mathcal{H}^{1}(c)+\alpha\,.$
Proof of Proposition 4.1.
If $\mathfrak{g}(\Gamma)=0$, the proposition holds trivially. In the following, we fix $g\in\mathbb{N}^{+}$ and only consider the case where $\mathfrak{g}(\Gamma)\in\mathbb{N}^{+}$.
It suffices to prove that for any $g^{\prime}\in\{1,2,\dots,g\}$ and $\varepsilon>0$, there exists $\delta^{\prime}(g,g^{\prime},\varepsilon)\in(0,\varepsilon)$ with the following property: For any smooth orientable closed surface $\Gamma$ containing disjoint closed disks $\{\overline{D_{i}}\}^{n}_{i=1}$ with piecewise smooth boundary, satisfying:
  1. (i)’
    $g^{\prime}\leq\mathfrak{g}(\Gamma)\leq g$,
  2. (ii)’
    $\mathcal{H}^{2}(\Gamma)<\delta^{\prime}(g,g^{\prime},\varepsilon)$,
  3. (iii)’
    $\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D_{i})<\delta^{\prime}(g,g^{\prime},\varepsilon)$,
there exist $g^{\prime}$ disjoint simple loops $\{c_{j}\}^{g^{\prime}}_{j=1}$ in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$ such that:
  1. (1)’
    Each $c_{j}$ has length $\mathcal{H}^{1}(c_{j})\leq\varepsilon$.
  2. (2)’
    $\{[c_{j}]\}^{g^{\prime}}_{j=1}$ is linearly independent in $H_{1}(\Gamma;\mathbb{Z}_{2})$.
We prove this existence result inductively on $g^{\prime}$.
For $g^{\prime}=1$, we define
$\delta^{\prime}(g,1,\varepsilon)\mathrel{\mathop{\ordinarycolon}}=\min\left(\frac{\varepsilon}{2},\frac{\varepsilon^{2}}{4C^{2}}\right)\,,$
where $C$ is the universal constant given in Theorem 4.4. By Theorem 4.4, for any $\Gamma$ and $\{\overline{D_{i}}\}^{n}_{i=1}$ satisfying (i)’,  (ii)’ and  (iii)’ there exists a loop $c^{\prime}_{1}\subset\Gamma$ such that
$\mathcal{H}^{1}(c^{\prime}_{1})\leq C\sqrt{\mathcal{H}^{2}(\Gamma)}\leq\frac{\varepsilon}{2}\,,\quad[c^{\prime}_{1}]\neq 0\in H_{1}(\Gamma;\mathbb{Z}_{2})\,.$
By Lemma 4.6, $c^{\prime}_{1}$ is homologous to a finite union of disjoint simple loops $\widetilde{c}_{1}$ lying outside $\{\overline{D_{i}}\}^{n}_{i=1}$, and
$\mathcal{H}^{1}(\widetilde{c}_{1})<\delta^{\prime}(g,1,\varepsilon)+\mathcal{H}^{1}(c^{\prime}_{1})<\varepsilon\,,\quad[\widetilde{c}_{1}]=[c^{\prime}_{1}]\neq 0\in H_{1}(\Gamma;\mathbb{Z}_{2})\,.$
Hence, there exists at least one loop in $\widetilde{c}_{1}$, denoted by $c_{1}$, such that
$\mathcal{H}^{1}(c_{1})<\varepsilon\,,\quad[c_{1}]\neq 0\in H_{1}(\Gamma;\mathbb{Z}_{2})\,.$
Inductively, assume the existence of $\delta^{\prime}(g,g^{\prime},\varepsilon)$ for any $\varepsilon$ and $g^{\prime}=k\in\{1,2,\dots,g-1\}$. We define
$\delta^{\prime}(g,k+1,\varepsilon)\mathrel{\mathop{\ordinarycolon}}=\min\left(\frac{\varepsilon}{4},\frac{\varepsilon^{2}}{8C^{2}},\delta^{\prime}\left(g,k,\frac{\varepsilon}{8k}\right)\right)\,.$
By the inductive assumption, for any $\Gamma$ and $\{\overline{D_{i}}\}^{n}_{i=1}$ satisfying  (i)’,  (ii)’ and  (iii)’, there exist $k$ disjoint simple loops $\{c_{j}\}^{k}_{j=1}$ in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$ satisfying  (1)’ and (2)’ with $\varepsilon$ replaced by $\frac{\varepsilon}{8k}$.
[…]
(4.1) $\mathcal{H}^{2}(\Gamma^{\prime})\leq\mathcal{H}^{2}(\Gamma)+\frac{\varepsilon^{2}}{8C^{2}}<\delta^{\prime}(g,k+1,\varepsilon)+\frac{\varepsilon^{2}}{8C^{2}}\leq\frac{\varepsilon^{2}}{4C^{2}}\,.$
[…]
(4.2) $\sum^{n+2k}_{i=1}\mathcal{H}^{1}(\partial D_{i})=\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D_{i})+2k\frac{\varepsilon}{8k}<\delta^{\prime}(g,k+1,\varepsilon)+\frac{\varepsilon}{4}\leq\frac{\varepsilon}{2}\,.$
Applying Theorem 4.4 again to $\Gamma^{\prime}$, together with  (4.1), we obtain a loop $c^{\prime}_{k+1}\subset\Gamma^{\prime}$ with length
$\mathcal{H}^{1}(c^{\prime}_{k+1})\leq C\sqrt{\mathcal{H}^{2}(\Gamma^{\prime})}\leq\frac{\varepsilon}{2}\,.$
By Lemma 4.6 with (4.2), $c^{\prime}_{k+1}$ is homologous to a finite union of loops $\widetilde{c}_{k+1}\subset\Gamma^{\prime}\setminus\bigcup^{n+2k}_{i=1}\overline{D_{i}}$ with
$\mathcal{H}^{1}(\widetilde{c}_{k+1})<\mathcal{H}^{1}(c^{\prime}_{k+1})+\frac{\varepsilon}{2}\leq\varepsilon\,,\quad[\widetilde{c}_{k+1}]=[c^{\prime}_{k+1}]\neq 0\in H_{1}(\Gamma^{\prime};\mathbb{Z}_{2})\,.$
Hence, there exists a simple loop $c_{k+1}$ in $\widetilde{c}_{k+1}$ with $[c_{k+1}]\neq 0\in H_{1}(\Gamma^{\prime};\mathbb{Z}_{2})$, which implies that $[c_{k+1}]$ is linearly independent of $[c_{1}],[c_{2}],\dots,[c_{k}]$ in $H_{1}(\Gamma;\mathbb{Z}_{2})$. Notice that
$\Gamma^{\prime}\setminus\bigcup^{n+2k}_{i=1}\overline{D_{i}}=\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}\,,$
so $c_{k+1}$ is also a simple loop in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$. Therefore, we construct $\{c_{j}\}^{k+1}_{j=1}$ satisfying  (1)’ and (2)’.
This closes the induction, and the proposition holds with
$\delta(g,\varepsilon)\mathrel{\mathop{\ordinarycolon}}=\min_{1\leq g^{\prime}\leq g}\delta^{\prime}(g,g^{\prime},\varepsilon)\,.$
4.2. Simple short loops in surfaces with large area
In this subsection, we prove the non-degenerate case as follows.
Proposition 4.7 (Existence of homologically independent simple short loops II).
In an orientable closed Riemannian 333-manifold $(M,\mathbf{g})$, let $\Sigma$ be an embedded smooth orientable genus $g_{0}$ closed surface. For any $\varepsilon\in(0,1)$ and any integer $g_{1}\geq g_{0}$, there exists $\delta=\delta(M,\mathbf{g},\Sigma,g_{1},\varepsilon)\in(0,\varepsilon)$ with the following property.
In $(M,\mathbf{g})$, for any embedded smooth orientable closed surface $\Gamma$ containing a finite set of disjoint closed disks $\{\overline{D_{i}}\}^{n}_{i=1}$ with piecewise smooth boundary satisfying:
  1. (i)
    $g_{0}\leq\mathfrak{g}(\Gamma)\leq g_{1}$,
  2. (ii)
    $\mathbf{F}(\Gamma,\Sigma)\leq\delta$,
  3. (iii)
    $\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D_{i})<\delta$,
  4. (iv)
    $\sum^{n}_{i=1}\operatorname{diam}_{M}(\overline{D_{i}})<\delta$,
there exist $\mathfrak{g}(\Gamma)-g_{0}$ disjoint simple loops $\{c_{j}\}^{\mathfrak{g}(\Gamma)-g_{0}}_{j=1}$ in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$ such that:
  1. (1)
    Each $c_{j}$ has length $\mathcal{H}^{1}(c_{j})\leq\varepsilon$.
  2. (2)
    $\{[c_{j}]\}^{\mathfrak{g}(\Gamma)-g_{0}}_{j=1}$ is linearly independent in $H_{1}(\Gamma;\mathbb{Z}_{2})$.
In particular, $\Gamma\setminus\bigcup^{g}_{j=1}c_{j}$ is a genus $g_{0}$ surface with $2(\mathfrak{g}(\Gamma)-g_{0})$ boundary circles.
Remark 4.8.
By Lemma 2.6, for sufficiently small $\delta$ depending on $M,\mathbf{g}$ and $\Sigma$, if $\mathbf{F}(\Gamma,\Sigma)<\delta$, then $\mathfrak{g}({\Gamma})\geq\mathfrak{g}({\Sigma})$. Therefore, we can replace the condition “$g_{0}\leq\mathfrak{g}(\Gamma)\leq g_{1}$” by “$\mathfrak{g}(\Gamma)\leq g_{1}$” in the proposition.
We first prove a useful lemma which enables us to reduce to the case where $\Gamma$ is in a small neighborhood of $\Sigma$.
Lemma 4.9.
In an orientable closed Riemannian 333-manifold $(M,\mathbf{g})$, let $\Sigma$ be an embedded smooth orientable closed surface. For any $\varepsilon\in(0,1)$, there exists $\delta=\delta(M,\mathbf{g},\Sigma,\varepsilon)\in(0,\varepsilon/2)$ with the following property.
In $(M,\mathbf{g})$, for any embedded smooth orientable closed surface $\Gamma$ containing a finite set of disjoint closed disks $\{\overline{D_{i}}\}^{n}_{i=1}$ with piecewise smooth boundary satisfying:
  1. (i)
    $\mathbf{F}(\Gamma,\Sigma)\leq\delta$,
  2. (ii)
    $\sum^{n}_{i=1}\operatorname{diam}_{M}(\overline{D_{i}})<\delta$,
there exists $\eta\in(0,\varepsilon)$ such that:
  1. (1)
    $\exp^{\perp}\mathrel{\mathop{\ordinarycolon}}\Sigma\times[-\eta,\eta]\to M$ is a diffeomorphism to its image $N_{\eta}(\Sigma)\mathrel{\mathop{\ordinarycolon}}=\exp^{\perp}(\Sigma,(-\eta,\eta))$.
  2. (2)
    $\partial N_{\eta}(\Sigma)$ intersects $\Gamma$ transversally and $\partial N_{\eta}(\Sigma)\cap\bigcup^{n}_{i=1}\overline{D_{i}}=\emptyset$.
  3. (3)
    $\mathcal{H}^{1}(\partial N_{\eta}(\Sigma)\cap\Gamma)<\varepsilon$.
  4. (4)
    $\mathcal{H}^{2}(\Gamma\setminus N_{\eta}(\Sigma))<\varepsilon$.
Moreover, by taking $\eta$ even smaller, we can remove a small tubular neighborhood of $\partial N_{\eta}(\Sigma)\cap\Gamma$ in $\Gamma$, and attach disjoint embedded disks $\{D^{\prime}_{i}\}^{n^{\prime}}_{i=1}$ in $M$ to obtain two disjoint embedded smooth orientable closed surfaces $\Gamma_{1}$ and $\Gamma_{2}$ satisfying:
  1. (5)
    $\Gamma_{1}\subset N_{\eta}(\Sigma)$.
  2. (6)
    $\mathbf{F}(\Gamma_{1},\Sigma)<\varepsilon$.
  3. (7)
    $\mathcal{H}^{2}(\Gamma_{2})<\varepsilon$.
And $\{\overline{D_{i}}\}^{n_{\Gamma}}_{i=1}\mathrel{\mathop{\ordinarycolon}}=\{\overline{D_{i}}\}^{n}_{i=1}\cup\{\overline{D^{\prime}_{i}}\}^{n^{\prime}}_{i=1}$ is also a finite set of disjoint closed disks with piecewise smooth boundary, which satisfies:
  1. (8)
    $\sum^{n^{\prime}}_{i=1}\mathcal{H}^{1}(\partial D^{\prime}_{i})<\varepsilon/2$.
  2. (9)
    $\sum^{n^{\prime}}_{i=1}\operatorname{diam}_{M}(\overline{D^{\prime}_{i}})<\varepsilon/2$.
Proof.
Firstly, we choose $\eta_{0}=\eta_{0}(M,\mathbf{g},\Sigma)\in(0,1)$ such that
$\exp^{\perp}\mathrel{\mathop{\ordinarycolon}}\Sigma\times[-\eta_{0},\eta_{0}]\to M$
is a diffeomorphism to its image. We fix $\zeta\in(0,\min(\varepsilon,\eta_{0}))$ which will be determined later.
By the definition of $\mathbf{F}$-metric, we can choose $\delta\in(0,\zeta/10)$ such that for any $\Gamma$ and $\{\overline{D}_{j}\}^{n}_{j=1}$ satisfying (i) and (ii),
$\mathcal{H}^{2}(\Gamma\setminus N_{\zeta/2}(\Sigma))<\frac{\zeta^{2}}{10}\,.$
By the coarea formula and Sard’s theorem, there exists a measurable set $E\in(\zeta/2,\zeta)$ with $\mathcal{H}^{1}(E)>\zeta/5$ such that for any $r\in E$, $\Gamma$ intersects $\partial N_{r}(\Sigma)$ transversally and
$\mathcal{H}^{1}(\Gamma\cap\partial N_{r}(\Sigma))<\zeta<\varepsilon\,.$
By (ii), we can choose $\eta\in E$ such that $\partial N_{\eta}(\Sigma)\cap\bigcup^{n}_{i=1}\overline{D_{i}}=\emptyset$. These confirm (1)(2)(3) and (4).
By the Jordan-Schoenflies theorem and the isoperimetric inequality, Lemma A.4, there exist positive constants $\nu_{\Sigma}$ and $\delta_{\Sigma}$, depending on $M$, $\mathbf{g}$ and $\Sigma$, such that for any $r\in(-\eta_{0},\eta_{0})$ and any simple loop $c$ in $\partial N_{r}(\Sigma)$ with $\mathcal{H}^{1}(c)\leq\delta_{\Sigma}$, the loop $c$ bounds a disk $D\subset\partial N_{r}(\Sigma)$ with
$\mathcal{H}^{2}(D)\leq\nu_{\Sigma}\left(\mathcal{H}^{1}(c)\right)^{2}\leq\nu_{\Sigma}\delta_{\Sigma}\mathcal{H}^{1}(c)\,,\quad\operatorname{diam}_{M}(\overline{D})<\nu_{\Sigma}\mathcal{H}^{1}(c)\,.$
Therefore, if we require that
$\zeta<\min\left(\frac{\varepsilon}{2+4\nu_{\Sigma}},\delta_{\Sigma}\right)\,,$
then in $(M,\mathbf{g})$, choosing suitable $\delta$ and $\eta$ as before, we can remove a small tubular neighborhood of $\Gamma\cap\partial N_{\eta}(\Sigma)=\bigcup^{n^{\prime}}_{i=1}c^{\prime}_{i}$ and attach disks $\{D^{\prime}_{i}\}^{2n^{\prime}}_{i=1}$ near $\partial N_{\eta}(\Sigma)$ to obtain two disjoint embedded closed surfaces $\Gamma_{1}\subset N_{\eta}(\Sigma)$ and $\Gamma_{2}$ satisfying:
$\mathcal{H}^{2}(\Gamma_{2})$ $<\mathcal{H}^{2}(\Gamma\setminus N_{\eta}(\Sigma))+\sum^{n^{\prime}}_{i=1}\mathcal{H}^{2}(D^{\prime}_{i})$
$<\mathcal{H}^{2}(\Gamma\setminus N_{\eta}(\Sigma))+2\nu_{\Sigma}\delta_{\Sigma}\sum^{n^{\prime}}_{i=1}\mathcal{H}^{1}(\partial D^{\prime}_{i})$
$\leq(1+2\nu_{\Sigma}\delta_{\Sigma})\zeta$
$\leq\varepsilon\,,$
$\sum^{n^{\prime}}_{i=1}\mathcal{H}^{1}(\partial D^{\prime}_{j})\leq\zeta<\varepsilon/2\,,$
and
$\sum^{n^{\prime}}_{i=1}\operatorname{diam}_{M}(\overline{D^{\prime}_{i}})\leq\nu_{\Sigma}\sum^{n_{\Gamma}}_{i=1}\mathcal{H}^{1}(\partial D^{\prime}_{i})+\delta<\nu_{\Sigma}\zeta+\zeta<\varepsilon/2\,.$
These confirm (5)(7),  (8) and (9).
Finally, we require that
$\zeta<\frac{\delta(M,\mathbf{g})}{1+2\nu_{\Sigma}\delta_{\Sigma}}\,,$
where $\delta(M,\mathbf{g})$ is a constant from Lemma A.4. By the isoperimetric inequality again, $\Gamma_{2}$ bounds a 333-dimensional modulo 222 integral current $A$ such that
$\mathcal{H}^{3}(A)\leq\nu_{M}(\mathcal{H}^{2}(\Gamma_{2}))^{3/2}\leq\nu_{M}(1+2\nu\Sigma\delta_{\Sigma})^{3/2}\zeta^{3/2}\,,$
where $\nu_{M}$ is also a constant from Lemma A.4. Hence,
$\mathcal{F}([\Gamma],[\Gamma_{1}])\leq\nu_{M}(1+2\nu_{\Sigma}\delta_{\Sigma})^{3/2}\zeta^{3/2}\,,$
and
$|\mathcal{H}^{2}(\Gamma)-\mathcal{H}^{2}(\Gamma_{1})|\leq 2\mathcal{H}^{2}(\Gamma_{2})<2(1+2\nu_{\Sigma}\delta_{\Sigma})\zeta\,.$
It follows from Corollary A.2 that, if we choose $\zeta$ small enough, choosing suitable $\delta$ and $\eta$ as before, we have
$\mathbf{F}(\Gamma,\Gamma_{1})<\varepsilon/2\,,\quad\mathbf{F}(\Gamma,\Sigma)\leq\delta<\varepsilon/2\,.$
This confirms (6) of the proposition.
Note that the choice of $\zeta$ only depends on $M$, $\mathbf{g}$, $\Sigma$ and $\varepsilon$ and so does $\delta$. ∎
This lemma can be extended to punctate surfaces, which will be used in Appendix B.
Corollary 4.10.
In an orientable closed Riemannian 333-manifold $(M,\mathbf{g})$, let $\Sigma$ be an embedded smooth orientable closed surface. For any $\varepsilon\in(0,1)$, there exists $\delta=\delta(M,\mathbf{g},\Sigma,\varepsilon)\in(0,\varepsilon)$ with the following property.
In $(M,\mathbf{g})$, for any punctate surface $S$ with $\mathbf{F}(S,\Sigma)\leq\delta$, there exists $\eta\in(0,\varepsilon)$ such that:
  1. (1)
    $\exp^{\perp}\mathrel{\mathop{\ordinarycolon}}\Sigma\times[-\eta,\eta]\to M$ is a diffeomorphism to its image $N_{\eta}(\Sigma)\mathrel{\mathop{\ordinarycolon}}=\exp^{\perp}(\Sigma,(-\eta,\eta))$.
  2. (2)
    $\partial N_{\eta}(\Sigma)$ intersects $S$ transversally.
  3. (3)
    $\mathcal{H}^{1}(\partial N_{\eta}(\Sigma)\cap S)<\varepsilon$.
  4. (4)
    $\mathcal{H}^{2}(S\setminus N_{\eta}(\Sigma))<\varepsilon$.
Moreover, by taking $\eta$ even smaller, we can remove a small tubular neighborhood of $\partial N_{\eta}(\Sigma)\cap S$ in $S$, and attach disjoint embedded disks $\{D^{\prime}_{i}\}^{n^{\prime}}_{i=1}$ in $M$ to obtain two disjoint embedded smooth orientable closed surfaces $S_{1}$ and $S_{2}$ satisfying:
  1. (5)
    $S_{1}\subset N_{\eta}(\Sigma)$.
  2. (6)
    $\mathbf{F}(S_{1},\Sigma)<\varepsilon$.
  3. (7)
    $\mathcal{H}^{2}(S_{2})<\varepsilon$.
Proof.
For a punctate surface $S$, a punctate set $P$ for $S$ can be viewed as a finite set of degenerating disks. Hence, one can follow the same proof of Lemma 4.9 to conclude the corollary. ∎
By Lemma 4.9, the closed surface $\Gamma$ can be decomposed as two components $\Gamma_{1}$ and $\Gamma_{2}$. Since $\Gamma_{2}$ has small area, Proposition 4.1 can be applied to exhaust all short simple loops it contains. For $\Gamma_{1}$, which lies in a $\eta$-normal neighborhood of $\Sigma$, we choose a fine triangulation of $\Sigma$ and apply Proposition 4.1 again to $\Gamma_{1}$ in the $\eta$-normal neighborhood of each triangle. With small enough $\eta$, removing all those short simple loops in $\Gamma_{1}$ as described above yields a genus $g_{0}$ surface with boundary. The details are presented below.
Proof of Proposition 4.7.
If $g_{1}=g_{0}$, the proposition holds trivially. In the following, we fix $g_{1}>g_{0}\in\mathbb{N}$, an embedded orientable genus $g_{0}$ closed surface $\Sigma\subset(M,\mathbf{g})$ and $\varepsilon>0$ as in the proposition.
Step 0. Set up
Set
(4.3) $\delta_{0}\mathrel{\mathop{\ordinarycolon}}=\delta(g_{1},\varepsilon/2)\in(0,\varepsilon/2)$
from Proposition 4.1.
Fix a triangulation $\mathcal{T}$ of $\Sigma$ such that:
  1. (0.i)
    For any $\Delta\in\mathcal{T}^{(2)}$, $\mathcal{H}^{2}(\Delta)\in\left(0,\frac{\delta_{0}}{10}\right)$.
  2. (0.ii)
    For any $e\in\mathcal{T}^{(1)}$, $\mathcal{H}^{1}(e)\in\left(0,\frac{\delta_{0}}{10}\right)$.
Let $m$ denote the total number of cells in $\mathcal{T}$. Set
$\tau\mathrel{\mathop{\ordinarycolon}}=\frac{1}{1000m^{4}}\,.$
Let $X$ be a smooth vector field defined on $\Sigma$ such that for any $e\in\mathcal{T}^{(1)}$, $X|_{e}$ is nowhere zero and is transverse to $e$. $X$ generates an isotopy
$\varphi\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\to\operatorname{Diff}(\Sigma)$
with $\varphi(0)=\operatorname{Id}$. For convenience, we denote $\varphi(t)$ by $\varphi_{t}$. Then there exists $T>0$ such that for any $t\in[-2T,2T]$, $e\in\mathcal{T}^{(1)}$ and $\Delta\in\mathcal{T}^{(2)}$,
(4.4) $\frac{\mathcal{H}^{1}(\varphi_{t}(e))}{\mathcal{H}^{1}(e)}\in\left(1-\frac{\tau^{2}}{100},1+\frac{\tau^{2}}{100}\right)\,,$
and
(4.5) $\frac{\mathcal{H}^{2}(\varphi_{t}(\Delta))}{\mathcal{H}^{2}(\Delta)}\in\left(1-\frac{\tau^{2}}{100},1+\frac{\tau^{2}}{100}\right)\,.$
For each vertex $v\in\mathcal{T}^{(0)}$, since $T_{v}\Sigma$ is a 222-dimensional normed vector space, it is isometrically isomorphic to $\mathbb{R}^{2}$. The isomorphism induces a smooth $\mathbb{S}^{1}$ family of vectors at $v$:
$Y_{v}\mathrel{\mathop{\ordinarycolon}}\mathbb{S}^{1}\to\{Z\in T_{v}\Sigma\mathrel{\mathop{\ordinarycolon}}|Z|=1\}\,.$
This can be extended to a continuous family of vector fields on $\Sigma$
$Y\mathrel{\mathop{\ordinarycolon}}\mathbb{S}^{1}\to\Gamma(T\Sigma)$
such that for any vertex $v\in\mathcal{T}^{(0)}$ and any $\theta\in\mathbb{S}^{1}$, $Y(\theta,v)=Y_{v}(\theta)$. Therefore, for any small $\xi>0$, $Y$ generates a smooth family of diffeomorphisms
$\psi\mathrel{\mathop{\ordinarycolon}}\overline{\mathbb{D}_{2\xi}}\to\operatorname{Diff}(\Sigma)$
with $\psi(\mathbf{0})=\operatorname{Id}$, where $\overline{\mathbb{D}_{2\xi}}\subset\mathbb{R}^{2}$ is a closed disk centered at the origin with radius $2\xi$. Note that for any $v\in\mathcal{T}^{(0)}$, $\psi(\cdot,v)$ is a diffeomorphism to the geodesic ball $B_{\Sigma}(v,\xi)$. For simplicity, we denote $\psi(z)$ by $\psi_{z}$ and $\varphi_{t}\circ\psi_{z}$ by $\varphi_{z,t}$.
By (4.4) and (4.5), there exists a sufficiently small $\xi>0$ such that for any $z\in\overline{\mathbb{D}_{2\xi}}$, $t\in[-2T,2T]$, $e\in\mathcal{T}^{(1)}$ and $\Delta\in\mathcal{T}^{(2)}$,
(4.6) $\frac{\mathcal{H}^{1}(\varphi_{z,t}(e))}{\mathcal{H}^{1}(e)}\in\left(1-\frac{\tau^{2}}{10},1+\frac{\tau^{2}}{10}\right)\,,$
and
(4.7) $\frac{\mathcal{H}^{2}(\varphi_{z,t}(\Delta))}{\mathcal{H}^{2}(\Delta)}\in\left(1-\frac{\tau^{2}}{10},1+\frac{\tau^{2}}{10}\right)\,.$
In the following, we fix such $\xi>0$.
Next we choose $T>0$ even smaller such that for any $t\in[-2T,2T]$ and $v\in\mathcal{T}^{(0)}$, the map
(4.8) $F_{v,t}\mathrel{\mathop{\ordinarycolon}}\overline{\mathbb{D}_{\xi}}\to\overline{\mathbb{D}_{(1+\tau)\xi}},\quad z\mapsto(\psi(\cdot,v))^{-1}(\varphi_{t}(\psi(z,v)))$
is a diffeomorphism to its image with a bi-Lipschitz constant $\left(1+\frac{\tau^{2}}{100}\right)$ and smoothly depends on $t$. This is possible because $F_{v,0}=\operatorname{Id}$.
Finally, we choose $\zeta>0$ such that the normal exponential map
$\exp^{\perp}\mathrel{\mathop{\ordinarycolon}}\Sigma\times(-\zeta,\zeta)\to M$
is a diffeomorphism to its image $N_{\zeta}(\Sigma)$ with a bi-Lipschitz constant $1+\frac{\tau^{2}}{100}$. We denote by
$\pi\mathrel{\mathop{\ordinarycolon}}N_{\zeta}(\Sigma)\to\Sigma$
the corresponding closest-point projection.
For convenience, in the following, for any $v\in\mathcal{T}^{(0)}$, $e\in\mathcal{T}^{(1)}$, $\Delta\in\mathcal{T}^{(2)}$ (note that they are any closed sets), $z\in\overline{\mathbb{D}}^{2}_{2\xi}$ and $t\in[-2T,2T]$, we define:
  • the point $v_{z,t}\mathrel{\mathop{\ordinarycolon}}=\varphi_{z,t}(v)$;
  • the vertical line $\mathrm{I}_{v,z,t}\mathrel{\mathop{\ordinarycolon}}=\pi^{-1}(v_{z,t})$;
  • the solid cylinder $\mathrm{S}_{v}\mathrel{\mathop{\ordinarycolon}}=\bigcup_{z\in\mathbb{D}^{2}_{(1+\tau)\xi}}\mathrm{I}_{v,z,0}$;
  • the edge $e_{z,t}\mathrel{\mathop{\ordinarycolon}}=\varphi_{z,t}(e)$;
  • the rectangle $\mathrm{R}_{e,z,t}\mathrel{\mathop{\ordinarycolon}}=\pi^{-1}(e_{z,t})$;
  • the cuboid $\mathrm{Cub}_{e,z,t}\mathrel{\mathop{\ordinarycolon}}=\bigcup_{t\in(-T,T)}\mathrm{R}_{e,z,t}$;
  • the triangle $\Delta_{z,t}\mathrel{\mathop{\ordinarycolon}}=\varphi_{z,t}(\Delta)$;
  • the cylinder $\mathrm{Cyl}_{\Delta,z,t}\mathrel{\mathop{\ordinarycolon}}=\pi^{-1}(\partial\Delta_{z,t})$;
  • the prism $\mathrm{Pri}_{\Delta,z,t}\mathrel{\mathop{\ordinarycolon}}=\pi^{-1}(\Delta_{z,t})$.
Step 1. Existence of a good triangulation
We show that there exists $\delta=\delta_{1}(M,\mathbf{g},\varepsilon,g_{1},\Sigma)\in(0,1)$ with the following property:
For any embedded smooth orientable closed surface $\Gamma\subset N_{\zeta}(\Sigma)$ with $\mathfrak{g}(\Gamma)\leq g_{1}$, $\mathbf{F}(\Gamma,\Sigma)<\delta$ and any finite set of disjoint closed disks $\{\overline{D_{i}}\}^{n}_{i=1}$ in $\Gamma$ with
$\sum^{n}_{i=1}\operatorname{diam}_{M}(\overline{D_{i}})<\delta\,,$
there exist $z\in\mathbb{D}^{2}_{\xi}$ and $t\in(-T,T)$ such that:
  1. (1)
    For any $v\in\mathcal{T}^{(0)}$, $\mathrm{I}_{v,z,t}\pitchfork\Gamma$ is a single point in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$.
  2. (2)
    For any $\Delta\in\mathcal{T}^{(2)}$, $\mathrm{Cyl}_{\Delta,z,t}\pitchfork\Gamma=\bigcup^{n_{\Delta}}_{k=1}c_{\Delta,k}$ is a finite union of disjoint simple loops in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$ satisfying
    $\frac{\mathcal{H}^{1}(\mathrm{Cyl}_{\Delta,z,t}\cap\Gamma)}{\mathcal{H}^{1}(\partial\Delta_{z,t})}\in(1-\tau,1+\tau)\,.$
Moreover, for any $\Delta\in\mathcal{T}^{(2)}$, we can arrange $\{c_{\Delta,k}\}^{n_{\Delta}}_{k=1}$ in such an order that:
  1. (3)
    $c_{\Delta,1}$ satisfies $[c_{\Delta,1}]\neq 0\in H_{1}(\mathrm{Cyl}_{\Delta,z,t};\mathbb{Z}_{2})$ and
    $\frac{\mathcal{H}^{1}(c_{\Delta,1})}{\mathcal{H}^{1}(\partial\Delta_{z,t})}\in(1-\tau,1+\tau)\,;$
    we call it the long loop for $\Delta$.
  2. (4)
    Every $c_{\Delta,k}$ with $k\in\{2,\dots,n_{\Delta}\}$ satisfies $[c_{\Delta,k}]=0\in H_{1}(\mathrm{Cyl}_{\Delta,z,t};\mathbb{Z}_{2})$, $c_{\Delta,k}\cap\bigcup_{v\in\mathcal{T}^{(0)}}\mathrm{I}_{v,z,t}=\emptyset$, and
    $\frac{\sum^{n_{\Delta}}_{k=2}\mathcal{H}^{1}(c_{\Delta,k})}{\mathcal{H}^{1}(\partial\Delta_{z,t})}\in(0,\tau)\,;$
    we call them the short loops for $\Delta$.
In the following, at each step, we will choose $\delta$ to be successively smaller, depending only on all the objects defined in Step 0, which were chosen in terms of $M$, $\mathbf{g}$, $\varepsilon$, $g_{1}$ and $\Sigma$.
Step 1.1.
Choose small $\delta$ such that $\Sigma$ and $\Gamma$ are homologous to each other.
By the isoperimetric inequality, Lemma A.4, we choose $\delta$ small enough, depending on $\Sigma$ and $\zeta$, such that $\mathbf{F}(\Gamma,\Sigma)\leq\delta$ and $\Gamma\subset N_{\zeta}(\Sigma)$ imply $\Sigma$ and $\Gamma$ are homologous to each other in $N_{\zeta}(\Sigma)$ with $\mathbb{Z}_{2}$ coefficients.
This implies that for any $v\in\mathcal{T}^{(0)}$, $e\in\mathcal{T}^{(1)}$, $\Delta\in\mathcal{T}^{(2)}$, $z\in\mathbb{D}^{2}_{(1+\tau)\xi}$, $t\in(-T,T)$, and $\Gamma$ with $\mathbf{F}(\Gamma,\Sigma)\leq\delta$, if $\mathrm{I}_{v,z,0}$ intersects $\Gamma$ transversely, then
(4.9) $[\mathrm{I}_{v,z,0}\cap\Gamma]=[v_{z,0}]\neq 0\in H_{0}(\mathrm{I}_{v,z,0};\mathbb{Z}_{2})\,;$
if $\mathrm{R}_{e,z,t}$ intersects $\Gamma$ transversely, then
(4.10) $[\mathrm{R}_{e,z,t}\cap\Gamma]=[e_{z,t}]\neq 0\in H_{1}(\mathrm{R}_{e,z,t};\mathbb{Z}_{2})\,;$
if $\mathrm{Cyl}_{\Delta,z,t}$ intersects $\Gamma$ transversely, then
(4.11) $[\mathrm{Cyl}_{\Delta,z,t}\cap\Gamma]=[\partial\Delta_{z,t}]\neq 0\in H_{1}(\mathrm{Cyl}_{\Delta,z,t};\mathbb{Z}_{2})\,.$
Step 1.2
Choose smaller $\delta$ and $W_{\Gamma}\subset\mathbb{D}^{2}_{(1+\tau)\xi}$ of relatively large measure such that for any $v\in\mathcal{T}^{(0)}$ and $z\in W_{\Gamma}$, $I_{v,z,0}\pitchfork\Gamma$ at exactly a single point.
For any $v\in\mathcal{T}^{(0)}$, by the definition of $\psi$, we can define a $C^{1}$ function $\mathfrak{z}_{v}\mathrel{\mathop{\ordinarycolon}}\mathrm{S}_{v}\to\mathbb{D}^{2}_{(1+\tau)\xi}$ by
$\psi(\mathfrak{z}_{v}(x),v)=\pi(x)\,,$
and in $S_{v}$, the Jacobian
$J^{M}_{\mathfrak{z}_{v}}\mathrel{\mathop{\ordinarycolon}}=\sqrt{\det\left(\left(\nabla^{M}\mathfrak{z}^{i}_{v}\cdot\nabla^{M}\mathfrak{z}^{j}_{v}\right)_{i,j\in\{1,2\}}\right)}>0\,.$
We also write
$J^{\Gamma}_{\mathfrak{z}_{v}}\mathrel{\mathop{\ordinarycolon}}=\sqrt{\det\left(\left(\nabla^{\Gamma}\mathfrak{z}^{i}_{v}\cdot\nabla^{\Gamma}\mathfrak{z}^{j}_{v}\right)_{i,j\in\{1,2\}}\right)}\,,$
and
$J^{\Sigma}_{\mathfrak{z}_{v}}\mathrel{\mathop{\ordinarycolon}}=\sqrt{\det\left(\left(\nabla^{\Sigma}\mathfrak{z}^{i}_{v}\cdot\nabla^{\Sigma}\mathfrak{z}^{j}_{v}\right)_{i,j\in\{1,2\}}\right)}\,.$
It is clear that for any $x\in\Gamma\cap S_{v}$, $J^{\Gamma}_{\mathfrak{z}_{v}}\leq J^{M}_{\mathfrak{z}_{v}}$ and for any $x\in\Sigma\cap S_{v}$, $J^{\Sigma}_{\mathfrak{z}_{v}}=J^{M}_{\mathfrak{z}_{v}}$.
Therefore, by the coarea formula, for any $v\in\mathcal{T}^{(0)}$,
(4.12) $\int_{S_{v}\cap\Sigma}J^{M}_{\mathfrak{z}_{v}}=\int_{S_{v}\cap\Sigma}J^{\Sigma}_{\mathfrak{z}_{v}}=\int_{\mathbb{D}^{2}_{(1+\tau)\xi}}\mathcal{H}^{0}(\mathrm{I}_{v,z,0}\cap\Sigma)\mathrm{d}z=\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\,,$
and
(4.13) $\int_{S_{v}\cap\Gamma}J^{M}_{\mathfrak{z}_{v}}\geq\int_{S_{v}\cap\Gamma}J^{\Gamma}_{\mathfrak{z}_{v}}=\int_{\mathbb{D}^{2}_{(1+\tau)\xi}}\mathcal{H}^{0}(\mathrm{I}_{v,z,0}\cap\Gamma)\mathrm{d}z\,.$
Note that for any $v\in\mathcal{T}^{(0)}$, $\partial S_{v}\pitchfork\Sigma$, so we have
$\lim_{r\to 0}\sup_{v\in\mathcal{T}^{(0)}}|\Sigma|(N_{r}(\partial S_{v}))=0\,.$
By Lemma A.3, we can choose $\delta$ even smaller such that, $\mathbf{F}(\Gamma,\Sigma)\leq\delta$ implies $\sup_{v\in\mathcal{T}^{(0)}}\mathbf{F}_{S_{v}}(|\Gamma|,|\Sigma|)$ is so small that for any $v\in\mathcal{T}^{(0)}$,
$\int_{S_{v}\cap\Gamma}J^{M}_{\mathfrak{z}_{v}}\leq\left(1+\frac{\tau^{2}}{10}\right)\int_{S_{v}\cap\Sigma}J^{M}_{\mathfrak{z}_{v}}\,.$
This together with (4.12) and (4.13) implies that for any $v\in\mathcal{T}^{(0)}$,
(4.14) $\int_{\mathbb{D}^{2}_{(1+\tau)\xi}}\mathcal{H}^{0}(\mathrm{I}_{v,z,0}\cap\Gamma)\mathrm{d}z\leq\left(1+\frac{\tau^{2}}{10}\right)\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\,.$
By Sard’s theorem, for any $\Gamma$ with $\mathbf{F}(\Gamma,\Sigma)\leq\delta$, there exists an open subset $U_{\Gamma}\subset\mathbb{D}^{2}_{(1+\tau)\xi}$ with $\mathcal{H}^{2}(U_{\Gamma})=\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})$ such that for any $z\in U_{\Gamma}$, $\mathrm{I}_{v,z,0}$ intersects $\Gamma$ transversely. By (4.9), we have
$\mathcal{H}^{0}(\mathrm{I}_{v,z,0}\cap\Gamma)\geq 1\,.$
For any $v\in\mathcal{T}^{(0)}$, let $W_{\Gamma,v}\subset U_{\Gamma}$ be the set of any $z$ such that $\mathcal{H}^{0}(\mathrm{I}_{v,z,0}\cap\Gamma)=1$, which is also an open subset. It follows from (4.14) that
$\mathcal{H}^{2}(W_{\Gamma,v})+2\left(\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})-\mathcal{H}^{2}(W_{\Gamma,v})\right)\leq\left(1+\frac{\tau^{2}}{10}\right)\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\,,$
and thus,
$\mathcal{H}^{2}(W_{\Gamma,v})\geq\left(1-\frac{\tau^{2}}{10}\right)\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\,.$
Hence, we can define an open set
$W_{\Gamma}\mathrel{\mathop{\ordinarycolon}}=\bigcap_{v\in\mathcal{T}^{(0)}}W_{\Gamma,v}\,.$
Then for any $v\in\mathcal{T}^{(0)}$ and $z\in W_{\Gamma}$, $I_{v,z,0}$ intersects $\Gamma$ transversely at exactly a point. Since #$\#\mathcal{T}^{(0)}\leq m$, the set satisfies
$\mathcal{H}^{2}(W_{\Gamma})\geq\left(1-m\frac{\tau^{2}}{10}\right)\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})>\left(1-\frac{\tau}{100}\right)\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\,,$
Step 1.3
Choose $z_{\Gamma}\in\mathbb{D}^{2}_{\xi}$ and $E_{\Gamma}\subset(-T,T)$ of relatively large measure such that for any $v\in\mathcal{T}^{(0)}$ and $t\in E_{\Gamma}$, $I_{v,z_{\Gamma},t}\pitchfork\Gamma$ at exactly a single point.
Fix $\Gamma$ with $\mathbf{F}(\Gamma,\Sigma)\leq\delta$. Let $W^{\prime}_{\Gamma}\mathrel{\mathop{\ordinarycolon}}=W_{\Gamma}\cap\mathbb{D}^{2}_{\zeta}$, which satisfies
$\mathcal{H}^{2}(W^{\prime}_{\Gamma})\geq\mathcal{H}^{2}(\mathbb{D}^{2}_{\xi})-\frac{\tau}{100}\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\geq\left(1-3\tau\right)\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\,.$
For any $v\in\mathcal{T}^{(0)}$ and $z\in W^{\prime}_{\Gamma}$, since $W^{\prime}_{\Gamma}$ is open, the set
$E_{\Gamma,v,z}\mathrel{\mathop{\ordinarycolon}}=\{t\in(-T,T)\mathrel{\mathop{\ordinarycolon}}F_{v,t}(z)\in W^{\prime}_{\Gamma}\}\,,$
is also open, where $F_{v,t}$ is defined in (4.8). For any $v\in\mathcal{T}^{(0)}$, we have
$\int_{W^{\prime}_{\Gamma}}\mathcal{H}^{1}(E_{\Gamma,v,z})\mathrm{d}z$ $=\int_{W^{\prime}_{\Gamma}}\int^{T}_{-T}\mathbf{1}_{W^{\prime}_{\Gamma}}(F_{v,t}(z))\mathrm{d}t\mathrm{d}z$
$=\int^{T}_{-T}\int_{W^{\prime}_{\Gamma}}\mathbf{1}_{W^{\prime}_{\Gamma}}(F_{v,t}(z))\mathrm{d}z\mathrm{d}t$
$=\int^{T}_{-T}\mathcal{H}^{2}\left(W^{\prime}_{\Gamma}\cap F^{-1}_{v,t}(W^{\prime}_{\Gamma})\right)\mathrm{d}t\,.$
Since for any $t\in(-T,T)$ and $v\in\mathcal{T}^{(0)}$, $F_{v,t}\mathrel{\mathop{\ordinarycolon}}\mathbb{D}^{2}_{\xi}\to\mathbb{D}^{2}_{(1+\tau)\xi}$ is a diffeomorphism of bi-Lispchitz constant $1+\frac{\tau^{2}}{100}$, the Hausdorff measure
$\mathcal{H}^{2}\left(W^{\prime}_{\Gamma}\cap F^{-1}_{v,t}(W^{\prime}_{\Gamma})\right)$ $\geq\left(1-\frac{\tau^{2}}{10}\right)\mathcal{H}^{2}\left(F_{v,t}(W^{\prime}_{\Gamma})\cap W^{\prime}_{\Gamma}\right)$
$\geq\left(1-\frac{\tau^{2}}{10}\right)\left(\mathcal{H}^{2}(F_{v,t}(W^{\prime}_{\Gamma}))+\mathcal{H}^{2}(W^{\prime}_{\Gamma})-\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\right)$
$\geq\left(1-\frac{\tau^{2}}{10}\right)\left(\left(1-\frac{\tau^{2}}{10}\right)\mathcal{H}^{2}(W^{\prime}_{\Gamma})+\mathcal{H}^{2}(W^{\prime}_{\Gamma})-\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\right)$
$\geq(1-10\tau)\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\,.$
These imply that for any $v\in\mathcal{T}^{(0)}$, the set
$W^{\prime\prime}_{\Gamma,v}\mathrel{\mathop{\ordinarycolon}}=\left\{z\in W^{\prime}_{\Gamma}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}^{1}(E_{\Gamma,v,z})\geq 2T(1-\sqrt{\tau})\right\}\,,$
satisfies
$\mathcal{H}^{2}(W^{\prime\prime}_{\Gamma,v})+(1-\sqrt{\tau})(\mathcal{H}^{2}(W^{\prime}_{\Gamma})-\mathcal{H}^{2}(W^{\prime\prime}_{\Gamma,v}))$ $\geq\frac{\int_{W^{\prime}_{\Gamma}}\mathcal{H}^{1}(E_{\Gamma,v,z})\mathrm{d}z}{2T}$
$\geq(1-10\tau)\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})$
and thus,
$\mathcal{H}^{2}(W^{\prime\prime}_{\Gamma,v})\geq(1-10\sqrt{\tau})\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\,.$
As in Step 1.2, we define
$W^{\prime\prime}_{\Gamma}\mathrel{\mathop{\ordinarycolon}}=\bigcap_{v\in\mathcal{T}^{(0)}}W^{\prime\prime}_{\Gamma,v}\,,$
which satisfies
$\mathcal{H}^{2}(W^{\prime\prime}_{\Gamma})\geq(1-m\cdot 10\sqrt{\tau})\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})>0\,.$
In the following, we set $z_{\Gamma}\in W^{\prime\prime}_{\Gamma}$, and
$E_{\Gamma}\mathrel{\mathop{\ordinarycolon}}=\bigcap_{v\in\mathcal{T}^{(0)}}E_{\Gamma,v,z_{\Gamma}}\,.$
Then $E_{\Gamma}$ has measure $\mathcal{H}^{1}(E_{\Gamma})\geq 2T(1-m\sqrt{\tau})$. Moreover, by definition, for any $v\in\mathcal{T}^{(0)}$ and $t\in E_{\Gamma}$, $I_{v,z_{\Gamma},t}$ intersects $\Gamma$ at exactly a single point.
Step 1.4
Choose smaller $\delta$ and $\widetilde{E}_{\Gamma}\subset E_{\Gamma}$ of relatively large measure such that for any $e\in\mathcal{T}^{(1)}$ and $t\in\widetilde{E}_{\Gamma}$, $\mathcal{H}^{1}(\mathrm{R}_{e,z_{\Gamma},t}\cap\Gamma)/\mathcal{H}^{1}(e_{z_{\Gamma},t})\in\left(1-\frac{\tau^{2}}{10},1+\frac{\tau}{2}\right)$. Furthermore, the path $\gamma_{e,z,t}$ in $\mathrm{R}_{e,z,t}\cap\Gamma$ connecting $\pi^{-1}(\partial e_{z,t})$ is unique.
For any 111-cell $e\in\mathcal{T}^{(1)}$ and $z\in\mathbb{D}^{2}_{\xi}$, by the definition of $\varphi$ in Step 0, we can define a $C^{1}$ function $\mathfrak{t}_{e,z}\mathrel{\mathop{\ordinarycolon}}\mathrm{Cub}_{e,z}\to(-T,T)$ by
$\pi(x)\in e_{z,\mathfrak{t}_{e,z}(x)}\,.$
By the coarea formula, for each $e\in\mathcal{T}^{(1)}$, we have
[…]
and
[…]
Note that for any $e\in\mathcal{T}^{(1)}$ and $z\in\mathbb{D}^{2}_{\xi}$, since $\partial\mathrm{Cub}_{e,z}$ intersects $\Sigma$ transversely and deforms continuously with respect to $z$, we have
$\lim_{r\to 0}\sup_{e\in\mathcal{T}^{(1)},z\in\mathbb{D}^{2}_{\xi}}|\Sigma|(N_{r}(\partial\mathrm{Cub}_{e,z}))=0\,.$
By Lemma A.3 again, we can choose $\delta$ even smaller such that, $\mathbf{F}(\Gamma,\Sigma)\leq\delta$ implies $\sup_{e\in\mathcal{T}^{(1)},z\in\mathbb{D}^{2}_{\xi}}\mathbf{F}_{\mathrm{Cub}_{e,z}}(|\Gamma|,|\Sigma|)$ is sufficiently so small that for any $e\in\mathcal{T}^{(1)}$ and $z\in\mathbb{D}^{2}_{\xi}$,
$\int_{\mathrm{Cub}_{e,z}\cap\Gamma}|\nabla^{M}\mathfrak{t}_{e,z}|\leq\left(1+\frac{\tau^{2}}{10}\right)\int_{\mathrm{Cub}_{e,z}\cap\Sigma}|\nabla^{M}\mathfrak{t}_{e,z}|\,.$
This together with (4.15) and (4.16) leads to
[…]
for any $z\in\mathbb{D}^{2}_{\xi}$.
By Sard’s theorem, for any $z\in\mathbb{D}^{2}_{\xi}$ and $\Gamma$ with $\mathbf{F}(\Gamma,\Sigma)\leq\delta$, there exists an open set $O_{\Gamma,z}\subset(-T,T)$ with $\mathcal{H}^{2}(O_{\Gamma,z})=2T$, such that for any $t\in O_{\Gamma,z}$ and $e\in\mathcal{T}^{(1)}$, $\mathrm{R}_{e,z,t}$ intersects $\Gamma$ transversely. It follows from (4.10) that, $\mathrm{R}_{e,z,t}\cap\Gamma$ has at least one path $\gamma_{e,z,t}$ connecting the two components $\pi^{-1}(\partial e_{z,t})$. Since $\exp^{\perp}$ is $\left(1+\frac{\tau^{2}}{10}\right)$-biLipschitz,
(4.18) $\mathcal{H}^{1}(\mathrm{R}_{e,z,t}\cap\Gamma)\geq\mathcal{H}^{1}(\gamma_{e,z,t})\geq\frac{\mathcal{H}^{1}(e_{z,t})}{1+\frac{\tau^{2}}{10}}>\left(1-\frac{\tau^{2}}{10}\right)\mathcal{H}^{1}(e_{z,t})\,.$
For any $e\in\mathcal{T}^{(1)}$, $z\in\mathbb{D}^{2}_{\xi}$, and $\Gamma$ with $\mathbf{F}(\Gamma,\Sigma)\leq\delta$, we define
$\widetilde{E}_{\Gamma,e,z}\mathrel{\mathop{\ordinarycolon}}=\left\{t\in O_{\Gamma,z}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}^{1}(\mathrm{R}_{e,z,t}\cap\Gamma)\leq\left(1+\frac{\tau}{2}\right)\mathcal{H}^{1}(e_{z,t})\right\}\,.$
The inequality (4.18) implies that for any $t\in\widetilde{E}_{\Gamma,e,z}$, the path $\gamma_{e,z,t}$ in $\mathrm{R}_{e,z,t}\cap\Gamma$ connecting $\pi^{-1}(\partial e_{z,t})$ is unique. By (4.6) and (4.18), we have
$\int_{O_{\Gamma,z}}\mathcal{H}^{1}(\mathrm{R}_{e,z,t}\cap\Gamma)\mathrm{d}t\geq$ $\mathcal{H}^{1}(\widetilde{E}_{\Gamma,e,z})\left(1-\frac{\tau^{2}}{10}\right)^{2}\mathcal{H}^{1}(e)$
$\quad+(2T-\mathcal{H}^{1}(\widetilde{E}_{\Gamma,e,z}))\left(1+\frac{\tau}{2}\right)\left(1-\frac{\tau^{2}}{10}\right)\mathcal{H}^{1}(e)\,,$
and
$\int^{T}_{-T}\mathcal{H}^{1}(e_{z,t})\mathrm{d}t\leq 2T\left(1+\frac{\tau^{2}}{10}\right)\mathcal{H}^{1}(e)\,.$
These together with (4.17) imply that the Hausdorff measure of $\widetilde{E}_{\Gamma,e,z}$ satisfies
$\mathcal{H}^{1}(\widetilde{E}_{\Gamma,e,z})\geq 2T(1-\tau)\,.$
For any $\Gamma$ with $\mathbf{F}(\Gamma,\Sigma)<\delta$, we define
$\widetilde{E}_{\Gamma}\mathrel{\mathop{\ordinarycolon}}=\bigcap_{e\in\mathcal{T}^{(1)}}\widetilde{E}_{\Gamma,e,z_{\Gamma}}\cap E_{\Gamma}\,,$
which satisfies
$\mathcal{H}^{1}(\widetilde{E}_{\Gamma})\geq 2T(1-m\tau-m\sqrt{\tau})\geq 2T(1-\sqrt[4]{\tau})>0\,.$
Moreover, for any $e\in\mathcal{T}^{(1)}$ and $t\in\widetilde{E}_{\Gamma}$,
$\frac{\mathcal{H}^{1}(\mathrm{R}_{e,z_{\Gamma},t}\cap\Gamma)}{\mathcal{H}^{1}(e_{z_{\Gamma},t})}\in\left(1-\frac{\tau^{2}}{10},1+\frac{\tau}{2}\right)\,,$
and the path $\gamma_{e,z_{\Gamma},t}$ in $\mathrm{R}_{e,z_{\Gamma},t}\cap\Gamma$ connecting $\pi^{-1}(\partial e_{z_{\Gamma},t})$ is unique.
Step 1.5
Choose smaller $\delta$ and $F_{\Gamma,\{\overline{D_{i}}\}_{i}}\subset(-T,T)$ of relatively large measure such that for any $e\in\mathcal{T}^{(1)}$ and $t\in F_{\Gamma,\{\overline{D_{i}}\}_{i}}$, $\mathrm{R}_{e,z_{\Gamma},t}\cap\bigcup^{n}_{i=1}\overline{D_{i}}=\emptyset$.
By the definition of $\varphi_{z,t}$, we choose $\delta$ smaller such that $\sum^{n}_{i=1}\mathrm{diam}^{M}(\overline{D_{i}})<\delta$ implies that for any $e\in\mathcal{T}^{(1)}$ and $z\in\mathbb{D}^{2}_{\xi}$, there exists $F_{e,z,\{\overline{D_{i}}\}_{i}}\subset(-T,T)$ with
$\mathcal{H}^{1}(F_{e,z,\{\overline{D_{i}}\}_{i}})\geq 2T(1-\sqrt[4]{\tau})$
such that for any $t\in F_{e,z,\{\overline{D_{i}}\}_{i}}$,
$\mathrm{R}_{e,z,t}\cap\bigcup^{n}_{i=1}\overline{D_{i}}=\emptyset\,.$
Then we define
$F_{\Gamma,\{\overline{D_{i}}\}_{i}}\mathrel{\mathop{\ordinarycolon}}=\bigcap_{e\in\mathcal{T}^{(1)}}F_{e,z_{\Gamma},\{\overline{D_{i}}\}_{i}}\,,$
which has $\mathcal{H}^{1}(F_{\Gamma,\{\overline{D_{i}}\}_{i}})\geq 2T(1-m\sqrt[4]{\tau})$. Clearly, for any $e\in\mathcal{T}^{(1)}$ and $t\in F_{\Gamma,\{\overline{D_{i}}\}_{i}}$,
$\mathrm{R}_{e,z_{\Gamma},t}\cap\bigcup^{n}_{i=1}\overline{D_{i}}=\emptyset\,.$
Step 1.6
We complete the proof.
Fix an arbitrary $\Gamma$ with $\mathbf{F}(\Gamma,\Sigma)<\delta$ and disks $\{\overline{D_{i}}\}^{n}_{i=1}$ in $\Gamma$ with $\sum^{n}_{i=1}\mathrm{diam}^{M}(\overline{D_{i}})<\delta$. We define
$G_{\Gamma,\{\overline{D_{i}}\}_{i}}\mathrel{\mathop{\ordinarycolon}}=\widetilde{E}_{\Gamma}\cap F_{\{\overline{D_{i}}\}_{i},z_{\Gamma}}\,,$
which satisfies
$\mathcal{H}^{1}(G_{\Gamma,\{\overline{D_{i}}\}_{i}})\geq 2T(1-2m\sqrt[4]{\tau})>0\,.$
Therefore, we can pick $t_{\Gamma,\{\overline{D_{i}}\}}\in G_{\Gamma,\{\overline{D_{i}}\}_{i}}$.
For simplicity, in the following, we write $\bar{z}=z_{\Gamma}$ and $\bar{t}=t_{\Gamma,\{\overline{D_{i}}\}}$.
For any $v\in\mathcal{T}^{(0)}$, by Step 1.3 and Step 1.5, $I_{v,\bar{z},\bar{t}}\pitchfork\Gamma$ is a single point in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$. This confirms (1).
For any $\Delta\in\mathcal{T}^{(2)}$, assume that $\partial\Delta=e^{1}\cup e^{2}\cup e^{3}$. Then by Step 1.4 and Step 1.5, for any $e^{i}$, $\mathrm{R}_{e^{i},z,t}\pitchfork\Gamma$ lies in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$, and
$\frac{\mathcal{H}^{1}(\mathrm{R}_{e^{i},z,t}\cap\Gamma)}{\mathcal{H}^{1}({e^{i}}_{z,t})}\in(1-\tau/2,1+\tau/2)\,.$
This confirms (2).
Moreover, for any $\Delta\in\mathcal{T}^{(2)}$ with $\partial\Delta=e^{1}\cup e^{2}\cup e^{3}$, by Step 1.4, for each $e^{i}$, there exists a unique connected component in $\mathrm{R}_{e^{i},z,t}\cap\Gamma$ connecting the two components $\pi^{-1}(\partial e_{z,t})$. Together with (1), we can conclude that the other connected component in $\mathrm{R}_{e^{i},z,t}\cap\Gamma$ are loops away from $\pi^{-1}(\partial e_{z,t})$. Therefore, the cylinder $C_{\Delta_{z,t}}\pitchfork\Gamma=\bigcup^{n_{\Delta}}_{k=1}c_{\Delta,k}$ has a unique loop $c_{\Delta,1}$ with $[c_{\Delta,1}]\neq 0\in H_{1}(\mathrm{Cyl}_{\Delta,z,t})$ and
$\frac{\mathcal{H}^{1}(c_{\Delta,1})}{\mathcal{H}^{1}(\partial\Delta_{z,t})}\in(1-\tau/2,1+\tau/2)\,.$
Every other loop $c_{\Delta,k}$ with $k\geq 2$, satisfies $[c_{\Delta,k}]=0\in H_{1}(\mathrm{Cyl}_{\Delta,z,t})$, $c_{\Delta,k}\cap\bigcup_{v\in\mathcal{T}^{(0)}}\mathrm{I}_{v,z,t}=\emptyset$, and
$\frac{\sum^{n_{\Delta}}_{k=2}\mathcal{H}^{1}(c_{\Delta,k})}{\mathcal{H}^{1}(\partial\Delta_{z,t})}\in(0,\tau)\,.$
These confirm both (3) and  (4).
This completes the proof of Step 1.
Step 2. Special case
We prove the proposition with an extra assumption that $\Gamma\subset N_{\zeta}(\Sigma)$.
Let $\delta_{0}$ be as in Step 0 and $\delta_{1}=\delta_{1}(M,\mathbf{g},\varepsilon,g_{1},\Sigma)$ be as in Step 1.
Note that for any $\Delta\in\mathcal{T}^{(2)}$, $z\in\mathbb{D}^{2}_{\xi}$, $t\in(-T,T)$, since $\mathrm{Cyl}_{\Delta,z,t}\pitchfork\Sigma$ and deforms continuosly with respect to $z$ and $t$, we have
$\lim_{r\to 0}\sup_{\Delta\in\mathcal{T}^{(2)},z\in\mathbb{D}^{2}_{\xi},t\in(-T,T)}|\Sigma|(N_{r}(\mathrm{Cyl}_{\Delta,z,t}))=0\,.$
By Lemma A.3,  (0.i) and (4.7), there exists positive $\delta=\delta_{2}(M,\mathbf{g},\varepsilon,g_{1},\Sigma)\in(0,\min(\delta_{0}/3,\delta_{1}))$ such that for any $\Delta\in\mathcal{T}^{(2)}$, $z\in\mathbb{D}^{2}_{\xi}$, $t\in(-T,T)$ and $\Gamma\subset N_{\zeta}(\Sigma)$ with $\mathbf{F}(\Gamma,\Sigma)<\delta$, we have
(4.19) $\mathcal{H}^{2}(\mathrm{Pri}_{\Delta,z,t}\cap\Gamma)<\frac{\delta_{0}}{3}\,.$
Fix $\Gamma\subset N_{\zeta}(\Sigma)$ with $g_{0}\leq\mathfrak{g}(\Gamma)\leq g_{1}$ and $\mathbf{F}(\Gamma,\Sigma)<\delta$ and closed disks $\{\overline{D_{i}}\}^{n}_{i=1}$ in $\Gamma$ with $\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D_{i})<\delta$ and $\sum^{n}_{i=1}\operatorname{diam}_{M}(\overline{D_{i}})<\delta$. Let $z_{1}\mathrel{\mathop{\ordinarycolon}}=\bar{z}\in\mathbb{D}^{2}_{\xi}$ and $t_{1}\mathrel{\mathop{\ordinarycolon}}=\bar{t}\in(-T,T)$ be determined by Step 1.6. For any $\Delta\in\mathcal{T}^{(2)}$, by (0.ii),  (4.7) and Step 1 (2), we have
(4.20) $\mathcal{H}^{1}(\mathrm{Cyl}_{\Delta,z_{1},t_{1}}\cap\Gamma)<\frac{\delta_{0}}{3}<\frac{\varepsilon}{2}\,,$
and by (4.19)
(4.21) $\mathcal{H}^{2}(\mathrm{Pri}_{\Delta,z_{1},t_{1}}\cap\Gamma)<\frac{\delta_{0}}{3}\,.$
For any $\Delta\in\mathcal{T}^{(2)}$, by Step 1 (1) and (3), we know that for any $v\in\Delta^{(0)}$,
$\mathrm{I}_{v,z_{1},t_{1}}\cap\Gamma=\mathrm{I}_{v,z_{1},t_{1}}\cap c_{\Delta,1}\,.$
Therefore, every short loop $c_{\Delta,k}$ with $k\geq 2$ lies in the interior of $\mathrm{R}_{e,z_{1},t_{1}}$ for some $e\in\Delta^{(1)}$. In particular, every short loop for $\Delta\in\mathcal{T}^{(2)}$ coincides with exactly one short loop in another $\Delta^{\prime}\in\mathcal{T}^{(2)}$.
In the following, we regard $\Gamma$ as an intrinsic surface.
First, we cut $\Gamma$ along every loop $\bigcup_{\Delta\in\mathcal{T}^{(2)}}\{c_{\Delta,k}\}^{n_{\Delta}}_{k=1}$. Then for any $\Delta\in\mathcal{T}^{(2)}$,
$\Gamma_{\Delta}\mathrel{\mathop{\ordinarycolon}}=\Gamma\cap\mathrm{Pri}_{\Delta,z_{1},t_{1}}$
is a surface with boundary; the boundary consists of smooth short loops $\{c_{\Delta,k}\}^{n_{\Delta}}_{k=1}$ and a piecewise smooth long loop $c_{\Delta,1}$. We glue disks $\{D_{\Delta,k}\}^{n_{\Delta}}_{k=1}$ of arbitrary small areas to its boundary and obtain a smooth closed orientable surface $\Gamma^{\prime}_{\Delta}$. By (4.20),  (4.21) and the choice of $\delta$, we have
$\sum^{n_{\Delta}}_{k=1}\mathcal{H}^{1}(\partial D_{\Delta,k})+\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D_{i})<\delta_{0},\quad\mathcal{H}^{2}(\Gamma^{\prime}_{\Delta})<\delta_{0}\,.$
Hence, for any $\Delta\in\mathcal{T}^{(2)}$, since $\mathfrak{g}(\Gamma^{\prime}_{\Delta})\leq g_{1}$, applying Proposition 4.1 to $\Gamma^{\prime}_{\Delta}$, we obtain $\mathfrak{g}(\Gamma^{\prime}_{\Delta})$ many disjoint loops $\{c^{\prime}_{\Delta,k}\}^{\mathfrak{g}(\Gamma^{\prime}_{\Delta})}_{k=1}$ in $\Gamma^{\prime}_{\Delta}\setminus\left(\bigcup^{n}_{i=1}\overline{D_{i}}\cup\bigcup^{n_{\Delta}}_{k=1}\overline{D_{\Delta,k}}\right)$, each of which has
(4.22) $\mathcal{H}^{1}(c^{\prime}_{\Delta,k})\leq\eta/2\,,\quad c^{\prime}_{\Delta,k}\subset\Gamma_{\Delta}\setminus\bigcup^{n}_{i=1}\overline{D_{i}}\,.$
Figure 11. $\Gamma_{\Delta}$ and $\Gamma^{\prime}_{\Delta}$
Then for any $\Delta\in\mathcal{T}^{(2)}$, we perform surgeries on $\Gamma^{\prime}_{\Delta}$ along $\{c^{\prime}_{\Delta,k}\}^{\mathfrak{g}(\Gamma^{\prime}_{\Delta})}_{k=1}$, and remove $D_{\Delta,1}$ whose boundary is the long loop $c_{\Delta,1}$. By Proposition 4.1, the resulting surface is a genus 0 closed surface, consisting of multiple spheres and a triangle $\widetilde{\Delta}$ with boundary $c_{\Delta,1}$. By Step 1 (1) and (3), one can glue any triangles $\{\widetilde{\Delta}\}_{\Delta\in\mathcal{T}^{(2)}}$ in the same manner as $\mathcal{T}^{(2)}=\{\Delta\}$ to obtain a smooth closed surface $\widetilde{\Gamma}$, which is homeomorphic to $\Sigma$.
$\mathfrak{g}(\widetilde{\Gamma})=\mathfrak{g}(\Sigma)=g_{0}\,.$
Therefore, we can perform surgeries on $\Gamma$ along the disjoint loops
$\bigcup_{\Delta\in\mathcal{T}}\{c_{\Delta,k}\}^{n_{\Delta}}_{k=1}\cup\{c^{\prime}_{\Delta,k}\}^{\mathfrak{g}(\Gamma^{\prime}_{\Delta})}_{k=1}$
to obtain a smooth genus $g_{0}$ closed surface $\widetilde{\Gamma}$ and multiple spheres. There are exactly $g=\mathfrak{g}(\Gamma)-g_{0}$ disjoint simple loops $\{c_{j}\}^{g}_{j=1}$ from $\bigcup_{\Delta\in\mathcal{T}}\{c_{\Delta,k}\}^{n_{\Delta}}_{k=1}\cup\{c^{\prime}_{\Delta,k}\}^{\mathfrak{g}(\Gamma^{\prime}_{\Delta})}_{k=1}$, each of which lies in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$ such that
  1. (2.i)
    For each $j\in\{1,2,\dots,g\}$, by (4.20) and (4.22), $\mathcal{H}^{1}(c_{j})\leq\varepsilon$.
  2. (2.ii)
    $\{[c_{j}]\}^{g}_{j=1}$ is linearly independent in $H_{1}(\Gamma;\mathbb{Z}_{2})$.
Step 3. General case
Let $\delta=\delta_{3}(M,\mathbf{g},\Sigma,g_{1},\varepsilon)\mathrel{\mathop{\ordinarycolon}}=\delta(M,\mathbf{g},\Sigma,\min(\delta_{2},\zeta))$ as in Lemma 4.9. Then by Lemma 4.9, for $\Gamma$ and $\{\overline{D_{i}}\}^{n}_{i=1}$ satisfying all the conditions in the proposition, there exists $\eta\in(0,\zeta)$ we can cut $\Gamma\pitchfork\partial N_{\eta}(\Sigma)=\bigcup^{n^{\prime}_{\Gamma}}_{j=1}c^{\prime}_{j}$ and attach disks to obtain two disjoint orientable closed surfaces $\Gamma^{1}$ and $\Gamma^{2}$ satisfying:
  1. (2.i)
    $\Gamma^{1}\subset N_{\zeta}(\Sigma)$.
  2. (2.ii)
    $\mathbf{F}(\Gamma^{1},\Sigma)<\delta_{2}$.
  3. (2.iii)
    $\mathcal{H}^{2}(\Gamma^{2})<\delta_{2}<=\delta_{0}$.
  4. (2.iv)
    $\sum^{n_{\Gamma}}_{j=1}\mathcal{H}^{1}(\partial D_{j})<\delta_{2}$ and $\sum^{n_{\Gamma}}_{j=1}\operatorname{diam}_{M}(\overline{D_{j}})<\delta_{2}$, where $\{\overline{D_{j}}\}^{n_{\Gamma}}_{j=1}$ is as defined in Lemma 4.9.
Applying Step 2 to $\Gamma^{1}$ and Proposition 4.1 to $\Gamma^{2}$ with closed disks $\bigcup^{n_{\Gamma}}_{j=1}\{\overline{D_{j}}\}$, we obtain $g^{\prime}=\mathfrak{g}(\Gamma^{1})+\mathfrak{g}(\Gamma^{2})-g_{0}$ disjoint simple loops $\{\widetilde{c}_{j}\}^{g^{\prime}}_{j=1}$ in $\Gamma\setminus\bigcup^{n_{\Gamma}}_{j=1}\overline{D_{j}}$. Moreover, if we view $\Gamma$ as an intrinsic surface and perform surgeries on $\Gamma$ along $\{\widetilde{c}_{j}\}^{g^{\prime}}_{j=1}\cup\{c^{\prime}_{j}\}^{n_{\Gamma}}_{j=1}$, we obtain a genus $g_{0}$ closed surface and multiple spheres as in Step 2. Therefore, as in Step 2, we can conclude the existence of $g=\mathfrak{g}(\Gamma)-g_{0}$ disjoint simple loops from $\{\widetilde{c}_{j}\}^{g^{\prime}}_{j=1}\cup\{c^{\prime}_{j}\}^{n_{\Gamma}}_{j=1}$ satisfying the proposition. ∎
4.3. Small balls in punctate surfaces
In the previous two subsections, we’ve proved how to detect simple short loops in a smooth surface. However, neither Proposition 4.1 or Proposition 4.7 can be applied to every surface in our Simon-Smith family (Definition 2.7), as not every surface is smooth therein.
In fact, for any $L\in\mathbb{R}^{+}$, one can construct a genus 111 punctate surface $T$ of area 111 but any nontrivial simple loop in $S$ has length at least $L$; see Appendix C for the sketch of such a construction. In particular, Proposition 4.1 fails for a punctate surface of genus 111. Fortunately, one can still find a nontrivial simple loop of $T$ that lies in a geodesic ball of uniformly bounded radius.
To generalize this statement to all punctate surfaces of any genus, we first prove the following useful lemma, which asserts that finitely many (not necessarily disjoint) small geodesic balls can be covered by relatively larger and pairwise disjoint geodesic balls in a controlled manner.
Lemma 4.11.
In an orientable closed Riemannian 333-manifold $(M,\mathbf{g})$, for any $q,n\in\mathbb{N}^{+}$, $\zeta\in\left(0,\frac{\operatorname{injrad}(M,\mathbf{g})}{2n(3n)^{q}}\right)$, and a finite set of geodesic balls $\{B(p_{i},\zeta)\}^{q}_{i=1}$ (not necessarily disjoint), there exist disjoint geodesic balls $\{B(p^{\prime}_{i},\zeta^{\prime})\}^{q^{\prime}}_{i=1}$ such that:
  1. (1)
    $q^{\prime}\leq q$, $\zeta^{\prime}=(3n)^{q-q^{\prime}}\zeta$ and $\{p^{\prime}_{i}\}^{q^{\prime}}_{i=1}\subset\{p_{i}\}^{q}_{i=1}$.
  2. (2)
    $\bigcup^{q}_{i=1}B(p_{i},\zeta)\subset\bigcup^{q^{\prime}}_{i=1}B(p^{\prime}_{i},\zeta^{\prime})$.
  3. (3)
    The closed balls $\left\{\overline{B(p^{\prime}_{i},n\zeta^{\prime})}\right\}^{q^{\prime}}_{i=1}$ are disjoint from each other.
Proof.
If the closed balls $\left\{\overline{B(p_{i},n\zeta)}\right\}^{q}_{i=1}$ are disjoint from each other, then the lemma holds trivially with $q^{\prime}=q$, $p^{\prime}_{i}=p_{i}$ and $\zeta^{\prime}=\zeta$.
Otherwise, there exist two indices $k\neq l\in\{1,2,\dots,q\}$ such that $\overline{B(p_{k},n\zeta)}\cap\overline{B(p_{l},n\zeta)}\neq\emptyset$. Then, we have
$B(p_{k},\zeta)\cup B(p_{l},\zeta)\subset B(p_{k},3n\zeta)\,.$
Set $\zeta_{1}\mathrel{\mathop{\ordinarycolon}}=3n\zeta$ and $\{p_{1,i}\}^{q-1}_{i=1}\mathrel{\mathop{\ordinarycolon}}=\{p_{i}\}^{q}_{i=1}\setminus\{p_{k}\}$. Note that we have
$\bigcup^{q}_{i=1}B(p_{j},\zeta)\subset\bigcup^{q-1}_{i=1}B(p_{1,j},\zeta_{1})\,.$
Inductively, we define $\zeta_{l+1}=3n\zeta_{l}$ provided that $\left\{\overline{B(p_{i},n\zeta_{l})}\right\}^{q-l}_{i=1}$ are not pairwise disjoint. We also define $\{p_{l+1,i}\}^{q-l-1}_{i=1}$ as a subset of $\{p_{l,i}\}^{q-l}_{i=1}$ by removing the center of a ball that intersects another as above.
Since the number of balls strictly decreases in the induction, this process stops at defining $\zeta_{l}$ for some $l\in\{1,\dots,q-1\}$. Hence, we can take
$\zeta^{\prime}=\zeta_{l},\quad q^{\prime}=q-l,\quad\{p^{\prime}_{i}\}^{q^{\prime}}_{i=1}=\{p_{l,i}\}^{q-l}_{i=1}\,,$
and complete the proof. ∎
Given two punctate surface $S_{1},S_{2}\in\mathcal{S}(M)$ with punctate sets $P_{1}$ and $P_{2}$, we say that $S_{1}$ and $S_{2}$ intersect transversally if
$S_{1}\cap P_{2}=S_{2}\cap P_{1}=\emptyset,\quad(S_{1}\setminus P_{1})\pitchfork(S_{2}\setminus P_{2})=\emptyset.$
We denote their intersection by $S_{1}\pitchfork S_{2}$.
We show how to approximate a puncate surface by a smooth surface.
Lemma 4.12.
In an orientable closed Riemannian 333-manifold $(M,\mathbf{g})$, suppose that $\Sigma$ is either an empty set or an embedded smooth orientable closed surface. For any $q\in\mathbb{N}^{+}$ and $\varepsilon\in\left(0,\frac{\operatorname{injrad}(M,\mathbf{g})}{2}\right)$, there exists a constant $\delta=\delta(M,\mathbf{g},\Sigma,q,\varepsilon)\in(0,\varepsilon)$ with the following property.
For any punctate surface $S\in\mathcal{S}(M)$ with a punctate set $P$ such that
#$\#P\leq q\,,\quad\mathbf{F}(S,\Sigma)<\delta\,,$
there exist an integer $\widetilde{q}\leq q$, a real number $r\in(0,\varepsilon)$ and geodesic balls $\{B(p_{i},r)\}^{\widetilde{q}}_{i=1}$ in $(M,\mathbf{g})$ such that:
  1. (1)
    The closed balls $\{\overline{B(p_{i},r)}\}^{\widetilde{q}}_{i=1}$ are pairwise disjoint geodesic balls.
  2. (2)
    For each $i\in\{1,2,\dots,\widetilde{q}\}$, $\partial B(p_{i},r)\pitchfork\Sigma$ is a union of finitely many loops.
  3. (3)
    $S\setminus\bigcup^{\widetilde{q}}_{i=1}B(p_{i},r)$ is an embedded smooth surface with boundary and genus at most $\mathfrak{g}(S)$.
  4. (4)
    There exists an embedded smooth orientable closed surface $\Gamma$ such that
    1. (a)
      $\Gamma\setminus\bigcup^{\widetilde{q}}_{i=1}B(p_{i},r)=S\setminus\bigcup^{\widetilde{q}}_{i=1}B(p_{i},r)$.
    2. (b)
      $\mathbf{F}(\Gamma,\Sigma)<\varepsilon$.
    3. (c)
      $\Gamma\cap\bigcup^{\widetilde{q}}_{i=1}\overline{B(p_{i},r)}$ is a union of disjoint closed disks $\{\overline{D_{j}}\}^{n}_{j=1}$ satisfying
      $\sum^{n}_{j=1}\mathcal{H}^{1}(\partial D_{j})<\varepsilon\,,\quad\sum^{n}_{j=1}\operatorname{diam}(\overline{D_{j}})<\varepsilon\,.$
Proof.
Since $\Sigma$ is either an empty set or a smooth closed surface, there exists a positive constant $C_{\Sigma}>1$ depending only on $M,\mathbf{g}$ and $\Sigma$, such that for any $p\in M$ and $r\in(0,\operatorname{injrad}(M,\mathbf{g})/2)$,
$\mathcal{H}^{2}(\overline{B(p,r)}\cap\Sigma)\leq C_{\Sigma}r^{2}\,,\quad\mathcal{H}^{3}(\overline{B(p,r)})\leq C_{\Sigma}r^{3}\,.$
By the Jordan-Schoenflies theorem and Lemma A.4, there exists $\nu_{M}>0$, depending only on $(M,\mathbf{g})$, such that for any $p\in M$ and $r\in(0,\operatorname{injrad}(M,\mathbf{g})/2)$, every simple loop $c$ in $\partial B(p,r)$ bounds a disk $D\subset\partial B(p,r)$ with
(4.23) $\mathcal{H}^{2}(D)\leq\nu_{M}\left(\mathcal{H}^{1}(c)\right)^{2}\,.$
Let $\eta_{1}=\delta(M,\mathbf{g},\Sigma,\varepsilon/2)/2$ be defined as in Corollary A.2. We choose $\eta_{2}\in(0,\operatorname{injrad}(M,\mathbf{g})/4)$ such that
$(2\nu_{M}+1)\eta^{2}_{2}<\eta_{1}\,,\quad qC_{\Sigma}(\eta_{2})^{3}<\eta_{1}\,,\quad\text{and }2\eta_{2}<\varepsilon\,.$
Let $C_{q}>1$ be a constant depending only on $q$ and determined later. And we set $\eta_{3}=\frac{\eta_{2}}{8qC_{q}C_{\Sigma}}$.
By the definition of $\mathbf{F}$-metric, there exists $\delta\in(0,\varepsilon/2)$ such that for any $S\in\mathcal{S}(M)$ with $\mathbf{F}(S,\Sigma)\leq\delta$ and any $p\in M$,
$\mathcal{H}^{2}(\overline{B(p,\eta_{3})}\cap S)\leq 2\mathcal{H}^{2}(\overline{B(p,2\eta_{3})}\cap\Sigma)\leq 2C_{\Sigma}(2\eta_{3})^{2}=8C_{\Sigma}(\eta_{3})^{2}\,.$
In the sequel, we fix a punctate surface $S$ with a punctate set $P$ satisfying $\mathbf{F}(S,\Sigma)\leq\delta$ and #$\#P\leq q$.
By Sard’s theorem and the coarea formula, we can choose a constant $C_{q}>0$, only depending on $q$, and a radius $\widetilde{r}\in\left(\eta_{3}/3^{q},2\eta_{3}/3^{q}\right)$ such that, for any $p\in P$ and any $l\in\{1,3,3^{2},\dots,3^{q}\}$, $\partial B(p,l\widetilde{r})\pitchfork S$ is a finite union of disjoint simple loops and
$\mathcal{H}^{1}(\partial B(p,l\widetilde{r})\cap S)\leq C_{q}\mathcal{H}^{2}(\overline{B(p,2\eta_{3})}\cap\Sigma)/\eta_{3}<8C_{q}C_{\Sigma}\eta_{3}=\eta_{2}/q\,.$
By Lemma 4.11 with $n=1$, there exist $P^{\prime}\subset P$ and a radius $r\in\{\widetilde{r},3\widetilde{r},3^{2}\widetilde{r},\dots,3^{q}\widetilde{r}\}\subset(0,\eta_{3})$ such that $\Gamma=S\setminus\bigcup_{p^{\prime}\in P^{\prime}}B(p^{\prime},r)$ is a smooth surface with boundary satisfying
$\sum_{p^{\prime}\in P^{\prime}}\mathcal{H}^{1}(\partial B(p^{\prime},r)\cap S)<(\#P^{\prime})\eta_{2}/q\leq\eta_{2}\,,$
$P\subset\bigcup_{p^{\prime}\in P^{\prime}}B(p^{\prime},r)$ and $\{\overline{B(p^{\prime},r)}\}_{p^{\prime}\in P^{\prime}}$ are disjoint from each other. Clearly, we also have $\mathfrak{g}(\Gamma)\leq\mathfrak{g}(S)$. These confirms (1)(2) and (3).
Then by (4.23), for each $p^{\prime}\in P^{\prime}$, every simple loop $c$ in $\partial B(p^{\prime},r)\cap S$ bounds a disk $D\subset\partial B(p^{\prime},r)$ with
$\mathcal{H}^{2}(D)\leq\nu_{M}\left(\mathcal{H}^{1}(c)\right)^{2}\,.$
In particular, for each $p^{\prime}\in P^{\prime}$, near $B(p^{\prime},r)$, we can remove $B(p^{\prime},r)\cap S$ and attach finitely many disks $\{D_{p^{\prime},j}\}^{n_{p^{\prime}}}_{j=1}$ near $\partial B(p^{\prime},r)$, yielding a smooth closed surface $\Gamma$ of genus $\mathfrak{g}(S)$, confirming (4)(a).
Furthermore, we have
$\sum_{p^{\prime}\in P^{\prime}}\sum^{m_{p^{\prime}}}_{j=1}\mathcal{H}^{2}(D_{p^{\prime},j})$ $\leq 2\sum_{p^{\prime}\in P^{\prime}}\sum^{m_{p^{\prime}}}_{j=1}\nu_{M}\left(\mathcal{H}^{1}(\partial D_{p^{\prime},j})\right)^{2}$
$\leq 2\nu_{M}\left(\sum_{p^{\prime}\in P^{\prime}}\sum^{m_{p^{\prime}}}_{j=1}\mathcal{H}^{1}(\partial D_{p^{\prime},j})\right)^{2}$
$\leq 2\nu_{M}(\eta_{2})^{2}\,.$
Therefore, we have
$|\mathcal{H}^{2}(\Gamma)-\mathcal{H}^{2}(S)|$ $\leq\sum_{p^{\prime}\in P^{\prime}}\sum^{m_{p^{\prime}}}_{j=1}\mathcal{H}^{2}(D_{p^{\prime},j})+\sum_{p^{\prime}\in P^{\prime}}\mathcal{H}^{2}(\overline{B(p^{\prime},r)}\cap S)$
$<2\nu_{M}(\eta_{2})^{2}+8qC_{\Sigma}(\eta_{3})^{2}$
$<\left(2\nu_{M}+1\right)(\eta_{2})^{2}<\eta_{1}\,,$
and
$\mathcal{F}([\Gamma],[S])\leq\sum_{p^{\prime}\in P^{\prime}}\mathcal{H}^{3}(B(p^{\prime},r))\leq qC_{\Sigma}r^{3}<qC_{\Sigma}(\eta_{2})^{3}<\eta_{1}\,.$
By the choice of $\eta_{1}$ and Corollary A.2, we have
$\mathbf{F}(\Gamma,S)<\varepsilon/2\,,$
and thus,
$\mathbf{F}(\Gamma,\Sigma)<\varepsilon\,,$
confirming (4)(b).
Finally, we have
$\sum_{p^{\prime}\in P^{\prime}}\sum^{m_{p^{\prime}}}_{j=1}\mathcal{H}^{1}(\partial D_{p^{\prime},j})=\sum_{p^{\prime}\in P^{\prime}}\mathcal{H}^{1}(\partial B(p^{\prime},r^{\prime})\cap S)<\eta_{2}<\varepsilon\,,$
and
$\sum_{p^{\prime}\in P^{\prime}}\sum^{m_{p^{\prime}}}_{j=1}\operatorname{diam}_{M}(\overline{D_{p^{\prime},j}})\leq 2\sum_{p^{\prime}\in P^{\prime}}\sum^{m_{p^{\prime}}}_{j=1}\mathcal{H}^{1}(\partial D_{p^{\prime},j})<2\eta_{2}<\varepsilon\,,$
confirming (4)(c). ∎
Here is a generalization of Proposition 4.1 to punctate surfaces, as an application of the previous approximation lemma.
Lemma 4.13 (Existence of small balls I).
In an orientable closed Riemannian 333-manifold $(M,\mathbf{g})$, for any integers $q,N\in\mathbb{N}^{+}$ and any real number $\varepsilon\in\left(0,\operatorname{injrad}(M,\mathbf{g})/(2N)\right)$, there exists a constant $\delta=\delta(M,\mathbf{g},q,N,\varepsilon)>0$ with the following property.
For any punctate surface $S\in\mathcal{S}(M)$ with a punctate set $P$ such that
$\mathfrak{g}(S)+\#P\leq q\,,\quad\mathcal{H}^{2}(S)<\delta\,,$
there exist an integer $\widetilde{q}\leq q$, a real number $r\in(\frac{\varepsilon}{2(3N)^{q}},\varepsilon)$ and geodesic balls $\{B(p_{i},r)\}^{\widetilde{q}}_{i=1}$ in $(M,\mathbf{g})$ such that:
  1. (1)
    The closed balls $\{\overline{B(p_{i},Nr)}\}^{\widetilde{q}}_{i=1}$ are pairwise disjoint geodesic balls.
  2. (2)
    For each $i\in\{1,2,\dots,\widetilde{q}\}$ and $l\in\{1,2,\dots,N\}$, $\partial B\left(p_{i},lr\right)\pitchfork S$ is a union of finitely many disjoint loops.
  3. (3)
    $S\setminus\bigcup^{\widetilde{q}}_{i=1}B\left(p_{i},r\right)$ is a smooth genus 0 surface with boundary.
Proof.
We set
(4.24) $\eta\mathrel{\mathop{\ordinarycolon}}=\frac{\varepsilon}{2(3N)^{q}}\,.$
Let $S$ be a punctate surface $S$ with a punctate set $P$ and #$\#P\leq q$. By Sard’s theorem and the coarea formula, there exist a constant $C=C(q)$ and a radius $r\in\left(\frac{\eta}{2\cdot 3^{q}},\frac{\eta}{3^{q}}\right)$ such that for any $p\in P$ and $l\in\{1,3,3^{2},\dots,3^{q}\}$, $\partial B(p,lr)\pitchfork S$ is a finite union of loops and
$\sum_{p\in P}\mathcal{H}^{1}(\partial B(p,lr)\cap S)\leq C\frac{\mathcal{H}^{2}(S)}{\eta}\,.$
By Lemma 4.11, there exist $P^{\prime}\subset P$ and $r^{\prime}\in\{r,3r,3r^{2},\dots,3^{q}r\}\subset(\frac{\eta}{2\cdot 3^{q}},\eta)$ such that $\Gamma=S\setminus\bigcup_{p^{\prime}\in P^{\prime}}B(p^{\prime},r^{\prime})$ is a smooth surface with boundary satisfying
(4.25) $\sum_{p^{\prime}\in P^{\prime}}\mathcal{H}^{1}(\partial B(p^{\prime},r^{\prime})\cap S)\leq C\frac{\mathcal{H}^{2}(S)}{\eta}\,,\quad P\subset\bigcup_{p^{\prime}\in P^{\prime}}B(p^{\prime},r^{\prime})\,,$
and $g\mathrel{\mathop{\ordinarycolon}}=\mathfrak{g}(\Gamma)\leq\mathfrak{g}(S)$.
In the sequel, we regard $\Gamma$ as an intrinsic surface, rather than a surface embedded in $(M,\mathbf{g})$. We also let $\widetilde{\delta}=\delta(q,\varepsilon)$ be the constant as determined in Proposition 4.1.
By (4.25), taking $\delta\in(0,\widetilde{\delta}/2)$ sufficiently small, if we further assume that $\mathcal{H}^{2}(S)<\delta$, then we can attach disks $\{D_{i}\}^{n}_{i=1}$ with arbitrarily small area to the boundary of $S\setminus\bigcup_{p^{\prime}\in P^{\prime}}B(p^{\prime},r^{\prime})$ to obtain a smooth genus $g$ closed surface $\Sigma$, and
(4.26) $\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D_{i})\leq\widetilde{\delta},\quad\mathcal{H}^{2}(\Gamma)\leq 2\mathcal{H}^{2}(S)<2\delta<\widetilde{\delta}\,.$
Moreover, it follows from Proposition 4.1 with (4.26), there exists $g$ disjoint simple loops $\{c_{j}\}^{g}_{j=1}$ in $\Sigma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}\subset S$ such that
  1. (i)
    For each $j\in\{1,\dots,g\}$, each $c_{j}$ satisfies $\mathcal{H}^{1}(c_{j})\leq\eta$.
  2. (ii)
    $\{[c_{j}]\}^{g}_{j=1}$ is linearly independent in $H_{1}(\Sigma)$.
In particular, each $c_{j}$ is contained in some ball $B(p_{j},\eta)$. Applying Sard’s theorem again, there exists $\eta^{\prime}\in(\eta,2\eta)$ such that for any integer $l\in\{1,2,3,\dots,N(3N)^{q}\}$, each $B(p_{j},l\eta^{\prime})$ with $j\in\{1,2,\dots,g\}$ intersects $S$ transversally, and so does each $B(p^{\prime},l\eta^{\prime})$ with $p^{\prime}\in P^{\prime}$.
In the sequel, we denote $\{B(p,\eta^{\prime})\}_{p\in P}\cup\{B(p_{j},\eta^{\prime})\}^{g}_{j=1}$ by $\{B(p_{j},\eta^{\prime})\}^{q^{\prime}}_{j=1}$, where $q^{\prime}=g+\#P^{\prime}\leq q$. Applying Lemma 4.11 to $\{B(p_{j},\eta^{\prime})\}^{q^{\prime}}_{j=1}$ with $n=N$, we obtain $\widetilde{q}\leq q^{\prime}$, $\{p^{\prime}_{j}\}^{\widetilde{q}}_{j=1}\subset\{p_{j}\}^{q^{\prime}}_{j=1}$, and a radius $\widetilde{r}=(3N)^{q^{\prime}-\widetilde{q}}\eta^{\prime}\in(\frac{\varepsilon}{2(3N)^{q}},\varepsilon)$, such that the closed balls $\{\overline{B(p^{\prime}_{j},N\widetilde{r})}\}^{\widetilde{q}}_{j=1}$ are disjoint from each other. This confirms (1).
Finally, the transversality result  (2) follows from the choice of $\eta^{\prime}$, and the genus result  (3) follows from the choice of the balls. ∎
Similarly, we can extend Proposition 4.7 to punctate surfaces.
Lemma 4.14 (Existence of small balls II).
In a 333-dimensional orientable closed Riemannian 333-manifold $(M,\mathbf{g})$, for any $q,N\in\mathbb{N}^{+}$, $g_{0}\in\mathbb{N}$, $\mathscr{S}_{g_{0}}$ a compact subset of embedded smooth orientable genus $g_{0}$ closed surfaces into $(M,\mathbf{g})$ (endowed with smooth topology), $\varepsilon\in(0,\operatorname{injrad}(M,\mathbf{g})/(2N))$, there exists $\delta=\delta(M,\mathbf{g},\mathscr{S}_{g_{0}},q,N,\varepsilon)>0$ with the following property.
For any punctate surface $S\in\mathcal{S}(M)$ with punctate set $P$ satisfies
$\mathfrak{g}(\Sigma)-g_{0}+\#P\leq q$
and any smooth surface $\Sigma\in\mathscr{S}_{\mathfrak{g}_{0}}$ with $\mathbf{F}(S,\Sigma)\leq\delta$, there exist an integer $\widetilde{q}\leq q$, a real number $r\in\left(\frac{\varepsilon}{2(3N)^{q}},\varepsilon\right)$ and geodesic balls $\{B(p_{i},r)\}^{\widetilde{q}}_{i=1}$ in $(M,\mathbf{g})$ such that:
  1. (1)
    The closed balls $\{\overline{B(p_{i},Nr)}\}^{\widetilde{q}}_{i=1}$ are pairwise disjoint geodesic balls.
  2. (2)
    For each $i\in\{1,2,\dots,\widetilde{q}\}$ and $l\in\{1,2,\dots,N\}$, $\partial B(p_{i},lr)\pitchfork\Sigma$ is a union of finitely many loops.
  3. (3)
    $S\setminus\bigcup^{\widetilde{q}}_{i=1}B(p_{i},r)$ is a smooth genus $g_{0}$ surface with boundary.
Proof.
Step 1
We prove the lemma when $\mathscr{S}_{g_{0}}=\{\Sigma\}$ is a singleton set.
Let $\widetilde{\delta}_{1}(\eta)=\delta(M,\mathbf{g},\Sigma,q,\eta)\in(0,\eta)$ be as determined in Proposition 4.7. We set
$\eta_{1}$ :$\mathrel{\mathop{\ordinarycolon}}=\frac{\varepsilon}{2(3N)^{q}}\,,$
$\eta_{2}$ :$\mathrel{\mathop{\ordinarycolon}}=\widetilde{\delta}_{1}(\eta_{1})/2\,.$
Let $\widetilde{\delta}_{2}=\delta(M,\mathbf{g},\Sigma,q,\eta_{2})$ be as in Lemma 4.12. If $\mathbf{F}(S,\Sigma)<\delta$, then there exist a finite set $P^{\prime}\subset P$, a radius $r^{\prime}\in(0,\eta_{2})$ and an embedded smooth orientable closed surface $\Gamma$ and a finite set of closed disks $\{\overline{D}_{j}\}^{n}_{j=1}$ in $\Gamma$ such that:
  1. (1)
    $\Gamma\setminus\bigcup^{n}_{j=1}\overline{D_{j}}=S\setminus\bigcup_{p^{\prime}\in P^{\prime}}\overline{B(p^{\prime},r^{\prime})}$ and thus, $g_{1}\mathrel{\mathop{\ordinarycolon}}=\mathfrak{g}(\Gamma)\leq\mathfrak{g}(S)$.
  2. (2)
    $\mathbf{F}(\Gamma,S)\leq 2\eta_{2}$.
  3. (3)
    $\sum^{n}_{j=1}\operatorname{diam}_{M}(\overline{D_{j}})<\eta_{2}$.
  4. (4)
    $\sum^{n}_{j=1}\mathcal{H}^{1}(\overline{D_{j}})<\eta_{2}$.
It follows from Proposition 4.7 that there exist $g=g_{1}-\mathfrak{g}(\Sigma)$ disjoint simple loops $\{c_{j}\}^{g}_{k=1}$ in $\Gamma^{\prime}\setminus\bigcup^{n}_{j=1}\overline{D_{j}}\subset S$, which are homologically independent in $H_{1}(\Gamma^{\prime})$. Moreover, each $c_{k}$ satisfies
$\mathcal{H}^{1}(c_{k})\leq\eta_{1}\,,$
so there exists $\widetilde{p}_{k}\in M$ and $c_{k}\subset B(\widetilde{p}_{k},\eta_{1})$.
We set $P^{\prime\prime}\mathrel{\mathop{\ordinarycolon}}=P^{\prime}\cup\{\widetilde{p}_{k}\}^{g}_{k=1}$. By Sard’s Lemma, there exists $\widetilde{r}\in(\eta_{1},2\eta_{1})=\left(\frac{\eta}{4N(3N)^{q}},\frac{\eta}{2N(3N)^{q}}\right)$ such that for every $p\in P^{\prime\prime}$, $l\in\{1,2,\dots,N(3^{N})\}$,
$\partial B(p,l\widetilde{r})\pitchfork\Sigma$
is a union of finitely many loops. Then applying Lemma 4.11 again to #$\#P^{\prime\prime}$ balls
$\{B(p,\widetilde{r})\}_{p\in P^{\prime\prime}}$
with $n=N$, we obtain $\widetilde{q}\leq\#P^{\prime\prime}\leq q$, $r\in(\frac{\varepsilon}{2(3N)^{q}},\varepsilon)$ and disjoint geodesic balls $\{B(p_{i},r)\}^{\widetilde{q}}_{i=1}$ satisfying  (1)(2)(3) of the lemma. In particular, we have defined $\delta(M,\mathbf{g},\varepsilon,q,N,\{\Sigma\})\mathrel{\mathop{\ordinarycolon}}=\widetilde{\delta}_{2}$ for any embedded smooth orientable closed genus $g_{0}$ surface $\Sigma$.
Step 2
Let $\mathscr{S}_{g_{0}}$ be a compact set with respect to the smooth topology.
Since the smooth topology is finer than the topology induced by the $\mathbf{F}$-distance defined in Definition 2.4, $\mathscr{S}_{g_{0}}\subset\mathcal{S}$ is a compact set with respect to the $\mathbf{F}$-distance. For each $\Sigma\in\mathscr{S}_{g_{0}}$, we define
$B_{\Sigma}\mathrel{\mathop{\ordinarycolon}}=\left\{\Gamma\in\mathscr{S}_{g_{0}}\mathrel{\mathop{\ordinarycolon}}\mathbf{F}(\Sigma,\Gamma)<\frac{\delta(M,\mathbf{g},\{\Sigma\},q,N\varepsilon)}{4}\right\}\,,$
where $\delta(M,\mathbf{g},\{\Sigma\},q,N,\varepsilon)$ is already defined in Step 1. Then the open balls
$\{B_{\Sigma}\}_{\Sigma\in\mathscr{S}_{g_{0}}}$
is an open covering of $\mathscr{S}_{g_{0}}$. By the compactness, it has a finite subcovering
$\{B_{\Sigma_{l}}\}^{m}_{l=1}.$
Therefore, we can define
$\delta(M,\mathbf{g},\mathscr{S}_{g_{0}},q,N,\varepsilon)=\min_{l\in\{1,2,\dots,m\}}\frac{\delta(M,\mathbf{g},\{\Sigma_{l}\},q,N,\varepsilon)}{4}\,,$
and then  (1)(2)(3) of the lemma hold. ∎
5. Pinch-off processes and interpolations
Let $g_{0}\leq g_{1}$ be two natural numbers. Intuitively, for a Simon-Smith family of genus $\leq g_{1}$ sufficiently close to a smooth genus $g_{0}$ closed surface $\Sigma$, we show that one can pinch all the small necks and shrink all small connected components in a continuous way to obtain a Simon-Smith family of genus $g_{0}$. Note that by Lemma 2.6, each surface in the Simon-Smith family has genus at least $g_{0}$.
With ingredients in this section, we will be able to prove Theorem 2.21, and Proposition 2.22 and 2.23.
Proposition 5.1 (Pinch-off proposition I).
Let $(M,\mathbf{g})$ be an orientable closed Riemannian 333-manifold, $X$ be a cubical subcomplex of $I(m,k)$, $g_{0}\in\mathbb{N}$, $\mathscr{S}_{g_{0}}$ be a compact subset of embeddings of a genus $g_{0}$ surface into $M$ (endowed with smooth topology), $q\in\mathbb{N}$ and $\varepsilon\in(0,1)$. Then there exists $\delta=\delta(M,\mathbf{g},\mathscr{S}_{g_{0}},m,q,\varepsilon)\in(0,\varepsilon)$ with the following property.
If $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ is a Simon-Smith family of genus $\leq g_{1}$ with
$N(P_{\Phi})+g_{1}-g_{0}\leq q\,,$
and $\Sigma\mathrel{\mathop{\ordinarycolon}}X\to\mathscr{S}_{g_{0}}$ is a map such that for every $x\in X$,
$\mathbf{F}(\Phi(x),\Sigma(x))\leq\delta\,,$
then there exists a Simon-Smith family $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}(M)$ of genus $\leq g_{1}$ such that for any $x\in X$:
  1. (1)
    $H(0,x)=\Phi(x)$, $\mathfrak{g}(H(1,x))=g_{0}$.
  2. (2)
    For any $0\leq t\leq t^{\prime}\leq 1$,
    $\mathcal{H}^{2}(\Phi(x))-\varepsilon\leq\mathcal{H}^{2}(H(t,x))\leq\mathcal{H}^{2}(\Phi(x))+\varepsilon\,,$
    $\mathcal{F}([H(t,x)],[\Phi(x)])\leq\varepsilon\,,$
    and
    $\mathfrak{g}(H(t^{\prime},x))\leq\mathfrak{g}(H(t,x))\,.$
  3. (3)
    The Simon-Smith family $\{H(x,t)\}_{t\in[0,1]}$ is a pinch-off process.
Moreover, if $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ consists of punctate surfaces of small area, we can perform pinch-off processes to obtain a Simon-Smith family of genus 0. This can be viewed as a special case of the previous proposition by taking $g_{0}=0$ and ${}\mathscr{S}_{0}=\{\emptyset\}$.
Proposition 5.2 (Pinch-off proposition II).
Let $(M,\mathbf{g})$ be an orientable closed Riemannian 333-manifold, $X$ be a cubical subcomplex of $I(m,k)$, $q\in\mathbb{N}$ and $\varepsilon\in(0,1)$. Then there exists $\delta=\delta(M,\mathbf{g},q,\varepsilon)\in(0,\varepsilon)$ with the following property.
If $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ is a Simon-Smith family of genus $\leq g_{1}$ with
$N(P_{\Phi})+g_{1}\leq q\,,$
and
$\sup_{x\in X}\mathcal{H}^{2}(\Phi(x))\leq\delta\,,$
then there exists a Simon-Smith family $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}(M)$ of genus $\leq g_{1}$ such that for any $x\in X$:
  1. (1)
    $H(0,x)=\Phi(x)$, $\mathfrak{g}(H(1,x))=0$.
  2. (2)
    For any $0\leq t\leq t^{\prime}\leq 1$,
    $\mathcal{H}^{2}(H(t,x))\leq\mathcal{H}^{2}(\Phi(x))+\varepsilon\,,\quad\mathfrak{g}(H(t^{\prime},x))\leq\mathfrak{g}(H(t,x))\,.$
  3. (3)
    The Simon-Smith family $\{H(x,t)\}_{t\in[0,1]}$ is a pinch-off process.
5.1. Local deformations
Recall the following local pinch-off processes, Pinching and Shrinking, introduced in [CL24].
Lemma 5.3 (Local Deformation A: Pinching).
In an orientable closed Riemannian 333-manifold $(M,\mathbf{g})$, there exists a constant $C_{\mathbf{g}}$ with the following property.
Let $\Lambda>0$, $r\in\left(0,\mathrm{injrad}(M,\mathbf{g})/4\right)$ and $p\in M$. Suppose that $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ is a Simon-Smith family and $x_{0}\in X$ so that
$\mathcal{H}^{2}(\Phi(x_{0})\cap A(p;r,2r))\leq\Lambda r^{2}\,.$
Then for any $\eta>0$, there exist $s\in(r,2r)$, $\zeta\in(0,\min\{\Lambda r,s-r,2r-s\}/5)$, an integer $K\geq 0$, a neighborhood $O_{x_{0}}$ of $x_{0}$ in $X$, and a Simon-Smith family
$H\mathrel{\mathop{\ordinarycolon}}[0,1]\times O_{x_{0}}\to\mathcal{S}(M)\,,$
with the following properties.
  1. (1)
    For any $y\in O_{x_{0}}$, $\Sigma_{y}\mathrel{\mathop{\ordinarycolon}}=\Phi(y)\cap A(p;s-5\zeta,s+5\zeta)$ is a smooth surface, varying smoothly in $y$.
  2. (2)
    For any $s^{\prime}\in[s-4\zeta,s+4\zeta]$ and $y\in O_{x_{0}}$, $\Phi(y)$ intersects $\partial B(p,s^{\prime})$ transversally, and
    $\Phi(y)\cap\partial B(p,s^{\prime})=\mathrel{\mathop{\ordinarycolon}}\bigsqcup_{i=1}^{K}\gamma_{i}(s^{\prime},y)\,,\quad\sum_{i=1}^{K}\mathcal{H}^{1}(\gamma_{i}(s^{\prime},y))\leq 2\Lambda r\,,$
    where each $\gamma_{i}(s^{\prime},y)$ is a simple loop.
If $K=0$, then $H(t,y)\equiv\Phi(y)$ holds true for any $y\in O_{x_{0}}$ and $t\in[0,1]$; and if $K\geq 1$, then:
  1. (3)
    For any $y\in O_{x_{0}}$, $H(0,y)=\Phi(y)$.
  2. (4)
    There exist $0=t_{0}<t_{1}<t_{2}<\dots<t_{K}<t_{K+1}=1$ such that for any $0\leq i\leq K$, $t\in(t_{i},t_{i+1})$ and $y\in O_{x_{0}}$,
    $\Sigma_{t,y}\mathrel{\mathop{\ordinarycolon}}=H(t,y)\cap A(p;s-5\zeta,s+5\zeta)$
[…]
    $H(1,y)\cap A(p;s-\zeta,s+\zeta)=\emptyset\,.$
  3. (5)
    For any $i\in\{1,2,\dots,K\}$ and $y\in O_{x_{0}}$, near $t_{i}$, $H(t,y)$ is a surgery process via pinching the cylinder $\bigcup_{s^{\prime}\in(s-2\zeta,s+2\zeta)}\gamma_{i}(s^{\prime},y)$.
  4. (6)
    For any $y\in O_{x_{0}}$ and $t\in[0,1]$, we have
    $H(t,y)\setminus A(p;s-3\zeta,s+3\zeta)$ $=\Phi(y)\setminus A(p;s-3\zeta,s+3\zeta)\,,$
    $\mathcal{H}^{2}(\Phi(y)\cap A(p;s-3\zeta,s+3\zeta))-\eta$ $\leq\mathcal{H}^{2}(H(t,y)\cap A(p;s-3\zeta,s+3\zeta))$
    $\leq\mathcal{H}^{2}(\Phi(y)\cap A(p;s-3\zeta,s+3\zeta))+C_{\mathbf{g}}\Lambda^{2}r^{2}\,,$
    and
    $\mathcal{F}([\Phi(y))],[H(t,y)])\leq\eta\,.$
Remark 5.4.
For every $y\in O_{x_{0}}$, by (5), $\{H(t,y)\}_{t\in[0,1]}$ is a pinch-off process.
Sketch of proof.
[…]
Lemma 5.5 (Local Deformation B: Shrinking, [CL24, Lemma 8.4]).
In an orientable closed Riemannian 333-manifold $(M,\mathbf{g})$, suppose that $p\in M$ and $0<r^{-}<r^{+}<\mathrm{injrad}(M,\mathbf{g})/4$. Then there exists a smooth one-parameter family of maps $\{\mathcal{R}_{t}\mathrel{\mathop{\ordinarycolon}}M\to M\}_{t\in[0,1]}$ such that:
  1. (1)
    $\mathcal{R}_{t}=\mathrm{id}$ for any $t\in[0,1/4]$; $\mathcal{R}_{t}(B(p,r^{-}))=\{p\}$ for any $t\in[3/4,1]$.
  2. (2)
    $\mathcal{R}_{t}$ is a diffeomorphism for any $t\in[0,3/4)$.
  3. (3)
    $\mathcal{R}_{t}|_{M\setminus B(p,r^{+})}=\mathrm{id}$ and $\mathcal{R}_{t}(B(p,r^{\pm}))\subset B(p,r^{\pm})$ for any $t\in[0,1]$.
  4. (4)
    $\mathcal{R}_{t}|_{B(p,r^{-})}$ is 111-Lipschitz for any $t\in[0,1]$.
Lemma 5.6 (Local Deformation C: Successive shrinkings, [CL24, Lemma 8.5]).
In an orientable closed Riemannian 333-manifold $(M,\mathbf{g})$, for $N\in\mathbb{N}^{+}$, suppose that $\{A(p_{j};r_{j}^{-},r_{j}^{+})\}_{j=1}^{N}$ is a collection of pairwise disjoint annuli in $M$, where $0<r_{j}^{-}<r_{j}^{+}<\mathrm{injrad}(M,\mathbf{g})/4$. For any $j\in\{1,2,\dots,N\}$, let $\mathcal{R}^{j}_{t}\mathrel{\mathop{\ordinarycolon}}M\to M$ be a map from Lemma 5.5 with $p_{j},r_{j}^{\pm}$ in place of $p,r^{\pm}$. For simplicity, we denote
$A_{j}\mathrel{\mathop{\ordinarycolon}}=A(p_{j};r^{-}_{j},r^{+}_{j})\,,\quad B^{\pm}_{j}\mathrel{\mathop{\ordinarycolon}}=B(p_{j},r^{\pm}_{j})\,.$
Then the following hold:
  1. (1)
    There exists a permutation $\sigma\in\mathfrak{S}_{N}$ such that for any $1\leq i<j\leq N$, either $B_{\sigma(i)}^{+}\cap B_{\sigma(j)}^{+}=\emptyset$, or $B_{\sigma(i)}^{+}\subset B_{\sigma(j)}^{-}$. In this case, we call such a $\sigma$ admissible.
  2. (2)
    For any admissible $\sigma$, the map
    $\mathcal{R}^{\sigma(N)}_{t}\circ\mathcal{R}^{\sigma(N-1)}_{t}\circ\cdots\circ\mathcal{R}^{\sigma(2)}_{t}\circ\mathcal{R}^{\sigma(1)}_{t}\mathrel{\mathop{\ordinarycolon}}M\to M$
    is 111-Lipschitz on every connect component of $M\setminus\bigcup_{j=1}^{N}A_{j}$; and it is independent of the choice of admissible $\sigma$.
5.2. Combinatorial arguments
Recall the following combinatorial arguments in [CL24], inspired by [Pit81, Lemma 4.8, Proposition 4.9] and [DR18, Lemma 5.7].
Lemma 5.7 (Combinatorial Argument I, [CL24, §8.3]).
For any positive integers $m\in\mathbb{N}^{+}$ and $q\in\mathbb{N}^{+}$, there exists $N=N(m,q)\in\mathbb{N}^{+}$ with the following property.
Given any $k\in\mathbb{N}^{+}$ and a cubical complex $I(m,k)$, suppose that $\{\mathcal{F}_{\sigma}\}_{\sigma\in I(m,k)}$ is a family of collections of open sets in a closed manifold $M$ assigned to each cell $\sigma$ of $I(m,k)$, where each collection $\mathcal{F}_{\sigma}$ has the form
$\mathcal{F}_{\sigma}=(\mathcal{O}_{\sigma,1},\mathcal{O}_{\sigma,2},\dots,\mathcal{O}_{\sigma,q})$
and each $\mathcal{O}_{\sigma,i}=\{U^{1}_{\sigma,i},U^{2}_{\sigma,i},\dots,U^{N}_{\sigma,i}\}$ satisfies
(5.1) $\operatorname{dist}(U^{r}_{\sigma,i},U^{s}_{\sigma,i})\geq 2\min\{\mathrm{diam}(U^{r}_{\sigma,i}),\mathrm{diam}(U^{s}_{\sigma,i})\}\,,$
for all $\sigma\in I(m,k)$, $i\in\{1,2,\dots,q\}$ and $r,s\in\{1,2,\dots,N\}$ with $r\neq s$. Then we can extract a family of open sets
$\{(U_{\sigma,1},U_{\sigma,2},\dots,U_{\sigma,q})\}_{\sigma\in I(m,k)}$
where $U_{\sigma,i}\in\mathcal{O}_{\sigma,i}$ such that
$\operatorname{dist}(U_{\sigma,i}\cap U_{\tau,j})>0\,,$
whenever $\sigma,\tau\in I(m,k)$, $(\sigma,i)\neq(\tau,j)$ and $\sigma,\tau$ are faces of a common cell $\gamma$ of $I(m,k)$.
Moreover, this is also true if the number of collections in each $\mathcal{F}_{\sigma}$ is no greater than $q$.
Lemma 5.8 (Combinatorial argument II, [CL24, §8.3]).
In a closed Riemannian manifold $(M,\mathbf{g})$, given any integers $q,N\in\mathbb{N}^{+}$ and any $R_{0}\in(0,\mathrm{injrad}(M,\mathbf{g})/4)$, if $P$ is a finite set of at most $q$ points, then there exists a radius $R\in(5^{-2Nq^{2}}R_{0},5^{-2N}R_{0})$ such that for every $p\in P$,
$A(p;R,5^{2N}R)\cap P=\emptyset\,.$
5.3. Proof of Proposition 5.1 and Proposition 5.2
The proof will proceed in three steps.
Step 1. Set up
Let $C_{\mathbf{g}}$ be as in Lemma 5.3 and $N\mathrel{\mathop{\ordinarycolon}}=N(m,q)$ as in Lemma 5.7.
In Proposition 5.1, since $\mathscr{S}_{g_{0}}$ is a compact set of embedded smooth closed surfaces, without loss of generality, we assume that
$\varepsilon<\min\left\{\mathrm{injrad}(M,\mathbf{g}),1\right\}\,,$
such that for any $r\in(0,\varepsilon/1000)$,
$\sup_{p\in M}\mathcal{H}^{3}(B(p,2r))<r^{2}\,.$
and
(5.2) $\sup_{\Sigma\in\mathscr{S}_{g_{0}},\ p\in M}\left\{\mathcal{H}^{2}\left(\Sigma\cap B(p,2r)\right)\right\}\leq 50r^{2}\,.$
We set
$\varepsilon_{1}\mathrel{\mathop{\ordinarycolon}}=\frac{\varepsilon}{10000Nq\cdot 3^{m}\max(C_{\mathbf{g}},1)}\,,$
and
$\varepsilon_{2}\mathrel{\mathop{\ordinarycolon}}=\frac{\varepsilon_{1}}{2(3N)^{q}q^{2}\cdot 5^{2N}}\,.$
By (5.2), we choose $\delta_{0}=\delta_{0}(\mathscr{S}_{g_{0}})>0$ such that for any $S\in\mathcal{S}(M)$, if $\mathbf{F}(S,\Sigma)<\delta_{0}$ for some $\Sigma\in\mathscr{S}_{g_{0}}$, then for any $r\in(\varepsilon_{2},\varepsilon_{1})$,
(5.3) $\sup_{p\in M}\left\{\mathcal{H}^{2}\left(S\cap B(p,2r)\right)\right\}\leq 100r^{2}\,.$
and in particular,
(5.4) $\sup_{p\in M}\left\{\mathcal{H}^{2}\left(S\cap A(p;r,2r)\right)\right\}\leq 100r^{2}\,.$
Note that these inequalities also hold for Proposition 5.2.
In the following, we choose $\delta_{1}\mathrel{\mathop{\ordinarycolon}}=\delta(M,\mathbf{g},q,N,\varepsilon_{2})$ as defined in Lemma 4.13 and $\delta_{2}\mathrel{\mathop{\ordinarycolon}}=\delta(M,\mathbf{g},\mathscr{S}_{g_{0}},q,N,\varepsilon_{2})$ in Lemma 4.14. Set
$\delta\mathrel{\mathop{\ordinarycolon}}=\min\{\delta_{0},\delta_{1},\delta_{2}\}\,.$
For a Simon-Smith family $\Phi$ as in Proposition 5.1 or Proposition 5.2, by Lemma 4.13 or Lemma 4.14 with $N=1$ therein, for any $x\in X$, there exist $r_{x}\in(\varepsilon_{2},2(3N)^{q}\varepsilon_{2})$ and a finite set of balls $\{B(p_{x,i},r_{x})\}^{q_{x}}_{i=1}$ such that
  1. (i)
    $q_{x}\leq q$;
  2. (ii)
    $\Sigma\setminus\bigcup^{q_{x}}_{i=1}B(p_{x,i},r_{x})$ is a smooth surface of genus $g_{0}$ with boundary.
Note that when $\Phi$ is a family as in Proposition 5.2, $g_{0}=0$.
We write
$\widetilde{P}(x)\mathrel{\mathop{\ordinarycolon}}=\{p_{x,i}\}^{q_{x}}_{i=1}$
and for every $l\in\{1,2,\dots,N\}$,
$r_{x,l}\mathrel{\mathop{\ordinarycolon}}=5^{2(l-1)}r_{x}\,.$
By (5.3) and (5.4), for every $1\leq l\leq N$, $x\in X$ and $p\in\widetilde{P}(x)$, since $r_{x,l}\in(\varepsilon_{2},\varepsilon_{1})$, we have
(5.5) $\mathcal{H}^{2}\left(\Phi(x)\cap B(p,2r_{x,l})\right)\leq 100r_{x,l}^{2}\,,$
and
(5.6) $\mathcal{H}^{2}\left(\Phi(x)\cap A(p;r_{x,l},2r_{x,l})\right)\leq 100r_{x,l}^{2}\,.$
Thus, applying Lemma 5.3 with $\eta=\varepsilon/2$ to $\Phi(x)$ in $A(p;r_{x,l},2r_{x,l})$ with $\Lambda=100$, we obtain the following data:
  1. (1)
    $s_{x,l,p}\in(r_{x,l},2r_{x,l})$,
  2. (2)
    $\zeta_{x,l,p}\in(0,\min\{100r_{x,l},s_{x,l}-r_{x,l},2r_{x,l}-s_{x,l}\}/5)$,
  3. (3)
    a neighborhood $O_{x,l,p}\subset X$ of $x$,
  4. (4)
    and a Simon-Smith family
    $H_{x,l,p}\mathrel{\mathop{\ordinarycolon}}[0,1]\times O_{x,l,p}\to\mathcal{S}(M)\,.$
In addition, for each $x\in X$, we fix a neighborhood of $x$
$O_{x}\subset\bigcap_{l\in\{1,2,\dots,N\},\ p\in\widetilde{P}(x)}O_{x,l,p}\,,$
such that $y\mapsto\Phi(y)\setminus\bigcup_{p\in\widetilde{P}(x)}B(p,r_{x})$ is continuous in the smooth topology for $y\in O_{x}$.
Step 2. Refinement
[…]
Then we refine $X$ and assign to each cell $\sigma$ of $X$ $x_{\sigma}\in\{x_{j}\}$ such that: Every cell $\tau$ of $X$ is a subset of $O_{x_{\sigma}}$ provided that $\sigma$ and $\tau$ are faces of some cell $\gamma$ of $X$. Note that for each $\sigma$, the number of such $\tau$ is no more than $5^{m}$.
By Lemma 5.7 with
$\mathcal{F}_{\sigma}=(\{A(p;r_{x_{\sigma},i},2r_{x_{\sigma},i})\}^{N}_{i=1})_{p\in\widetilde{P}(x_{\sigma})}\,,$
we obtain, associated with each $\sigma$, a collection of annuli denoted by
$(A_{\sigma,p}\mathrel{\mathop{\ordinarycolon}}=A(p,r_{\sigma,p},2r_{\sigma,p})_{p\in\widetilde{P}(x_{\sigma})}$
such that whenever $\sigma$ and $\tau$ are faces of some cell $\gamma$ of $X$ and for any $p\in\widetilde{P}(x_{\sigma})$ and $p^{\prime}\in\widetilde{P}(x_{\tau})$, we have
$A_{\sigma,p}\cap A_{\tau,p^{\prime}}=\emptyset\,,$
unless $(\sigma,p)=(\tau,p^{\prime})$.
Step 3. Construction of $H$
For each cell $\sigma$ and $p\in\widetilde{P}(x_{\sigma})$, for convenience, we denote the data associated with each $A_{\sigma,p}$, constructed at the end of Step 1, as follows:
  1. (1)
    $s_{\sigma,p}\in(r_{\sigma,p},2r_{\sigma,p})$,
  2. (2)
    $\zeta_{\sigma,p}\in(0,\min\{s_{\sigma,p}-r_{\sigma,p},2r_{\sigma,p}-s_{\sigma,p}\}/5)$,
  3. (3)
    a neighborhood $O_{\sigma}\mathrel{\mathop{\ordinarycolon}}=O_{x_{\sigma}}$,
  4. (4)
    and a Simon-Smith family
    $H_{\sigma,p}\mathrel{\mathop{\ordinarycolon}}[0,1]\times O_{\sigma}\to\mathcal{S}(M)\,.$
We also denote
$r_{\sigma,p}^{\pm}\mathrel{\mathop{\ordinarycolon}}=s_{\sigma,p}\pm\zeta_{\sigma,p}\,,\quad B_{\sigma,p}^{\pm}\mathrel{\mathop{\ordinarycolon}}=B_{r_{\sigma,p}^{\pm}}(p)\,,\quad\hat{A}_{\sigma,p}\mathrel{\mathop{\ordinarycolon}}=\overline{B_{\sigma,p}^{+}\setminus B_{\sigma,p}^{-}}\subset A_{\sigma,p}\,.$
After the refinement in Step 2, suppose that $X$ is a cubical subcomplex of the unit cube $I(m,k^{\prime})$. For each cell $\sigma$ of $X$ and for each $x\in X$, we define
$\mathbf{d}_{\sigma}(x)\mathrel{\mathop{\ordinarycolon}}=\min\left\{\frac{2\|x-c_{\sigma}\|_{\ell_{\infty}}}{3^{-k^{\prime}}},1\right\}\,,$
to be the normalized $l^{\infty}$ distant function to $\sigma$, where $c_{\sigma}$ is the center of $\sigma$.
For each $x\in X$, we define $H(x,t)$, $0\leq t\leq 1/2$, using pinching surgeries as described in Lemma 5.3: For every cell $\sigma$ of $X$ and every $p\in\widetilde{P}(x_{\sigma})$,
$H(t,x)\cap A_{\sigma,p}=H_{\sigma,p}\left(\min\{8t(1-\mathbf{d}_{\sigma}(x)),1\},x\right)\cap A_{\sigma,p}$
and $H(t,x)=\Phi(x)$ outside $\bigcup_{\sigma\in X,p\in\widetilde{P}(x_{\sigma})}A_{\sigma,p}$.
The map has the following properties for every $x\in X$:
  • (i)
    The set
    $Z_{x}\mathrel{\mathop{\ordinarycolon}}=\{\sigma\mathrel{\mathop{\ordinarycolon}}\mathbf{d}_{\sigma}(x)<1\}$
    satisfies that every cell $\sigma$ in $Z_{x}$ is a face of the cell $\gamma$, which is the smallest cell containing $x$. By Step 2, the corresponding annuli $\{A_{\sigma,p}\}_{\sigma\in Z_{x},p\in\widetilde{P}(x_{\sigma})}$ are pairwise disjoint, and the number is no greater than $3^{m}\cdot q$. In particular, the map $H(x,t)$ is well-defined.
  • (ii)
    By Lemma 5.3 (4) and (5), $t\mapsto\mathfrak{g}(H(x,t))$ is non-increasing and $t\mapsto H(x,t)$ is a pinch-off process.
  • (iii)
    By (5.6) and Lemma 5.3 (6), for every $t\in[0,1/2]$,
    $\mathcal{H}^{2}(H(t,x))$ $\leq\mathcal{H}^{2}(\Phi(x))+(3^{m}\cdot q)\cdot C_{\mathbf{g}}100^{2}\varepsilon^{2}_{1}$
    $\leq\mathcal{H}^{2}(\Phi(x))+\varepsilon/2\,,$
    $\mathcal{H}^{2}(H(t,x))\geq\mathcal{H}^{2}(\Phi(x))-\varepsilon/2\,,$
    and
    $\mathcal{F}([H(t,x)],[\Phi(x)])\leq\varepsilon/2\,;$
    in fact, for every $x\in X$, $\sigma\in Z_{x}$ and $p\in\widetilde{P}(x)$,
    (5.7) $\mathcal{H}^{2}(H(t,x)\cap B^{+}_{\sigma,p}(p_{\sigma,p}))\leq 100(r^{+}_{\sigma,p})^{2}+\varepsilon/(100q\cdot 3^{m})<\varepsilon/(10q\cdot 3^{m})\,.$
  • (iv)
    If $\mathbf{d}_{\sigma}(x)\leq 3/4$, then for each $p\in\widetilde{P}(x_{\sigma})$, in $A_{\sigma,p}$, $H(1/2,x)=H_{\sigma,p}(1,x)$, and thus by Lemma 5.3 (4), we know that
    $H(1/2,x)\cap\hat{A}_{\sigma,p}=\emptyset\,.$
Finally, for each $t\in[1/2,1]$ and each $x\in X$, we define $H(x,t)$ using the shrinking process as described in Lemma 5.5:
For each cell $\sigma$ of $X$ and $p\in\widetilde{P}(x_{\sigma})$, we let $\mathcal{R}_{(\sigma,p),t}\mathrel{\mathop{\ordinarycolon}}S^{3}\to S^{3}$ be the one-parameter family of shrinking deformation from Lemma 5.5 with $r^{-}=r^{-}_{\sigma,p}$ and $r^{+}=r^{+}_{\sigma,p}$.
Then for each $x\in X$, we label $\{(\sigma,p)\mid\sigma\in Z_{x},\ p\in P^{\prime}(x_{\sigma})\}$ as
$\{(\sigma_{i},p_{i})\}_{1\leq i\leq N^{\prime}}$
such that for every $1\leq i<j\leq N^{\prime}$, either $B_{\sigma_{i},p_{i}}^{+}\cap B_{\sigma_{j},p_{j}}^{+}=\emptyset$, or $B_{\sigma_{i},p_{i}}^{+}\subset B_{\sigma_{j},p_{j}}^{-}$, where $N^{\prime}\leq 3^{m}\cdot q$. The existence of such labeling follows from Lemma 5.6 (1). Then, we denote for simplicity,
^$\hat{\mathcal{R}}^{(i)}_{t,x}\mathrel{\mathop{\ordinarycolon}}=\mathcal{R}_{(\sigma_{i},p_{i}),(2t-1)(1-\mathbf{d}_{\sigma_{i}}(x))}\,.$
Note that
  • By Lemma 5.5 (1), if $\mathbf{d}_{\sigma_{i}}(x)\geq 3/4$ or $t\in[1/2,5/8]$, then ^$\hat{\mathcal{R}}^{i}_{t,x}=\mathrm{id}$; when $\mathbf{d}_{\sigma_{i}}(x)<3/4$, by the construction above, $H(1/2,x)\cap\hat{A}_{\sigma_{i},p_{i}}=\emptyset$.
  • if $\mathbf{d}_{\sigma_{i}}(x)\leq 1/4$, then ^$\hat{\mathcal{R}}^{(i)}_{1,x}(B_{\sigma_{i},p_{i}}^{-})\subset\{p_{i}\}$.
Now for $t\in[1/2,1]$ and $x\in X$, we define
$H(t,x)\mathrel{\mathop{\ordinarycolon}}=\hat{\mathcal{R}}^{(N^{\prime})}_{t,x}\circ\cdots\circ\hat{\mathcal{R}}^{(2)}_{t,x}\circ\hat{\mathcal{R}}^{(1)}_{t,x}\left(H(x,1/2)\right)\,.$
Intuitively, $\{H(t,x)\}_{t\in[1/2,1]}$ is obtained by shrinking some connected components of $H(1/2,x)$ to points. Therefore, for every $x\in X$, $t\mapsto\mathfrak{g}(H(t,x))$ is non-increasing, and $\{H(t,x)\}_{t\in[1/2,1]}$ is a pinch-off process. This confirms statement (3) of both propositions.
By Lemma 5.5 (3), (4) and Lemma 5.6 (2), we see that $\mathcal{H}^{2}(H(t,x))$ is non-increasing for $t\in[1/2,1]$. In particular, for every $t\in[1/2,1]$,
$\mathcal{H}^{2}(H(t,x))\leq\mathcal{H}^{2}(\Phi(x))+\varepsilon\,.$
Also, by the estimate 5.7, one also have
$\mathcal{H}^{2}(H(t,x))\geq\mathcal{H}^{2}(H(1/2,x))-\varepsilon/(2\cdot 3^{q}\cdot m)\cdot N^{\prime}\geq\mathcal{H}^{2}(\Phi(x))-\varepsilon\,,$
and
$\mathcal{F}([H(t,x)],[\Phi(x)])$ $\leq\mathcal{F}([H(t,x)],[H(1/2,x)])+\mathcal{F}([H(1/2,x)],[\Phi(x)])$
$\leq\varepsilon/2+N^{\prime}(\varepsilon_{1})^{2}$
$\leq\varepsilon.$
due to the essence of shrinking $N^{\prime}$ balls. This confirms statement (2) of both propositions.
Furthermore, by the construction above, $\mathfrak{g}(H(1,x))=g_{0}$. This verifies statement (1) of both propositions and completes the proof.
5.4. Proof of Theorem 2.21 and Proposition 2.22
Equipped with the ingredients above, the proof of Theorem 2.21 proceeds similarly to [CL24, §8.5]: The details are included in Appendix D.
As for the proof of Proposition 2.22, we suppose that $\operatorname{dmn}(X)$ is a cubical subcomplex of $I(m,k)$. Let $q=N(P_{\Phi})+g$, and $\delta=\delta(\varepsilon/2,q,m)$ as in Proposition 5.2.
Since $\mathbf{L}(\Lambda(\Phi))=0$, there exists $\Phi^{\prime}\in\Lambda(\Phi)$ with
$\sup_{x\in X}\mathcal{H}^{2}(\Phi^{\prime}(x))<\delta\,.$
Then, the proposition follows from applying Proposition 5.2 on $\Phi^{\prime}$.
5.5. Proof of Proposition 2.23
Following the proof of Proposition 5.1 and Proposition 5.2, we can similarly prove Proposition 2.23 in three steps. This argument is much simpler, as it does not involve area control, and the family is “continuous in the Simon-Smith sense” rather than merely with respect to the $\mathbf{F}$-metric.
Step 1. Set up
Let $q=N(P_{\Phi})$ and $m$ be the ambient dimension of $X$, and $N\mathrel{\mathop{\ordinarycolon}}=N(m,q)$ as in Lemma 5.7. Fix
$\varepsilon_{1}\mathrel{\mathop{\ordinarycolon}}=\frac{\min\{\mathrm{injrad}(M,\mathbf{g}),1\}}{3^{m}}\,,\quad\varepsilon_{2}\mathrel{\mathop{\ordinarycolon}}=\frac{\varepsilon_{1}}{3^{q}5^{2qN}5^{2N}}\,.$
For every $x\in K$, let $P_{\Phi}(x)$ denote a punctate set associated with $\Phi(x)$. By Sard’s theorem and Lemma 4.11 with $n=5^{2N}$, there exist $q_{x}\leq q$, $\widetilde{P}(x)\mathrel{\mathop{\ordinarycolon}}=\{p_{x,i}\}^{q_{x}}_{i=1}\subset P_{\Phi}(x)$ and $r_{x}\in(\varepsilon_{2},3^{q}5^{2qN}\varepsilon_{2})$ such that:
  1. (1)
    $\{\overline{B(p_{x,i},5^{2N}r_{x})}\}$ are disjoint from each other;
  2. (2)
    For every $k\in\{1,2,\cdots,5^{2N}\}$, $\Phi(x)\setminus\bigcup^{q_{x}}_{i=1}B(p_{x,i},kr_{x})$ is an embedded smooth surface with boundary and genus at most $g_{0}$.
For every $l\in\{1,2,\cdots,N\}$, we write
$r_{x,l}\mathrel{\mathop{\ordinarycolon}}=5^{2(l-1)}r_{x}\,.$
For every $x\in K$, $p\in\widetilde{P}(x)$ and $l\in\{1,2,\cdots,N\}$, applying Lemma 5.3 with $\eta=1$ to $\Phi(x)$ in $A(p;r_{x,l},2r_{x,l})$ with suitable $\Lambda$, we obtain the following data:
  1. (1)
    $s_{x,l,p}\in(r_{x,l},2r_{x,l})$,
  2. (2)
    $\zeta_{x,l,p}\in(0,\min\{100r_{x,l},s_{x,l}-r_{x,l},2r_{x,l}-s_{x,l}\}/5)$,
  3. (3)
    a neighborhood $O_{x,l,p}\subset X$ of $x$,
  4. (4)
    and a Simon-Smith family
    $H_{x,l,p}\mathrel{\mathop{\ordinarycolon}}[0,1]\times O_{x,l,p}\to\mathcal{S}(M)\,.$
In addition, for each $x\in K$, we fix a neighborhood of $x$
$O_{x}\subset\bigcap_{l\in\{1,2,\dots,N\},\ p\in\widetilde{P}(x)}O_{x,l,p}\,,$
such that $y\mapsto\Phi(y)\setminus\bigcup_{p\in\widetilde{P}(x)}B(p,r_{x})$ is continuous in the smooth topology for $y\in O_{x}$; in particular, each $\Phi(y)\setminus\bigcup_{p\in\widetilde{P}(x)}B(p,r_{x})$ has genus $\leq g_{0}$.
Step 2. Refinement
Since $K$ is compact and $\{O_{x}\}_{x\in K}$ is an open cover, there exists a finite subcover
$\{O_{x_{j}}\}_{j\in J}\,.$
Up to refinement of $X$, there exists a cubical subcomplex $Z$ of $X$ such that
$K\subset\operatorname{int}(Z)\subset Z\subset\bigcup_{j\in J}O_{x_{j}}\,.$
For simplicity, we denote $\Phi|_{Z}$ by $\Psi$, $O_{x_{j}}\cap Z$ by $U_{x_{j}}$ for every $j\in J$.
Then we refine $Z$ and assign to each cell $\sigma$ of $Z$ an $x_{\sigma}\in\{x_{j}\}_{j\in J}$ such that: Every cell $\tau$ of $Z$ is a subset of $U_{x_{\sigma}}$ provided that $\sigma$ and $\tau$ are faces of a common cell $\gamma$ of $X$.
By Lemma 5.7 with
$\mathcal{F}_{\sigma}=\left(\{A(p;r_{x_{\sigma},l},2r_{x_{\sigma},l}\}^{N}_{l=1}\right)_{p\in\widetilde{P}(x_{s}igma)}\,,$
we obtain, associated with each cell $\sigma$ in $Z$, a collection of annuli denoted by
$\left(A_{\sigma,p}\mathrel{\mathop{\ordinarycolon}}=A(p;r_{\sigma,p},2r_{\sigma,p})\right)_{p\in\widetilde{P}(x_{\sigma})}$
such that whenever $\sigma$ and $\tau$ are faces of a common cell of $Z$ and for any $p\in\widetilde{P}(x_{\sigma})$ and $p^{\prime}\in\widetilde{P}(x_{\tau})$, we have
$A_{\sigma,p}\cap A_{\tau,p^{\prime}}=\emptyset,$
unless $(\sigma,p)=(\tau,p^{\prime})$.
Step 3. Construction of $H$
Following the same construction as in Step 3 of the proof of Proposition 5.1 and Proposition 5.2, without keeping track of the area estimates, one obtain
$H\mathrel{\mathop{\ordinarycolon}}[0,1]\times Z\to\mathcal{S}(M)$
such that $\Psi^{\prime}\mathrel{\mathop{\ordinarycolon}}=H(1)$ is a Simon-Smith family of genus $\leq g_{0}$, and $H$ is a deformation via pinch-off processes from $\Psi$ to $\Psi^{\prime}$.
6. Obtaining an optimal family via repetitive min-max
The goal of this section is to prove Theorem 3.1. In fact, the methodology of the proof is already given in [CL24, §3.2 - 3.4]. But note here we consider arbitrary genus, because we now have the interpolation theorem (Theorem 2.21) for arbitrary genus.
First, we will repetitively run min-max using the given family $\Phi$, and obtain a number of families. Then, we assemble these families to obtain the desired “optimal family” $\Phi^{\prime}$, whose domain is a union of parts $D_{0},D_{1},\dots,D_{l}$. After that, we conclude the proof of Theorem 3.1.
6.1. Repetitive min-max: $k$-th stage
In this section, we will repetitively run min-max. But before that, let us fix some constants.
  • First, we choose some $r^{\prime}>0$ such that $r^{\prime}<r$, where $r$ is given in Theorem 3.1.
  • Let $d_{0}$ be the minimum value of $\mathbf{F}(\Sigma,\Sigma^{\prime})$ where $\Sigma,\Sigma^{\prime}$ are any orientable embedded minimal surfaces of genus $\leq g$ in $(M,\mathbf{g})$: Recall there are only finitely many orientable embedded minimal surfaces of genus $\leq g$ by assumption. By choosing a smaller value for $r^{\prime}$, let us assume $r^{\prime}<d_{0}$.
  • By further shrinking $r^{\prime}$, by Lemma 2.6, we can assume that for each of the finitely many orientable embedded minimal surface $\Sigma$ of genus $\leq g$, if $S\in\mathcal{S}(M)$ is such that $\mathbf{F}(S,\Sigma)<r^{\prime}$, then $\mathfrak{g}(S)\geq\mathfrak{g}(\Sigma)$.
Remark 6.1.
Throughout §6, a superscript k or subscript k would indicate that the object concerned is constructed from running the min-max process for the $k$-th time.
We will recursively describe the process. Let $k=1,2,\dots$. Right before the $k$-th stage of min-max, we would have on our hands a subcomplex $X^{k-1}$ of $X$ (up to refinement), and a Simon-Smith family
$\Phi^{k-1}\mathrel{\mathop{\ordinarycolon}}X^{k-1}\to\mathcal{S}_{\leq g}.$
(Before the first time we run min-max, i.e. $k=1$, we of course have $X^{0}\mathrel{\mathop{\ordinarycolon}}=X$ and $\Phi^{0}\mathrel{\mathop{\ordinarycolon}}=\Phi$.)
We apply the min-max theorem, Theorem 2.16, to $\Phi^{k-1}$ and $r^{\prime}>0$. Suppose the width $\mathbf{L}(\Lambda(\Phi^{k-1}))$ is positive, then we obtain a varifold $V^{k}$ in $\mathcal{W}^{k}\mathrel{\mathop{\ordinarycolon}}=\mathcal{W}_{\mathbf{L}(\Lambda(\Phi^{k-1})),\leq g}$ (defined in Definition 2.15) satisfying Theorem 2.16 (2). If the width is zero instead, then this $k$-th stage of min-max would be actually the last stage, which will be discuss in the subsequent section. So let us assume the width is positive at this stage.
Since $(M,\mathbf{g})$ has positive Ricci curvature, the Frankel property holds. Consequently, every varifold in $\mathcal{W}^{k}$ is associated with a connected surface. Also, we know that $\operatorname{spt}(V^{k})$ must be orientable, for otherwise it would be one-sided (as $M$ is orientable) and its double cover would be stable (by Theorem 2.16 (2) (b)), contradicting the fact that $M$ has positive Ricci curvature. As a result, by Theorem 2.16 (2), $V^{k}$ is of multiplicity one. Furthermore, we know the following:
Lemma 6.2.
Fix an ambient 333-manifold $(M,\mathbf{g})$ that satisfies the assumption in Theorem 3.1. Suppose that when we apply the min-max theorem (Theorem 2.16) to some $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}_{\leq g}$, we know the width $L=\mathbf{L}(\Lambda(\Phi))$ is positive and the set $\mathcal{W}_{L,\leq g}$ contains some element $V$ of multiplicity one. Then we have the following.
  1. (1)
    If $\operatorname{spt}V$ has genus $g$, then every varifold in $\mathcal{W}$ is associated with a multiplicity one, orientable, minimal surface of genus $g$.
  2. (2)
[…]
In particular, $\mathcal{W}$ consists of finitely many varifolds, each associated with a multiplicity one orientable minimal surface.
Proof.
Recall that we defined $\mathcal{A}$ as the set of all orientable embedded minimal surface of genus $\leq g-1$, and non-orientable embedded minimal surface of genus $\leq g+1$. Note that:
  • By the assumptions in Theorem 3.1, given any orientable minimal surface $\Gamma$ of genus $g$, numbers in the set
    $\{\operatorname{area}(\Gamma^{\prime})\mathrel{\mathop{\ordinarycolon}}\Gamma^{\prime}\in\mathcal{A}\}\cup\{\operatorname{area}(\Gamma)\}$
    are linearly independent over $\mathbb{Z}$.
  • On the other hand, we know that every element of $\mathcal{W}$ must satisfy (2.1), and thus must either be: (A) a multipicity one orientable genus $g$ surface, or (B) an element in $\mathcal{A}$ possibly with multiplicities.
[…]
The finiteness of $\mathcal{W}$ follows from the assumptions in Theorem 3.1 that there are only finitely many orientable embedded minimal surfaces of genus $\leq g$. ∎
Using this lemma, there are two cases regarding $V^{k}\in\mathcal{W}^{k}\mathrel{\mathop{\ordinarycolon}}=\mathcal{W}_{\mathbf{L}(\Lambda(\Phi^{k-1})),\leq g}$:
  • Case 1.
    $\mathcal{W}^{k}$ consists of finitely many elements, each associated with a multiplicity one, orientable, embedded minimal surface of genus $g$: Denote them as $T^{k}_{1},\dots,T^{k}_{n_{k}}\in\mathcal{S}_{g}$;
  • Case 2.
    $\mathcal{W}^{k}=\{V^{k}\}$, and $V^{k}$ is associated with a multiplicity one, orientable, embedded minimal surface of genus $\leq g-1$, which must be non-degenerate as assumed in Theorem 3.1. Let $T^{k}\in\mathcal{S}_{\leq g-1}$ be the associated element,
Let us analyze them separately.
6.1.1. Case 1
Applying Theorem 2.17 to $\Phi^{k-1}\mathrel{\mathop{\ordinarycolon}}X^{k-1}\to\mathcal{S}(M)$ and $r^{\prime}>0$, we can obtain some $\widetilde{\Phi}^{k}\in\Lambda(\Phi^{k-1})$ and $\delta_{k}>0$ such that
$\mathcal{H}^{2}(\widetilde{\Phi}^{k}(x))\geq\mathbf{L}(\Lambda(\Phi^{k-1}))-\delta_{k}\implies[\widetilde{\Phi}^{k}(x)]\in\mathbf{B}^{\mathbf{F}}_{r^{\prime}}([\mathcal{W}^{k}]).$
Hence, we can choose some subcomplex $\hat{D}^{k}$ of $X^{k-1}$ (after refining $X^{k-1}$), which we call a “cap”, such that
$\mathcal{H}^{2}(\widetilde{\Phi}^{k}(x))\geq\mathbf{L}(\Lambda(\Phi^{k-1}))-\delta_{k}/2\implies x\in\hat{D}^{k},$
and
$x\in\hat{D}^{k}\implies[\widetilde{\Phi}^{k}(x)]\in\mathbf{B}^{\mathbf{F}}_{r^{\prime}}([\mathcal{W}^{k}]).$
In addition, for future purpose, let us choose a slightly larger subcomplex $D^{k}\supset\hat{D}^{k}$ in $X^{k-1}$ (after refinement) such that $D^{k}$ also satisfies these two properties, and that there exists a smooth cut-off function $\eta_{k}\mathrel{\mathop{\ordinarycolon}}X\to[0,1]$ such that $\eta_{k}=1$ on $\overline{X^{\backslash}D^{k}}$ while $\eta_{k}=0$ on $\hat{D}^{k}$.
Since we assumed $r^{\prime}<d_{0}$, by further refining $X^{k-1}$, we can assume that $D^{k}$ can be decomposed into a disjoint union of subcomplexes $D^{k}_{1},\dots,D^{k}_{n_{k}}$ such that: For each $j=1,\dots,n_{k}$,
$\widetilde{\Phi}^{k}(D^{k}_{j})\subset\mathbf{B}^{\mathbf{F}}_{r^{\prime}}(T^{k}_{j})\subset\mathbf{B}^{\mathbf{F}}_{r}(T^{k}_{j})\,.$
In fact, since we defined $r^{\prime}$ to be less than the constant given by Lemma 2.6 (applied to any $T^{k}_{j}$), we know that every member of $\widetilde{\Phi}^{k}|_{\hat{D}^{k}}$ is of genus $g$. Hence, we in fact have
$\widetilde{\Phi}^{k}(D^{k}_{j})\subset\mathcal{S}_{g}\cap\mathbf{B}^{\mathbf{F}}_{r}(T^{k}_{j}).$
In addition, recalling by definition $\widetilde{\Phi}^{k}$ is in the same Simon-Smith isotopy class as $\Phi^{k-1}$, we let $H^{k}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X^{k-1}\to\mathcal{S}_{\leq g}$ be such that $H^{k}(0,\cdot)=\Phi^{k-1}$, $H^{k}(1,\cdot)=\widetilde{\Phi}^{k}$, and $t\mapsto H^{k}(t,y)$ is a smooth deformation for every $y$.
6.1.2. Case 2
In this case, we let $T^{k}\in\mathcal{S}_{\leq g-1}$ be the element associated to the varifold $V^{k}$, and apply Theorem 2.21 to $\Phi^{k-1}\mathrel{\mathop{\ordinarycolon}}X^{k-1}\to\mathcal{S}(M)$ and $r$. We obtain:
  • a Simon-Smith family of genus $\leq g$
    $\widetilde{\Phi}^{k}\mathrel{\mathop{\ordinarycolon}}X^{k-1}\to\mathcal{S}_{\leq g}\,,$
  • a Simon-Smith family of genus $\leq g$
    $H^{k}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X^{k-1}\to\mathcal{S}_{\leq g},$
  • a constant $\delta_{k}>0$,
satisfying the following:
  • $H^{k}(0,\cdot)=\Phi^{k-1}$, $H^{1}(1,\cdot)=\widetilde{\Phi}^{k}$,
  • If $\mathcal{H}^{2}(\widetilde{\Phi}^{k}(x))\geq\mathbf{L}(\Lambda(\Phi^{k-1}))-\delta_{k}$, then $\widetilde{\Phi}^{k}(x)\in\mathcal{S}_{\mathfrak{g}(T^{k})}\cap\mathbf{B}^{\mathbf{F}}_{r}(T^{k})$.
  • $t\mapsto H^{k}(t,x)$ is a pinch-off process for each $x\in X^{k-1}$.
Clearly, we can choose some subcomplex $\hat{D}^{k}\subset X^{k-1}$, which we call a “cap”, such that
$\mathcal{H}^{2}(\widetilde{\Phi}^{k}(x))\geq\mathbf{L}(\Lambda(\Phi^{k-1}))-\delta_{k}/2\implies x\in\hat{D}^{k},$
and
$x\in\hat{D}^{k}\implies\widetilde{\Phi}^{k}(x)\in\mathcal{S}_{\mathfrak{g}(T^{k})}\cap\mathbf{B}^{\mathbf{F}}_{r}(T^{k}).$
As in Case 1, for future use, we choose a slightly larger subcomplex $D^{k}\supset D^{k}$ in $X^{k-1}$ (after refinement) such that $D^{k}$ also satisfies these two properties, and that there exists a smooth cut-off function $\eta_{k}\mathrel{\mathop{\ordinarycolon}}X\to[0,1]$ such that $\eta_{k}=1$ on $\overline{X^{\backslash}D^{k}}$ while $\eta_{k}=0$ on $\hat{D}^{k}$.
6.1.3. Family for running the $(k+1)$-th min-max
Now, we have described the objects we obtain in the $k$-th stage of min-max, in both Case 1 and 2. To enter the next stage, we need a new family. Namely, we define
$X^{k}\mathrel{\mathop{\ordinarycolon}}=\overline{X^{k-1}\backslash\hat{D}^{k}},\quad\textrm{ and }\quad\Phi^{k}\mathrel{\mathop{\ordinarycolon}}=\widetilde{\Phi}^{k}|_{X^{k}}.$
Note, by how we defined $\hat{D}^{k}$, it is guaranteed that the new width $\mathbf{L}(\Lambda(\Phi^{k}))$ is strictly less than $\mathbf{L}(\Lambda(\Phi^{k-1}))$. Combined with the assumption that $(M,\mathbf{g})$ has only finitely many orientable embedded minimal surfaces of genus $\leq g$, our repetitive min-max process must terminate in finitely many steps. Specifically, for some $K\geq 2$, when we apply the min-max process for the $K$-th time using the family $\Phi^{K-1}$, the width $\mathbf{L}(\Lambda(\Phi^{K-1}))$ is zero.
6.2. Last stage
We now have that the width $\mathbf{L}(\Lambda(\Phi^{K-1}))$ is zero. Thus, using Theorem 2.22, we have:
  • a Simon-Smith family of genus 0, $\Phi^{K}\mathrel{\mathop{\ordinarycolon}}X^{K-1}\to\mathcal{S}_{0}$,
  • a Simon-Simon family of genus $\leq g$, $H^{K}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X^{K-1}\to\mathcal{S}_{\leq g}$,
such that:
  • $H^{K}(0,\cdot)=\Phi^{K-1}$ and $H^{K}(1,\cdot)=\Phi^{K}$.
  • For all $x\in X^{K-1}$, $t\mapsto H^{K}(t,x)$ is a pinch-off process.
6.3. An optimal family $\Phi^{\prime}$
We have already run the min-max process $K$ times. In this subsection, we will use the families defined above to reconstruct a new Simon-Smith family $\Phi^{\prime}$ of genus $\leq g$, with parameter space $X$, and a deformation via pinch-off processes, $H$, from $\Phi$ to $\Phi^{\prime}$. Readers may refer to Figure 12 for a schematic of the families constructed.
Figure 12. In the first picture, the red region is $X^{D-1}$, the blue is $D^{K-1}$, and the green is $D^{K-2}$. In the second picture, the gray surface (with suitable smoothening) denotes $\Xi^{K-1}$. In the last picture, the gray surface denotes $\Xi^{K-2}$.
  • First, we take last piece, $\Phi^{K}$, which has domain $X^{K-1}=\overline{\widetilde{X}^{K-2}\backslash\hat{D}^{K-1}}$. We “lift” $\Phi^{K}$ on “bridge region” $\overline{D^{K-1}\backslash\hat{D}^{K-1}}$, using the deformation $H^{K}$ and the cut-off $\eta_{K-1}$. More precisely, we consider the family
    $H^{K}(\eta_{K-1}(x),x),\quad x\in X^{K-1}.$
    Note, for $x$ at the “boundary” $X^{K-1}\cap\hat{D}^{K-1}$, the expression above becomes $H^{K}(0,x)=\Phi^{K-1}(x)$.
  • As a result, this family can be glued with the family $\widetilde{\Phi}^{K-1}|_{\hat{D}^{K-1}}$ along $X^{K-1}\cap\hat{D}^{K-1}$. Then, the resulting family, denoted $\Xi^{K-1}$, has domain $X^{K-2}$. Now, it is easy to see that there exists a deformation via pinch-off processes from $\widetilde{\Phi}^{K-1}$ to $\Xi^{K-1}$. Recall we have also the deformation via pinch-off processes $H^{K-1}$ from $\Phi^{K-2}$ to $\widetilde{\Phi}^{K-1}$. Thus, let us concatenate these two processes, to obtain a deformation via pinch-off processes
    $G^{K-1}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X^{K-2}\to\mathcal{S}_{\leq g}$
    such that $G^{K-1}(0,\cdot)=\Phi^{K-2}$ and $G^{K-1}(1,\cdot)=\Xi^{K-1}$.
  • Now, we “lift” $\Xi^{K-1}$ on the “bridge region” $\overline{D^{K-2}\backslash\hat{D}^{K-2}}$, using the deformation $G^{K-1}$ and the cut-off $\eta_{K-2}$. More precisely, We consider the family
    $G^{K-1}(\eta_{K-2}(x),x),\quad x\in X^{K-2}.$
    Note, for $x$ at the “boundary” $X^{K-2}\cap\hat{D}^{K-2}$, the expression above becomes $G^{K-1}(0,x)=\Phi^{K-2}(x)$.
  • As a result, this family can be glued with the family $\widetilde{\Phi}^{K-2}|_{\hat{D}^{K-2}}$ along $X^{K-2}\cap\hat{D}^{K-2}$. The resulting family, denoted $\Xi^{K-2}$ has domain $X^{K-3}$. Now, as before, we know there exists a deformation via pinch-off processes
    $G^{K-2}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X^{K-3}\to\mathcal{S}_{\leq g}$
    such that $G^{K-2}(0,\cdot)=\Phi^{K-3}$ and $G^{K-1}=\Xi^{K-2}$.
We repeat this process. At the final step, we would be gluing some family of the form
$G^{2}(\eta_{1}(x),x),\quad x\in X^{1}$
to the family $\widetilde{\Phi}^{1}|_{\hat{D}^{1}}$, and obtain a family $\Xi^{1}$ with domain $X^{1}\cup\hat{D}^{1}=X$. Now, we define the desired “optimal family” $\Phi^{\prime}$ to be $\Xi^{1}$. By construction, it is easy to see that there exists some deformation via pinch-off processes
$G^{1}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}_{\leq g}$
such that $G^{1}(0,\cdot)=\Phi$ and $G^{1}(1,\cdot)=\Phi^{\prime}$. But we should really denote this process by $H$ instead, as it is the one required by Theorem 3.1.
6.4. To conclude the proof
Now, for simplicity, let us collect all the minimal surfaces $T^{k}$’s and $T^{k}_{j}$’s together, and relabel them as $\Gamma_{1},\Gamma_{2},\dots,\Gamma_{l}$. At the same time, we relabel their respective associated caps $D^{k}$’s and $D^{k}_{j}$’s as $D_{1},D_{2},\dots,D_{l}$. Moreover, we define $D_{0}\mathrel{\mathop{\ordinarycolon}}=X^{K-1}\backslash D^{K-1}$, on which by definition $\Phi^{\prime}=\Phi^{K}$, which has genus 0.
Let us finish the proof of Theorem 3.1. Item (1) is already explained above, For item (3), we fix an $i=1,..,l$, and suppose the cap $D_{i}$ was defined in the $k$-th stage of min-max. Then, by how we defined the caps $D^{k}$ and $D^{k}_{j}$, we have $\widetilde{\Phi}^{k}|_{D_{i}}\in\mathcal{S}_{\mathfrak{g}(\Gamma_{i})}\cap\mathbf{B}^{\mathbf{F}}_{r}(\Gamma_{i})$, where $\Gamma_{i}$ is the minimal surface associated with the cap $D_{i}$. It is easy to observe that, by construction, the family $\Phi|_{D_{i}}$ can be deformed via pinch-off processes to become $\widetilde{\Phi}^{k}|_{D_{i}}$, which in turn can be deformed via pinch-off processes to become $\Phi^{\prime}|_{D_{i}}$, and both deformations appear as subfamilies in $H$. This has two immediate consequences. First, item (3) follows immediately. Second, $\Phi^{\prime}|_{D_{i}}$ must be of genus $\leq\mathfrak{g}(\Gamma_{i})$, Finally, we know $\Phi^{\prime}|_{D_{0}}$ is of genus 0 because $\Phi^{\prime}|_{D_{0}}$ is by construction equal to $\Phi^{K}$, which has genus 0. So we have item (2) too, This finishes the proof of Theorem 3.1.
Remark 6.3.
Above, we concatenated multiple pinch-off processes together. Namely, the ending surface of one process is the beginning of another. If the ending surface of the first process has a cone point, this cone point would be preserved in the second process. This explains why in Definition 2.18 (2) we allow the cone to be left intact.
7. Construction of the family $\Psi$
The goal of this section is to construct the 777-parameter family $\Psi$ and prove Theorem 3.2. The form of $\Psi$ has been defined in (3.1), and it remains to specify the cut-off function $\rho$ and study the genus and singularity of $\Psi$.
Let us introduce the coordinate system we would use on $S^{3}$. First, we view $S^{3}$ as the set of points $x\in\mathbb{R}^{4}$ of unit distance to the origin. Let $C\subset S^{3}$ be a great circle given by $C\mathrel{\mathop{\ordinarycolon}}=\{x\in S^{3}\mathrel{\mathop{\ordinarycolon}}x_{3}=x_{4}=0\}$. We parametrize $S^{3}$ by
(7.1) $\mathbf{D}\times\mathbb{S}^{1}\to S^{3},\quad(x,\alpha)\mapsto(x_{1},x_{2},\varsigma(x)\cos\alpha,\varsigma(x)\sin\alpha)\,,$
where $\varsigma(x)\mathrel{\mathop{\ordinarycolon}}=\sqrt{1-x_{1}^{2}-x_{2}^{2}}$, $\mathbf{D}$ is the closed unit disk in $\mathbb{R}^{2}$, $\mathbb{S}^{1}\mathrel{\mathop{\ordinarycolon}}=\mathbb{R}/2\pi\mathbb{Z}$. For later reference, we also let $\mathbf{D}_{\kappa}(b)$ be the closed disc of radius $\kappa$ centered at $b$. Note that (7.1) is a diffeomorphism from $\operatorname{int}(\mathbf{D})\times\mathbb{S}^{1}$ to $S^{3}\setminus C$, and restricting to a smooth $\mathbb{S}^{1}$ fibration on $\partial\mathbf{D}\times\mathbb{S}^{1}$, mapping $\{z\}\times\mathbb{S}^{1}$ to a point on $C$. Note also that here we use the symbol $\bf D$ in order to distinguish it from the $\mathbb{D}$ in $\mathrm{dmn}(\Psi)={\mathbb{RP}}^{5}\times\mathbb{D}$.
This section is organized as follows. We first study in section 7.1 zero sets of the form
$\{x\in S^{3}\mathrel{\mathop{\ordinarycolon}}a_{0}+a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3}+a_{4}x_{4}+a_{5}x_{1}x_{2}=0\}.$
This corresponds to an ${\mathbb{RP}}^{5}$ family of cycles which (if nonempty) are either punctate surfaces with genus 0, or union of two transverse round sphere. Section 7.2 contains some abstract lemmas discussing zeros of functions on $\mathbb{S}^{1}$, which helps to understand in Lemma 7.6 the genus and other geometric behavior of the zero set for a special class of function of form (3.1). Using them, we specify the choice of cut-off $\rho$ in section 7.3 and discuss basic properties of the family $\Psi$ in Lemma 7.7. Finally, Theorem 3.2 is proved in section 7.4 based on Lemma 7.7, an extension Lemma 7.11 of embedded loops and a purely topology Lemma 7.12 on Hopf links.
7.1. The canonical ${\mathbb{RP}}^{5}$ family
[…]
$\widetilde{\Phi}_{5}(a)$ :=$\mathrel{\mathop{\ordinarycolon}}=\{x\in\mathbb{R}^{4}\mathrel{\mathop{\ordinarycolon}}a_{0}+a_{1}x_{1}+a_{2}x_{2}+a_{3}x_{3}+a_{4}x_{4}+a_{5}x_{1}x_{2}=0\}\subset\mathbb{R}^{4}\,;$
$\Phi_{5}(a)$ :$\mathrel{\mathop{\ordinarycolon}}=S^{3}\cap\widetilde{\Phi}_{5}(a)\subset S^{3}\,.$
[…]
As pointed out before, $\Phi_{5}$ gives a 5-sweepout in the Almgren-Pitts setting, but it is not a Simon-Smith family. However, we still have genus bound when $\Phi_{5}(a)$ is not singular:
Lemma 7.1.
Let
[…]
Then,
  1. (i)
    for every $a\in A_{\operatorname{sing}}$, $\Phi_{5}(a)$ is the union of two round (but not necessarily equatorial) spheres that intersect transversely along the circle $\{x\in S^{3}\mathrel{\mathop{\ordinarycolon}}x_{1}=-a_{2},x_{2}=-a_{1}\}$;
  2. (ii)
    for a dense subset of $a\in{\mathbb{RP}}^{5}\setminus A_{\operatorname{sing}}$, $\widetilde{\Phi}_{5}(a)$ is transverse to $S^{3}$, and hence $\Phi_{5}(a)$ is smooth;
  3. (iii)
    for every $a\in{\mathbb{RP}}^{5}\setminus A_{\operatorname{sing}}$, $\widetilde{\Phi}_{5}(a)$ is transverse to $S^{3}$ at all but at most 999 points; and $\mathfrak{g}(\Phi_{5}(a))=0$.
Proof.
(i) follows by noticing that when $a\in A_{\operatorname{sing}}$, $\widetilde{\Phi}_{5}(a)=\{(x_{1}+a_{2})(x_{2}+a_{1})=0\}$ is the union of two orthogonal hyperplanes $\{x_{1}=-a_{2}\}$ and $\{x_{2}=-a_{1}\}$ whose intersection transversely intersect $S^{3}$ along the circle described in (i).
To prove (ii) and (iii), consider an arbitrary $a\in{\mathbb{RP}}^{5}\setminus A_{\operatorname{sing}}$. If $a_{5}=0$, then $\widetilde{\Phi}_{5}(a)$ is either empty or an affine hyperplane in $\mathbb{R}^{4}$, hence (iii) holds for such $a$; if not, then we write $a=[a_{0}\mathrel{\mathop{\ordinarycolon}}\dots\mathrel{\mathop{\ordinarycolon}}a_{4}\mathrel{\mathop{\ordinarycolon}}1]$. If $a_{0}=a_{1}a_{2}$ and $a_{3}=a_{4}=0$, then by $a\notin A_{\operatorname{sing}}$, we have $a_{1}^{2}+a_{2}^{2}\geq 1$, hence $\Phi_{5}(a)$ is the union of at most two 2-spheres (possibly degenerated to points), either disjoint or tangent to each other, which also meets the assertion in (iii). For the remaining case,
$\widetilde{\Phi}_{5}(a)=\{x\in\mathbb{R}^{4}\mathrel{\mathop{\ordinarycolon}}(x_{1}+a_{2})(x_{2}+a_{1})+(a_{3}x_{3}+a_{4}x_{4})+a_{0}-a_{1}a_{2}=0\}\,.$
where $(a_{3},a_{4},a_{0}-a_{1}a_{2})\neq(0,0,0)$. Hence there exists $\mathcal{R}\in\mathrm{SO}(4)$, $b\in\mathbb{R}^{4}$ and $\lambda>0$ such that
$\eta(\widetilde{\Phi}_{5}(a))=\begin{cases}\{x\in\mathbb{R}^{4}\mathrel{\mathop{\ordinarycolon}}x_{1}x_{2}=1\},&\ \text{ if }(a_{3},a_{4})=(0,0)\,,\\ \{x\in\mathbb{R}^{4}\mathrel{\mathop{\ordinarycolon}}x_{1}x_{2}=x_{3}\},&\ \text{ if }(a_{3},a_{4})\neq(0,0)\,;\end{cases}$
where $\eta(x)=\eta_{\mathcal{R},b,\lambda}(x)\mathrel{\mathop{\ordinarycolon}}=\lambda\cdot\mathcal{R}x+b$ is a diffeomorphism of $\mathbb{R}^{4}$. Also notice that $\eta(S^{3})=\partial B^{4}_{\lambda}(b)$, hence $\eta(\Phi_{5}(a))=(P\times\mathbb{R})\cap\partial B^{4}_{\lambda}(b)$, where $P$ is the smooth surface described in the following Lemma 7.2.
Write $b=(\breve{b},b_{4})\in\mathbb{R}^{3}\times\mathbb{R}$, $\varrho_{\breve{b}}(\breve{x})\mathrel{\mathop{\ordinarycolon}}=|\breve{x}-\breve{b}|^{2}$ is a function on $\mathbb{R}^{3}$. Notice that $P\times\mathbb{R}$ intersects non-transversely with $\partial B^{4}_{\lambda}(b)$ at $x=(\breve{x},x_{4})$ if and only if $x_{4}=b_{4}$ and $\breve{b}-\breve{x}\perp T_{\breve{x}}P$, i.e. $\varrho_{\breve{b}}(\breve{x})=\lambda^{2}$ and $\breve{x}$ is a critical point of $\varrho_{\breve{b}}|_{P}$. Also notice that if $\lambda^{2}$ is a regular value of $\varrho_{\breve{b}}|_{P}$, then the smooth surface $(P\times\mathbb{R})\cap\partial B_{\lambda}^{4}(b)$ is homeomorphic to the double of the smooth surface with boundary $P\cap\overline{B^{3}_{\lambda}(\breve{b})}$.
Therefore, (ii) follows by Sard Theorem, as the regular value of $\varrho_{\breve{b}}|_{P}$ is dense; (iii) follows from Lemma 7.2 below, where the genus control is immediate when the intersection of $\widetilde{\Phi}_{5}(a)$ and $S^{3}$ is transverse, since the double of finitely many disks is finitely many spheres; when the intersection is not transverse, the genus control follows from the approximation of regular $\Phi(a_{j})$ guaranteed by (ii) and the lower-semi-continuity of genus $\mathfrak{g}$ under $C^{\infty}_{loc}$ convergence away from finitely many point. ∎
Lemma 7.2.
Let $P$ be one of the following two smooth surfaces in $\mathbb{R}^{3}$
$\{x\in\mathbb{R}^{3}\mathrel{\mathop{\ordinarycolon}}x_{1}x_{2}=1\}\,,\qquad\{x\in\mathbb{R}^{3}\mathrel{\mathop{\ordinarycolon}}x_{1}x_{2}=x_{3}\}\,.$
Then for every $b\in\mathbb{R}^{3}$, let $\varrho_{b}(x)\mathrel{\mathop{\ordinarycolon}}=|x-b|^{2}$, we have,
  1. (a)
    $\varrho_{b}|_{P}$ has at most 999 critical points;
  2. (b)
    if $r^{2}$ is a regular value of $\varrho_{b}|_{P}$, then $\overline{B_{r}^{3}(b)}\cap P$ is diffeomorphic to a finite collection (possibly empty) of disjoint disks.
Proof.
We only focus on the case when $P=\{x\in\mathbb{R}^{3}\mathrel{\mathop{\ordinarycolon}}x_{1}x_{2}=x_{3}\}$, the other surface can be treated in a similar (actually even simpler) way.
For every $x\in P$, $x$ is a critical point of $\varrho_{b}|_{P}$ if and only if $2(x_{1}-b_{1},x_{2}-b_{2},x_{3}-b_{3})=\nabla\varrho_{b}(x)\perp T_{x}P$, which is equivalent to
(7.2) $\begin{cases}x_{1}-b_{1}=-x_{2}(x_{3}-b_{3})\,,\\ x_{2}-b_{2}=-x_{1}(x_{3}-b_{3})\,.\end{cases}$
When $(x_{3}-b_{3})^{2}\neq 1$, (7.2) has solution
$x_{1}=\frac{b_{1}-b_{2}(x_{3}-b_{3})}{1-(x_{3}-b_{3})^{2}}\,,$ $x_{2}=\frac{b_{2}-b_{1}(x_{3}-b_{3})}{1-(x_{3}-b_{3})^{2}}$
Plugging back to the defining equation $x_{1}x_{2}=x_{3}$ for $P$ gives a 555-th order polynomial equation in $x_{3}$, hence has at most 555 solutions for $(x_{1},x_{2},x_{3})$.
When $x_{3}-b_{3}=\pm 1$, by (7.2),
$x_{1}\pm x_{2}=b_{1}=\pm b_{2}\,,$
Combining with $x_{1}x_{2}=x_{3}=b_{3}\pm 1$ gives at most 444 solutions for $(x_{1},x_{2},x_{3})$. In summary, there are at most 5+4=95+4=95 + 4 = 9 solutions $x$ to $x\in P$ and $\nabla\varrho_{b}\perp T_{x}P$, which proves (a).
To prove (b), we can first take a smooth perturbation of $\varrho_{b}|_{P}$ to get a Morse function $\bar{\varrho}\in C^{\infty}(P)$ such that $\operatorname{spt}(\bar{\varrho}-\varrho_{b}|_{P})$ is compact and that $\bar{\varrho}^{-1}(\mathbb{R}_{\leq r^{2}})$ is diffeomorphic to $\varrho_{b}^{-1}(\mathbb{R}_{\leq r^{2}})\cap P=\overline{B_{r}^{3}(b)}\cap P$. Also, since the second fundamental form $A_{P}$ of $P$ is everywhere indefinite,
$\nabla^{2}(\varrho_{b}|_{P})=\nabla^{2}\varrho_{b}+(\nabla^{\perp}\varrho_{b})\cdot A_{P}=2\,\mathrm{id}_{T_{*}P}+(\nabla^{\perp}\varrho_{b})\cdot A_{P}\,,$
has at least 111 strictly positive eigenvalue everywhere on $P$. Thus we can further ask the smooth perturbation $\bar{\varrho}$ to also have no local maximum, i.e. any critical point of $\bar{\varrho}$ has Morse index 0 or 111. Therefore, the first betti number function
$\lambda\mapsto\dim H_{1}(\bar{\varrho}^{-1}(\mathbb{R}_{\leq\lambda}))$
is monotone non-decreasing in regular value $\lambda$ of $\bar{\varrho}$. While since when $\lambda\gg 1$, $\bar{\varrho}^{-1}(\mathbb{R}_{\leq\lambda})=\varrho_{b}^{-1}(\mathbb{R}_{\leq\lambda})\cap P=\overline{B^{3}_{\sqrt{\lambda}}(b)}\cap P$ is a single disk, we conclude that
$\dim H_{1}(\overline{B_{r}^{3}(b)}\cap P)=\dim H_{1}(\bar{\varrho}^{-1}(\mathbb{R}_{\leq r^{2}}))=0\,.$
Also, since $P$ is a graph over the $x_{1}x_{2}$-plane and hence diffeomorphic to $\mathbb{R}^{2}$, as a compact subdomain with first betti number 0, $\overline{B_{r}^{3}(b)}\cap P$ is diffeomorphic to a finite disjoint union of disks. ∎
7.2. Some abstract lemmas
We list some useful abstract Lemmas in this subsection for later application. The first time readers are encouraged to skip the proof of this subsection.
For $f\in C^{\infty}(\mathbb{S}^{1};\mathbb{R})$, $\alpha\in f^{-1}(0)$. Define
$\operatorname{ord}_{f}(\alpha)\mathrel{\mathop{\ordinarycolon}}=\sup\{l\in\mathbb{Z}_{\geq 1}\mathrel{\mathop{\ordinarycolon}}f(\alpha)=f^{(1)}(\alpha)=\dots=f^{(l-1)}(\alpha)=0\}\,.$
Also for $E\subset\mathbb{S}^{1}$, let
$\operatorname{ord}(f;E)\mathrel{\mathop{\ordinarycolon}}=\sum_{\alpha\in f^{-1}(0)\cap E}\operatorname{ord}_{f}(\alpha)\,,$ $\mathcal{Z}(f;E)\mathrel{\mathop{\ordinarycolon}}=\sum_{\alpha\in f^{-1}(0)\cap E}\operatorname{ord}_{f}(\alpha)\cdot\alpha\,;$
and we may omit $E$ if $E=\mathbb{S}^{1}$. Note that since $\alpha\in\mathbb{S}^{1}=\mathbb{R}/2\pi\mathbb{Z}$, $\mathcal{Z}(f;E)$ also takes value in $\mathbb{R}/2\pi\mathbb{Z}$ provided $0<\operatorname{ord}(f;E)<+\infty$; when $\operatorname{ord}(f;E)=0$, we set $\mathcal{Z}(f;E)=0\text{ mod}\,2\pi$; when $\operatorname{ord}(f;E)=+\infty$, we set $\mathcal{Z}(f;E)\mathrel{\mathop{\ordinarycolon}}=\infty$.
Lemma 7.3.
Suppose $f\in C^{\infty}(\mathbb{S}^{1})$, $\alpha\in f^{-1}(0)$ with $\operatorname{ord}_{f}(\alpha)\leq N$; $f_{j}\in C^{\infty}(\mathbb{S}^{1})$ $C^{N}$-converges to $f$ as $j\to\infty$, $U_{j}$ is a smaller and smaller closed interval containing $\alpha$ and shrinking to $\alpha$ as $j\to\infty$. Then for $j\gg 1$,
$\operatorname{ord}(f_{j};U_{j})\leq\operatorname{ord}_{f}(\alpha)\,.$
Proof.
Notice that by intermediate value theorem, any closed interval $U\subset\mathbb{S}^{1}$ ended by two zeros of $f$ contains a zero of $f^{\prime}$ in its interior. Hence inductively we have,
$\operatorname{ord}(f;U)\leq\operatorname{ord}(f^{(1)};U)+1\leq\dots\leq\operatorname{ord}(f^{(N)};U)+N\,.$
Suppose now for contradiction $\operatorname{ord}(f_{j};U_{j})>\operatorname{ord}_{f}(\alpha)=\mathrel{\mathop{\ordinarycolon}}N^{\prime}$ for infinitely $j$, then after passing to a subsequence, for every $j$,
$\operatorname{ord}(f_{j}^{(N^{\prime})},U_{j})\geq\operatorname{ord}(f_{j},U_{j})-N^{\prime}>0\,,$
i.e. $\exists\,\alpha_{j}\in U_{j}$ such that $f_{j}^{(N^{\prime})}(\alpha_{j})=0$. Take $j\to\infty$, we get $\alpha_{j}\to\alpha$ and $f^{(N^{\prime})}(\alpha)=0$. This forces $\operatorname{ord}_{f}(\alpha)\geq N^{\prime}+1$ which is a contradiction. ∎
For every integer $k>0$, a trigonometric polynomial of degree $k$ is a function on $\mathbb{S}^{1}$
(7.3) $T(\alpha)\mathrel{\mathop{\ordinarycolon}}=b_{0}+b_{1}\cos(\alpha+\theta_{1})+\dots b_{k}\cos(k\alpha+\theta_{k})\,,$
where $b_{j}\in\mathbb{R}$, $b_{k}\neq 0$, $\theta_{j}\in\mathbb{R}/2\pi\mathbb{Z}$. (We use the convention that 0-function has degree 1-1- 1.) We let
$\|T\|\mathrel{\mathop{\ordinarycolon}}=\left(b_{0}^{2}+b_{1}^{2}+\dots+b_{k}^{2}\right)^{1/2}\,.$
And let $\mathcal{T}_{k}$ be the space of all trigonometric polynomial of degree $\leq k$.
Lemma 7.4.
For every nonzero $T\in\mathcal{T}_{k}$, $\operatorname{ord}(T)\leq 2k$; and if the equality holds with $T$ of the form (7.3), then $\mathcal{Z}(T)=-2\theta_{k}$.
Proof.
Without loss of generality $b_{k}\neq 0$. Notice that $\alpha\in T^{-1}(0)$ if and only if $z\mathrel{\mathop{\ordinarycolon}}=e^{i\alpha}$ is a root of the polynomial with multiplicity $\operatorname{ord}_{T}(\alpha)$:
$q_{T}(z)\mathrel{\mathop{\ordinarycolon}}=z^{k}\left(b_{0}+\frac{b_{1}}{2}(e^{i\theta_{1}}z+e^{-i\theta_{1}}z^{-1})+\dots+\frac{b_{k}}{2}(e^{i\theta_{k}}z^{k}+e^{-i\theta_{k}}z^{-k})\right)$
[…]
Corollary 7.5.
For every integer $k>0$, there exists $\delta_{1}(k)>0$ such that if $f\in C^{\infty}(\mathbb{S}^{1})$ and $T\in\mathcal{T}_{k}$ has the form (7.3) such that
$\|f-T\|_{C^{2k}}\leq\delta_{1}\|T\|\,,$
then $\operatorname{ord}(f)\leq 2k$, and if in fact $\operatorname{ord}(f)=2k$, then
$b_{k}\neq 0\,,$ $\operatorname{dist}_{\mathbb{S}^{1}}(\mathcal{Z}(f),-2\theta_{k})\leq 1000^{-1}\,.$
Proof.
Argue by contradiction and apply Lemma 7.3 and 7.4. ∎
Lemma 7.6.
For every $b=(b_{1},b_{2})\in\mathbf{D}$ and $\kappa>0$ such that $|b|+2\kappa<1$, there exists $\delta_{2}(\kappa)\in(0,\kappa/4)$ with the following significant.
Suppose $F\in C^{\infty}(\mathbf{D}\times\mathbb{S}^{1};\mathbb{R})$, $\mathbf{x}\in C^{\infty}(\mathbb{S}^{1};\mathbf{D}_{\kappa/4}(b))$ satisfies
  1. (a)
[…]
    $|a_{0}-a_{1}a_{2}|+|a_{1}+b_{2}|+|a_{2}+b_{1}|+|a_{3}|+|a_{4}|<\delta_{2}\,;$
  2. (b)
    $f(\alpha)\mathrel{\mathop{\ordinarycolon}}=F(\mathbf{x}(\alpha),\alpha)$ has $\operatorname{ord}(f)<+\infty$;
  3. (c)
    in $\mathbf{D}_{\kappa}(b)\times\mathbb{S}^{1}$, $\nabla_{x}F(x,\alpha)=0$ if and only if $x=\mathbf{x}(\alpha)$; and
    $\left|\nabla^{2}_{x}F-\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\right|<\delta_{2}\,.$
Then
  1. (1)
    Under the parametrization (7.1), $\Sigma$ is a smooth surface in $S^{3}$ with isolated singularities in $\{(\mathbf{x}(\alpha),\alpha)\mathrel{\mathop{\ordinarycolon}}\alpha\in f^{-1}(0),\operatorname{ord}_{f}(\alpha)\geq 2\}$.
  2. (2)
    For every $\alpha\in\mathbb{S}^{1}$,
    $F(\cdot,\alpha)\begin{cases}>0\ \text{ on }\{\mathbf{x}(\alpha)+(t,t)\in\mathbf{D}\mathrel{\mathop{\ordinarycolon}}0\neq t\in\mathbb{R}\}\,,&\ \text{ if }f(\alpha)\geq 0\,;\\ <0\ \text{ on }\{\mathbf{x}(\alpha)+(t,-t)\in\mathbf{D}\mathrel{\mathop{\ordinarycolon}}0\neq t\in\mathbb{R}\}\,,&\ \text{ if }f(\alpha)\leq 0\,;\end{cases}$
  3. (3)
    the genus of $\Sigma$ satisfies
    $\mathfrak{g}(\Sigma)\leq\max\left\{0,\ \frac{1}{2}\#\{\alpha\in f^{-1}(0)\mathrel{\mathop{\ordinarycolon}}\operatorname{ord}_{f}(\alpha)\text{ is odd}\}-1\right\}\,.$
    and equality holds if every zero of $f$ has order 111.
Proof.
To prove (1), by (a) with $\delta_{2}\ll 1$, $\Sigma=\Phi_{5}(a)$ is smooth outside $\mathbf{D}_{\kappa}(b)\times\mathbb{S}^{1}$. Hence it suffices to show that for every $(x,\alpha)\in\Sigma\cap\mathbf{D}_{\kappa}(b)\times\mathbb{S}^{1}$,
(7.4) $\nabla F(x,\alpha)=0$$\Leftrightarrow$ $x=\mathbf{x}(\alpha)\ \text{ and }\ \operatorname{ord}_{f}(\alpha)\geq 2\,.$
To see this, note that by (c), $\nabla_{x}F(x,\alpha)=0$ if and only if $x=\mathbf{x}(\alpha)$; while by definition of $f$,
$f^{\prime}(\alpha)=\nabla_{x}F(\mathbf{x}(\alpha),\alpha)\cdot\mathbf{x}^{\prime}(\alpha)+\partial_{\alpha}F(\mathbf{x}(\alpha),\alpha)=\partial_{\alpha}F(\mathbf{x}(\alpha),\alpha)\,.$
Combining these proves (7.4).
To prove (2), for $0\neq v=(v_{1},v_{2})\in\mathbb{R}^{2}$ such that $\mathbf{x}(\alpha)+v\in\mathbf{D}_{\kappa}(b)$, let $h(t)\mathrel{\mathop{\ordinarycolon}}=F(\mathbf{x}(\alpha)+tv,\alpha)$. Note that by (b) and (c),
$h(0)=f(\alpha)\,,$ $h^{\prime}(0)=0\,,$ $|h^{\prime\prime}(t)-2v_{1}v_{2}|\leq\delta_{2}|v|^{2}\,.$
Hence ±$\pm h^{\prime\prime}(t)>0$ if $v=(t,\pm t)$. Combined this with the fact
[…]
proves (2) in $\mathbf{D}_{\kappa}(b)$. While by (a) with $\delta_{2}\ll 1$ and a compactness argument, in $(\mathbf{D}\setminus\mathbf{D}_{\kappa}(b))\times\mathbb{S}^{1}$, $\Sigma=\Phi_{5}(a)$ doesn’t intersect $\{(\mathbf{x}(\alpha)+(t,\pm t),\alpha)\in\mathbf{D}\times\mathbb{S}^{1}\mathrel{\mathop{\ordinarycolon}}0\neq t\in\mathbb{R},\alpha\in\mathbb{S}^{1}\}$, so the sign of $F(\cdot,\alpha)$ on the corresponding line segments are not changing outside $\mathbf{D}_{\kappa}(b)$.
To prove (3), we first further assume that every zero of $f$ has order 111 (hence $\Sigma$ is smooth everywhere by (1)). Consider the Killing vector field $\partial_{\alpha}$ on $\mathbf{D}\times\mathbb{S}^{1}$ (which is also a killing vector field on the round $S^{3}$ under parametrization (7.1)). For $(x_{\circ},\alpha_{\circ})\in\mathbf{D}_{\kappa}(b)\times\mathbb{S}^{1}\cap\Sigma$, the orthogonal projection $\partial_{\alpha}^{\Sigma}$ of $\partial_{\alpha}$ onto $T\Sigma$ is zero at $(x_{\circ},\alpha_{\circ})$ if and only if $\nabla F(x_{\circ},\alpha_{\circ})\mathbin{\!/\mkern-5.0mu/\!}\partial_{\alpha}$, or equivalently, $\nabla_{x}F(x_{\circ},\alpha_{\circ})=0$, which is equivalent to $x_{\circ}=\mathbf{x}(\alpha_{\circ})$ (and $\alpha_{\circ}\in f^{-1}(0)$) by (c). Moreover, near such $(x_{\circ},\alpha_{\circ})$, view $\alpha$ as a single valued function on $\mathbf{D}\times\mathbb{S}^{1}$, $\partial_{\alpha}^{\Sigma}=\nabla(\alpha|_{\Sigma})$, and if we set $\partial_{\alpha}|_{(x_{\circ},\alpha_{\circ})}=c_{\circ}\nabla F(x_{\circ},\alpha_{\circ})$ for some nonzero constant $c_{\circ}\in\mathbb{R}$, then at $(x_{\circ},\alpha_{\circ})$,
$\nabla^{2}(\alpha|_{\Sigma})=\nabla^{2}\alpha|_{T\Sigma\otimes T\Sigma}+\vec{A}_{\Sigma}\cdot\partial_{\alpha}$ $=c_{\circ}\vec{A}_{\Sigma}\cdot\nabla F(x_{\circ},\alpha_{\circ})$
$=c_{\circ}(\nabla^{2}(F|_{\Sigma})-\nabla^{2}F(x_{\circ},\alpha_{\circ})|_{T\Sigma\otimes T\Sigma})$
$=-c_{\circ}\nabla_{x}^{2}F(x_{\circ},\alpha_{\circ})$
is indefinite and nondegenerate by (c) if $\delta_{2}$ is chosen small. Hence, the Poincaré-Hopf index of the vector field $\partial_{\alpha}^{\Sigma}$ at its zero $(x_{\circ},\alpha_{\circ})$ is 1-1- 1.
[…]
$\chi(\Sigma)=\chi(\Sigma^{\prime})=4+(-1)\cdot\#f^{-1}(0)\,.$
(3) then follows from the fact that $\Sigma^{\prime}$ has at most two connected component, and is connected when $f^{-1}(0)\neq\emptyset$.
Finally, when $f$ has zeros of higher order, let $Z_{ev}$ and $Z_{od}$ be the set of zeros of $f$ with even order and odd order respectively. We can construct another smooth function $h\in C^{\infty}(\mathbb{S}^{1})$ such that
  • $h^{-1}(0)=Z_{od}$ and each zero of $h$ has order 111;
  • $fh\geq 0$ everywhere on $\mathbb{S}^{1}$.
Then, let $\xi\in C^{\infty}_{c}(\mathbf{D}_{\kappa}(b),[0,1])$ be a cut-off such that $\xi|_{\mathbf{D}_{\kappa/2}(b)}=1$, we can consider
$F_{\epsilon}(x,\alpha)\mathrel{\mathop{\ordinarycolon}}=F(x,\alpha)+\epsilon\xi(x)h(\alpha)\,.$
It is easy to check that when $0<\epsilon\ll 1$, $F_{\epsilon}$ also satisfies (a)-(c) with $f$ replaced by $f_{\epsilon}\mathrel{\mathop{\ordinarycolon}}=f+\epsilon h$. By the construction above, $f_{\epsilon}$ only has zeros of order 111, and $f_{\epsilon}^{-1}(0)=Z_{od}$. Therefore, as $F_{\epsilon}^{-1}(0)$ locally smoothly converges to $\Sigma$ in the regular part, we have
$\mathfrak{g}(\Sigma)\leq\liminf_{\epsilon\searrow 0}\mathfrak{g}(F_{\epsilon}^{-1}(0))=\max\left\{0,\ \frac{1}{2}\#Z_{od}-1\right\}\,.$
7.3. Construction of $\Psi$
Let $\mathbb{D}$ denote the closed 2-dimensional unit disc for part of the domain of the 777-parameter family $\Psi$. We will often use polar coordinate $z=re^{i\theta}$ for $\mathbb{D}$. Recall we parametrize $S^{3}$ by (7.1). Under this parametrization, the family $\Psi\mathrel{\mathop{\ordinarycolon}}{\mathbb{RP}}^{5}\times\mathbb{D}\to\mathcal{S}(S^{3})$ takes the form
$\Psi(a,z)\mathrel{\mathop{\ordinarycolon}}=\{(x,\alpha)\in\mathbf{D}\times\mathbb{S}^{1}\mathrel{\mathop{\ordinarycolon}}F_{a,z}(x,\alpha)=0\}\,,$
[…]
$F_{a,z}(x,\alpha)$ :$\mathrel{\mathop{\ordinarycolon}}=a_{0}+a_{1}x_{1}+a_{2}x_{2}+\varsigma(x)(a_{3}\cos\alpha+a_{4}\sin\alpha)$
$\;\;\;+a_{5}\Big{[}x_{1}x_{2}+\rho(a,x)\big{(}r\cos(\theta+2\alpha)+(1-r)\cos(3\alpha)\big{)}\Big{]}\,.$
The cut-off function $\rho\mathrel{\mathop{\ordinarycolon}}{\mathbb{RP}}^{5}\times S^{3}\to[0,1]$ is chosen as follows: let $\zeta\in C^{\infty}(\mathbb{R})$ be a fixed non-increasing cut-off such that $\zeta|_{\mathbb{R}_{\leq 1/2}}=1$, $\zeta|_{\mathbb{R}_{\geq 1}}=0$,
  1. (I)
[…]
    $\rho(a,x)\mathrel{\mathop{\ordinarycolon}}=$ $\ \zeta\left(64\cdot\frac{(x_{1}+a_{2})^{2}+(x_{2}+a_{1})^{2}}{\left(1-a_{1}^{2}-a_{2}^{2}\right)^{2}}\right)\cdot$
    $\ \zeta\left(\frac{a_{3}^{2}+a_{4}^{2}+(a_{0}-a_{1}a_{2})^{2}}{\delta(a_{1}^{2}+a_{2}^{2})}\right)\delta(a_{1}^{2}+a_{2}^{2})$
    where $\delta\in C^{\infty}(\mathbb{R};[0,1])$ is 0 outside (1,1)(-1,1)( - 1 , 1 ), positive and very small in (1,1)(-1,1)( - 1 , 1 ), to be specified in Lemma 7.7 below.
  2. (II)
    When $a$ does not take the form above, we set $\rho(a,x)=0$.
Note that
  • $\rho$ is a well-defined continuous function on ${\mathbb{RP}}^{5}\times S^{3}$ and is $C^{\infty}$ in $x$ variable;
  • for $a\in{\mathbb{RP}}^{5}$ that is not in a small neighborhood of $A_{\operatorname{sing}}$, $\rho(a,\cdot)\equiv 0$, hence (under the parametrization 7.1) by Lemma 7.1, $\Psi(a,z)=\Phi_{5}(a)$ is a surface with at most 999 singular points and genus 0;
  • when $\rho(a,x)\neq 0$,
    $(x_{1}+a_{2})^{2}+(x_{2}+a_{1})^{2}\leq\frac{1}{64}(1-a_{1}^{2}-a_{2}^{2})^{2}<\frac{1}{16}\left(1-\sqrt{a_{1}^{2}+a_{2}^{2}}\right)^{2}\,,$
    hence $|x|\leq(1+3\sqrt{a_{1}^{2}+a_{2}^{2}}\ )/4<1$. As a consequence, $\Psi(a,z)$ agrees with $\Phi_{5}(a)$ in a small neighborhood of $\partial\mathbf{D}\times\mathbb{S}^{1}$ (whose image under parametrzation (7.1) is a neighborhood of $C$ in $S^{3}$). We shall establish the smoothness of $\Psi(a,z)$ away from at most 333 points in the interior of $\mathbf{D}\times\mathbb{S}^{1}$, which then implies the regularity of $\Psi(a,z)$ in the whole $S^{3}$;
  • when $a\in A_{\operatorname{sing}}$, $\rho(a,x)=\delta(a_{1}^{2}+a_{2}^{2})$ is a nonzero constant for $(x,\alpha)$ close to the singular circle of $\Phi_{5}(a)$. Heuristically, this means $\Psi(a,z)$ is a desingularization of $\Phi_{5}(a)$ using some trigonometric polynomials in $\alpha$ variable, whose number of zeros bounds the genus of $\Psi(a,z)$ as in Lemma 7.6.
Lemma 7.7.
[…]
  1. (a)
    there exists $\mathbf{x}_{a}\in C^{\infty}(\mathbb{S}^{1};\mathbf{D}_{\kappa(a)/4}(-a_{2},-a_{1}))$ such that
    $|\mathbf{x}_{a}(\alpha)-(-a_{2},-a_{1})|\leq C(|a_{3}|+|a_{4}|)\,$
    for some constant $C=C(a^{2}_{1}+a^{2}_{2})>0$; and that for $(x,\alpha)\in\mathbf{D}_{\kappa(a)}(-a_{2},-a_{1})\times\mathbb{S}^{1}$,
    $\nabla_{x}F_{a,z}(x,\alpha)=0\,,$$\Leftrightarrow$ $x=\mathbf{x}_{a}(\alpha)\,.$
  2. (b)
    Let $f_{a,z}\in C^{\infty}(\mathbb{S}^{1})$ be given by $f_{a,z}(\alpha)\mathrel{\mathop{\ordinarycolon}}=F_{a,z}(\mathbf{x}(\alpha),\alpha)$. Then
    #$\#f_{a,z}^{-1}(0)\leq\begin{cases}4\,,&\ \text{ if }|z|=1\,;\\ 6\,,&\ \text{ if }|z|<1\,.\end{cases}$
[…]
    $\operatorname{dist}_{\mathbb{S}^{1}}(\mathcal{Z}(f_{a,z}),-2\theta)\leq 1000^{-1}\,.$
  3. (c)
    For every $z\in\mathbb{D}$, $\Psi(a,z)$ is a smooth surface in $S^{3}$ away from at most 333 points, with genus estimate,
    $\mathfrak{g}(\Psi(a,z))\leq\begin{cases}1\,,&\ \text{ if }|z|=1\,;\\ 2\,,&\ \text{ if }|z|<1\,.\end{cases}$
  4. (d)
[…]
    $\Omega_{\Psi}\mathrel{\mathop{\ordinarycolon}}=\{(a,z)\in{\mathbb{RP}}^{5}\times\partial\mathbb{D}\mathrel{\mathop{\ordinarycolon}}\rho(a,\cdot)\neq 0,\ \mathfrak{g}(\Psi(a,z))=1\}\,.$
[…]
    • $\beta^{\pm}_{a,z}\subset\{\pm F_{a,z}>0\}$, and $\beta^{+}_{a,z}\cup\beta^{-}_{a,z}$ is a Hopf link in $S^{3}$ (see Definition 7.8);
    • $\Theta^{\pm}\mathrel{\mathop{\ordinarycolon}}(a,z)\mapsto\beta^{\pm}_{a,z}$ are continuous maps over $\Omega_{\Psi}$ into $\Gamma(S^{3})$, the space of (unoriented) embedded $S^{1}$ in $S^{3}$ (see Appendix E.2);
    • if $\gamma\mathrel{\mathop{\ordinarycolon}}S^{1}\to\Omega_{\Psi}$ is a loop homologous to $\gamma_{\circ}$ in $H_{1}({\mathbb{RP}}^{5}\times\partial\mathbb{D},\mathbb{Z}_{2})$, then
      $(\Theta^{\pm}\circ\gamma)_{*}[S^{1}]\neq 0,\quad\text{ in }\ H_{1}(\Gamma(S^{3});\mathbb{Z}_{2})\,.$
    • if $\gamma\mathrel{\mathop{\ordinarycolon}}S^{1}\to\Omega_{\Psi}$ is a loop homologous to 0 in $H_{1}({\mathbb{RP}}^{5}\times\partial\mathbb{D},\mathbb{Z}_{2})$, then
      $(\Theta^{\pm}\circ\gamma)_{*}[S^{1}]=0,\quad\text{ in }\ H_{1}(\Gamma(S^{3});\mathbb{Z}_{2})\,.$
Definition 7.8.
Recall that $S^{3}\mathrel{\mathop{\ordinarycolon}}=\{x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}=1\}\subset\mathbb{R}^{4}.$ Suppose we have two embedded, non-intersecting loops $\alpha$ and $\beta$ in $S^{3}$. We say that $\alpha\cup\beta$ is a Hopf link if it is isotopic to the standard link
$\{x_{1}=x_{2}=0\}\cup\{x_{3}=x_{4}=0\}.$
Proof of Lemma 7.7.
To prove (a), in the following $\Lambda_{i}>1$ will be absolute constants changing from line to line, and $C(a^{2}_{1}+a^{2}_{2})>0$ also changes from line to line. Recall that
$F_{a,z}(x,\alpha)=(x_{1}+a_{2})(x_{2}+a_{1})$ $+(a_{0}-a_{1}a_{2})+\varsigma(x)(a_{3}\cos\alpha+a_{4}\sin\alpha)$
$+\rho(a,x)(r\cos(\theta+2\alpha)+(1-r)\cos(3\alpha))$
where $\rho(a,\cdot)$ is constant for $x\in\mathbf{D}_{\kappa(a)/4}(-a_{2},-a_{1})$, $\rho(a,\cdot)=0$ for $x\notin\mathbf{D}_{\kappa(a)}(-a_{2},-a_{1})$ and in general,
$|\nabla_{x}\rho|\leq\Lambda_{1}(1-a_{1}^{2}-a_{2}^{2})^{-2}\cdot\delta(a_{1}^{2}+a_{2}^{2})\,.$
Also since $\rho(a,\cdot)\neq 0$, we have
(7.5) $a_{3}^{2}+a_{4}^{2}+(a_{0}-a_{1}a_{2})^{2}\leq\delta(a_{1}^{2}+a_{2}^{2})\,.$
And in $\mathbf{D}_{\kappa(a)}(-a_{2},-a_{1})$,
$|\varsigma|+|\nabla\varsigma|+|\nabla^{2}\varsigma|\leq\Lambda_{2}\ \kappa(a)^{-4}\,.$
Hence, in $\mathbf{D}_{\kappa(a)}(-a_{2},-a_{1})\setminus\mathbf{D}_{\kappa(a)/4}(-a_{2},-a_{1})$,
$\nabla_{x}F_{a,z}(x,\alpha)$ $=(x_{2}+a_{1},x_{1}+a_{2})+O\left(|a_{3}|+|a_{4}|+\delta(a_{1}^{2}+a_{2}^{2})\right)\kappa(a)^{-4}$
$=(x_{2}+a_{1},x_{1}+a_{2})+O\left(\delta(a_{1}^{2}+a_{2}^{2})^{1/2}\right)\kappa(a)^{-4}$
0$\neq$ 0≠ 0
as long as $\delta(a_{1}^{2}+a_{2}^{2})\leq\Lambda_{3}^{-1}\kappa(a)^{10}$. While in $\mathbf{D}_{\kappa(a)/4}(-a_{2},-a_{1})$,
$\nabla_{x}F_{a,z}(x,\alpha)$ $=(x_{2}+a_{1},\ x_{1}+a_{2})+(a_{3}\cos\alpha+a_{4}\sin\alpha)\nabla\varsigma(x)\,;$
$\nabla_{x}F_{a,z}(-a_{2},-a_{1},\alpha)$ $=O(|a_{3}|+|a_{4}|)\kappa(a)^{-4}=O(\delta(a_{1}^{2}+a_{2}^{2})^{1/2})\kappa(a)^{-4}\,;$
(7.6) $\nabla^{2}_{x}F_{a,z}(x,\alpha)=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}+O(|a_{3}|+|a_{4}|)\kappa(a)^{-4}\,.$
Hence consider
$\varpi\mathrel{\mathop{\ordinarycolon}}(y_{1},y_{2})\mapsto-\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\nabla_{x}F_{a,z}(-a_{2}+y_{1},-a_{1}+y_{2},\alpha)+(y_{1},y_{2})$
By taking $\delta(a_{1}^{2}+a_{2}^{2})\leq\Lambda_{4}^{-1}\kappa(a)^{10}$, we can see that $\varpi$ has small Lipschitz constant and is a contraction map from $\mathbf{D}_{\kappa(a)/4}$ to itself, and by Banach Fix Point Theorem, it has a unique fixed point (which corresponds to zero of $\nabla_{x}F_{a,z}$) with length no more than
$|\varpi(0,0)|+\|\nabla\varpi\|_{C^{0}(\mathbf{D}_{\kappa(a)/4})}\leq C(a_{1}^{2}+a_{2}^{2})(|a_{3}|+|a_{4}|)\,.$
This finishes the proof of (a). Moreover, since $\nabla_{x}F_{a,z}(\mathbf{x}_{a}(\alpha),\alpha)=0$, by chain rule we have,
(7.7) $\|\mathbf{x}_{a}-(-a_{2},-a_{1})\|_{C^{6}(\mathbb{S}^{1})}\leq C(a_{1}^{2}+a_{2}^{2})(|a_{3}|+|a_{4}|)\,.$
To prove (b), notice that
$T_{a,z}(\alpha)$ :$\mathrel{\mathop{\ordinarycolon}}=F_{a,z}(-a_{2},-a_{1},\alpha)$
$\ =(a_{0}-a_{1}a_{2})+\varsigma(-a_{2},-a_{1})(a_{3}\cos\alpha+a_{4}\sin\alpha)$
$\ \ +\rho(a,(-a_{2},-a_{1}))(r\cos(\theta+2\alpha)+(1-r)\cos(3\alpha))$
is a trigonometric polynomial of degree 3$\leq$ 3≤ 3 for every $(a,z)$, and it is of degree 222 if $z\in\partial\mathbb{D}$. And since $\rho(a,\cdot)$ is a constant in $\mathbf{D}_{\kappa(a)/4}$,
$f_{a,z}(\alpha)-T_{a,z}(\alpha)$ $=(\mathbf{x}_{a}(\alpha)_{1}+a_{2})(\mathbf{x}_{a}(\alpha)_{2}+a_{1})$
$+(\varsigma(\mathbf{x}_{a}(\alpha))-\varsigma(-a_{2},-a_{1}))(a_{3}\cos\alpha+a_{4}\sin\alpha)\,.$
Therefore
$\|f_{a,z}-T_{a,z}\|_{C^{6}(\mathbb{S}^{1})}$ $\leq C(a_{1}^{2}+a_{2}^{2})\Big{(}\|\mathbf{x}_{a}-(-a_{2},-a_{1})\|_{C^{6}(\mathbb{S}^{1})}^{2}$
$\;\;\qquad\qquad+\;(|a_{3}|+|a_{4}|)\|\mathbf{x}_{a}-(-a_{2},-a_{1})\|_{C^{6}(\mathbb{S}^{1})}\Big{)}$
$\leq C(a_{1}^{2}+a_{2}^{2})(|a_{3}|+|a_{4}|)^{2}$
$\leq\bar{C}(a_{1}^{2}+a_{2}^{2})(|a_{3}|+|a_{4}|)\|T_{a,z}\|\,.$
Hence by taking $\delta(a_{1}^{2}+a_{2}^{2})\ll 1$ such that $\bar{C}(a_{1}^{2}+a_{2}^{2})(|a_{3}|+|a_{4}|)\leq\delta_{1}(3)$ from Corollary 7.5, (b) follows directly by Corollary 7.5.
To prove (c) and (d), by (7.5) and (7.6), we can take $\delta(a_{1}^{2}+a_{2}^{2})$ further small such that assumptions (a)-(c) of Lemma 7.6 hold for $F_{a,z},\ (-a_{2},-a_{1}),\ \kappa(a),\ \mathbf{x}_{a}$ in place of $F,\ b,\ \kappa,\ \mathbf{x}$ therein. Hence, (c) follows immediately from (b) proved above together with Lemma 7.6 (1) and (3). Moreover, if $z=e^{i\theta}\in\partial\mathbb{D}$ and $a=[0\mathrel{\mathop{\ordinarycolon}}\dots\mathrel{\mathop{\ordinarycolon}}0\mathrel{\mathop{\ordinarycolon}}1]$, then $\nabla_{x}F_{a,z}(0,0)=(0,0)$ and hence $\mathbf{x}_{a}(\alpha)=(0,0)$, which implies
[…]
$\breve{\beta}_{a,z}^{+}\mathrel{\mathop{\ordinarycolon}}=\left\{\big{(}\mathbf{x}(\omega_{i}(a,z))+(t,t),\ \omega_{i}(a,z)\big{)}\mathrel{\mathop{\ordinarycolon}}t\in\mathbb{R},\ i\in\{1,3\}\right\}\cap\mathbf{D}\times\mathbb{S}^{1}$
is the union of two line segments contained in $\{F_{a,z}>0\}$, which can be extended to a loop $\beta_{a,z}^{+}$ in $\{F_{a,z}>0\}\subset S^{3}$ by connecting the closer pairs of end points of it in the great circle $C\subset S^{3}$ (which is the image of $\partial\mathbb{D}\times\mathbb{S}^{1}$ under parametrization 7.1). Similarly,
$\breve{\beta}_{a,z}^{-}\mathrel{\mathop{\ordinarycolon}}=\left\{\big{(}\mathbf{x}(\omega_{i}(a,z))+(t,-t),\ \omega_{i}(a,z)\big{)}\mathrel{\mathop{\ordinarycolon}}t\in\mathbb{R},\ i\in\{2,4\}\right\}\cap\mathbf{D}\times\mathbb{S}^{1}$
[…]
$\breve{\beta}^{+}_{a,z}(\tau)$ =$=\left\{\big{(}(1-\tau)\mathbf{x}(\omega_{i}(a,z))+(t,t),\ \omega_{i}(a,z)\big{)}\mathrel{\mathop{\ordinarycolon}}t\in\mathbb{R},\ i\in\{1,3\}\right\}\cap\mathbf{D}\times\mathbb{S}^{1}\,;$
$\breve{\beta}^{-}_{a,z}(\tau)$ =$=\left\{\big{(}(1-\tau)\mathbf{x}(\omega_{i}(a,z))+(t,-t),\ \omega_{i}(a,z)\big{)}\mathrel{\mathop{\ordinarycolon}}t\in\mathbb{R},\ i\in\{2,4\}\right\}\cap\mathbf{D}\times\mathbb{S}^{1}\,.$
Hence $\beta^{+}_{a,z}\cup\beta^{-}_{a,z}$ also forms a Hopf link.
We are left to prove the third and fourth bullet point of (d). We only focus on the third bullet point below, as the fourth follows by similar argument. Let $\gamma\mathrel{\mathop{\ordinarycolon}}S^{1}\mapsto\Omega_{\Psi}$ be a loop homologous to $\gamma_{\circ}$ in $H_{1}({\mathbb{RP}}^{5}\times\partial\mathbb{D})$, set $\gamma(s)=(a(s),e^{-i\theta(s)})$ where $\theta\mathrel{\mathop{\ordinarycolon}}S^{1}\to\mathbb{R}/2\pi\mathbb{Z}$, then $\theta$ has odd degree. Viewing $S^{1}=[0,1]/\{0,1\}$, this mean $\theta$ lifts to a map $\widetilde{\theta}\mathrel{\mathop{\ordinarycolon}}[0,1]\to\mathbb{R}$ such that
(7.8) $\widetilde{\theta}(1)-\widetilde{\theta}(0)\ \text{ is an odd integer multiple of }2\pi\,.$
While recall that $(a,z)\mapsto f^{-1}_{a,z}(0)$ is continuous over $\Omega_{\Psi}$, there exists a continuous lifting of zeros:
$\widetilde{\alpha}_{i}\mathrel{\mathop{\ordinarycolon}}[0,1]\to\mathbb{R},\quad i=1,\dots,4$
such that if we denote by $\alpha_{i}(s)\mathrel{\mathop{\ordinarycolon}}=\widetilde{\alpha}_{i}(s)\ \text{mod }2\pi\in\mathbb{S}^{1}$, then $\forall\,s\in[0,1]$,
$\widetilde{\alpha}_{1}(s)<\widetilde{\alpha}_{2}(s)<\widetilde{\alpha}_{3}(s)<\widetilde{\alpha}_{4}(s)<\widetilde{\alpha}_{1}(s)+2\pi\,;$
(7.9) $f_{a(s),z(s)}^{-1}(0)=\{\alpha_{i}(s)\mathrel{\mathop{\ordinarycolon}}i=1,\dots,4\}\,.$
We thus have
(7.10) $\sum_{i=1}^{4}\widetilde{\alpha}_{i}(0)=\sum_{i=1}^{4}\widetilde{\alpha}_{i}(1)\quad\text{mod }2\pi\,,$
since they projects to the sum of zeros of the same function on $\mathbb{S}^{1}$. On the other hand, after possibly subtracting $\widetilde{\alpha}_{i}$ by a common integer multiple of $2\pi$, we have,
(7.11) $\operatorname{dist}_{\mathbb{R}}\left(\sum_{i=1}^{4}\widetilde{\alpha}_{i}(s),-2\widetilde{\theta}(s)\right)=\operatorname{dist}_{\mathbb{S}^{1}}\left(\sum_{i=1}^{4}\alpha_{i}(s),-2\theta(s)\right)\leq 1000^{-1}\,.$
for every $s\in[0,1]$, where the last inequality follows from (b) proved above. Combining (7.8), (7.10) and (7.11), we conclude that
$\sum_{i=1}^{4}\widetilde{\alpha}_{i}(1)-\sum_{i=1}^{4}\widetilde{\alpha}_{i}(0)=4m\pi$
for some odd integer $m$. Combined with (7.9), this forces (under convention $\alpha_{i}(s)\mathrel{\mathop{\ordinarycolon}}=\alpha_{i-4}(s)$ if $i>4$.)
$\alpha_{i}(1)=\alpha_{i+2}(0),\quad i=1,\dots,4\,.$
This means, when $s$ goes around $S^{1}$, $\Theta^{\pm}\circ\gamma(s)=\beta^{\pm}_{a(s),z(s)}$ both reverse their orientation. Hence neither of $\Theta^{\pm}\circ\gamma\mathrel{\mathop{\ordinarycolon}}S^{1}\to\Gamma(S^{3})$ lifts to a map $S^{1}\to\widetilde{\Gamma}(S^{3})$. By Lemma E.2 and Hurewicz Theorem,
$(\Theta^{\pm}\circ\gamma)_{*}[S^{1}]\neq 0,\quad\text{ in }\ H_{1}(\Gamma(S^{3});\mathbb{Z}_{2})\,.$
7.4. Proof of Theorem 3.2.
The construction of $\Psi$ has been carried out in section 7.3 with the choice of $\delta$ fixed by Lemma 7.7.
The assertion that $\Psi$ takes value in the set of punctate surfaces of genus 2$\leq$ 2≤ 2 as well as (3) of Theorem 3.2 follows from (c) of Lemma 7.7 and the second bullet point above Lemma 7.7.
Theorem 3.2 (1) follows from Lemma 7.1 by noticing that in the construction of $\Psi$, if $a\in{\mathbb{RP}}^{5}$ with $a_{5}=0$, then $a\notin A_{\operatorname{sing}}$ and $\Psi(a,z)=\Phi_{5}(a)$.
Theorem 3.2 (2) follows by noticing that for every $z\in\mathbb{D}$,
$\breve{\omega}(s)\mathrel{\mathop{\ordinarycolon}}=([s/(1-s^{2})\mathrel{\mathop{\ordinarycolon}}1\mathrel{\mathop{\ordinarycolon}}0\mathrel{\mathop{\ordinarycolon}}0\mathrel{\mathop{\ordinarycolon}}0\mathrel{\mathop{\ordinarycolon}}0],z),\quad s\in(-1,1)$
extends to a nontrivial loop $\omega\subset{\mathbb{RP}}^{5}\times\{z\}$ and
$s\mapsto\Psi(\omega(s))=\Phi_{5}(\omega(s))=\{x\in S^{3}\mathrel{\mathop{\ordinarycolon}}x_{1}+s/(1-s^{2})=0\}$
forms a 111-sweepout in the Almgren-Pitts sense. Since $[\omega]$ generates $H^{*}({\mathbb{RP}}^{5}\times\{z\};\mathbb{Z}_{2})$ and clearly $\Psi$ is continuous under flat topology, $\Psi|_{{\mathbb{RP}}^{5}\times\{z\}}$ is a 555-sweepout in the Almgren-Pitts sense.
To prove (4) of Theorem 3.2, let $\gamma\subset{\mathbb{RP}}^{5}\times\partial\mathbb{D}$ be the 111-cycle specified therein. First note that $\gamma$ is the sum of loops in a small neighborhood of $A_{\operatorname{sing}}\times\partial\mathbb{D}\subset{\mathbb{RP}}^{5}\times\partial\mathbb{D}$ which deformation retracts to $[0\mathrel{\mathop{\ordinarycolon}}\dots\mathrel{\mathop{\ordinarycolon}}0\mathrel{\mathop{\ordinarycolon}}1]\times\partial\mathbb{D}$, hence the number of connected component of $\gamma$ that also homologous to $[0\mathrel{\mathop{\ordinarycolon}}\dots\mathrel{\mathop{\ordinarycolon}}0\mathrel{\mathop{\ordinarycolon}}1]\times\partial\mathbb{D}$ in $H_{1}({\mathbb{RP}}^{5}\times\partial\mathbb{D},\mathbb{Z}_{2})$ is odd. Hence by (d) of Lemma 7.7, $\gamma$ is supported in $\Omega_{\Psi}\subset{\mathbb{RP}}^{5}\times\partial\mathbb{D}$ and
(7.12) $\Theta^{\pm}_{*}[\gamma]\neq 0\;\textrm{ in }\;H_{1}(\Gamma(S^{3});\mathbb{Z}_{2}).$
To proceed the proof of Theorem 3.2 (4), let us suppose by contradiction that there exists a Simon-Smith family $\hat{\Phi}\mathrel{\mathop{\ordinarycolon}}W\to\mathcal{S}_{1}$ having the properties described in Theorem 3.2 (4). We claim that we can assume without loss of generality that $\hat{\Phi}$ maps into the set $\mathcal{U}$ of “unknotted elements of $\mathcal{S}_{1}(S^{3})$”. Namely, we first define the following.
Definition 7.9.
[…]
Now, we can assume, by subtracting some components from $W$, that every component of $W$ has some boundary. Hence, each point in $W$ can be connected by some path to $\partial W\cong\gamma$, on which we know $\hat{\Phi}|_{\partial W}=\Psi|_{\gamma}$ maps into $\mathcal{U}$ (by Lemma 7.7 (d)). As a result, we know by the proposition below that the whole $\hat{\Phi}$ must map into $\mathcal{U}$.
Proposition 7.10.
Let $\Phi\mathrel{\mathop{\ordinarycolon}}[0,1]\to\mathcal{S}_{1}(S^{3})$ be a Simon-Smith family such that $\Phi(0)\in\mathcal{U}$. Then $\Phi(t)\in\mathcal{U}$ for every $t\in[0,1]$.
The proof of this proposition is purely topological, and is included in Appendix E.1. In conclusion, we have that $\Phi\mathrel{\mathop{\ordinarycolon}}W\to\mathcal{U}$ is such that:
  • $\partial W\cong\gamma$.
  • $\Phi|_{\partial W}=\Psi|_{\gamma}$.
  • There is a consistent choice, for every $x\in W$, of inside and outside regions for $\Phi(x)$.
However, we have the following abstract extension lemma which (together with Lemma 7.7 (d), which gives the existence of the map $\Upsilon_{\pm}$ on $\partial W$ for $\hat{\Phi}$) implies that using such map $\hat{\Phi}\mathrel{\mathop{\ordinarycolon}}W\to\mathcal{U}$ one can construct an extension $\bar{\Upsilon}\mathrel{\mathop{\ordinarycolon}}W\to\Gamma(S^{3})$ of $\Theta^{+}|_{\gamma}$. This implies that $\Theta^{+}_{*}[\gamma]=0$ in $H_{1}(\Gamma(S^{3});\mathbb{Z}_{2})$, which contradicts (7.12).
Lemma 7.11.
Suppose $W$ is a compact simplicial 222-complex, $\Phi\mathrel{\mathop{\ordinarycolon}}W\to\mathcal{U}\subset\mathcal{S}_{1}(S^{3})$ is a Simon-Smith family of unknotted punctate surfaces of genus 111 (see Definition 7.9), and $U_{\pm}\mathrel{\mathop{\ordinarycolon}}W\to\{\text{open subset of }S^{3}\}$ is continuous map in flat topology such that
$S^{3}\setminus\Phi(q)=U_{+}(q)\sqcup U_{-}(q)\,,$ $\partial U_{\pm}(q)=\Phi(q)\,.$
And suppose $\Upsilon_{\pm}\mathrel{\mathop{\ordinarycolon}}\partial W\to\Gamma(S^{3})$ are continuous maps such that
$\Upsilon_{\pm}(q)\subset U_{\pm}(p)\,,$ $\Upsilon_{+}(q)\cup\Upsilon_{-}(q)\ \text{ is a Hopf link in }\,S^{3}\,.$
Then there exists a continuous map $\bar{\Upsilon}\mathrel{\mathop{\ordinarycolon}}W\to\Gamma(S^{3})$ such that $\bar{\Upsilon}|_{\partial W}=\Upsilon_{+}$.
Hence, to finish the proof of Theorem 3.2 (4), it suffices to prove the above lemma.
Proof.
The extension will be constructed cell by cell.
Claim 1
There exist a subdivision of $W$ (not relabel) and a map associates to each 222-cell $\sigma$ of $W$ a Hopf link $\beta_{+}(\sigma)\cup\beta_{-}(\sigma)$ such that,
  1. (i)
    for every $q\in\partial W$ and $q^{\prime}\in W$ in the same 222-cell of $q$, we have
    $\Upsilon_{\pm}(q)\subset U_{\pm}(q^{\prime})\,;$
  2. (ii)
    for every 222-cell $\sigma$ such that $\sigma\cap\partial W\neq\emptyset$, there exists $q_{\sigma}\in\sigma\cap\partial W$ such that
    $\beta_{\pm}(\sigma)=\Upsilon_{\pm}(q_{\sigma})$
  3. (iii)
    for every 222-cell $\sigma$ and point $q$ in $\sigma$ or its adjacent 222-cell, we have
    $\beta_{\pm}(\sigma)\subset U_{\pm}(q)\,.$
Proof of Claim 1..
By the definition of $\mathcal{U}$, for every $q\in W$, there exists a Hopf link $\check{\beta}_{+}(q)\cup\check{\beta}_{-}(q)\subset S^{3}$ such that $\check{\beta}_{\pm}(q)\subset U_{\pm}(q)$ (of course, $q\mapsto\check{\beta}_{\pm}(q)$ need not be a continuous map), and when $q\in\partial W$, we can just choose $\check{\beta}_{\pm}(q)=\Upsilon_{\pm}(q)$. By the closedness of $\Phi$ (see Definition 2.7 (2)), there exists $\delta_{q}>0$ such that $\check{\beta}_{\pm}(q)\subset U_{\pm}(q^{\prime})$ for every $q^{\prime}\in B_{\delta_{q}}(q)\subset W$. Since $q\mapsto\beta_{\pm}(q)$ is continuous on $\partial W$, we can even ask $\delta_{\partial W}\mathrel{\mathop{\ordinarycolon}}=\inf_{q\in\partial W}\delta_{q}>0$.
We can then choose of finite subset $J\subset W$ such that $\{B_{\delta_{q}/10}(q)\}_{q\in J}$ forms a cover of $W$. Hence, there is a subdivision of $W$ (not relabeled) such that
  • every 222-cell $\sigma$ of $W$ is contained in $B_{\delta_{q(\sigma)}/4}(q(\sigma))$ for some $q(\sigma)\in J$;
  • every 222-cell $\sigma$ that intersects $\partial W$ is contained in $B_{\delta_{\partial W}/4}(\partial W)$.
For every 222-cell $\sigma$ of $W$, define
$\beta_{\pm}(\sigma)\mathrel{\mathop{\ordinarycolon}}=\begin{cases}\Upsilon_{\pm}(q),&\ \text{ if }\sigma\cap\partial W\neq\emptyset,\ q\text{ is some point on }\sigma\cap\partial W\,;\\ \beta_{\pm}(q(\sigma)),&\ \text{ if }\sigma\cap\partial W=\emptyset\,.\end{cases}$
It is easy to verify from the definition that (i) - (iii) in Claim 1 hold. ∎
Now we are ready to construct the extension $\bar{\Upsilon}$.
Step 1. 0-cells
For every 0-cell $q\in W$, define [^3] The choice is not unique, any such choice works for later purpose.
$\bar{\Upsilon}(q)\mathrel{\mathop{\ordinarycolon}}=\begin{cases}\Upsilon_{+}(q),&\ \text{ if }q\in\partial W\,;\\ \beta_{+}(\sigma),&\ \text{ if }q\notin\partial W,\ \text{ pick a $2$-cell }\sigma\text{ that contains }q\,.\end{cases}$
Claim 2
For every 0-cell $q_{\circ}\in W$, if $\sigma$ is an arbitrary 222-cell containing $q_{\circ}$ and $q\in W$ lies in some 111-cell containing $q_{\circ}$, then
$\bar{\Upsilon}(q_{\circ})\subset U_{+}(q)\,,$ $\beta_{-}(\sigma)\subset U_{-}(q)\,;$
and $\bar{\Upsilon}(q_{\circ})\cup\beta_{-}(\sigma)$ forms a Hopf link.
Proof of Claim 2..
First note that $q$ is contained either in $\sigma$ or in its adjacent 222-cells, by (iii) of Claim 1, $\beta_{\pm}(\sigma)\subset U_{\pm}(q)$ forms a Hopf link.
If $q_{\circ}\in\partial W$, (i) of Claim 1 implies that
$\bar{\Upsilon}(q_{\circ})=\Upsilon_{+}(q_{\circ})\subset U_{+}(q),$ $\Upsilon_{-}(q_{\circ})\subset U_{-}(q)$
and they also form a Hopf link. Then by Lemma 7.12, $\bar{\Upsilon}(q_{\circ})\cup\beta_{-}(\sigma)$ also forms a Hopf link.
If $q_{\circ}\notin\partial W$, let $\sigma_{\circ}\ni q_{\circ}$ be the 222-cell such that $\bar{\Upsilon}(q_{\circ})=\beta_{+}(\sigma_{\circ})$. Then $q$ also lies in either $\sigma_{\circ}$ or its adjacent 222-cells, by (iii) of Claim 1, $\beta_{\pm}(\sigma_{\circ})\subset U_{\pm}(q)$ forms a Hopf link. Again by Lemma 7.12, $\bar{\Upsilon}(q_{\circ})\cup\beta_{-}(\sigma)$ also forms a Hopf link. ∎
Step 2. 111-cells
For every 111-cell $e\subset\partial W$, define $\bar{\Upsilon}|_{e}=\Upsilon_{+}|_{e}$. Note that by (i), (ii) of Claim 1, for every 222-cell $\sigma$ containing $e$ and every $q\in e$, we have $\bar{\Upsilon}(q)\cap\beta_{-}(\sigma)=\emptyset$.
Now if $e\not\subset\partial W$, let $q_{1},q_{2}\in W$ be the end points of $e$ (which are 0-cells of $W$) and $q\in W$ be the midpoint of $e$. Claim 1 implies that
  • $\bar{\Upsilon}(q_{i})\subset U_{+}(q)$,
  • for every 222-cell $\sigma^{\prime}$ containing $e$, $\beta_{-}(\sigma^{\prime})\subset U_{-}(q)$,
  • $\bar{\Upsilon}(q_{i})\cup\beta_{-}(\sigma^{\prime})$ forms a Hopf link.
Then by Lemma 7.12, there exists a continuous isotopy $\bar{\Upsilon}|_{e}\mathrel{\mathop{\ordinarycolon}}e\to\Gamma(S^{3})$ extending $\bar{\Upsilon}|_{q_{i}}$, such that $\bar{\Upsilon}(q^{\prime})\cup\beta_{-}(\sigma^{\prime})$ forms a Hopf link for every $q^{\prime}\in e$ and every 222-cell $\sigma^{\prime}$ containing $e$.
Step 3. 222-cells
For every 222-cell $\sigma$ of $W$, Step 2 produces a map $\bar{\Upsilon}|_{\partial\sigma}\mathrel{\mathop{\ordinarycolon}}\partial\sigma\to\Gamma(S^{3})$ which agrees with $\Upsilon_{+}$ on $\partial W$ when $\partial\sigma\cap\partial W\neq\emptyset$, such that $\bar{\Upsilon}(q)\cup\beta_{-}(\sigma)$ forms a Hopf link for every $q^{\prime}\in\partial\sigma$. Since the linking number is never zero, this forces $\bar{\Upsilon}|_{\partial\sigma}\mathrel{\mathop{\ordinarycolon}}\partial\sigma\to\Gamma(S^{3})$ lifts to a map $\partial\sigma\to\widetilde{\Gamma}(S^{3})$, and then by Lemma E.2, there exists a continous map $\bar{\Upsilon}\mathrel{\mathop{\ordinarycolon}}\sigma\to\Gamma(S^{3})$ extending $\bar{\Upsilon}|_{\partial\sigma}$. This finishes the construction of $\bar{\Upsilon}$ on $W$. ∎
The following Lemma used in the proof above is purely topological.
Lemma 7.12.
Let $T\in\mathcal{S}_{1}(S^{3})$ be a punctate surface of genus 111, $U_{\pm}$ be the two open regions bounded by $T$ (i.e. $S^{3}\setminus T=U_{+}\sqcup U_{-}$ such that $\partial U_{\pm}=T$). Suppose for $i=0,1$, $\beta_{i}^{\pm}\subset U_{\pm}$ are embedded loops such that $\beta_{i}^{+}\cup\beta_{i}^{-}$ forms a Hopf link. Then there exists an isotopy $t\mapsto\beta^{\pm}(t)$ of embedded loops over $t\in[0,1]$ such that
  • $\beta^{\pm}(i)=\beta^{\pm}_{i}$, $i=0,1$;
  • for every $t_{\pm}\in[0,1]$, $\beta^{\pm}(t_{\pm})\subset U_{\pm}$ and $\beta^{+}(t_{+})\cup\beta^{-}(t_{-})$ is also a Hopf link.
Proof.
By subtracting sufficiently small balls centered at the punctate set and gluing in discs, we can replace $T$ by some smooth surface of genus 111 that bounds domains which still satisfy the assumptions for the same $\beta_{i}^{\pm}$, hence we may assume without loss of generality that $T$ is smooth everywhere. Also, by the Schoenflies Theorem in $S^{3}$, every connected component of $U_{\pm}$ whose boundary is a 222-sphere must be contractible, hence can’t contain either of $\beta_{i}^{\pm}$. By inductively subtracting all the inner-most or outer-most spherical component (which still doesn’t affect the assumptions on $\beta_{i}^{\pm}$), we may assume without loss of generality that $T$ is diffeomorphic to a smooth tori and $U_{\pm}$ are connected.
By van Kampen Theorem, one of the homomorphism $\iota^{\pm}\mathrel{\mathop{\ordinarycolon}}\pi_{1}(T)\to\pi_{1}(U_{\pm})$ induced by natural inclusion is not injective, say $\iota^{+}$ isn’t injective. Then by Dehn’s Lemma, there exists an embedded nontrivial loop $\gamma\subset T$ which bounds an embedded disc $D\subset U_{+}$. When $\epsilon\ll 1$, after smoothing the corner, $\bar{U}_{+}\setminus B_{\epsilon}(D)$ is a smooth domain with boundary diffeomorphic to $S^{2}$, hence by Schoenflies Theorem in $S^{3}$, $\bar{U}_{+}\setminus B_{\epsilon}(D)$ (after smoothing out corner) is diffeomorphic to $B^{3}$, hence $\bar{U}_{+}$ is diffeomorphic to $B^{3}$ attached with a 111-handle, which is a solid torus $D^{2}\times S^{1}$. In summary, $U_{+}$ is diffeomorphic to the tubular neighborhood of some knot $K\subset U_{+}\subset S^{3}$.
Claim
For every $i=0,1$ and every map from the disc $f\mathrel{\mathop{\ordinarycolon}}D\to\bar{U}_{+}\setminus\beta_{i}^{+}$ such that $f(\partial D)\subset\partial U_{+}=T$, we have $f_{*}[\partial D]=0$ in $H_{1}(T)$.
Proof of Claim..
[…]
$0=H_{2}(S^{3})\to H_{1}(T)\to H_{1}(U_{-})\oplus H_{1}(U_{+})\to H_{1}(S^{3})=0\,,$
there exists a loop $\lambda\subset T$ such that $[\lambda]\in H_{1}(T)$ is mapping to $(0,[K])\in H_{1}(U_{-})\oplus H_{1}(U_{+})$. Hence, $m_{i}\lambda$ is homologous to $\beta_{i}^{+}$ in $U_{+}$, and there is some map $\Sigma\to U_{-}$ from surface $\Sigma$ such that $\partial\Sigma$ is mapping to $\lambda$. Therefore,
±$\pm 1=\operatorname{link}(\beta_{i}^{-},\beta_{i}^{+})=\text{ algebraic intersection of }\beta_{i}^{-}\text{ and }m_{i}\Sigma\,.$
Thus $m_{i}=\pm 1$.
Now since $f(D)\cap\beta^{+}_{i}=\emptyset$, the algebraic intersection of $f(D)$ and $K$ is also 0. This forces $f_{*}[\partial D]=0$ in $H_{1}(T)$ by Poincaré duality. ∎
Now that with this Claim, by Hurewicz Theorem, the natural inclusion $T\hookrightarrow\bar{U}_{+}\setminus\beta_{i}^{+}$ induces an injective map $\pi_{1}(T)\to\pi_{1}(U_{+}\setminus\beta_{i}^{+})$, hence by van Kampen Theorem, the homomorphism induced by inclusion
$\pi_{1}(U_{-}\setminus\beta_{i}^{-})\to\pi_{1}(S^{3}\setminus(\beta_{i}^{+}\cup\beta_{i}^{-}))$
is also injective. Since $\beta_{i}^{+}\cup\beta_{i}^{-}$ forms a Hopf link, the right hand side is an abelian group, and so is the left hand side. While since $U_{+}$ is diffeomorphic to a tubular neighborhood of $K$, we have
$\pi_{1}(S^{3}\setminus(K\cup\beta_{i}^{-}))=\pi_{1}(U_{-}\setminus\beta_{i}^{-})$
is also an abelian group. By [XZ23, Remark 1.7] (see also [Kaw96, Section 6.3]), $K\cup\beta_{i}^{-}$ is also a Hopf link, and then $K$ is an unknot, forcing $U^{-}=S^{3}\setminus\bar{U}_{+}$ to be also diffeomorphic to a tubular neighborhood of a (unknotted) embedded loop. Therefore, for each $i$, there exist $\varepsilon>0$ and an embedded disc $D_{i}\subset S^{3}$ with boundary $\beta_{i}^{-}$ such that $D_{i}\cap B_{\varepsilon}(K)$ is an $\varepsilon$-disc $D_{i}^{\prime}$ whose boundary is a meridian of $K$. We can then construct isotopy from $\beta_{0}^{-}$ to $\partial D_{0}^{\prime}$ in $D_{0}$, from $\partial D_{0}^{\prime}$ to $\partial D_{1}^{\prime}$ in $\partial B_{\varepsilon}(K)$ and from $\partial D_{1}^{\prime}$ to $\partial D_{1}=\beta_{1}^{-}$ in $D_{1}$. Since all these isotopies are not intersecting $B_{\varepsilon/2}(K)$, they can be concatenated and deformed to an isotopy from $\beta_{0}^{-}$ to $\beta_{1}^{-}$ in $U_{-}$. In summary, there exists an isotopy $[0,1]\ni t\mapsto\beta^{-}(t)\subset U_{-}$ such that $\beta^{-}(i)=\beta^{-}_{i}$ for every $i=0,1$. The same process also produces an isotopy $[0,1]\mapsto\beta^{+}(t)\subset U_{+}$ such that $\beta^{+}(i)=\beta^{+}_{i}$ for every $i=0,1$. These isotopies satisfies the desired properties. ∎
Appendix A Some geometric measure theory results
In an $N$-dimensional closed Riemannian manifold $(M,\mathbf{g})$, let $G_{n}(M)$ denote its Grassmanian (unoriented) $n$-plane bundle. Given two $n$-dimensional varifolds $V,W\in\mathcal{V}(M)$, the distance function $\mathbf{F}$ is defined by
$\mathbf{F}(V,W)\mathrel{\mathop{\ordinarycolon}}=\sup\{V(f)-W(f)\mathrel{\mathop{\ordinarycolon}}f\in C(G_{n}(M)),|f|\leq 1,\operatorname{Lip}(f)\leq 1\}\,;$
for a Borel set $B\subset M$, we also define
$\mathbf{F}_{B}(V,W)\mathrel{\mathop{\ordinarycolon}}=\mathbf{F}(V\llcorner G_{n}(B),W\llcorner G_{n}(B))\,,$
where $G_{n}(B)$ denotes the restriction of $G_{n}(M)$ to the base $B$.
Let us first recall a result by Pitts relating the $\mathbf{F}$-norm and the $\mathcal{F}$-norm on the space of modulo 222 integral currents.
Lemma A.1 ([Pit81, 68].).
In a closed $N$-dimensional Riemannian manifold $(M,\mathbf{g})$, given a sequence of modulo 222 integral currents $\{S_{i}\}^{\infty}_{i=1}$ in $\mathbf{I}_{n}(M;\mathbb{Z}_{2})$ and $S\in\mathbf{I}_{n}(M;\mathbb{Z}_{2})$,
$\lim_{i\to\infty}\mathbf{F}(S_{i},S)=0\,,$
if and only if
$\lim_{i\to\infty}\mathcal{F}(S_{i},S)=0,\quad\lim_{i\to\infty}\mathbf{M}(S_{i})=\mathbf{M}(S)\,.$
A direct consequence of this lemma is the following quantitative corollary.
Corollary A.2.
In a closed $N$-dimensional Riemannian manifold $(M,\mathbf{g})$, for any $S\in\mathcal{S}(M)$ and $\varepsilon>0$, there exists $\delta=\delta(M,\mathbf{g},S,\varepsilon)>0$ such that for any $S^{\prime}\in\mathcal{S}(M)$, if
$\mathbf{F}(S^{\prime},S)<\delta\,,$
then
$\mathcal{F}([S^{\prime}],[S])<\varepsilon,\quad|\mathcal{H}^{2}(S^{\prime})-\mathcal{H}^{2}(S)|<\varepsilon\,;$
similarly, if
$\mathcal{F}([S^{\prime}],[S])<\delta,\quad|\mathcal{H}^{2}(S^{\prime})-\mathcal{H}^{2}(S)|<\delta\,,$
then
$\mathbf{F}(S^{\prime},S)<\varepsilon\,.$
Proof.
We argue by contradiction.
Suppose for the sake of contradiction that there exists a sequence $\{S_{i}\}^{\infty}_{i=1}$ in $\mathcal{S}(M)$ such that $\mathbf{F}(S_{i},S)<\frac{1}{i}$, but $\mathcal{F}([S_{i}],[S])\geq\varepsilon$. Then we have
$\lim_{i\to\infty}\mathbf{F}(S_{i},S)=0,\quad\liminf_{i\to\infty}\mathcal{F}([S_{i}],[S])>\varepsilon\,,$
contradicting Lemma A.1.
The other directions can be proved similarly. ∎
Lemma A.3.
In an $N$-dimensional closed Riemannian manifold $(M,\mathbf{g})$, for any $V\in\mathcal{V}_{n}(M)$, $\varepsilon>0$ and any family of Borel sets $\{B_{\alpha}\}_{\alpha\in I}$ such that
$\lim_{r\to 0}\sup_{\alpha\in I}V(\overline{N_{r}(\partial B_{\alpha})})=0\,,$
where $\overline{N_{r}(\partial B_{\alpha})}=\{x\in M\mathrel{\mathop{\ordinarycolon}}\operatorname{dist}_{M}(x,\partial B_{\alpha})\leq r\}$, there exists a constant $\delta=\delta(M,\mathbf{g},V,\varepsilon,\{B_{\alpha}\}_{\alpha\in I})$ satisfying the following property.
For any $W\in\mathcal{V}_{n}(M)$, if $\mathbf{F}(V,W)<\delta$, then
$\sup_{\alpha\in I}\mathbf{F}_{B_{\alpha}}(V,W)\leq\varepsilon\,.$
Proof.
Since $\lim_{r\to 0}\sup_{\alpha\in I}V(\overline{N_{r}(\partial B_{\alpha})})=0$, we can pick $r_{0}\in(0,1)$ such that $\sup_{\alpha\in I}V(\overline{N_{r_{0}}(\partial B_{\alpha})})<\varepsilon/10$.
We set
$\delta\mathrel{\mathop{\ordinarycolon}}=\frac{r_{0}\varepsilon}{1000}\,.$
Firstly, we claim that
$\sup_{\alpha\in I}W(\overline{N_{r_{0}/2}(\partial B_{\alpha})})<\varepsilon/5\,.$
Suppose by contradiction that there exists $\alpha_{0}\in I$ such that $W(\overline{N_{r_{0}/2}(\partial B_{\alpha_{0}})})\geq\varepsilon/5$. We can define a cut-off function $h\mathrel{\mathop{\ordinarycolon}}M\to[0,1]$ with $h\equiv 1$ in $\overline{N_{r_{0}/2}(\partial B_{\alpha_{0}})}$, $g\equiv 0$ outside $\overline{N_{r_{0}}(\partial B_{\alpha_{0}})}$ and $\operatorname{Lip}(g)\leq\frac{3}{r_{0}}$. Then we have
$\frac{\varepsilon}{10}>V(\overline{N_{r_{0}}(\partial B_{\alpha})})\geq V(g)\geq W(g)-\mathbf{F}(V,W)\operatorname{Lip}(g)\geq\frac{\varepsilon}{5}-\frac{\delta r_{0}}{3}\geq\frac{\varepsilon}{10}\,.$
a contradiction.
Then fix an arbitrary $\alpha\in I$. For any Lipschitz function $f\mathrel{\mathop{\ordinarycolon}}M\to\mathbb{R}$ with $\sup|f|\leq 1$ and $\operatorname{Lip}(f)\leq 1$, define
$\widetilde{f}(x)=\begin{cases}f(x),&x\in B_{\alpha}\\ 0,&\operatorname{dist}_{M}(x,\partial B_{\alpha})\geq r_{0}\,.\end{cases}$
We can extend $\widetilde{f}$ to a function defined on $M$ with Lipschitz constant $\frac{1}{r_{0}}$ and $L^{\infty}$-norm 111, still denoted by $\widetilde{f}$. Then we have
$\quad W\llcorner G_{n}(B_{\alpha})(f)-V\llcorner G_{n}(B_{\alpha})(f)$
$=W(\widetilde{f})-V(\widetilde{f})$
$\quad-W\llcorner G_{n}(N_{r_{0}}(\partial B_{\alpha})\setminus B_{\alpha})(\widetilde{f})+V\llcorner G_{n}(N_{r_{0}}(\partial B_{\alpha})\setminus B_{\alpha})(\widetilde{f})$
$\leq|W(\widetilde{f})-V(\widetilde{f})|+\|W\|(N_{r_{0}}(\partial B_{\alpha})\setminus B_{\alpha})+\|V\|(N_{r_{0}}(\partial B_{\alpha})\setminus B_{\alpha})$
$\leq\mathbf{F}(W,V)\operatorname{Lip}(\widetilde{f})+\|W\|(N_{r_{0}}(\partial B_{\alpha})\setminus B_{\alpha})+\|V\|(N_{r_{0}}(\partial B_{\alpha})\setminus B_{\alpha})$
$\leq\mathbf{F}(W,V)\frac{1}{r_{0}}+\varepsilon/5<\varepsilon\,.$
This implies that $\mathbf{F}_{B_{\alpha}}(V,W)\leq\varepsilon$.
Since $\alpha\in I$ is arbitrary, we conclude that $\sup_{\alpha\in I}\mathbf{F}_{B_{\alpha}}(V,W)\leq\varepsilon$. ∎
Lemma A.4 (Isoperimetric inequality, [FF60, 6.2]).
In a closed $N$-dimensional Riemannian manifold $(M,\mathbf{g})$, there exists $\delta=\delta(M,\mathbf{g})>0$ and $\nu>0$ such that, if $T\in\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$ and $\mathcal{F}(T)<\delta$, there exists $S\in\mathbf{I}_{n+1}(M;\mathbb{Z}_{2})$ with
$\partial S=T,\quad\mathbf{M}(S)\leq\nu\mathbf{M}(T)^{\frac{n+1}{n}}\,.$
Appendix B A genus inequality
Lemma B.1.
Let $\Sigma$ be a smooth closed oriented surface, $S\subset\Sigma\times[0,1]$ be an embedded oriented smooth surface homologous to $\Sigma\times\{0\}$ in $H_{2}(\Sigma\times[0,1],\mathbb{Z}_{2})$. Then
$\mathfrak{g}(S)\geq\mathfrak{g}(\Sigma)\,.$
Proof.
Without loss of generality, we assume $\Sigma$ and $S$ are connected and $S\cap\Sigma\times\{0\}=\emptyset$. Consider the projection map onto the first factor $\varpi\mathrel{\mathop{\ordinarycolon}}\Sigma\times[0,1]\to\Sigma$. Since $S$ is homologous to $\Sigma\times\{0\}$, we claim that the mod 222 degree of $\pi\mathrel{\mathop{\ordinarycolon}}=\varpi|_{S}$ is nonzero. To see this, let $x\in\Sigma$ be a regular value of $\pi$, $D\subset\Sigma\times[0,1]$ be the domain such that $\partial[D]=[S]-[\Sigma\times\{0\}]$ in $\mathbf{I}_{2}(\Sigma\times[0,1],\mathbb{Z}_{2})$. Then $D\cap\varpi^{-1}(x)$ is a finite union of line segment whose end points contains exactly one point in $\Sigma\times\{0\}$ and hence odd number of points in $S$, which coincides with $\pi^{-1}(S)$.
Now we shall show that $\pi^{*}\mathrel{\mathop{\ordinarycolon}}H^{1}(\Sigma;\mathbb{Z}_{2})\to H^{1}(S;\mathbb{Z}_{2})$ is injective, which immediately implies the genus inequality. Let $\alpha\in\ker\pi^{*}\cap H^{1}(\Sigma;\mathbb{Z}_{2})$ be arbitrary, for every $\beta\in H^{1}(\Sigma;\mathbb{Z}_{2})$,
$\langle\alpha\cup\beta,[\Sigma]\rangle=\langle\alpha\cup\beta,\pi_{*}[S]\rangle=\langle\pi^{*}\alpha\cup\pi^{*}\beta,[S]\rangle=0\,.$
By the nondegeneracy of cup product, it follows that $\alpha=0$. ∎
Proof of Lemma 2.6.
Let $\eta$ be defined as in Corollary 4.10, and $\varepsilon=\delta(M,\mathbf{g})$ as in Lemma A.4. We also let $r=\delta(M,\mathbf{g},\Sigma,\varepsilon)$ as defined in Corollary 4.10.
By Corollary 4.10, there exist a punctate surface $S^{\prime}$ such that:
  1. (i)
    $S^{\prime}\subset N_{\eta}(\Sigma)$.
  2. (ii)
    $\mathcal{F}(S^{\prime}-\Sigma)\leq\mathbf{F}(S^{\prime},\Sigma)<\varepsilon$.
  3. (iii)
    $\mathfrak{g}(S^{\prime})\leq\mathfrak{g}(S)$.
By (i),  (ii) and Lemma A.4, $S^{\prime}$ and $\Sigma$ are homologous in $H_{2}(N_{\eta}(\Sigma),\mathbb{Z}_{2})$.
Since $N_{\eta}(\Sigma)=\exp^{\perp}(\Sigma,(-\eta,\eta))$ is diffeomorphic to $\Sigma\times(0,1)$, by  (iii) and Lemma B.1, we can conclude that
$\mathfrak{g}(S)\geq\mathfrak{g}(S^{\prime})\geq\mathfrak{g}(\Sigma)\,.$
Appendix C Punctate surfaces with arbitrarily long loops
In this section, we show that for all $L\in\mathbb{R}^{+}$, there exists a punctate surface $S$ of genus 111 in $\mathbb{R}^{3}$ such that its area is 111 but any nontrivial loop in $S$ has length at least $L$.
[…]
Figure 13. A slit torus
It is easy to see that every nontrivial loop in the slit torus has length more than $L$ while the slit torus still has area 111. However, note that this is not a punctate surface in $\mathbb{R}^{3}$, as it is not embedded in $\mathbb{R}^{3}$ and the singular set is a slit instead of a finite set of points.
Next, in $\mathbb{R}^{3}$, we can choose a sufficiently small embedded torus with one point removed, denoted by $T_{2}$, and a smooth embedding $\Phi\mathrel{\mathop{\ordinarycolon}}T_{1}\to\mathbb{R}^{3}$ which maps the slit in $T_{1}$ to the point in $T_{2}$. Moreover, we require that $\Phi$ is a strictly short map in the sense that its Lipschitz constant is less than 111; see Figure 14.
Figure 14. A short embedding
Finally, one can follow Nash-Kuiper’s spiral-strain construction [Kui55] to deform $\Phi$ and, without taking the limit, to obtain a smooth almost isometry $\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}T_{1}\to\mathbb{R}^{3}$ such that $T^{\prime}_{2}\mathrel{\mathop{\ordinarycolon}}=\Phi^{\prime}(T_{1})$ is still a torus with a point removed. One can also verify that $T^{\prime}_{2}$ is a punctate surface we need. The details are left to the readers.
Appendix D Proof of Theorem 2.21
In this section, we use Proposition 5.1 and the following local min-max theorem to finish the proof of Theorem 2.21. The proof closely follows that in [CL24, § 8.5], with a slight improvement: we no longer modify the parameter space. For completeness, we include the details below.
Theorem D.1 (Local min-max theorem [MN21, Theorem 6.1]).
[…]
  1. (1)
    $F_{0}=\operatorname{Id}$, $F_{v}=F^{-1}_{v}$ for all $v\in\overline{\mathbb{B}}^{k}$.
  2. (2)
    $\|F_{v}-\operatorname{Id}\|_{C^{1}}<\beta$ for all $v\in\overline{\mathbb{B}}^{k}$.
  3. (3)
    The function
    $A^{\Sigma}\mathrel{\mathop{\ordinarycolon}}\overline{\mathbb{B}}^{k}\to[0,\infty],v\mapsto\mathcal{H}^{2}((F_{v})_{\#}\Sigma)\,,$
    is strictly concave.
  4. (4)
    For every $T\in\mathcal{Z}_{2}(M;\mathbb{Z}_{2})$ with $\mathcal{F}(T,[\Sigma])<\varepsilon_{0}$, we have
    $\max_{v\in\overline{\mathbb{B}}^{k}}\mathbf{M}((F_{v})_{\#}T)\geq\mathcal{H}^{2}(\Sigma)$
    with equality only if $[\Sigma]=(F_{v})_{\#}T$ for some $v\in\overline{\mathbb{B}}^{k}$.
We proceed the proof in three steps.
Step 1. Set up
Since $\mathcal{W}_{L,\leq g}(M,\mathbf{g})$ consists of a varifold associated with a non-degenerate multiplicity-one minimal surface $\Gamma$, we define
$\delta_{0}\mathrel{\mathop{\ordinarycolon}}=\delta(M,\mathbf{g},\Gamma,r/10),$
where the function $\delta(\cdot,\cdot,\cdot,\cdot)$ is as defined in Corollary A.2.
Let $\varepsilon_{0}>0$ and $\{F_{v}\}_{v\in\overline{\mathbb{B}}^{k}}$ be the smooth family associated with $\Gamma$ given by Theorem D.1 such that:
  • (i)
    For every $S_{1},S_{2}\in\mathcal{S}(M)$ and $v\in\overline{\mathbb{B}}^{k}$,
    $\mathbf{F}(F_{v}(S_{1}),F_{v}(S_{2}))\leq 2\mathbf{F}(S_{1},S_{2})\,,$
    and
    $\mathbf{F}(F_{v}(S_{1}),S_{1}))\leq\frac{\delta_{0}}{10L}\mathcal{H}^{2}(S_{1})\,.$
We can choose smaller $\varepsilon_{0}>0$ such that:
  • (ii)
    For every $S\in\mathcal{S}(M)$ with $\mathbf{F}(S,\Gamma)<\varepsilon_{0}$, the function
    $A^{T}\mathrel{\mathop{\ordinarycolon}}\overline{\mathbb{B}}^{k}\to[0,\infty],v\mapsto\mathcal{H}^{2}(F_{v}(S))\,,$
    is strictly concave with a unique maximum in $\mathbb{B}^{k}_{1/2}$.
  • (iii)
    We have
    $\varepsilon_{1}\mathrel{\mathop{\ordinarycolon}}=\min_{v\in\partial\mathbb{B}^{k}}(\mathcal{H}^{2}(\Gamma)-\mathcal{H}^{2}(F_{v}(\Gamma)))>0\,.$
[…]
$\delta_{1}\mathrel{\mathop{\ordinarycolon}}=\delta(M,\mathbf{g},\mathscr{S},m+1,N_{P}(\Phi)+g,\min(\delta_{0},\varepsilon_{0},\varepsilon_{1})/10)$
as defined in Proposition 5.1.
Step 2. Initial Simon-Smith family
Let $r_{0}\mathrel{\mathop{\ordinarycolon}}=\delta_{1}/10$. It follows from Theorem 2.17 with $r=r_{0}$ that there exist $\eta\in(0,\delta_{0}/100)$ and $\Phi_{1}\in\Lambda(\Phi)$ such that
$\mathcal{H}^{2}(\Phi_{1}(x))\geq L-4\eta\implies[\Phi_{1}(x)]\in\mathbf{B}^{\mathbf{F}}_{r_{0}}([\Gamma])\,.$
In particular, there exists a Simon-Smith family
$H_{1}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}(M)$
such that $H_{1}(0,\cdot)=\Phi$ and $H_{1}(1,\cdot)=\Phi_{1}$. In fact, for every $x\in X$, $H_{1}(\cdot,x)$ is a pinch-off process, because it is induced by a one-parameter group of diffeomorphisms.
We refine the cubical subcomplex $X$ so that, for each cell $\sigma$ of $X$, exactly one of the following conditions holds:
  • There exists a point $x_{0}\in\sigma$ such that $\mathcal{H}^{2}(\Phi_{1}(x_{0}))\geq L-2\eta$; Moreover, in this case, for every cell $\sigma^{\prime}$ with $\sigma\cap\sigma^{\prime}\neq\emptyset$ and $x\in\sigma^{\prime}$, $\mathcal{H}^{2}(\Phi_{1}(x))\geq L-4\eta$.
  • For every $x\in\sigma$, $\mathcal{H}^{2}(\Phi_{1}(x))<L-2\eta$.
We denote the side length of each cell in $X$ by $\ell$. Let $Y$ be the smallest cubical subcomplex of $X$ containing all cells that satisfy the first condition, and let $\widetilde{Y}$ be the cubical subcomplex of $X$ whose underlying set is
$\{x\in X\mathrel{\mathop{\ordinarycolon}}d(x,Y)\leq\ell\}\,.$
Let $Z\mathrel{\mathop{\ordinarycolon}}=\overline{\widetilde{Y}\setminus Y}$, which is compact. Clearly, for every $z\in Z$,
$L-4\eta\leq\mathcal{H}^{2}(\Phi_{1}(z))\leq L-2\eta\,.$
Step 3. Deformation
For every $z\in Z$, let $A^{z}\mathrel{\mathop{\ordinarycolon}}\overline{\mathbb{B}}^{k}\to[0,\infty)$ be the function
$A^{z}(v)\mathrel{\mathop{\ordinarycolon}}=\mathcal{H}^{2}(F_{v}(\Phi_{1}(z)))\,,$
By (ii) of Step 1, $A^{z}$ is strictly concave and has a unique maximum $m(z)\in\mathbb{B}^{k}_{1/2}$. As shown in [CL24, §2], the Simon-Smith family $\Phi_{1}$ is continuous in the $\mathbf{F}$-metric, so the function
$m\mathrel{\mathop{\ordinarycolon}}Z\to\mathbb{B}^{k}_{1/2}$
is also continuous.
It follows from Theorem D.1 (4) and $\mathcal{H}^{2}(\Phi_{1}(z))<L$, that $m(z)\neq 0$ for every $z\in Z$. Hence, there exists $\alpha>0$ such that
$\alpha\leq|m(z)|<1/2\,.$
Consider the one-parameter flow $\{\phi^{z}(\cdot,t)\}_{t\geq 0}\subset\operatorname{Diff}(\overline{\mathbb{B}}^{k})$ generated by
$v\mapsto-(1-|v|^{2})\nabla A^{z}(v)\,.$
For every $v\in\overline{\mathbb{B}}^{k}\setminus m(y)$, $t\mapsto A^{z}(\phi^{z}(v,t))$ is decreasing, and the limit $\lim_{t\to\infty}\varphi^{z}(v,t)\in\partial\mathbb{B}^{k}$. In particular, by (iii) of Step 1, we have
$A^{z}(0)-\lim_{t\to\infty}A^{z}(\phi^{z}(v,t))\geq\varepsilon_{1}\,.$
By the compactness of $Z$, we can choose $T_{0}>0$ such that for every $z\in Z$, $t\mapsto A^{z}(\phi^{z}(v,t))$ is decreasing along $[0,T_{0}]$ and
$A^{z}(0)-A^{z}(\phi^{z}(v,T_{0}))\geq\varepsilon_{1}/2\,.$
Now, let us define a bump function $\xi\mathrel{\mathop{\ordinarycolon}}[0,\infty)\to[0,1]$ by
$\xi(s)\mathrel{\mathop{\ordinarycolon}}=\begin{cases}3s/\ell&s\in[0,\ell/3)\\ 1&s\in[\ell/3,2\ell/3)\\ 3-3s/\ell&s\in[2\ell/3,\ell)\\ 0&s\in[\ell,\infty)\,.\end{cases}$
Then we define a new Simon-Smith family $\Phi_{2}\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ by
$\Phi_{2}(x)\mathrel{\mathop{\ordinarycolon}}=F_{\phi^{x}(0,\xi(d(x,Y))\cdot T_{0})}(\Phi_{1}(x))\,.$
We also define a deformation $H_{2}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}(M)$ from $\Phi_{1}$ to $\Phi_{2}$ by
$H_{2}(t,x)$ $=F_{\phi^{x}(0,t\cdot\xi(d(x,Y))\cdot T_{0})}(\Phi_{1}(x))\,.$
Note that when $x\notin Z$, we abuse the notation by setting $F_{\phi^{x}(0,0)}=\operatorname{Id}$.
Similarly, for every $x\in X$, $H_{2}(\cdot,x)$ is also induced by an isotopy, and thus, a pinch-off process. In particular, we have
(D.1) $N_{P}(\Phi)=N_{P}(\Phi_{1})=N_{P}(\Phi_{2})\,.$
Furthermore, $\Phi_{2}$ has the following properties: For every $x\in X$,
  • (i)
    if $d(x,Y)>0$,
    $\mathcal{H}^{2}(\Phi_{2}(x))\leq L-2\eta\,;$
  • (ii)
    if $d(x,Y)\leq 2\ell/3$, by (i) of Step 1,
    $\mathbf{F}(\Phi_{2}(x),\mathscr{S})\leq 2\mathbf{F}(\Phi_{1}(x),\Gamma)<2r_{0}<\delta_{1}\,,$
    $\mathbf{F}(\Phi_{2}(x),\Gamma)$ $\leq\mathbf{F}(\Phi_{2}(x),\mathscr{S})+\mathbf{F}(\mathscr{S},\Gamma)$
    $<\delta_{1}+\delta_{0}/10$
    $\leq\delta_{0}/5,$
    and thus,
    $\mathcal{F}([\Phi_{2}(x)],[\Gamma])\leq\delta_{0}/10\,.$
  • (iii)
    if $d(x,Y)\in[\ell/3,2\ell/3]$,
    $\mathcal{H}^{2}(\Phi_{2}(x))\leq L-2\eta-\varepsilon_{1}/2\,.$
Step 4. Interpolation
Let $Y^{\prime}\mathrel{\mathop{\ordinarycolon}}=\{x\in X\mathrel{\mathop{\ordinarycolon}}d(x,Y)\leq 2\ell/3\}$. By Step 3 (ii), we can apply Proposition 5.1 to $\Phi_{2}|_{Y^{\prime}}$ and obtain a deformation
$H_{3,Y^{\prime}}\mathrel{\mathop{\ordinarycolon}}[0,1]\times Y^{\prime}\to\mathcal{S}(M)\,.$
Note that by the choice of $\delta$, we can assume $\varepsilon=\min(\delta_{0},\varepsilon_{0},\varepsilon_{1})/10$ in Proposition 5.1.
Consequently, we can define a Simon-Smith family $H_{3}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}(M)$ as
$H_{3}(t,x)=\begin{cases}H_{3,Y^{\prime}}(t,x)&d(x,Y)\leq\ell/3\\ H_{3,Y^{\prime}}\left((2-3d(x,Y)/\ell)t,x\right)&x\in Y^{\prime}\\ \Phi_{2}(x),&x\notin Y^{\prime}\,.\end{cases}$
We denote $H_{3}(1,\cdot)\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ by $\Phi_{3}$.
By Proposition 5.1 (1), for every $x\in X$ with $d(x,Y)\leq\ell/3$,
$\mathfrak{g}(\Phi_{3}(x))=\mathfrak{g}(H_{3}(1,x))=\mathfrak{g}(\Gamma)\,,$
and
$\mathcal{H}^{2}(\Phi_{3}(x))=\mathcal{H}^{2}(H_{3,Y^{\prime}}(1,x))\leq\mathcal{H}^{2}(\Phi_{2}(x))+\delta_{0}/10\,.$
By Proposition 5.1 (2), Step 3 (ii) and (iii), for every $x\in X$ with $d(x,Y)\in[\ell/3,2\ell/3]$, we have
$\mathcal{H}^{2}(\Phi_{3}(x))=\mathcal{H}^{2}\left((2-3d(x,Y)/\ell)t,x\right)\leq\mathcal{H}^{2}(\Phi_{2}(x))+\varepsilon_{1}/10\leq L-2\eta\,.$
By Step 3 (i), for every $x\in X$ with $d(x,Y)\geq 2\ell/3$,
$\mathcal{H}^{2}(\Phi_{3}(x))=\mathcal{H}^{2}(\Phi_{2}(x))\leq L-2\eta\,.$
Therefore, for every $x\in X$ with $\mathcal{H}^{2}(\Phi_{3}(x))\geq L-\eta$, we have $d(x,Y)\leq\ell/3$, which implies:
  • (i)
    $\mathfrak{g}(\Phi_{3}(x))=\mathfrak{g}(\Gamma)$;
  • (ii)
    $|\mathcal{H}^{2}(\Phi_{3}(x))-L|\leq\delta_{0}$;
  • (iii)
    $\mathcal{F}([\Phi_{3}(x)],[\Gamma])\leq\mathcal{F}([\Phi_{3}(x)],[\Phi_{2}(x)])+\mathcal{F}([\Phi_{2}(x)],[\Gamma])<\delta_{0}\,.$
In particular, by (ii), (iii) and Corollary A.2, we can conclude that
$\mathbf{F}(\Phi_{3}(x),\Gamma)<r\,.$
Moreover, by Proposition 5.1 (3), for every $x\in X$, $\{H_{3}(t,x)\}_{t\in[0,1]}$ is also a pinch-off process.
Finally, we can set $\Phi^{\prime}=\Phi_{3}$, which satisfies (1) of the theorem. The deformation map
$H\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}(M)\,,$
is simply the concatenation of $H_{1}$, $H_{2}$ and $H_{3}$ defined above. Clearly, it is a Simon-Smith family of genus $\leq g$ satisfying (2)(a) and (2)(b) of the theorem.
This completes the proof.
Appendix E Some topological facts
E.1. Unknotted elements in $\mathcal{S}_{1}(S^{3})$
In this section, we prove Proposition 7.10.
[…]
Lemma E.1.
Let $S\in\mathcal{S}_{1}(S^{3})$ and let $S^{\prime}$ be obtained from $S$ by performing a neck-pinch surgery, such that $S^{\prime}$ still has genus 111. Then $S\in\mathcal{U}$ if and only if $S^{\prime}\in\mathcal{U}$.
Proof.
For the “if” part, by definition there exist some loops $\alpha\subset\mathrm{in}(S^{\prime})$ and $\beta\subset\mathrm{out}(S^{\prime})$ that form a Hopf link. When we undo the neck-pinch surgery, we can first deform the two loops to avoid the surgery region, so that after gluing back in the 1-handle to obtain $S$, we still have $\alpha\subset\mathrm{in}(S)$ and $\beta\subset\mathrm{out}(S)$. So $S\in\mathcal{U}$.
To prove the “only if” part, we first choose some loops $\alpha\subset\mathrm{in}(S)$ and $\beta\subset\mathrm{out}(S)$ that form a Hopf link. Let $\gamma\subset S$ be the smooth embedded loop along which the surgery is performed, and $D$ be the smooth embedded disc it bounds such that the surgery process involves gluing back two discs close to $D$. Without loss of generality, we choose an inward direction for $S$, and assume that the interior of $D$ lies in $\mathrm{in}(S)$.
Step 1: First, by definition, $S$ has at most finitely many singularities. We let $B_{1},\dots,B_{n}\subset S^{3}$ be sufficiently small balls centered at these singularities such that:
  • Each $B_{i}$ avoids $\alpha$ and $\beta$.
  • $S\backslash\cup^{n}_{i=1}B_{i}$ still has genus 1.
  • Each $\partial B_{i}$ intersects $S$ on smooth part, and transversely.
  • Each $B_{i}$ avoids some neighborhood of $D$ such that the surgery process would leave each $B_{i}$ intact.
For each ball $B_{i}$, the intersection $S\cap\partial B_{i}$ consists of a finite union of smooth loops $\{\gamma_{k}\}_{k}$. We then perform neck-pinch surgery along these loops one-by-one: Namely, whenever such a loop $\gamma_{k}$ bounds a disc on $\partial B_{i}$ that does not contain any other loop $\gamma_{k^{\prime}}$, we remove a short cylinder piece from $S$ (whose geometry is roughly $\gamma_{k}\times[0,\epsilon]$) and then glue back two discs (each being close to the disc on $\partial B_{i}$ bounded by $\gamma_{k}$). We perform such surgery operation for all $\gamma_{k}$ and for all balls $B_{i}$, and then remove all components lying within $\cup_{i}B_{i}$, and denote the resulting surface at the end by $\widetilde{S}$. Note that by assuming the balls $B_{i}$ to be sufficiently small, $\widetilde{S}$ is a smooth genus one surface: This follows from $\mathfrak{g}(S)=1$ and the definition of genus for elements in $\mathcal{S}(M)$ in Definition 2.5.
Step 2: Now, by the way we chose the balls $B_{i}$, we would still have $\alpha\subset\mathrm{in}(\widetilde{S})$ and $\beta\subset\mathrm{out}(\widetilde{S})$. Hence, since $\alpha\cup\beta$ is a Hopf link, $\widetilde{S}$ is a genus one Heegaard surface. Then, using the assumption $\mathfrak{g}(S^{\prime})=1$, we know that the surface obtained from $\widetilde{S}$ by performing a surgery along $\gamma$, using the disc $D$, has genus one. Hence, the disc $D$ can be “pushed” (i.e. deformed via smooth embeddings) within $\overline{\mathrm{in}(\widetilde{S})}$ to become a disc $\widetilde{D}\subset\widetilde{S}$, while keeping the boundary $\gamma$ fixed. Then, we can use the reverse of this pushing procedure to deform $\alpha$, and obtain a loop $\widetilde{\alpha}\subset\mathrm{in}(\widetilde{S})$ that does not intersect $D$. As a result, when we perform surgery on $\widetilde{S}$ along $\gamma$ using the disc $D$, the loops $\widetilde{\alpha}$ and $\beta$ would be left intact. Call this surface $\widetilde{S}^{\prime}$
Step 3: Finally, on the new surface $\widetilde{S}^{\prime}$, we undo the procedure described in Step 1. Namely, for $\widetilde{S}^{\prime}$, we put back the connected components discarded, and undo all the neck-pinch surgeries (near the balls $B_{i}$), by gluing back the 1-handles, so that we obtain the surface $S^{\prime}$ with singularities: This $S^{\prime}$ is precisely the surface of interest in Lemma E.1. It is easy to see that $\alpha\subset\mathrm{in}(S^{\prime})$ and $\beta\subset\mathrm{out}(S^{\prime})$, and $\alpha\cup\beta$ is a Hopf link, we have $S^{\prime}\in\mathcal{U}$
We now prove Proposition 7.10. Suppose by contradiction there exists some $\Phi(t)\notin\mathcal{U}$. Let
$T\mathrel{\mathop{\ordinarycolon}}=\inf\{t\in[0,1]\mathrel{\mathop{\ordinarycolon}}\Phi(t)\notin\mathcal{U}\}.$
First, we claim that $\Phi(T)\notin\mathcal{U}$. Indeed, if $\Phi(T)\in\mathcal{U}$, then we can find embedded loops $\alpha\subset\mathrm{in}(\Phi(T))$ and $\beta\subset\mathrm{out}(\Phi(T))$ such that they form a Hopf link: Note that we can consistently choose an inward direction for each $\Phi(t)$. Then using the definition of Simon-Smith family, for any $t$ sufficiently close to $T$, we have $\alpha\subset\mathrm{in}(\Phi(t))$ and $\beta\subset\mathrm{out}(\Phi(t))$, and thus $\Phi(t)\in\mathcal{U}$ for such $t$. This contradicts the definition of $T$.
As a result, $T>0$, since $\Phi(0)\in\mathcal{U}$ by assumption. Take $t_{1},t_{2},\dots\uparrow T.$ By the definition of Simon-Smith family, we can choose finitely many small open balls $B_{j}\subset S^{3}$ such that $\Phi(t_{i})\to\Phi(T)$ smoothly outside $B\mathrel{\mathop{\ordinarycolon}}=\cup_{j}B_{j}$. Without loss of generality, we can assume also that:
  • The boundary of each $B_{j}$ intersects each $\Phi(t_{i})$ on the smooth part only, and such intersection is transverse, for each $t_{i}$.
[…]
Now, we fix $t_{i}$. We will modify the surface $\Phi(t_{i})$ such that it avoids $\partial B$ as follows. For each ball $B_{j}$, the intersection $\Phi(t_{i})\cap\partial B_{j}$ consists of a finite union of smooth loops $\{\gamma_{k}\}_{k}$. We then perform neck-pinch surgery along these loops one-by-one: Namely, whenever such a loop $\gamma_{k}$ bounds a disc on $\partial B_{j}$ that does not contain any other loop $\gamma_{k^{\prime}}$, we remove a short cylinder piece from $\Phi(t_{i})$ (whose geometry is roughly $\gamma_{k}\times[0,\epsilon]$) and then glue back two discs (each roughly looks like the disc on $\partial B_{j}$ bounded by $\gamma_{k}$). We perform such surgery operation for all $\gamma_{k}$, for all ball $B_{j}$, and denote the resulting surface at the end by ${\Phi(t_{i})^{\prime}}$. As a result, ${\Phi(t_{i})^{\prime}}$ avoids $\partial B$, and is smooth outside $B$.
[…]
This finishes the proof of Proposition 7.10.
E.2. Topology of the space of embedded circles
For a closed manifold $M$, we denote by
$\Gamma(M)$ :=$\mathrel{\mathop{\ordinarycolon}}=\{\text{unoriented embedded loops in }M\}\,;$
$\widetilde{\Gamma}(M)$ :=$\mathrel{\mathop{\ordinarycolon}}=\{\text{oriented embedded loops in }M\}\,.$
And let $\mathrm{p}\mathrel{\mathop{\ordinarycolon}}\widetilde{\Gamma}(M)\to\Gamma(M)$ be the orientation forgetting map (which is a double cover). We shall need the following consequence of a Theorem by Hatcher.
Lemma E.2.
$\pi_{1}(\Gamma(S^{3}))=\mathbb{Z}_{2}\,.$
And a loop $\gamma\mathrel{\mathop{\ordinarycolon}}\mathbb{S}^{1}\to\Gamma(S^{3})$ is trivial in $\pi_{1}(\Gamma(S^{3}))$ if and only if it lifts to a loop $\widetilde{\gamma}\mathrel{\mathop{\ordinarycolon}}\mathbb{S}^{1}\to\widetilde{\Gamma}(S^{3})$.
Proof.
We view = $S^{3}$ as the unit sphere in $\mathbb{R}^{4}$, let
$\Gamma_{\circ}(S^{3})$ :=$\mathrel{\mathop{\ordinarycolon}}=\{\text{unoriented great circles in }S^{3}\}\,;$
$\widetilde{\Gamma}_{\circ}(S^{3})$ :=$\mathrel{\mathop{\ordinarycolon}}=\{\text{oriented great circles in }S^{3}\}=\mathrm{p}^{-1}(\Gamma_{\circ}(S^{3}))\,.$
By [Hat83, Appendix (7)], $\Gamma(S^{3})$ has a deformation retract to $\Gamma_{\circ}(S^{3})$, which naturally lifts to a deformation retract of $\widetilde{\Gamma}(S^{3})$ to $\widetilde{\Gamma}_{\circ}(S^{3})$. The lemma follows immediately since $\widetilde{\Gamma}_{\circ}(S^{3})$ is the space of oriented 222-planes in $\mathbb{R}^{4}$, which is diffeomorphic to $S^{2}\times S^{2}$ and hence is connected and simply connected. ∎
E.3. A Lemma about cap product
In this section we prove Lemma 3.9. We consider the following diagram.
[…]
Here, all four horizontal arrows are given by cap products, and all the vertical arrows are induced by inclusions: Note that the inclusion
$(C,C\cap A)\hookrightarrow(C,W_{1}\cap(A\cup B))$
(for the middle arrow in the middle column) and
$(C,W_{2}\cap(A\cup B))\hookrightarrow(C,C\cap B)$
(for the middle arrow in the right column) are well-defined because $W_{2}\cap A=\emptyset$ by assumption. Moreover, this diagram satisfies the naturality property for relative cap products (see [Hat02, p.241]).
To prove Lemma 3.9, we need to “pull $[C]\frown\omega$ backward” along this chain of arrows induced by inclusions
(E.2) $H_{p-q}(W_{2},W_{2}\cap B)\rightarrow H_{p-q}(C,W_{2}\cap(A\cup B))\rightarrow H_{p-q}(C,C\cap B)\rightarrow H_{p-q}(X,B).$
More precisely, we need to find some element $\theta\in H_{p-q}(W_{2},W_{2}\cap B)$ whose pushforward under this three maps is $[C]\frown\omega$. Hence, there are three steps.
Step 1: Observe that $[C]\frown\omega$ has $[C]\frown\omega|_{C}\in H_{p-q}(C,C\cap B)$ (in the latter expression we treat $[C]\in H_{p}(C,C\cap(A\cup B))$) as a preimage, using the naturality for the first two rows of (E.1).
Step 2: Then, noting $C\cap A=W_{1}\cap A\subset W_{1}\subset C$, we consider the exact sequence for this triple:
(E.3) $H^{q}(C,W_{1})\to H^{q}(C,C\cap A)\to H^{q}(W_{1},C\cap A).$
Since the pullback $\omega|_{W_{1}}$ of $\omega$ under the inclusion $(W_{1},W_{1}\cap A)\hookrightarrow(X,A)$ is zero by assumption, we know that $\omega|_{C}$ gets mapped to zero under the map $H^{q}(C,C\cap A)\to H^{q}(W_{1},C\cap A)$ (recall $C\cap A=W_{1}\cap A$). Hence, using the exact sequence (E.3), we know that under the map $H^{q}(C,W_{1})\to H^{q}(C,C\cap A)$, there exists some $\mu$ that gets mapped to $\omega|_{C}$. Thus, there exists some $\mu^{\prime}$ that gets mapped to $\omega|_{C}$ under the map $H^{q}(C,W_{1}\cap(A\cup B))\to H^{q}(C,C\cap A)$ (namely, we define $\mu^{\prime}$ as the image of $\mu$ under the bottom arrow in the middle column of (E.1)). Hence, using the second and the third row of (E.1) and by naturality, we know $[C]\frown\omega|_{C}$ has $[C]\frown\mu^{\prime}$ as a preimage, under the map
$H_{p-q}(C,W_{2}\cap(A\cup B))\rightarrow H_{p-q}(C,C\cap B).$
Step 3: Finally, based on the inclusions $W_{2}\cap B\subset W_{2}\subset C$, we consider the exact sequence
$H_{p-q}(W_{2},W_{2}\cap B)\to H_{p-q}(C,W_{2}\cap B)\to H_{p-q}(C,W_{2}).$
Note the middle term here is the same as $H_{p-q}(C,W_{2}\cap(A\cup B))$ as $W_{2}\cap A=\emptyset$. Thus, in order to obtain the desired $\theta\in H_{p-q}(W_{2},W_{2}\cap B)$ that gets mapped to $[C]\frown\mu^{\prime}$ under the first arrow of (E.2), it suffices to show that $[C]\frown\mu^{\prime}$ gets mapped to 0 under the map
(E.4) $H_{p-q}(C,W_{2}\cap B)\to H_{p-q}(C,W_{2}).$
But now note that this map (E.4) is precisely the map at the bottom of the right column of (E.1). So we can deduce that $[C]\frown\mu^{\prime}$ would indeed get mapped to 0 from these facts:
  • By definition $\mu$ is sent to $\mu^{\prime}$ under the map $H^{q}(C,W_{1})\to H^{q}(C,W_{1}\cap(A\cup B))$.
  • $H_{p}(C,C)=0$.
  • The last two rows of (E.1) satisfy the naturality for cup product.
This completes the proof of Lemma 3.9.