### v2

# Min-max theory and minimal surfaces with prescribed genus

Abstract.
We establish a general min-max type theorem that produces minimal surfaces with prescribed genus in 3-manifolds with positive Ricci curvature. An important intermediate step is to show that, in a generic metric with positive Ricci curvature, any family of smooth embedded surfaces, possibly with finitely many singularities, can be deformed into a certain “topologically optimal family”.
Results in this paper will be crucial to our program on the construction of multiple minimal surfaces with prescribed genus in 3-spheres via topological methods [chu2025arbitraryGenus, chuLiWang2025fourGenus2].
1. Introduction
In recent years we have seen a rapid development on the construction of minimal surfaces from min-max theory. The Almgren-Pitts min-max theory has lead to the proof of Yau’s conjecture on the existence of infinitely many minimal surfaces in every 3-manifold [MN17, IMN18, ChodoshMantoulidis2020AC, Zho20, Son23]. Another line of research is to construct of minimal surfaces of prescribed genus. It was conjectured by Yau (1982) that every 3-sphere has at least 4 minimal spheres, and by White that every 3-sphere has least 5 minimal tori. Wang–Zhou [WangZhou23FourMinimalSpheres] verified Yau’s conjecture for generic metrics and metrics with positive Ricci curvature. In [chuLi2024fiveTori], the first two authors verified White’s conjecture for metrics with positive Ricci curvature. Regarding more works on the construction of minimal submanifolds of controlled topological types under general metrics, see also [LS47, Str84, GJ86, Grayson89, Whi91, Zho16, HK19, bettiolPiccione2023bifurcationsCliffordTorus, HK23, Ko23a, Ko23b, bettiolPiccione2024nonplanarMinSpheres, LiWang2024NineTori, liWangYao2025minimalLensSpace].
The above conjectures lead to the following question:
Question 1.1.
In any given Riemannian 3-manifold $M$, how can one construct embedded minimal surfaces of a prescribed genus?
In this paper, we reduce this problem to the purely topological problem of constructing suitable “non-trivial” families of surfaces in $M$. Results in this paper are instrumental to our program on constructing minimal surfaces with prescribed genus in 3-spheres – namely, the existence of minimal surfaces of arbitrary genus in Ricci-positive 3-spheres [chu2025arbitraryGenus], and the existence of four genus 2 minimal surfaces in Ricci-positive 3-spheres [chuLiWang2025fourGenus2] [^1] The current paper and [chuLiWang2025fourGenus2] together supersede [chuLiWang2025genus2PartI]..
Here is the setting. In a closed Riemannian $3$-manifold $M$, we let $\mathcal{S}(M)$ be the set of all surfaces smoothly embedded in $M$ that possibly have finitely many singularities. For some technical reasons, we will focus on surfaces that are orientable, finite-area, and separate $M$ into two regions. Note that the notion of genus extends naturally to these singular surfaces, so we may let $\mathcal{S}_{\leq g}(M)\subset\mathcal{S}(M)$ be the set of all singular surfaces of genus $\leq g$; readers may refer to §2 for the precise definitions. Then, our main theorem below essentially states that if there is a family $\Phi$ of singular surfaces with genus $\leq g$ that detects some non-trivial relative structure of the pair $(\mathcal{S}_{\leq g}(M),\mathcal{S}_{\leq g-1}(M))$, then $\Phi$ can be used to produce a minimal surface of genus $g$; see Figure 1 for a schematic illustration.
Theorem 1.2.
Let $M$ be a closed orientable Riemannian $3$-manifold with positive Ricci curvature, and $g$ be a positive integer. Suppose there exists a Simon–Smith family
$\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}_{\leq g}(M)$
that cannot be deformed via pinch-off processes into a map into $\mathcal{S}_{\leq g-1}(M)$. Then $M$ admits an orientable, embedded minimal surface of genus $g$ with area at most $\max_{x\in X}\operatorname{area}(\Phi(x))$.
Figure 1. This schematic diagram shows the set $\mathcal{S}_{\leq g}(M)$ of all surfaces with genus at most $g$, possibly with singularities. The gray cap below represents $\mathcal{S}_{\leq g-1}(M)$, while the blue path represents the image of a map $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}_{\leq g}(M)$, detecting some non-trivial relative structure of the pair $(\mathcal{S}_{\leq g}(M),\mathcal{S}_{\leq g-1}(M))$.
The definitions of “Simon–Smith family” and “deformation via pinch-off processes” are given in Definition 2.7 and Definition 2.20, respectively. Note that we will not equip $\mathcal{S}_{\leq g}(M)$ with a topology. Instead, the notion of the Simon–Smith family includes mild conditions on the map $\Phi$ to ensure the applicability of the Simon–Smith min-max theory. A pinch-off process refers to a deformation process for elements in $\mathcal{S}(M)$ that consists of any combination of the following three procedures: (1) isotopy, (2) neck-pinch surgery, and (3) shrinking connected components to points (see Figure 2). In particular, pinch-off processes do not increase genus.
Figure 2. This is an example of pinch-off process. It has one neck-pinch surgery, and one connected component shrunk to a point.
Remark 1.3.
Regarding the assumption with positive Ricci curvature, a primary purpose is to guarantee the Frankel property, that every embedded minimal surface is connected. Without this, heuristically, it is possible that while the pair $(\mathcal{S}_{\leq g}(M),\mathcal{S}_{\leq g-1}(M))$ has some non-trivial relative structure, the “genus gg” minimal surface produced is however just the union of multiple minimal surfaces of lower genus whose total genus sums to $g$. Furthermore, positive Ricci curvature enables the application of the multiplicity one theorem by Wang–Zhou [wangZhou2022higherMulti], which will play a crucial role in our argument.
Theorem 1.2 is a consequence of the following theorem, which establishes the existence of a family that is “topologically optimal”.
Theorem 1.4 (Existence of a topologically optimal family).
Let $(M,\mathbf{g})$ be a closed orientable Riemannian $3$-manifold with positive Ricci curvature, and $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ be a Simon–Smith family of genus $\leq g$, and the min-max width $L\mathrel{\mathop{\ordinarycolon}}=\mathbf{L}(\Lambda(\Phi))>0$. Assume Condition A for $(M,\mathbf{g})$, $g$ and $L$ (which automatically holds if $\mathbf{g}$ is a generic metric with positive Ricci curvature [^2] This follows easily from White’s bumpy metric theorem [Whi91, white2017bumpy], Choi–Schoen’s compactness theorem [ChoiSchoen1985compactness] and Marques–Neves [MN21, Proposition 8.5]). Then for any $r>0$, there exist:
  • a deformation via pinch-off processes, $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}_{\leq g}(M)$,
  • distinct, orientable, multiplicity one, embedded minimal surfaces $\Gamma_{1},\Gamma_{2},\dots,\Gamma_{l}$ of genus $\leq g$ and area $\leq L$,
  • subcomplexes $D_{0},D_{1},\dots,D_{l}$ of $X$ (after refinement) that cover $X$,
such that:
  1. (1)
    $H(0,\cdot)=\Phi$.
  2. (2)
    The family $\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}=H(1,\cdot)$ satisfies that each member of $\Phi^{\prime}|_{D_{i}}$ has genus at most that of $\Gamma_{i}$, for each i=1,,li=1,...,l, while each member of $\Phi^{\prime}|_{D_{0}}$ has genus 0.
  3. (3)
    For each $i=1,\dots,l$, there is a smooth function $f_{i}\mathrel{\mathop{\ordinarycolon}}D_{i}\to[0,1]$ such that for every $x\in D_{i}$, $H(f_{i}(x),x)$ has the same genus as $\Gamma_{i}$ and is $r$-close to $\Gamma_{i}$ in the $\mathbf{F}$-metric.
Below is the definition of Condition A. Let $\mathcal{O}_{g,\leq L}(M,\mathbf{g})$ (resp. $\mathcal{N}_{g,\leq L}(M,\mathbf{g})$) be the set of all orientable (resp. non-orientable) embedded minimal surfaces in $(M,\mathbf{g})$ with genus $g$ and area $\leq L$. Analogously, we define the sets $\mathcal{O}_{\leq g,\leq L}(M,\mathbf{g})$ and $\mathcal{N}_{\leq g,\leq L}(M,\mathbf{g})$ in which the genus of the minimal surfaces concerned is $\leq g$.
Condition A.
For a closed Riemannian 3-manifold $(M,\mathbf{g})$, a positive integer $g$, and a real number $L>0$, the following hold:
  • The set $\mathcal{O}_{\leq g,\leq L}(M,\mathbf{g})$ is finite.
  • Every element in $\mathcal{O}_{\leq g-1,\leq L}(M,\mathbf{g})$ is non-degenerate.
  • For every $\Gamma\in\mathcal{O}_{g,\leq L}(M,\mathbf{g})$, numbers in the set
    $\{\operatorname{area}(\Gamma^{\prime})\mathrel{\mathop{\ordinarycolon}}\Gamma^{\prime}\in\mathcal{O}_{\leq g-1,\leq L}(M,\mathbf{g})\cup\mathcal{N}_{\leq g+1,\leq L}(M,\mathbf{g})\}\cup\{\operatorname{area}(\Gamma)\}$
    are linearly independent over $\mathbb{Z}$.
The above theorem builds upon numerous important and foundational results in Simon–Smith min-max theory [Smith82, CD03, DP10, Ket19, Zho20, MN21, WangZhou23FourMinimalSpheres]. Theorem 1.4 will be crucial for the enumerative min-max theorem in our work [chuLiWang2025fourGenus2], which constructs multiple minimal surfaces of prescribed genus.
1.1. Main ideas
To derive Theorem 1.2 using Theorem 1.4, suppose by contradiction that no embedded genus $g$ minimal surface has area at most $\max\mathcal{H}^{2}\circ\Phi$. This implies, by Choi-Schoen compactness [ChoiSchoen1985compactness], that for some $\delta_{0}>0$, no orientable embedded genus $g$ minimal surface has area at most $\max\mathcal{H}^{2}\circ\Phi+\delta_{0}$. Thus, using [MN21, Proposition 8.5] and Choi-Schoen compactness [ChoiSchoen1985compactness], it is possible to perturb the metric $\mathbf{g}$ to $\mathbf{g}$$^{\prime}$ such that $(M,{\mathbf{g}}^{\prime})$ satisfies the assumptions in Theorem 1.4 and have no genus $g$ minimal surface.
Hence, applying Theorem 1.4 to $\Phi$ in $(M,\mathbf{g}^{\prime})$, we obtain an optimal family $\Phi^{\prime}$. In particular, all the minimal surfaces $\Gamma_{1},\dots,\Gamma_{l}$ given by Theorem 1.4 must have genus $\leq g-1$. Hence, by item (2) in Theorem 1.4, the family $\Phi^{\prime}$ actually maps into $\mathcal{S}_{\leq g-1}(M)$. So we have deformed $\Phi$ to become a map into $\mathcal{S}_{\leq g-1}(M)$. This contradicts the assumption on $\Phi$, finishing the proof of Theorem 1.2.
As for Theorem 1.4, the optimal family $\Phi^{\prime}$ is constructed using a repeated application of Simon–Smith min-max theory. Starting from $\Psi$ and running min-max theory, we obtain a family that detects some minimal surfaces. We remove the portion of this family that is in the neighborhood of minimal surfaces, and then repeat this process; this must terminate in finitely many steps, and we then assemble the families created to obtain $\Psi^{\prime}$ and the deformation $H$. The strategy of repetitive min-max was first employed by the first two authors in the construction of 5 minimal tori in Ricci-positive 3-spheres [chuLi2024fiveTori]. The key advancement in the current paper is the consideration of arbitrary genus: This poses significant difficulty in the interpolation step (Theorem 2.21) of deforming a family of surfaces, via pinch-off processes, to one of lower genus (this interpolation step is the reason why it is necessary to consider surfaces with singularities: See [chuLi2024fiveTori, Example 2.6]).
Roughly speaking, to decrease the genus of all members in a family of (singular) surfaces, we need to locate in each member certain numbers of short loops, and then pinch all of them. This pinching process needs to be done continuously across the family. However, sometimes it is impossible to choose the family of short loops continuously. In addition, for proving Theorem 1.4, it is necessary to ensure that during the pinching, the area of all members do not increase by much. These make the problems of choosing short loops and pinching very delicate. See the beginning of §5 for the main strategy.
1.2. Organization
We first introduce the necessary terminologies in §2. Then in §3 we prove Theorem 1.2 assuming Theorem 1.4, and in §4 we prove Theorem 1.2 using a repetitive min-max procedure. An essential step in the construct of optimal family is the interpolation from a family of surfaces to one with lower genus using pinch-off processes. In §5, we locate short loops in members of the family for pinching, and finally in §6 we carry out the interpolation.
Acknowledgment
We would like to thank André Neves for his constant support. This material is based upon work supported by the NSF Grant DMS-1928930, while the first author was in residence at the SLMath. The first and the second author were partially supported by the AMS-Simons travel grant.
Part I Main arguments
2. Preliminaries
2.1. Notations
In the sequel, all simplicial complexes and cubical complexes are assumed to be finite; all chains, homology groups, and cohomology groups are assumed to have $\mathbb{Z}_{2}$-coefficients, unless specified otherwise. In general, we use $M$ to denote a closed Riemannian manifold.
  • $\mathbf{I}_{n}(M;\mathbb{Z}_{2})$: the set of integral $n$-dimensional currents in $M$ with $\mathbb{Z}_{2}$-coefficients.
  • $\mathcal{Z}_{n}(M;\mathbb{Z}_{2})\subset\mathbf{I}_{n}(M;\mathbb{Z}_{2})$: the subset that consists of elements $T$ such that $T=\partial Q$ for some $Q\in\mathbf{I}_{n+1}(M;\mathbb{Z}_{2})$ (such $T$ are also called flat $k$-cycles).
  • $\mathcal{Z}_{n}(M;\nu;\mathbb{Z}_{2})$ with $\nu=\mathcal{F},\mathbf{F},\mathbf{M}$: the set $\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$ equipped with the three topologies given corresponding respectively to the flat norm $\mathcal{F}$, the $\mathbf{F}$-metric, and the mass norm $\mathbf{M}$ (see [Pit81] and the survey paper [MN20]). For the flat norm, there are two definitions that, by the isoperimetric inequality, would induce the same topology:
    $\mathcal{F}(T)\mathrel{\mathop{\ordinarycolon}}=\inf\{\mathbf{M}(P)+\mathbf{M}(Q)\mathrel{\mathop{\ordinarycolon}}T=P+\partial Q\},\quad\textrm{ and }\quad\mathcal{F}(T)\mathrel{\mathop{\ordinarycolon}}=\inf\{\mathbf{M}(Q)\mathrel{\mathop{\ordinarycolon}}T=\partial Q\}.$
    In this paper, we will use the second definition.
  • $\mathcal{V}_{n}(M)$: the closure, in the varifold weak topology, of the space of $n$-dimensional rectifiable varifolds in $M$.
  • $\mathcal{C}(M)$: the space of Caccioppoli sets in $M$, equipped with the metric induced by the Hausdorff measure of the symmetric difference.
  • $\partial^{*}\Omega$: the reduced boundary of $\Omega\in\mathcal{C}(M)$.
  • $\nu_{\Omega}$: the inward pointing normal of $\Omega\in\mathcal{C}(M)$.
  • $\mathinner{\!\left\lVert V\right\rVert}$: the Radon measure induced on $M$ by $V\in\mathcal{V}_{n}(M)$.
  • $|T|$: the varifold in $\mathcal{V}_{n}(M)$ induced by a current $T\in\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$, or a countable $n$-rectifiable set $T$. In the same spirit, given a map $\Phi$ into $\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$, the associated map into $\mathcal{V}_{n}(M)$ is denoted by $|\Phi|$.
  • $\operatorname{spt}(\cdot)$: the support of a current or a measure.
  • $[W]$: the $\mathbb{Z}_{2}$-current induced by $W$, if $W$ is a countably $2$-rectifiable set with $\mathcal{H}^{2}(W)<\infty$. In the same spirit, given a map $f$ that outputs countably $2$-rectifiable sets, the associated map into the space $\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$ is denoted by $[f]$.
  • $\mathcal{W}]\mathrel{\mathop{\ordinarycolon}}=\{[\Sigma_{i}]\}\subset\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$ for a set $\mathcal{W}$ of varifolds $\{V_{i}\}\subset\mathcal{V}_{n}(M)$, each associated with a countably $n$-rectifiable set $\Sigma_{i}$.
  • $\mathbf{B}^{\nu}_{\varepsilon}(\cdot)$: the open $\varepsilon$-neighborhood of an element or a subset of the space $\mathcal{Z}_{n}(M;\nu;\mathbb{Z}_{2})$.
  • $\mathbf{B}^{\mathbf{F}}_{\varepsilon}(\cdot)$: the open $\varepsilon$-neighborhood of an element or a subset of $\mathcal{V}_{n}(M)$ under the $\mathbf{F}$-metric.
  • $\Gamma^{\infty}(M)$: the set of smooth Riemannian metrics on $M$.
  • $B_{r}(p)$: the open $r$-neighborhood of a point $p$.
  • $B_{r}(\{p_{1},\dots,p_{n}\})\mathrel{\mathop{\ordinarycolon}}=\bigcup_{i}B_{r}(p_{i})$.
  • $\mathfrak{g}(S)$: the genus of a surface $S$.
  • $\mathrm{in}(S)$: the inside, open region of a surface $S$.
  • $\mathrm{out}(S)$: the outside, open region of a surface $S$.
For an $m$-dimensional cube $I^{m}=\mathbb{R}^{m}\cap\{x\mathrel{\mathop{\ordinarycolon}}0\leq x_{i}\leq 1,i=1,2,\dots,m\}$, we can give it cubical complex structures as follows.
  • $I(1,j)$: the cubical complex on $I\mathrel{\mathop{\ordinarycolon}}=[0,1]$ whose $1$-cells and 0-cells are respectively
[…]
  • $I(m,j)$: the cubical complex structure
    $I(m,j)=I(1,j)\otimes\dots\otimes I(1,j)\;\;(m\textrm{ times})$
    on $I^{m}$. $\alpha=\alpha_{1}\otimes\cdots\otimes\alpha_{m}$ is a $q$-cell of $I(m,j)$ if and only if each $\alpha_{i}$ is a cell in $I(1,j)$ and there are exactly $q$ 11-cells. A cell $\beta$ is a face of a cell $\alpha$ if and only if $\beta\subset\alpha$ as sets.
We call $X\subset I(m,j)$ a cubical subcomplex of $I(m,j)$ if every face of a cell in $X$ is also a cell in $X$. For convenience, we also call $X$ a cubical complex without referring to the ambient cube.
We denote by $|X|$ the underlying space of $X$. For the sake of simplicity, we will also consider a complex and its underlying space as identical unless there is ambiguity.
Given a cubical subcomplex $X$ of some $I(m,j)$, for $j^{\prime}>j$, one can refine $X$ to a cubical subcomplex
$X(j^{\prime})\mathrel{\mathop{\ordinarycolon}}=\{\sigma\in I(m,j^{\prime})\mathrel{\mathop{\ordinarycolon}}\sigma\cap|X|\neq\emptyset\}$
of $I(m,j^{\prime})$. For the sake of convenience, we will denote the refined cubical subcomplex by $X$ unless there is ambiguity.
2.2. Surface with singularities
Let $M$ be a closed orientable 3-manifold, equipped with a Riemannian metric (though all notions in §2.2, except those involving the $\mathbf{F}$-metric, will not depend on the metric of $M$). In this paper, we will consider surfaces with finite area and finitely many singularities, which possibly include some isolated points too:
Definition 2.1 (Punctate surface).
Let $M$ be a closed orientable 3-manifold. A closed subset $S\subset M$ is a punctate surface in $M$ provided that:
  1. (1)
    The 2-dimensional Hausdorff measure $\mathcal{H}^{2}(S)$ is finite.
  2. (2)
    There exists a (possibly empty) finite set $P\subset S$ such that $S\setminus P$ is a smooth, orientable, embedded surface.
  3. (3)
[…]
We denote by $\mathcal{S}(M)$ the set of all punctate surfaces in $M$.
[…]
Definition 2.2 (Punctate set).
Let $S\in\mathcal{S}(M)$. For any finite set $P$ such that $S\setminus P$ is a smooth embedded surface, we call $P$ a punctate set for $S$.
Figure 3. This picture shows an element $S\in\mathcal{S}(M)$: It is a closed set, which contains a (black) smooth surface part and also the red points. The smallest possible punctate set for $S$ is given by the red points. Note $S$ has genus $1$.
Remark 2.3.
As shown in [chuLi2024fiveTori, §2], every element $S\in\mathcal{S}(M)$ can be associated uniquely to an integral 2-current $S]\in\mathbf{I}_{2}(M;\mathbb{Z}_{2})$ with $\partial[S]=0$. In other words, $S]\in\mathcal{Z}_{2}(M;\mathbb{Z}_{2})$.
Hence, we can borrow the $\mathbf{F}$-metric on $\mathcal{Z}_{2}(M;\mathbb{Z}_{2})$ and impose it on $\mathcal{S}(M)$.
Definition 2.4.
We define $\mathbf{F}\mathrel{\mathop{\ordinarycolon}}\mathcal{S}(M)\times\mathcal{S}(M)\to\mathbb{R}$ by
$\mathbf{F}(S_{1},S_{2})\mathrel{\mathop{\ordinarycolon}}=\mathbf{F}([S_{1}],[S_{2}])=\mathcal{F}([S_{1}],[S_{2}])+\mathbf{F}(|S_{1}|,|S_{2}|).$
One can check that $\mathbf{F}$ is a pseudometric on $\mathcal{S}(M)$ (two members of $\mathcal{S}(M)$ differing just by finitely many points would have zero $\mathbf{F}$-distance).
Now, for any $S\in\mathcal{S}(M)$, let $S_{\operatorname{sing}}$ be the set of non-smooth points of $S$. Then by Sard’s theorem, there exists a set $E\subset(0,\infty)$ of full measure such that for every $r\in E$, $S\setminus B_{r}(S_{\operatorname{sing}})$ is a compact smooth surface with boundary, where $B_{r}(P)$ denotes the union of the open $r$-balls centered at the points in $P$. Moreover, by [CM12, Chapter 1 §2.1 Lemma 1.5], for any $r_{1}>r_{2}>0$ in $E$, we have the following genus inequality
$\mathfrak{g}(S\setminus B_{r_{1}}(S_{\operatorname{sing}}))\leq\mathfrak{g}(S\setminus B_{r_{2}}(S_{\operatorname{sing}}))\,.$
Hence, the limit
$\lim_{r\in E,r\to 0}\mathfrak{g}(S\setminus B_{r}(S_{\operatorname{sing}}))\in\mathbb{N}\cup\{\infty\}$
always exists.
Definition 2.5 (Genus of a punctate surface).
Let $S\in\mathcal{S}(M)$, and let $E$ be the set of $r>0$ such that $S\setminus B_{r}(S_{\operatorname{sing}})$ is a smooth surface with boundary. Then we define the genus of $S$ by
$\mathfrak{g}(S)\mathrel{\mathop{\ordinarycolon}}=\lim_{r\in E,r\to 0}\mathfrak{g}(S\setminus B_{r}(S_{\operatorname{sing}}))\,.$
Moreover, the set of $S\in\mathcal{S}(M)$ with genus $g$ is denoted $\mathcal{S}_{g}(M)$, or simply $\mathcal{S}_{g}$; the set of $S\in\mathcal{S}(M)$ with genus $\leq g$ is denoted $\mathcal{S}_{\leq g}(M)$, or simply $\mathcal{S}_{\leq g}$.
Figure 3 is an example of a genus one punctate surface. Moreover, it is easy to check that given any $S\in\mathcal{S}(M)$, for any punctate set $P$, if we let $E$ be the set of $r>0$ such that $S\setminus B_{r}(P)$ is a smooth surface with boundary, then $\lim_{r\in E,r\to 0}\mathfrak{g}(S\setminus B_{r}(P))=\mathfrak{g}(S)$.
We finish this section with a lemma relating genus and the $\mathbf{F}$-pseudometric, which is proved in Appendix B.
Lemma 2.6.
Let $(M,\mathbf{g})$ be a closed Riemannian $3$-manifold, and $\Sigma$ be an orientable smooth embedded surface in $M$. Then there exists $r>0$ such that for any $S\in\mathcal{S}(M)$ with $\mathbf{F}(S,\Sigma)<r$, we have $\mathfrak{g}(S)\geq\mathfrak{g}(\Sigma)$.
2.3. Simon–Smith family
Definition 2.7 (Simon–Smith family).
Let $X$ be a cubical or simplicial complex. A map $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ is called a Simon–Smith family with parameter space $X$ if the following hold.
  1. (1)
    The map $x\mapsto\mathcal{H}^{2}(\Phi(x))$ is continuous.
  2. (2)
    (Closedness) The family $\Phi$ is “closed” in the sense that $\{(x,p)\in X\times M\mathrel{\mathop{\ordinarycolon}}p\in\Phi(x)\}$ is a closed subset of $X\times M$.
  3. (3)
    ($C^{\infty}$-continuity away from singularities) For each $x\in X$, we can choose a finite set $P_{\Phi}(x)\subset\Phi(x)$ such that:
[…]
    • For any $x_{0}\in X$ and within any open set $U\subset\subset M\backslash P_{\Phi}(x_{0})$, $\Phi(x)\to\Phi(x_{0})$ smoothly whenever $x\to x_{0}$.
    • $N(P_{\Phi})\mathrel{\mathop{\ordinarycolon}}=\sup_{x\in X}|P_{\Phi}(x)|<\infty.$
Example 2.8.
The set $P_{\Phi}(x)$ might contain more points than the non-smooth points of $\Phi(x)$. Consider a family $\Phi\mathrel{\mathop{\ordinarycolon}}[0,1]\to\mathcal{S}(M)$, where $\Phi(t)$ is a smooth torus for $t<1$, $\Phi(1)$ is a smooth sphere, and as $t\to 1$, the handle of the torus becomes smaller and eventually vanishes at some smooth point $p$ on the sphere $\Phi(1)$. In this case, $P_{\Phi}(1)$ contains the point $p$.
Remark 2.9.
In previous literature, Simon–Smith families are often allowed to contain certain subsets of measure zero as members. In this paper, to simplify the notations, we choose not to do that. But we note that all results can be modified in obvious ways to allow Simon–Smith families to contain closed subsets of measure zero. See [chuLi2024fiveTori] for such a modification.
Definition 2.10.
A Simon–Smith family $\Phi$ is of genus $\leq g$ if each $\Phi(x)$ has genus $\leq g$.
As shown in [chuLi2024fiveTori, §2], every Simon–Smith family is $\mathbf{F}$-continuous.
Definition 2.11 (Simon–Smith isotopy).
Two Simon–Smith families $\Phi,\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ are called (Simon–Smith) isotopic to each other if there exists a continuous map
$\varphi\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\operatorname{Diff}^{\infty}(M)$
such that:
  1. (1)
    $\varphi(0,x)=\operatorname{Id}$ for all $x\in X$,
  2. (2)
    $\varphi(1,x)(\Phi(x))=\Phi^{\prime}(x)$ for all $x\in X$,
where $\textrm{Diff}^{\infty}(M)$ denotes the group of diffeomorphisms on $M$ equipped with the $C^{\infty}$ topology. Moreover, the set of all families isotopic to a Simon–Smith family $\Phi$ is called the (Simon–Smith) isotopy class associated to $\Phi$, denoted by $\Lambda(\Phi)$.
Remark 2.12.
  • In Definition 2.7, for any choice of sets $P_{\Phi}(x)$ for $\Phi$, and for any isotopy $\varphi$, the sets $P_{\Phi^{\prime}}(x)\mathrel{\mathop{\ordinarycolon}}=\varphi(1,x)\circ P_{\Phi}(x)$ would satisfy the conditions for the family $\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}=\varphi(1,\cdot)\circ\Phi\in\Lambda(\Phi)$. Consequently, we may assume that $\sup_{x}|P_{\Phi}(x)|<\infty$ is a constant within the isotopy class $\Lambda(\Phi)$.
  • If $\Phi$ is of genus $\leq g$, then so is every $\Phi^{\prime}\in\Lambda(\Phi)$.
2.4. Min-max theorems
Definition 2.13 (Simon–Smith width).
Given an isotopy class $\Lambda$ of Simon–Smith families, its width is defined as
$\mathbf{L}(\Lambda)\mathrel{\mathop{\ordinarycolon}}=\inf_{\Phi\in\Lambda}\sup_{x\in X}\mathcal{H}^{2}(\Phi(x))\,.$
Definition 2.14 (Minimizing sequence and min-max sequence).
A sequence $\{\Phi_{i}\}$ in $\Lambda$ is called minimizing if
$\lim_{i\to\infty}\sup_{x\in X}\mathcal{H}^{2}(\Phi_{i}(x))=\mathbf{L}(\Lambda)\,.$
If $\{\Phi_{i}\}$ is a minimizing sequence in $\Lambda$ and $\{x_{i}\}\subset X$ is such that $\mathcal{H}^{2}(\Phi_{i}(x_{i}))\to\mathbf{L}(\Lambda)\,$, then $\{\Phi_{i}(x_{i})\}$ is called a min-max sequence. For a minimizing sequence $\{\Phi_{i}\}$, we define its critical set $\mathbf{C}(\{\Phi_{i}\})$ to be the set of all subsequential varifold-limit of its min-max sequences:
$\mathbf{C}(\{\Phi_{i}\})\mathrel{\mathop{\ordinarycolon}}=\{V=\lim_{j}|\Phi_{i_{j}}(x_{j})|\mathrel{\mathop{\ordinarycolon}}x_{j}\in X,\|V\|(M)=\mathbf{L}(\Lambda)\}\,.$
And $\{\Phi_{i}\}$ is called pulled-tight if every varifold in $\mathbf{C}(\{\Phi_{i}\})$ is stationary.
Definition 2.15.
Let $(M,{\mathbf{g}})$ be a closed orientable Riemannian $3$-manifold and $L>0$.
  1. (1)
    Denote by $\mathcal{W}_{L}(M,{\mathbf{g}})$ the set of all varifolds $W\in\mathcal{V}_{2}(M)$, with $\|W\|(M)=L$, of the form
    (2.1) $W=m_{1}|\Gamma_{1}|+\cdots+m_{l}|\Gamma_{l}|\,$,
    where $m_{1},\dots,m_{l}$ are positive integers, and $\Gamma_{1},\dots,\Gamma_{l}$ are disjoint smooth, connected, embedded minimal surfaces in $(M,{\mathbf{g}})$, with $m_{i}$ even if $\Gamma_{i}$ is non-orientable (see [Ket19, Remark 1.4]).
  2. (2)
    Denote by $\mathcal{W}_{L,\leq g}(M,{\mathbf{g}})$ the set of all varifolds $W\in\mathcal{W}_{L}(M,{\mathbf{g}})$ (of the form (2.1)) satisfying
    (2.2) $\sum_{j\in I_{O}}m_{j}\mathfrak{g}(\Gamma_{j})+\frac{1}{2}\sum_{j\in I_{N}}m_{j}(\mathfrak{g}(\Gamma_{j})-1)\leq g\,$,
    where $\Gamma_{j}$ is orientable if $j\in I_{O}$ and non-orientable if $j\in I_{N}$.
For simplicity, when the ambient manifold $(M,{\mathbf{g}})$ is clear, we may omit it in the notations, and use the notations $\mathcal{W}_{L}$ and $\mathcal{W}_{L,\leq g}$, respectively.
Theorem 2.16 (Simon–Smith min-max theorem).
Let $(M,\mathbf{g})$ be a closed orientable Riemannian $3$-manifold. Suppose a Simon–Smith family $\Phi$ of genus $\leq g$ has $L\mathrel{\mathop{\ordinarycolon}}=\mathbf{L}(\Lambda(\Phi))>0$. Then for any $r>0$, there exists a minimizing sequence $\{\Phi_{i}\}$ in $\Lambda(\Phi)$ such that:
  1. (1)
    The sequence $\{\Phi_{i}\}$ is pulled-tight and $\mathbf{C}(\{\Phi_{i}\})\cap\mathcal{W}_{L,\leq g}\neq\emptyset\,.$
  2. (2)
    There exists some
    $W=m_{1}|\Gamma_{1}|+\cdots+m_{l}|\Gamma_{l}|\in\mathbf{C}(\{\Phi_{i}\})\cap\mathcal{W}_{L,\leq g},$
    under the notation in Definition 2.15 (1), such that:
    1. (a)
      If $\Gamma_{j}$ is unstable and two-sided, then $m_{j}=1$.
    2. (b)
      If $\Gamma_{j}$ is one-sided, then its connected double cover is stable and $m_{j}$ is even.
  3. (3)
    Furthermore, there exists $\eta>0$ such that for all sufficiently large $i$,
    $\mathcal{H}^{2}(\Phi_{i}(x))\geq L-\eta\implies|\Phi_{i}(x)|\in\mathbf{B}^{\mathbf{F}}_{r}(\mathcal{W}_{L,\leq g})\,.$
The above theorem builds upon the foundational results  [Smith82, CD03, DP10, Ket19, MN21, WangZhou23FourMinimalSpheres], and was proven in [chuLi2024fiveTori]: See Theorem 2.16 therein (we can guarantee that $m_{i}$ is even if $\Gamma_{i}$ is one-sided [Ket19, Remark 1.4]). Note that since $M$ is orientable, a surface is two-sided if and only if it is orientable.
The following theorem, proven in [chuLi2024fiveTori, §3], is also useful. Roughly speaking, if all min-max minimal surfaces are of multiplicity one, then there exists some family where in the “large-area region” we can guarantee closeness to the minimal surfaces in the $\mathbf{F}$-norm for currents.
Theorem 2.17 ($\mathbf{F}$-closeness).
Let $(M,{\mathbf{g}})$ be a closed orientable Riemannian $3$-manifold, and $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ be a Simon–Smith family of genus $\leq g$ such that $L\mathrel{\mathop{\ordinarycolon}}=\mathbf{L}(\Lambda(\Phi))>0$. Suppose that $\mathcal{W}_{L,\leq g}$ consists of only finitely many varifolds, each induced by a multiplicity one embedded minimal surface. Then for any $r>0$, there exists $\eta>0$ and $\Phi^{\prime}\in\Lambda(\Phi)$ such that
$\mathcal{H}^{2}(\Phi^{\prime}(x))\geq L-\eta\implies[\Phi^{\prime}(x)]\in\mathbf{B}^{\mathbf{F}}_{r}([\mathcal{W}_{L,\leq g}])\,.$
2.5. Pinch-off process and interpolation
In this section, we state an important interpolation result, Theorem 2.21, an improvement of Theorem 2.17. Namely, given a family $\Phi$ that produces a multiplicity one, non-degenerate min-max minimal surface $\Gamma$, we can construct a “better” family $\Phi^{\prime}\in\Lambda(\Phi)$ such that on the large-area region, each element has the same genus as $\Gamma$. This imposes a topological conditions on the large-area region, compared to Theorem 2.17. This is achieved by performing pinch-off processes, continuously across the whole region of punctate surfaces of large area in $\Phi^{\prime\prime}$.
Definition 2.18.
In a closed 3-manifold $M$, a collection $\{\Gamma(t)\}_{t\in[a,b]}$ of elements in $\mathcal{S}(M)$ is called a pinch-off process if the following holds. Regarding the parameter space $[a,b]$ as a time interval, there exist finitely many spacetime points $(t_{1},p_{1}),\dots,(t_{n},p_{n})$ in $(a,b]\times M$, with each $p_{i}\in\Gamma(t_{i})$, such that each $(t,p)\in[a,b]\times M$ is of one of the following types.
If $(t,p)$ is not equal to any $(t_{i},p_{i})$, then:
  1. (1)
    There exists a closed neighborhood $J\subset[a,b]$ around $t$ and a ball $U\subset M$ around $p$ such that $\{\Gamma(t)\}_{t\in J}$ deforms by an isotopy within $U$.
If $(t,p)$ is equal to some $(t_{i},p_{i})$, then there exists a closed neighborhood $J\subset[a,b]$ around $t$ and a ball $U\subset M$ around $p$ such that the family $\{\Gamma(t^{\prime})\cap U\}_{t^{\prime}\in J}$ of surfaces is of one of the following types (Note it is possible that $t=b$.):
  1. (2)
    Surgery : The initial surface deforms by an isotopy for $t^{\prime}<t$, but then at time $t$ becomes, up to diffeomorphism, a double cone with $p$ as the cone point (as shown in the second picture in Figure 2). Afterwards, it either remains a double cone or splits into two smooth discs: In either case it is allowed to deform by another isotopy.
  2. (3)
    Shrinking into points: For each $t^{\prime}\in J$, $\Gamma(t^{\prime})$ does not intersect $\partial U$. Moreover, for the family $\{\Gamma(t^{\prime})\cap U\}_{t^{\prime}\in J}$, the initial surface deforms by isotopy in $U$ when $t^{\prime}<t$, but then becomes just the point $p$ at time $t$, and this point moves smoothly after time $t.$
Note, all pinch-off processes are Simon–Smith families. In addition, in (3), since $\Gamma(t^{\prime})$ does not intersect $\partial U$, the part we are shrinking must be a union of several connected components. The reason for allowing the double cone to remain after time $t$ in case (2) will be addressed in Remark 4.3. Now, the lemma below follows directly from the definition of the pinch-off process.
Lemma 2.19.
Along a pinch-off process, genus is non-increasing.
Definition 2.20.
Let $\Phi,\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ be Simon–Smith families. Suppose we have a Simon–Smith family $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}(M)$ such that:
  • $H(0,\cdot)=\Phi$ and $H(1,\cdot)=\Phi^{\prime}$.
  • For each $x\in X$, $t\mapsto H(t,x)$ is a pinch-off process.
Then $H$ is called a deformation via pinch-off processes from $\Phi$ to $\Phi^{\prime}$.
We now state a crucial interpolation theorem.
Theorem 2.21 (Interpolation).
In a closed orientable Riemannian $3$-manifold $(M,{\mathbf{g}})$, let $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ be a Simon–Smith family of genus $\leq g$ with $X$ a cubical complex. Assume that $L\mathrel{\mathop{\ordinarycolon}}=\mathbf{L}(\Lambda(\Phi))>0$, and $\mathcal{W}_{L,\leq g}=\{|\Gamma|\}$ where $\Gamma$ is a nondegenerate embedded closed minimal surface. Then for any $r>0$, there exists a Simon–Smith family $\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}_{\leq g}$ with the following properties:
  1. (1)
    There exists an $\eta>0$ such that for all $x\in X$ with $\mathcal{H}^{2}(\Phi^{\prime}(x))\geq L-\eta$, we have $\Phi^{\prime}(x)\in\mathcal{S}_{\mathfrak{g}(\Gamma)}\cap\mathbf{B}^{\mathbf{F}}_{r}(\Gamma).$
  2. (2)
    $\Phi$ can be deformed via pinch-off processes to $\Phi^{\prime}$.
We postpone the proof of Theorem 2.21 to Section 6, which details the pinch-off constructions, and Appendix D. As a special case, when $L=0$, the whole family can be deformed via pinch-off processes to surfaces of genus 0:
Proposition 2.22.
In a closed orientable Riemannian $3$-manifold $(M,{\mathbf{g}})$, suppose that $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ is a Simon–Smith family of genus $\leq g$ with $X$ a cubical complex and $\mathbf{L}(\Lambda(\Phi))=0$. Then $\Phi$ can be deformed via pinch-off processes into some Simon–Smith family $\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}_{0}$ of genus 0.
The above proposition is a direct corollary of Proposition 6.2; see §6.4.
Finally, we finish this section with a proposition useful in proving the enumerative min-max theorem in [chuLiWang2025fourGenus2], which is also proved in §6
Proposition 2.23.
Given a closed orientable Riemannian $3$-manifold $(M,{\mathbf{g}})$, and integers $0\leq g_{0}<g_{1}$, suppose that $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ is a Simon–Smith family of genus $\leq g_{1}$ with $X$ a cubical complex. Let $K$ be an arbitrary compact subset in $X$, and assume $\Phi|_{K}$ maps into $\mathcal{S}_{\leq g_{0}}$. Then there exists a subcomplex $Z\subset X$ (after refinement) whose interior contains $K$ such that $\Phi|_{Z}$ can be deformed via pinch-off processes to become some Simon–Smith family of genus $\leq g_{0}$.
3. Proof of Theorem 1.2 via Theorem 1.4
Assuming Theorem 1.4, let us prove Theorem 1.2 by contradiction: Suppose that no embedded genus $g$ minimal surface has area at most $\max\mathcal{H}^{2}\circ\Phi$. This implies, by Choi-Schoen compactness [ChoiSchoen1985compactness], that for some $\delta_{0}>0$, no orientable embedded genus $g$ minimal surface has area at most $\max\mathcal{H}^{2}\circ\Phi+\delta_{0}$. Thus, using [MN21, Proposition 8.5] and Choi-Schoen compactness [ChoiSchoen1985compactness], we can perturb the metric $\mathbf{g}$ to $\mathbf{g}$$^{\prime}$ such that $(M,{\mathbf{g}}^{\prime})$ has the following properties:
  1. (1)
    It has positive Ricci curvature.
  2. (2)
    All embedded minimal surfaces are non-degenerate.
  3. (3)
    The areas of the embedded minimal surfaces are linearly independent over $\mathbb{Z}$.
  4. (4)
    There does not exist an orientable embedded minimal surface of genus $g$ and has area at most $\max\mathcal{H}^{2}_{\mathbf{g}^{\prime}}\circ\Phi+\delta_{0}/2$.
It is easy to see that now $(M,\mathbf{g}^{\prime})$ satisfy the assumptions in Theorem 1.4. We claim that applying Theorem 1.4 (to be proven in §4) to $(M,\mathbf{g}^{\prime})$, $g$ and $\Phi$, with $r=1$, leads to a contradiction and thus prove Theorem 1.2. Indeed, using the defining properties (1) - (3) of $(M,\mathbf{g}^{\prime})$, the assumptions of Theorem 1.4 are satisfied. So we can apply Theorem 1.4 and obtain a Simon–Smith family $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}_{\leq g}$ that satisfies the properties in Theorem 1.4. However, recall that by property (4) of $(M,\mathbf{g}^{\prime})$, there does not exist an orientable embedded minimal surface of genus $g$ that has area at most $\max\mathcal{H}^{2}_{\mathbf{g}^{\prime}}\circ\Phi$. As a result, all those minimal surfaces $\Gamma_{1},\dots,\Gamma_{l}$ produced in Theorem 1.4 must have genus $\leq g-1$. By item (3) of the theorem, $\Phi^{\prime}|_{D_{i}}$ maps into $\mathcal{S}_{\leq g-1}$ for each $i=1,\dots,l$. Moreover, $\Phi^{\prime}|_{D_{0}}$ has genus 0, so we know actually the whole family $\Phi^{\prime}$ maps into $\mathcal{S}_{\leq g-1}$. This contradicts the assumption of Theorem 1.2 that $\Phi$ cannot be deformed via pinch-off processes to become a map into $\mathcal{S}_{\leq g-1}$. Therefore, we have proven Theorem 1.2.
4. Proof of Theorem 1.4: Repetitive min-max
The goal of this section is to prove Theorem 1.4.
First, we will repetitively run min-max using the given family $\Phi$, and obtain a number of families. Then, we assemble these families to obtain the desired “optimal family” $\Phi^{\prime}$, whose domain is a union of parts $D_{0},D_{1},\dots,D_{l}$. The min-max construction together with our interpolation theorem, Theorem 2.21, leads to the conclusions of Theorem 1.4.
4.1. Repetitive min-max: $k$-th stage
In this section, we will repetitively run min-max. But before that, let us fix some constants.
  • First, we choose a $r^{\prime}>0$ such that $r^{\prime}<r$, where $r$ is given in Theorem 1.4.
  • Let $d_{0}$ be the minimum value of $\mathbf{F}(\Sigma,\Sigma^{\prime})$ where $\Sigma,\Sigma^{\prime}$ are any two distinct orientable embedded minimal surfaces of genus $\leq g$ in $(M,\mathbf{g})$: Recall there are only finitely many orientable embedded minimal surfaces of genus $\leq g$ by assumption. By choosing a smaller value for $r^{\prime}$, let us assume $r^{\prime}<d_{0}$.
  • By further shrinking $r^{\prime}$, by Lemma 2.6, we can assume that for each of the finitely many orientable embedded minimal surfaces $\Sigma$ of genus $\leq g$, if $S\in\mathcal{S}(M)$ is such that $\mathbf{F}(S,\Sigma)<r^{\prime}$, then $\mathfrak{g}(S)\geq\mathfrak{g}(\Sigma)$.
Remark 4.1.
Throughout §4, a superscript k or subscript k would indicate that the object concerned is constructed from running the min-max process for the $k$-th time.
We will recursively describe the process. Let $k=1,2,\dots$. Right before the $k$-th stage of min-max, we would have on our hands a subcomplex $X^{k-1}$ of $X$ (up to refinement), and a Simon–Smith family $\Phi^{k-1}\mathrel{\mathop{\ordinarycolon}}X^{k-1}\to\mathcal{S}_{\leq g}.$ (Before the first time we run min-max, i.e., $k=1$, we of course have $X^{0}\mathrel{\mathop{\ordinarycolon}}=X$ and $\Phi^{0}\mathrel{\mathop{\ordinarycolon}}=\Phi$.)
For each $k\in\mathbb{N}^{+}$, we apply the min-max theorem, Theorem 2.16, to $\Phi^{k-1}$ and $r^{\prime}>0$. If the width $\mathbf{L}(\Lambda(\Phi^{k-1}))$ is positive, then we obtain a varifold $V^{k}$ in $\mathcal{W}^{k}\mathrel{\mathop{\ordinarycolon}}=\mathcal{W}_{\mathbf{L}(\Lambda(\Phi^{k-1})),\leq g}$ (defined in Definition 2.15) satisfying Theorem 2.16 (2); Otherwise, this $k$-th stage of min-max would actually be the last stage, which will be discussed in the subsequent section. So let us assume the width is positive at this stage.
Since $(M,\mathbf{g})$ has positive Ricci curvature, the Frankel property holds. Consequently, every varifold in $\mathcal{W}^{k}$ is associated with a connected surface. Also, we know that $\operatorname{spt}(V^{k})$ must be orientable, for otherwise it would be one-sided (as $M$ is orientable) and its double cover would be stable (by Theorem 2.16 (2) (b)), contradicting the fact that $M$ has positive Ricci curvature. As a result, by Theorem 2.16 (2), $V^{k}$ is of multiplicity one. Furthermore, we know the following:
Lemma 4.2.
Fix an ambient $3$-manifold $(M,\mathbf{g})$ that satisfies the assumption in Theorem 1.4. Suppose that when we apply the min-max theorem (Theorem 2.16) to some $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}_{\leq g}$, the width $L=\mathbf{L}(\Lambda(\Phi))$ is positive and the set $\mathcal{W}_{L,\leq g}$ contains some element $V$ of multiplicity one. Then we have the following.
  1. (1)
    If $\operatorname{spt}V$ has genus $g$, then every varifold in $\mathcal{W}_{L,\leq g}$ is associated with a multiplicity one, orientable, minimal surface of genus $g$.
  2. (2)
    If $\operatorname{spt}V$ has genus <g<g, then $\mathcal{W}_{L,\leq g}=\{V\}$.
In particular, $\mathcal{W}_{L,\leq g}$ consists of finitely many varifolds, each associated with a multiplicity one orientable minimal surface.
Proof.
Notice:
  • By the assumptions in Theorem 1.4, given any orientable minimal surface $\Gamma$ of genus $g$, numbers in the set $\{\operatorname{area}(\Gamma^{\prime})\mathrel{\mathop{\ordinarycolon}}\Gamma^{\prime}\in\mathcal{O}_{\leq g-1,\leq L}(M,\mathbf{g})\cup\mathcal{N}_{\leq g+1,\leq L}(M,\mathbf{g})\}\cup\{\operatorname{area}(\Gamma)\}$ are linearly independent over $\mathbb{Z}$.
  • On the other hand, we know that every element of $\mathcal{W}_{L,\leq g}$ must satisfy (2.1), and thus must either be: (A) a multiplicity one orientable genus $g$ surface, or (B) an element in $\mathcal{O}_{\leq g-1,\leq L}(M,\mathbf{g})\cup\mathcal{N}_{\leq g+1,\leq L}(M,\mathbf{g})$ possibly with multiplicities.
Now, to prove item (1), we assume $\operatorname{spt}V$ has genus $g$ and is orientable. Then, combining the two bullet points above, we know Case (B) is impossible. So we have proven item (1). As for item (2), if $\operatorname{spt}V$ has genus <g<g (so $\operatorname{spt}V\in\mathcal{O}_{\leq g-1,\leq L}(M,\mathbf{g})\cup\mathcal{N}_{\leq g+1,\leq L}(M,\mathbf{g})$), we can similarly see from the above two bullet points that $\mathcal{W}_{L,\leq g}$ can only consist of the element $V$.
The finiteness of $\mathcal{W}$ follows from the assumptions in Theorem 1.4 that there are only finitely many orientable embedded minimal surfaces of genus $\leq g$. ∎
Using this lemma, there are two cases regarding $V^{k}\in\mathcal{W}^{k}\mathrel{\mathop{\ordinarycolon}}=\mathcal{W}_{\mathbf{L}(\Lambda(\Phi^{k-1})),\leq g}$:
  • Case 1.
    $\mathcal{W}^{k}$ consists of finitely many elements, each associated with a multiplicity one, orientable, embedded minimal surface of genus $g$; Denote them as $T^{k}_{1},\dots,T^{k}_{n_{k}}\in\mathcal{S}_{g}$;
  • Case 2.
    $\mathcal{W}^{k}=\{V^{k}\}$, and $V^{k}$ is associated with a multiplicity one, orientable, embedded minimal surface of genus $\leq g-1$, which must be non-degenerate as assumed in Theorem 1.4. Let $T^{k}\in\mathcal{S}_{\leq g-1}$ be the associated element,
Let us analyze them separately.
4.1.1. Case 1
Applying Theorem 2.17 to $\Phi^{k-1}\mathrel{\mathop{\ordinarycolon}}X^{k-1}\to\mathcal{S}(M)$ and $r^{\prime}>0$, we can obtain some $\widetilde{\Phi}^{k}\in\Lambda(\Phi^{k-1})$ and $\delta_{k}>0$ such that
$\mathcal{H}^{2}(\widetilde{\Phi}^{k}(x))\geq\mathbf{L}(\Lambda(\Phi^{k-1}))-\delta_{k}\implies\widetilde{\Phi}^{k}(x)\in\mathbf{B}^{\mathbf{F}}_{r^{\prime}}(\mathcal{W}^{k}).$
Hence, we can choose some subcomplex $\hat{E}^{k}$ of $X^{k-1}$ (after refining $X^{k-1}$), which we call a “cap”, such that:
  • $\mathcal{H}^{2}(\widetilde{\Phi}^{k}(x))\geq\mathbf{L}(\Lambda(\Phi^{k-1}))-\delta_{k}/2\implies x\in\hat{E}^{k},$
  • $x\in\hat{E}^{k}\implies\widetilde{\Phi}^{k}(x)\in\mathbf{B}^{\mathbf{F}}_{r^{\prime}}(\mathcal{W}^{k}).$
In addition, for future purpose, let us choose a slightly larger subcomplex $E^{k}\supset\hat{E}^{k}$ in $X^{k-1}$ (after refinement) such that $E^{k}$ also satisfies these two properties, and that there exists a smooth cut-off function $\eta_{k}\mathrel{\mathop{\ordinarycolon}}X\to[0,1]$ such that $\eta_{k}=1$ on $\overline{X\backslash E^{k}}$ while $\eta_{k}=0$ on $\hat{E}^{k}$.
Since $r^{\prime}<d_{0}$, $E^{k}$ is a disjoint union of subcomplexes $E^{k}_{1},\dots,E^{k}_{n_{k}}$ such that, for each $j=1,\dots,n_{k}$, there exists $T^{k}_{j}\in\mathcal{W}^{k}$ satisfying $\widetilde{\Phi}^{k}(E^{k}_{j})\subset\mathbf{B}^{\mathbf{F}}_{r^{\prime}}(T^{k}_{j})\subset\mathbf{B}^{\mathbf{F}}_{r}(T^{k}_{j})\,.$ In fact, since $r^{\prime}$ is less than the constant given by Lemma 2.6 (applied to any $T^{k}_{j}$), we know that every member of $\widetilde{\Phi}^{k}|_{E^{k}}$ is of genus $g$. Hence, we have $\widetilde{\Phi}^{k}(E^{k}_{j})\subset\mathcal{S}_{g}\cap\mathbf{B}^{\mathbf{F}}_{r}(T^{k}_{j}).$ In addition, recalling by definition $\widetilde{\Phi}^{k}$ is in the same Simon–Smith isotopy class as $\Phi^{k-1}$, we let $H^{k}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X^{k-1}\to\mathcal{S}_{\leq g}$ be such that $H^{k}(0,\cdot)=\Phi^{k-1}$, $H^{k}(1,\cdot)=\widetilde{\Phi}^{k}$, and $t\mapsto H^{k}(t,y)$ is a smooth deformation for every $y$.
4.1.2. Case 2
In this case, we let $T^{k}\in\mathcal{S}_{\leq g-1}$ be the element associated to the varifold $V^{k}$, and apply Theorem 2.21 to $\Phi^{k-1}\mathrel{\mathop{\ordinarycolon}}X^{k-1}\to\mathcal{S}(M)$ and $r$. We obtain:
  • a Simon–Smith family of genus $\leq g$, $\widetilde{\Phi}^{k}\mathrel{\mathop{\ordinarycolon}}X^{k-1}\to\mathcal{S}_{\leq g}\,$,
  • a Simon–Smith family of genus $\leq g$, $H^{k}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X^{k-1}\to\mathcal{S}_{\leq g},$
  • a constant $\delta_{k}>0$,
satisfying the following:
  • $H^{k}(0,\cdot)=\Phi^{k-1}$, $H^{1}(1,\cdot)=\widetilde{\Phi}^{k}$,
  • If $\mathcal{H}^{2}(\widetilde{\Phi}^{k}(x))\geq\mathbf{L}(\Lambda(\Phi^{k-1}))-\delta_{k}$, then $\widetilde{\Phi}^{k}(x)\in\mathcal{S}_{\mathfrak{g}(T^{k})}\cap\mathbf{B}^{\mathbf{F}}_{r}(T^{k})$.
  • $t\mapsto H^{k}(t,x)$ is a pinch-off process for each $x\in X^{k-1}$.
Clearly, we can choose some subcomplex $\hat{E}^{k}\subset X^{k-1}$, which we call a “cap”, such that
  • $\mathcal{H}^{2}(\widetilde{\Phi}^{k}(x))\geq\mathbf{L}(\Lambda(\Phi^{k-1}))-\delta_{k}/2\implies x\in\hat{E}^{k},$
  • $x\in\hat{E}^{k}\implies\widetilde{\Phi}^{k}(x)\in\mathcal{S}_{\mathfrak{g}(T^{k})}\cap\mathbf{B}^{\mathbf{F}}_{r}(T^{k}).$
As in Case 1, for future use, we choose a slightly larger subcomplex $E^{k}\supset\hat{E}^{k}$ in $X^{k-1}$ (after refinement) such that $E^{k}$ also satisfies these two properties, and that there exists a smooth cut-off function $\eta_{k}\mathrel{\mathop{\ordinarycolon}}X\to[0,1]$ such that $\eta_{k}=1$ on $\overline{X\setminus E^{k}}$ while $\eta_{k}=0$ on $\hat{E}^{k}$.
4.1.3. Family for running the $(k+1)$-th min-max
[…]
4.2. The last stage
We now have that the width $\mathbf{L}(\Lambda(\Phi^{K-1}))$ is zero. Thus, using Theorem 2.22, we have:
  • a Simon–Smith family of genus 0, $\Phi^{K}\mathrel{\mathop{\ordinarycolon}}X^{K-1}\to\mathcal{S}_{0}$,
  • a Simon-Simon family of genus $\leq g$, $H^{K}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X^{K-1}\to\mathcal{S}_{\leq g}$,
such that:
  • $H^{K}(0,\cdot)=\Phi^{K-1}$ and $H^{K}(1,\cdot)=\Phi^{K}$.
  • For all $x\in X^{K-1}$, $t\mapsto H^{K}(t,x)$ is a pinch-off process.
4.3. An optimal family $\Phi^{\prime}$
In this subsection, we will use the families defined above to reconstruct a new Simon–Smith family $\Phi^{\prime}$ of genus $\leq g$, with parameter space $X$, and a deformation via pinch-off processes, $H$, from $\Phi$ to $\Phi^{\prime}$. Readers may refer to Figure 4 for a schematic of the families constructed.
Figure 4. In the first picture, the red region is $X^{K-1}$, the blue is $E^{K-1}$, and the green is $E^{K-2}$. In the second picture, the gray surface (with suitable smoothening) denotes $\Xi^{K-1}$. In the last picture, the gray surface denotes $\Xi^{K-2}$.
[…]
    $H^{K}(\eta_{K-1}(x),x),\quad x\in X^{K-1}.$
    Note, for $x$ at the “boundary” $X^{K-1}\cap\hat{E}^{K-1}$, the expression above becomes $H^{K}(0,x)=\Phi^{K-1}(x)$.
  • As a result, this family can be glued with the family $\widetilde{\Phi}^{K-1}|_{\hat{E}^{K-1}}$ along $X^{K-1}\cap\hat{E}^{K-1}$. Then, the resulting family, denoted $\Xi^{K-1}$, has domain $X^{K-2}$. Now, it is easy to see that there exists a deformation via pinch-off processes from $\widetilde{\Phi}^{K-1}$ to $\Xi^{K-1}$. Recall we have also the deformation via pinch-off processes $H^{K-1}$ from $\Phi^{K-2}$ to $\widetilde{\Phi}^{K-1}$. Thus, let us concatenate these two processes, to obtain a deformation via pinch-off processes
    $G^{K-1}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X^{K-2}\to\mathcal{S}_{\leq g}$
    such that $G^{K-1}(0,\cdot)=\Phi^{K-2}$ and $G^{K-1}(1,\cdot)=\Xi^{K-1}$.
[…]
    $G^{K-1}(\eta_{K-2}(x),x),\quad x\in X^{K-2}.$
    Note, for $x$ at the “boundary” $X^{K-2}\cap\hat{E}^{K-2}$, the expression above becomes $G^{K-1}(0,x)=\Phi^{K-2}(x)$.
  • As a result, this family can be glued with the family $\widetilde{\Phi}^{K-2}|_{\hat{E}^{K-2}}$ along $X^{K-2}\cap\hat{E}^{K-2}$. The resulting family, denoted $\Xi^{K-2}$ has domain $X^{K-3}$. Now, as before, we know there exists a deformation via pinch-off processes
    $G^{K-2}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X^{K-3}\to\mathcal{S}_{\leq g}$
    such that $G^{K-2}(0,\cdot)=\Phi^{K-3}$ and $G^{K-1}=\Xi^{K-2}$.
We repeat this process. At the final step, we would be gluing some family of the form $G^{2}(\eta_{1}(x),x)$, $x\in X^{1}$ to the family $\widetilde{\Phi}^{1}|_{\hat{E}^{1}}$, and obtain a family $\Xi^{1}$ with domain $X^{1}\cup\hat{E}^{1}=X$. Now, we define the desired “optimal family” $\Phi^{\prime}$ to be $\Xi^{1}$. By construction, it is easy to see that there exists some deformation via pinch-off processes $G^{1}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}_{\leq g}$ such that $G^{1}(0,\cdot)=\Phi$ and $G^{1}(1,\cdot)=\Phi^{\prime}$.
In the following, we will denote $G^{1}$ by $H$ instead, to align with the notation in Theorem 1.4.
4.4. Completion of the Proof
For simplicity, we now collect all the minimal surfaces $T^{k}$ and $T^{k}_{j}$ together, and relabel them $\Gamma_{1},\Gamma_{2},\dots,\Gamma_{l},$ but in an order that reverses the order they were obtained in the min-max process (i.e. if $i\leq j$, then $\Gamma_{i}$ was obtained in a later stage of min-max than $\Gamma_{j}$). Note they all must have area $\leq\mathbf{L}(\Lambda(\Phi))$. At the same time, we relabel their corresponding associated caps $E^{k}$ and $E^{k}_{j}$ as $\widetilde{D}_{1},\widetilde{D}_{2},\dots,\widetilde{D}_{l}.$ We also set $\widetilde{D}_{0}\mathrel{\mathop{\ordinarycolon}}=X^{K-1}\,.$ Finally, for each $i=0,1,2,\cdots,l$, we define
$D_{i}\mathrel{\mathop{\ordinarycolon}}=\overline{\widetilde{D}_{i}\setminus\bigcup^{l}_{j=i+1}\widetilde{D}_{j}}\,.$
In other words, if $\widetilde{D}_{i}$ was obtained at the $k$-th stage of min-max, then we remove from it all the “bridge parts” coming from the $k$-th stage or earlier (see Figure 4). Moreover, if some $D_{i}$ is empty, we remove $D_{i}$ from the collection $\{D_{j}\}$ and the corresponding $\Gamma_{i}$ from the set $\{\Gamma_{j}\}$, and then relabel the remaining sets accordingly.
Let us finish the proof of Theorem 1.4. Item (1) is already explained at the end of the previous section. We fix an $i=1,\dots,l$, and suppose the cap $\widetilde{D}_{i}$ was defined in the $k$-th stage of min-max. Then, by how we defined the caps $E^{k}$ and $E^{k}_{j}$, we have $\widetilde{\Phi}^{k}|_{\widetilde{D}_{i}}\in\mathcal{S}_{\mathfrak{g}(\Gamma_{i})}\cap\mathbf{B}^{\mathbf{F}}_{r}(\Gamma_{i})$, where $\Gamma_{i}$ is the minimal surface associated with the cap $\widetilde{D}_{i}$. Note that $D_{i}$ is constructed by removing the “bridge parts” above $\widetilde{D}_{i}$, and thus, the family $\Phi|_{D_{i}}$ can be deformed via pinch-off processes to become $\widetilde{\Phi}^{k}|_{D_{i}}$, which in turn can be deformed via pinch-off processes to become $\Phi^{\prime}|_{D_{i}}$, and both deformations appear as subfamilies in $H$. Then, item (3) follows immediately. Moreover, this also imply that $\Phi^{\prime}|_{D_{i}}$ must be of genus $\leq\mathfrak{g}(\Gamma_{i})$, which is the first part of item (2). Finally, $\Phi^{\prime}|_{D_{0}}$ is exactly $\Phi^{K}|_{D_{0}}$, so $\Phi^{\prime}|_{D_{0}}$ is of genus 0. Thus, item (2) is true. This finishes the proof of Theorem 1.4.
Remark 4.3.
In the above construction, we concatenated multiple pinch-off processes together. Namely, the ending surface of one process is the beginning of another. If the ending surface of the first process has a cone point, this cone point would be preserved in the second process. This explains why in Definition 2.18 (2) we need to allow the cone to be left intact.
5. Existence of homologically independent simple short loops
Let $(M,\mathbf{g})$ be an orientable closed Riemannian $3$-manifold. We let $\Gamma$ and $\Sigma$ denote smooth embedded orientable closed surfaces in $M$, and $S$ denote a punctate surface in $\mathcal{S}(M)$. The main goal of these sections is to locate homologically independent simple short loops in a smooth surface and small balls covering nontrivial topology in a punctate surface.
Recall that the systole of a surface $\Gamma$, denoted by $\operatorname{sys}\pi_{1}(\Gamma)$, is the length of the least length of a noncontractible loop in $\Gamma$. In [Gro92], Gromov introduced the notion of $\mathbb{Z}_{2}$ homological systole of $\Gamma$, denoted by $\operatorname{sys}H_{1}(\Gamma;\mathbb{Z}_{2})$, defined as the least length of a loop in $\Gamma$, which represents a non-trivial class in $H_{1}(\Gamma;\mathbb{Z}_{2})$. The remarkable systolic inequality by Loewner and Gromov states that if the area of $\Gamma$ is small, then the $\mathbb{Z}_{2}$ homological systole is also small; see Theorem 5.4. By the standard minimization process in the calculus of variations, the systole can be realized by a simple loop —specifically, simple closed geodesic.
In general, if the genus $\mathfrak{g}(\Gamma)>1$ and the area $\mathcal{H}^{2}(\Gamma)$ is small, one expects to find $\mathfrak{g}(\Gamma)$ homologically independent simple short loops in $\Gamma$. This is closely related to Bers’ constant for hyperbolic surfaces [Bers_74_degenerating_surf, Bers_85_inequality_surf]. A natural idea is to perform a neck-pinch surgery along the detected loop in $M$ realizing the systole and then employ an inductive argument; see Proposition 5.1 in §5.1
In §5.2, we show that in $(M,\mathbf{g})$, for $g_{1}\geq g_{0}\in\mathbb{N}$, a smooth embedded orientable closed surface $\Gamma$ of genus $g_{1}$ must contain $g_{1}-g_{0}$ disjoint simple short loops which are homologically independent, provided that it is close to a smooth embedded orientable closed surface $\Sigma$ of genus $g_{0}$ with respect to the $\mathbf{F}$-metric.
If $\mathfrak{g}(\Gamma)>\mathfrak{g}(\Sigma)$, and $\Gamma$ lies in a tubular neighborhood $N_{\eta}(\Sigma)$ of $\Sigma$, then the map $f\mathrel{\mathop{\ordinarycolon}}\Gamma\hookrightarrow N_{\eta}(\Sigma)\to\Sigma$ induces a non-injective map $f_{*}\mathrel{\mathop{\ordinarycolon}}\pi_{1}(\Gamma)\to\pi_{1}(\Sigma)\,.$ It follows from Gabai’s resolution of the so-called Simple Loop Conjecture [Gabai_85_simple_loop, Theorem 2.1] that there exists a non-contractible simple closed curve $\alpha\subset\Gamma$ such that $f|_{\alpha}$ is homotopically trivial. In particular, together with Dehn’s lemma, $\alpha$ bounds a compressing disk in $M$. However, this topological result does not give us any control on the length of the loop, and our result can be regarded as a quantitative and homological generalization of this simple loop conjecture.
Heuristically, if $\Gamma$ and $\Sigma$ are sufficiently close with respect to the $\mathbf{F}$-metric, a large portion $\Gamma^{\prime}$ of $\Gamma$ is homeomorphic to $\Sigma$ removing finitely many disks. Hence, $\Gamma\setminus\Gamma^{\prime}$ has small area, and one may appeal to Proposition 5.1 to detect multiple simple short loops which lie outside $\Gamma^{\prime}$, provided that its boundary also has small length. The technical details will be presented in the proof of Proposition 5.7, where we also need to deal with the case where some thin parts, such as filigrees, of $\Gamma$ are not inside a tubular neighborhood of $\Sigma$.
Next, in §5.3, we extend this result to a genus $g_{1}$ punctate surface. However, as illustrated in Appendix C, one cannot expect these loops to have small lengths; instead, they are contained in $3$-balls of small radius, i.e., small balls.
5.1. Simple short loops in surfaces with small area
If the surface $\Gamma$ is an embedded surface, then in general, the simple short loop realizing the systole might be knotted in $M$, and in this case, there is no compressing disk to perform the surgery in the ambient manifold. Therefore, in this subsection, $\Gamma$ is always regarded as an intrinsic surface so that we can carry out the inductive argument.
Proposition 5.1 (Existence of homologically independent simple short loops I).
For any $g\in\mathbb{N}$ and $\varepsilon>0$, there exists $\delta=\delta(g,\varepsilon)\in(0,\varepsilon)$ with the following property. For any smooth orientable closed surface $\Gamma$ containing a finite set of disjoint closed disks $\{\overline{D_{i}}\}^{n}_{i=1}$ in $\Gamma$ with piecewise smooth boundary satisfying
$\mathfrak{g}(\Gamma)\leq g\,,\quad\mathcal{H}^{2}(\Gamma)<\delta\,,\quad\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D_{i})<\delta\,$,
there exist $\mathfrak{g}(\Gamma)$ disjoint simple loops $\{c_{j}\}^{\mathfrak{g}(\Gamma)}_{j=1}$ in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$ such that:
  1. (1)
    Each $c_{j}$ has length $\mathcal{H}^{1}(c_{j})\leq\varepsilon$.
  2. (2)
    $\{[c_{j}]\}^{\mathfrak{g}(\Gamma)}_{j=1}$ is linearly independent in $H_{1}(\Gamma;\mathbb{Z}_{2})$.
In particular, $\Gamma\setminus\bigcup^{\mathfrak{g}(\Gamma)}_{j=1}c_{j}$ is a genus 0 surface with boundary.
Remark 5.2.
In this proposition, as well as Proposition 5.7, we specify a finite set of disks that serves as forbidden regions for the loops. This technical assumption allows us to apply the propositions to punctate surfaces by removing neighborhoods of their punctate sets.
Remark 5.3.
By Corollary A.2, in $(M,\mathbf{g})$, small $\mathcal{H}^{2}(\Gamma)$ is equivalent to the assumption that $\mathbf{F}(\Gamma,\emptyset)$ is small. Hence, if we regard $\emptyset$ as a degenerate genus 0 surface, then the condition $\mathcal{H}^{2}(\Gamma)<\delta$ can be replaced by $\mathbf{F}(\Gamma,\emptyset)<\delta$ and the proposition is a special case of Proposition 5.7 with $\Sigma=\emptyset$.
Note that without the forbidden regions condition, this follows immediately from a quantitative estimate on the lengths of pants decomposition [Balacheff_Parlier_Sabourau_12_short_loop_decomp, Proposition 6.3]. In the following, we prove it by applying the systolic inequality inductively and a technical result that allows loops to bypass the forbidden regions.
Theorem 5.4 (Loewner torus inequality [Pu52, Section 3], Gromov’s homological systolic inequality [Gro92, Section 2.C]).
Given a smooth orientable closed surface $\Gamma$ of positive genus, there exists a universal constant $C>0$ such that
$\operatorname{sys}H_{1}(\Gamma;\mathbb{Z}_{2})\mathrel{\mathop{\ordinarycolon}}=\inf\left\{\mathcal{H}^{1}(c)\mathrel{\mathop{\ordinarycolon}}[c]\neq 0\in H_{1}(\Gamma;\mathbb{Z}_{2})\right\}\leq C\sqrt{\mathcal{H}^{2}(\Gamma)}\,.$
In particular, there exists a simple closed geodesic $c\in\Gamma$ such that $c]\neq 0\in H_{1}(\Gamma;\mathbb{Z}_{2})$ and $\mathcal{H}^{1}(c)\leq C\sqrt{\mathcal{H}^{2}(\Gamma)}$.
Remark 5.5.
In [Gro92], Gromov proved an optimal estimate that there exists a universal constant $\widetilde{C}>0$ such that for any $\Gamma$ of genus $g\geq 2$, $\operatorname{sys}H_{1}(\Gamma;\mathbb{Z}_{2})\leq\widetilde{C}{\log g}\sqrt{\mathcal{H}^{2}(\Gamma)/g}\,.$ This optimal result is not needed in our paper.
Lemma 5.6 (Bypassing the forbidden regions).
Let $\alpha$ be a positive real number. Suppose that in a smooth orientable closed surface $\Gamma$, there exists a simple loop $c\subset\Gamma$ and a finite set of disjoint closed disks $\{\overline{D_{i}}\}^{n}_{i=1}$ in $\Gamma$ with piecewise smooth boundary satisfying $\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D_{i})<\alpha\,.$ Then there exists a finite union of disjoint simple loops $\widetilde{c}$ in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$ such that $\widetilde{c}]=[c]$ in $H_{1}(\Gamma;\mathbb{Z}_{2})$ and $\mathcal{H}^{1}(\widetilde{c})<\mathcal{H}^{1}(c)+\alpha\,.$
Proof.
Since each disk has piecewise smooth boundary, by Sard’s theorem, for each disk $D_{i}$, we can choose a larger disk $D^{\prime}_{i}\supset\overline{D_{i}}$ such that $\{\overline{D^{\prime}_{i}}\}^{n}_{i=1}$ are also pairwise disjoint, $c$ intersects $\partial D^{\prime}_{i}$ transversally, and $\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D^{\prime}_{i})<\alpha\,.$
Since each $D^{\prime}_{i}$ is contractible, $c$ separates $D^{\prime}_{i}$ into two sets $E^{\prime}_{i,1}$ and $E^{\prime}_{i,2}$ such that $\partial E^{\prime}_{i,1}\cap D^{\prime}_{i}=\partial E^{\prime}_{i,2}\cap D^{\prime}_{i}=c\cap D^{\prime}_{i}\,.$ We define (see figure 5)
$\widetilde{c}=(c\setminus\bigcup^{n}_{i=1}(\partial E^{\prime}_{i,1}\cap D^{\prime}_{i}))\cup\bigcup^{n}_{i=1}(\partial D^{\prime}_{i}\cap\partial E^{\prime}_{i,1})\,.$
Figure 5. Constructing $\widetilde{c}$ from $c$
Clearly, by the transversality, $\widetilde{c}$ is a union of disjoint simple loops in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$ satisfying $\widetilde{c}]=[c]+\partial[cup_{i}E^{\prime}_{i,1}]=[c]$ in $H_{1}(\Gamma;\mathbb{Z}_{2})$, and
$\mathcal{H}^{1}(\widetilde{c})\leq\mathcal{H}^{1}(c)+\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D^{\prime}_{i}\cap\partial E^{\prime}_{i,1})\leq\mathcal{H}^{1}(c)+\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D^{\prime}_{i})<\mathcal{H}^{1}(c)+\alpha\,.$
Proof of Proposition 5.1.
If $\mathfrak{g}(\Gamma)=0$, the proposition holds trivially. In the following, we fix $g\in\mathbb{N}^{+}$ and only consider the case where $\mathfrak{g}(\Gamma)\in\mathbb{N}^{+}$.
It suffices to prove that for any $g^{\prime}\in\{1,2,\dots,g\}$ and $\varepsilon>0$, there exists $\delta^{\prime}(g,g^{\prime},\varepsilon)\in(0,\varepsilon)$ with the following property: For any smooth orientable closed surface $\Gamma$ containing disjoint closed disks $\{\overline{D_{i}}\}^{n}_{i=1}$ with piecewise smooth boundary, satisfying:
  1. (i)’
    $g^{\prime}\leq\mathfrak{g}(\Gamma)\leq g$,
  2. (ii)’
    $\mathcal{H}^{2}(\Gamma)<\delta^{\prime}(g,g^{\prime},\varepsilon)$,
  3. (iii)’
    $\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D_{i})<\delta^{\prime}(g,g^{\prime},\varepsilon)$,
there exist $g^{\prime}$ disjoint simple loops $\{c_{j}\}^{g^{\prime}}_{j=1}$ in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$ such that:
  1. (1)’
    Each $c_{j}$ has length $\mathcal{H}^{1}(c_{j})\leq\varepsilon$.
  2. (2)’
    $\{[c_{j}]\}^{g^{\prime}}_{j=1}$ is linearly independent in $H_{1}(\Gamma;\mathbb{Z}_{2})$.
We prove this existence result inductively on $g^{\prime}$. For $g^{\prime}=1$, we define $\delta^{\prime}(g,1,\varepsilon)\mathrel{\mathop{\ordinarycolon}}=\min({\varepsilon}/{2},{\varepsilon^{2}}/({4C^{2}}))\,$, where $C$ is the universal constant given in Theorem 5.4. By Theorem 5.4, for any $\Gamma$ and $\{\overline{D_{i}}\}^{n}_{i=1}$ satisfying (i)’,  (ii)’ and  (iii)’ there exists a loop $c^{\prime}_{1}\subset\Gamma$ such that
$\mathcal{H}^{1}(c^{\prime}_{1})\leq C\sqrt{\mathcal{H}^{2}(\Gamma)}\leq\frac{\varepsilon}{2}\,,\quad[c^{\prime}_{1}]\neq 0\in H_{1}(\Gamma;\mathbb{Z}_{2})\,.$
By Lemma 5.6, $c^{\prime}_{1}$ is homologous to a finite union of disjoint simple loops $\widetilde{c}_{1}$ lying outside $\{\overline{D_{i}}\}^{n}_{i=1}$, and
$\mathcal{H}^{1}(\widetilde{c}_{1})<\delta^{\prime}(g,1,\varepsilon)+\mathcal{H}^{1}(c^{\prime}_{1})<\varepsilon\,,\quad[\widetilde{c}_{1}]=[c^{\prime}_{1}]\neq 0\in H_{1}(\Gamma;\mathbb{Z}_{2})\,.$
Hence, there exists at least one loop in $\widetilde{c}_{1}$, denoted by $c_{1}$, such that $\mathcal{H}^{1}(c_{1})<\varepsilon\,$, and $c_{1}]\neq 0\in H_{1}(\Gamma;\mathbb{Z}_{2})\,.$
Inductively, assume the existence of $\delta^{\prime}(g,g^{\prime},\varepsilon)$ for any $\varepsilon$ and $g^{\prime}=k\in\{1,2,\dots,g-1\}$. We define
$\delta^{\prime}(g,k+1,\varepsilon)\mathrel{\mathop{\ordinarycolon}}=\min({\varepsilon}/{4},{\varepsilon^{2}}/({8C^{2}}),\delta^{\prime}(g,k,{\varepsilon}/({8k}))\,)\,.$
By the inductive assumption, for any $\Gamma$ and $\{\overline{D_{i}}\}^{n}_{i=1}$ satisfying  (i)’,  (ii)’ and  (iii)’, there exist $k$ disjoint simple loops $\{c_{j}\}^{k}_{j=1}$ in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$ satisfying  (1)’ and (2)’ with $\varepsilon$ replaced by $\frac{\varepsilon}{8k}$.
We perform a sequence of surgeries on $\Gamma$ along $\{c_{j}\}^{k}_{j=1}$—cut open along these disjoint loops and attach $2k$ disks, denoted by $\{D_{i}\}^{n+2k}_{i=n+1}$, of area small enough such that the resulting smooth orientable closed surface $\Gamma^{\prime}$ of positive genus has
(5.1) $\mathcal{H}^{2}(\Gamma^{\prime})\leq\mathcal{H}^{2}(\Gamma)+\frac{\varepsilon^{2}}{8C^{2}}<\delta^{\prime}(g,k+1,\varepsilon)+\frac{\varepsilon^{2}}{8C^{2}}\leq\frac{\varepsilon^{2}}{4C^{2}}\,.$
Moreover, since each $c_{i}$ has $\mathcal{H}^{1}(c_{i})\leq\frac{\varepsilon}{8k}$, the $(n+2k)$ disks $\{D_{i}\}^{n+2k}_{i=1}$ in $\Gamma^{\prime}$ satisfy
(5.2) $\sum^{n+2k}_{i=1}\mathcal{H}^{1}(\partial D_{i})=\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D_{i})+2k\frac{\varepsilon}{8k}<\delta^{\prime}(g,k+1,\varepsilon)+\frac{\varepsilon}{4}\leq\frac{\varepsilon}{2}\,.$
Applying Theorem 5.4 again to $\Gamma^{\prime}$, together with  (5.1), we obtain a loop $c^{\prime}_{k+1}\subset\Gamma^{\prime}$ with length
$\mathcal{H}^{1}(c^{\prime}_{k+1})\leq C\sqrt{\mathcal{H}^{2}(\Gamma^{\prime})}\leq{\varepsilon}/{2}\,.$
By Lemma 5.6 with (5.2), $c^{\prime}_{k+1}$ is homologous to a finite union of loops $\widetilde{c}_{k+1}\subset\Gamma^{\prime}\setminus\bigcup^{n+2k}_{i=1}\overline{D_{i}}$ with
$\mathcal{H}^{1}(\widetilde{c}_{k+1})<\mathcal{H}^{1}(c^{\prime}_{k+1})+\frac{\varepsilon}{2}\leq\varepsilon\,,\quad[\widetilde{c}_{k+1}]=[c^{\prime}_{k+1}]\neq 0\in H_{1}(\Gamma^{\prime};\mathbb{Z}_{2})\,.$
Hence, there exists a simple loop $c_{k+1}$ in $\widetilde{c}_{k+1}$ with $c_{k+1}]\neq 0\in H_{1}(\Gamma^{\prime};\mathbb{Z}_{2})$, which implies that $c_{k+1}]$ is linearly independent of $c_{1}],[c_{2}],\dots,[c_{k}]$ in $H_{1}(\Gamma;\mathbb{Z}_{2})$. Notice that
$\Gamma^{\prime}\setminus\bigcup^{n+2k}_{i=1}\overline{D_{i}}=\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}\,$,
so $c_{k+1}$ is also a simple loop in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$. Therefore, we construct $\{c_{j}\}^{k+1}_{j=1}$ satisfying  (1)’ and (2)’.
This closes the induction, and the proposition holds with $\delta(g,\varepsilon)\mathrel{\mathop{\ordinarycolon}}=\min_{1\leq g^{\prime}\leq g}\delta^{\prime}(g,g^{\prime},\varepsilon)\,.$
5.2. Simple short loops in surfaces with large area
In this subsection, we prove the non-degenerate case as follows.
Proposition 5.7 (Existence of homologically independent simple short loops II).
In an orientable closed Riemannian $3$-manifold $(M,\mathbf{g})$, let $\Sigma$ be an embedded smooth orientable genus $g_{0}$ closed surface. For any $\varepsilon\in(0,1)$ and any integer $g_{1}\geq g_{0}$, there exists $\delta=\delta(M,\mathbf{g},\Sigma,g_{1},\varepsilon)\in(0,\varepsilon)$ with the following property. In $(M,\mathbf{g})$, for any embedded smooth orientable closed surface $\Gamma$ containing a finite set of disjoint closed disks $\{\overline{D_{i}}\}^{n}_{i=1}$ with piecewise smooth boundary satisfying:
  1. (i)
    $g_{0}\leq\mathfrak{g}(\Gamma)\leq g_{1}$,
  2. (ii)
    $\mathbf{F}(\Gamma,\Sigma)\leq\delta$,
  3. (iii)
    $\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D_{i})<\delta$,
  4. (iv)
    $\sum^{n}_{i=1}\operatorname{diam}_{M}(\overline{D_{i}})<\delta$,
there exist $\mathfrak{g}(\Gamma)-g_{0}$ disjoint simple loops $\{c_{j}\}^{\mathfrak{g}(\Gamma)-g_{0}}_{j=1}$ in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$ such that:
  1. (1)
    Each $c_{j}$ has length $\mathcal{H}^{1}(c_{j})\leq\varepsilon$.
  2. (2)
    $\{[c_{j}]\}^{\mathfrak{g}(\Gamma)-g_{0}}_{j=1}$ is linearly independent in $H_{1}(\Gamma;\mathbb{Z}_{2})$.
In particular, $\Gamma\setminus\bigcup^{g}_{j=1}c_{j}$ is a genus $g_{0}$ surface with $2(\mathfrak{g}(\Gamma)-g_{0})$ boundary circles.
Remark 5.8.
By Lemma 2.6, for sufficiently small $\delta$ depending on $M,\mathbf{g}$ and $\Sigma$, if $\mathbf{F}(\Gamma,\Sigma)<\delta$, then $\mathfrak{g}({\Gamma})\geq\mathfrak{g}({\Sigma})$. Therefore, we can replace the condition $g_{0}\leq\mathfrak{g}(\Gamma)\leq g_{1}$” by $\mathfrak{g}(\Gamma)\leq g_{1}$” in the proposition.
We first prove a useful lemma which enables us to reduce to the case where $\Gamma$ is in a small neighborhood of $\Sigma$.
Lemma 5.9.
In an orientable closed Riemannian $3$-manifold $(M,\mathbf{g})$, let $\Sigma$ be an embedded smooth orientable closed surface. For any $\varepsilon\in(0,1)$, there exists $\delta=\delta(M,\mathbf{g},\Sigma,\varepsilon)\in(0,\varepsilon/2)$ with the following property. In $(M,\mathbf{g})$, for any embedded smooth orientable closed surface $\Gamma$ containing a finite set of disjoint closed disks $\{\overline{D_{i}}\}^{n}_{i=1}$ with piecewise smooth boundary satisfying:
  1. (i)
    $\mathbf{F}(\Gamma,\Sigma)\leq\delta$,
  2. (ii)
    $\sum^{n}_{i=1}\operatorname{diam}_{M}(\overline{D_{i}})<\delta$,
there exists $\eta\in(0,\varepsilon)$ such that:
  1. (1)
    $\exp^{\perp}\mathrel{\mathop{\ordinarycolon}}\Sigma\times[-\eta,\eta]\to M$ is a diffeomorphism to its image $\exp^{\perp}(\Sigma\times[-\eta,\eta])$; We denote $N_{\eta}(\Sigma)\mathrel{\mathop{\ordinarycolon}}=\exp^{\perp}(\Sigma\times(-\eta,\eta))$.
  2. (2)
    $\partial N_{\eta}(\Sigma)$ intersects $\Gamma$ transversally and $\partial N_{\eta}(\Sigma)\cap\bigcup^{n}_{i=1}\overline{D_{i}}=\emptyset$.
  3. (3)
    $\mathcal{H}^{1}(\partial N_{\eta}(\Sigma)\cap\Gamma)<\varepsilon$.
  4. (4)
    $\mathcal{H}^{2}(\Gamma\setminus N_{\eta}(\Sigma))<\varepsilon$.
Moreover, by taking $\delta$ even smaller, we can remove a small tubular neighborhood of $\partial N_{\eta}(\Sigma)\cap\Gamma$ in $\Gamma$, and attach disjoint embedded disks $\{D^{\prime}_{i}\}^{n^{\prime}}_{i=1}$ in $M$ to obtain two disjoint embedded smooth orientable closed surfaces $\Gamma_{1}$ and $\Gamma_{2}$ satisfying:
  1. (5)
    $\Gamma_{1}\subset N_{\eta}(\Sigma)$.
  2. (6)
    $\mathbf{F}(\Gamma_{1},\Sigma)<\varepsilon$.
  3. (7)
    $\mathcal{H}^{2}(\Gamma_{2})<\varepsilon$.
And $\{\overline{D_{i}}\}^{n_{\Gamma}}_{i=1}\mathrel{\mathop{\ordinarycolon}}=\{\overline{D_{i}}\}^{n}_{i=1}\cup\{\overline{D^{\prime}_{i}}\}^{n^{\prime}}_{i=1}$ is also a finite set of disjoint closed disks with piecewise smooth boundary, which satisfies:
  1. (8)
    $\sum^{n^{\prime}}_{i=1}\mathcal{H}^{1}(\partial D^{\prime}_{i})<\varepsilon/2$.
  2. (9)
    $\sum^{n^{\prime}}_{i=1}\operatorname{diam}_{M}(\overline{D^{\prime}_{i}})<\varepsilon/2$.
Proof.
Firstly, we choose $\eta_{0}=\eta_{0}(M,\mathbf{g},\Sigma)\in(0,1)$ such that $\exp^{\perp}\mathrel{\mathop{\ordinarycolon}}\Sigma\times[-\eta_{0},\eta_{0}]\to M$ is a diffeomorphism to its image. We fix $\zeta\in(0,\min(\varepsilon,\eta_{0}))$ which will be determined later.
By the definition of $\mathbf{F}$-metric, we can choose $\delta\in(0,\zeta/10)$ such that for any $\Gamma$ and $\{\overline{D}_{j}\}^{n}_{j=1}$ satisfying (i) and (ii), $\mathcal{H}^{2}(\Gamma\setminus N_{\zeta/2}(\Sigma))<\zeta^{2}/10.$ By the coarea formula and Sard’s theorem, there exists a measurable set $E\in(\zeta/2,\zeta)$ with $\mathcal{H}^{1}(E)>\zeta/5$ such that for any $r\in E$, $\Gamma$ intersects $\partial N_{r}(\Sigma)$ transversally and $\mathcal{H}^{1}(\Gamma\cap\partial N_{r}(\Sigma))<\zeta<\varepsilon\,.$ By (ii), we can choose $\eta\in E$ such that $\partial N_{\eta}(\Sigma)\cap\bigcup^{n}_{i=1}\overline{D_{i}}=\emptyset$. These confirm (1)(2)(3) and (4).
By the Jordan-Schoenflies theorem and the isoperimetric inequality, Lemma A.4, there exist positive constants $\nu_{\Sigma}$ and $\delta_{\Sigma}$, depending on $M$, $\mathbf{g}$ and $\Sigma$, such that for any $r\in(-\eta_{0},\eta_{0})$ and any simple loop $c$ in $\partial N_{r}(\Sigma)$ with $\mathcal{H}^{1}(c)\leq\delta_{\Sigma}$, the loop $c$ bounds a disk $D\subset\partial N_{r}(\Sigma)$ with
$\mathcal{H}^{2}(D)\leq\nu_{\Sigma}(\mathcal{H}^{1}(c))^{2}\leq\nu_{\Sigma}\delta_{\Sigma}\mathcal{H}^{1}(c)\,,\quad\operatorname{diam}_{M}(\overline{D})<\nu_{\Sigma}\mathcal{H}^{1}(c)\,.$
Therefore, if we require that $\zeta<\min\left(\frac{\varepsilon}{2+4\nu_{\Sigma}},\delta_{\Sigma}\right)\,$, then in $(M,\mathbf{g})$, choosing suitable $\delta$ and $\eta$ as before, we can remove a small tubular neighborhood of $\Gamma\cap\partial N_{\eta}(\Sigma)=\bigcup^{n^{\prime}}_{i=1}c^{\prime}_{i}$ and attach disks $\{D^{\prime}_{i}\}^{2n^{\prime}}_{i=1}$ near $\partial N_{\eta}(\Sigma)$ to obtain two disjoint embedded closed surfaces $\Gamma_{1}\subset N_{\eta}(\Sigma)$ and $\Gamma_{2}$ satisfying:
$\mathcal{H}^{2}(\Gamma_{2})<\mathcal{H}^{2}(\Gamma\setminus N_{\eta}(\Sigma))+\sum^{n^{\prime}}_{i=1}\mathcal{H}^{2}(D^{\prime}_{i})<\mathcal{H}^{2}(\Gamma\setminus N_{\eta}(\Sigma))+2\nu_{\Sigma}\delta_{\Sigma}\sum^{n^{\prime}}_{i=1}\mathcal{H}^{1}(\partial D^{\prime}_{i})\leq(1+2\nu_{\Sigma}\delta_{\Sigma})\zeta\leq\varepsilon\,$,
$\sum^{n^{\prime}}_{i=1}\mathcal{H}^{1}(\partial D^{\prime}_{i})\leq\zeta<\varepsilon/2\,,\quad\textrm{ and }\quad\sum^{n^{\prime}}_{i=1}\operatorname{diam}_{M}(\overline{D^{\prime}_{i}})\leq\nu_{\Sigma}\sum^{n^{\prime}}_{i=1}\mathcal{H}^{1}(\partial D^{\prime}_{i})+\delta<\nu_{\Sigma}\zeta+\zeta<\varepsilon/2\,.$
These confirm (5)(7),  (8) and (9).
Finally, we require that $\zeta<\frac{\delta(M,\mathbf{g})}{1+2\nu_{\Sigma}\delta_{\Sigma}}\,$, where $\delta(M,\mathbf{g})$ is a constant from Lemma A.4. By the isoperimetric inequality again, $\Gamma_{2}$ bounds a $3$-dimensional modulo $2$ integral current $A$ such that
$\mathcal{H}^{3}(A)\leq\nu_{M}(\mathcal{H}^{2}(\Gamma_{2}))^{3/2}\leq\nu_{M}(1+2\nu_{\Sigma}\delta_{\Sigma})^{3/2}\zeta^{3/2}\,$,
where $\nu_{M}$ is also a constant from Lemma A.4. Hence,
$\mathcal{F}([\Gamma],[\Gamma_{1}])\leq\nu_{M}(1+2\nu_{\Sigma}\delta_{\Sigma})^{3/2}\zeta^{3/2}\,$,
and
$|\mathcal{H}^{2}(\Gamma)-\mathcal{H}^{2}(\Gamma_{1})|\leq 2\mathcal{H}^{2}(\Gamma_{2})<2(1+2\nu_{\Sigma}\delta_{\Sigma})\zeta\,.$
It follows from Corollary A.2 that, if we choose $\zeta$ small enough, choosing suitable $\delta$ and $\eta$ as before, we have $\mathbf{F}(\Gamma,\Gamma_{1})<\varepsilon/2$ and $\mathbf{F}(\Gamma,\Sigma)\leq\delta<\varepsilon/2\,.$ This confirms (6) of the proposition. Note that the choice of $\zeta$ only depends on $M$, $\mathbf{g}$, $\Sigma$ and $\varepsilon$ and so does $\delta$. ∎
This lemma can be extended to punctate surfaces, which will be used in Appendix B.
Corollary 5.10.
In an orientable closed Riemannian $3$-manifold $(M,\mathbf{g})$, let $\Sigma$ be an embedded smooth orientable closed surface. For any $\varepsilon\in(0,1)$, there exists $\delta=\delta(M,\mathbf{g},\Sigma,\varepsilon)\in(0,\varepsilon)$ with the following property. In $(M,\mathbf{g})$, for any punctate surface $S$ with $\mathbf{F}(S,\Sigma)\leq\delta$, there exists $\eta\in(0,\varepsilon)$ such that:
  1. (1)
    $\exp^{\perp}\mathrel{\mathop{\ordinarycolon}}\Sigma\times[-\eta,\eta]\to M$ is a diffeomorphism to its image $\exp^{\perp}(\Sigma\times[-\eta,\eta])$. We denote $N_{\eta}(\Sigma)\mathrel{\mathop{\ordinarycolon}}=\exp^{\perp}(\Sigma\times(-\eta,\eta))$.
  2. (2)
    $\partial N_{\eta}(\Sigma)$ intersects $S$ transversally.
  3. (3)
    $\mathcal{H}^{1}(\partial N_{\eta}(\Sigma)\cap S)<\varepsilon$.
  4. (4)
    $\mathcal{H}^{2}(S\setminus N_{\eta}(\Sigma))<\varepsilon$.
Moreover, by taking $\eta$ even smaller, we can remove a small tubular neighborhood of $\partial N_{\eta}(\Sigma)\cap S$ in $S$, and attach disjoint embedded disks $\{D^{\prime}_{i}\}^{n^{\prime}}_{i=1}$ in $M$ to obtain two disjoint embedded punctate orientable closed surfaces $S_{1}$ and $S_{2}$ satisfying:
  1. (5)
    $S_{1}\subset N_{\eta}(\Sigma)$.
  2. (6)
    $\mathbf{F}(S_{1},\Sigma)<\varepsilon$.
  3. (7)
    $\mathcal{H}^{2}(S_{2})<\varepsilon$.
Proof.
For a punctate surface $S$, a punctate set $P$ for $S$ can be viewed as a finite set of degenerating disks. Hence, one can follow the same proof of Lemma 5.9 to conclude the corollary. ∎
By Lemma 5.9, the closed surface $\Gamma$ can be decomposed as two components $\Gamma_{1}$ and $\Gamma_{2}$. Since $\Gamma_{2}$ has small area, Proposition 5.1 can be applied to exhaust all short simple loops it contains. For $\Gamma_{1}$, which lies in a $\eta$-normal neighborhood of $\Sigma$, we choose a fine triangulation of $\Sigma$ and apply Proposition 5.1 again to $\Gamma_{1}$ in the $\eta$-normal neighborhood of each triangle. With small enough $\eta$, removing all those short simple loops in $\Gamma_{1}$ as described above yields a genus $g_{0}$ surface with boundary. The details are presented below.
Proof of Proposition 5.7.
If $g_{1}=g_{0}$, the proposition holds trivially. In the following, we fix $g_{1}>g_{0}\in\mathbb{N}$, an embedded orientable genus $g_{0}$ closed surface $\Sigma\subset(M,\mathbf{g})$ and $\varepsilon>0$ as in the proposition.
Step 0. Set up
Set
(5.3) $\delta_{0}\mathrel{\mathop{\ordinarycolon}}=\delta(g_{1},\varepsilon/2)\in(0,\varepsilon/2)$
from Proposition 5.1. Fix a triangulation $\mathcal{T}$ of $\Sigma$ such that:
  1. (0.i)
    For any $\Delta\in\mathcal{T}^{(2)}$, $\mathcal{H}^{2}(\Delta)\in(0,{\delta_{0}}/{10})$.
  2. (0.ii)
    For any $e\in\mathcal{T}^{(1)}$, $\mathcal{H}^{1}(e)\in(0,{\delta_{0}}/{10})$.
Let $m$ denote the total number of cells in $\mathcal{T}$. Set $\tau\mathrel{\mathop{\ordinarycolon}}=\frac{1}{1000m^{4}}$.
Let $X$ be a smooth vector field defined on $\Sigma$ such that for any $e\in\mathcal{T}^{(1)}$, $X|_{e}$ is nowhere zero and is transverse to $e$. $X$ generates an isotopy $\varphi\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\to\operatorname{Diff}(\Sigma)$ with $\varphi(0)=\operatorname{Id}$. For convenience, we denote $\varphi(t)$ by $\varphi_{t}$. Then there exists $T>0$ such that for any $t\in[-2T,2T]$, $e\in\mathcal{T}^{(1)}$ and $\Delta\in\mathcal{T}^{(2)}$,
(5.4) $\frac{\mathcal{H}^{1}(\varphi_{t}(e))}{\mathcal{H}^{1}(e)}\in(1-\frac{\tau^{2}}{100},1+\frac{\tau^{2}}{100})\,\quad\textrm{ and }\quad\frac{\mathcal{H}^{2}(\varphi_{t}(\Delta))}{\mathcal{H}^{2}(\Delta)}\in(1-\frac{\tau^{2}}{100},1+\frac{\tau^{2}}{100})\,.$
For each vertex $v\in\mathcal{T}^{(0)}$, since $T_{v}\Sigma$ is a $2$-dimensional normed vector space, it is isometrically isomorphic to $\mathbb{R}^{2}$. The isomorphism induces a smooth $\mathbb{S}^{1}$ family of vectors at $v$:
$Y_{v}\mathrel{\mathop{\ordinarycolon}}\mathbb{S}^{1}\to\{Z\in T_{v}\Sigma\mathrel{\mathop{\ordinarycolon}}|Z|=1\}\,.$
This can be extended to a continuous family of vector fields on $\Sigma$, $Y\mathrel{\mathop{\ordinarycolon}}\mathbb{S}^{1}\to\Gamma(T\Sigma)$, such that for any vertex $v\in\mathcal{T}^{(0)}$ and any $\theta\in\mathbb{S}^{1}$, $Y(\theta,v)=Y_{v}(\theta)$. Therefore, for any small $\xi>0$, $Y$ generates a smooth family of diffeomorphisms $\psi\mathrel{\mathop{\ordinarycolon}}\overline{\mathbb{D}_{2\xi}}\to\operatorname{Diff}(\Sigma)$ with $\psi(\mathbf{0})=\operatorname{Id}$, where $\overline{\mathbb{D}_{2\xi}}\subset\mathbb{R}^{2}$ is a closed disk centered at the origin with radius $2\xi$. Note that for any $v\in\mathcal{T}^{(0)}$, $\psi(\cdot,v)$ is a diffeomorphism to the geodesic ball $B_{\Sigma}(v,\xi)$. For simplicity, we denote $\psi(z)$ by $\psi_{z}$ and $\varphi_{t}\circ\psi_{z}$ by $\varphi_{z,t}$.
By (5.4), there exists a sufficiently small $\xi>0$ such that for any $z\in\overline{\mathbb{D}_{2\xi}}$, $t\in[-2T,2T]$, $e\in\mathcal{T}^{(1)}$ and $\Delta\in\mathcal{T}^{(2)}$,
(5.5) $\frac{\mathcal{H}^{1}(\varphi_{z,t}(e))}{\mathcal{H}^{1}(e)}\in(1-{\tau^{2}}/{10},1+{\tau^{2}}/{10})\,$,
and
(5.6) $\frac{\mathcal{H}^{2}(\varphi_{z,t}(\Delta))}{\mathcal{H}^{2}(\Delta)}\in(1-{\tau^{2}}/{10},1+{\tau^{2}}/{10})\,.$
In the following, we fix such $\xi>0$.
Next we choose $T>0$ even smaller such that for any $t\in[-2T,2T]$ and $v\in\mathcal{T}^{(0)}$, the map
(5.7) $F_{v,t}\mathrel{\mathop{\ordinarycolon}}\overline{\mathbb{D}_{\xi}}\to\overline{\mathbb{D}_{(1+\tau)\xi}},\quad z\mapsto(\psi(\cdot,v))^{-1}(\varphi_{t}(\psi(z,v)))$
is a diffeomorphism to its image with a bi-Lipschitz constant $1+{\tau^{2}}/{100}$ and smoothly depends on $t$. This is possible because $F_{v,0}=\operatorname{Id}$. Finally, we choose $\zeta>0$ such that the normal exponential map $\exp^{\perp}\mathrel{\mathop{\ordinarycolon}}\Sigma\times(-\zeta,\zeta)\to M$ is a diffeomorphism to its image $N_{\zeta}(\Sigma)$ with a bi-Lipschitz constant $1+\frac{\tau^{2}}{100}$. We denote by $\pi\mathrel{\mathop{\ordinarycolon}}N_{\zeta}(\Sigma)\to\Sigma$ the corresponding closest-point projection.
For convenience, in the following, for any $v\in\mathcal{T}^{(0)}$, $e\in\mathcal{T}^{(1)}$, $\Delta\in\mathcal{T}^{(2)}$ (note that they are all closed sets), $z\in\overline{\mathbb{D}}^{2}_{2\xi}$ and $t\in[-2T,2T]$, we define:
  • the point $v_{z,t}\mathrel{\mathop{\ordinarycolon}}=\varphi_{z,t}(v)$;
  • the vertical line $\mathrm{I}_{v,z,t}\mathrel{\mathop{\ordinarycolon}}=\pi^{-1}(v_{z,t})$;
  • the solid cylinder $\mathrm{S}_{v}\mathrel{\mathop{\ordinarycolon}}=\bigcup_{z\in\mathbb{D}^{2}_{(1+\tau)\xi}}\mathrm{I}_{v,z,0}$;
  • the edge $e_{z,t}\mathrel{\mathop{\ordinarycolon}}=\varphi_{z,t}(e)$;
  • the rectangle $\mathrm{R}_{e,z,t}\mathrel{\mathop{\ordinarycolon}}=\pi^{-1}(e_{z,t})$;
  • the cuboid $\mathrm{Cub}_{e,z}\mathrel{\mathop{\ordinarycolon}}=\bigcup_{t\in(-T,T)}\mathrm{R}_{e,z,t}$;
  • the triangle $\Delta_{z,t}\mathrel{\mathop{\ordinarycolon}}=\varphi_{z,t}(\Delta)$;
  • the cylinder $\mathrm{Cyl}_{\Delta,z,t}\mathrel{\mathop{\ordinarycolon}}=\pi^{-1}(\partial\Delta_{z,t})$;
  • the prism $\mathrm{Pri}_{\Delta,z,t}\mathrel{\mathop{\ordinarycolon}}=\pi^{-1}(\Delta_{z,t})$.
Step 1. Existence of a good triangulation
We show that there exists $\delta=\delta_{1}(M,\mathbf{g},\varepsilon,g_{1},\Sigma)\in(0,1)$ with the following property: For any embedded smooth orientable closed surface $\Gamma\subset N_{\zeta}(\Sigma)$ with $\mathfrak{g}(\Gamma)\leq g_{1}$, $\mathbf{F}(\Gamma,\Sigma)<\delta$ and any finite set of disjoint closed disks $\{\overline{D_{i}}\}^{n}_{i=1}$ in $\Gamma$ with $\sum^{n}_{i=1}\operatorname{diam}_{M}(\overline{D_{i}})<\delta\,$, there exist $z\in\mathbb{D}^{2}_{\xi}$ and $t\in(-T,T)$ such that:
  1. (1)
    For any $v\in\mathcal{T}^{(0)}$, $\mathrm{I}_{v,z,t}\pitchfork\Gamma$ is a single point in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$.
  2. (2)
    For any $\Delta\in\mathcal{T}^{(2)}$, $\mathrm{Cyl}_{\Delta,z,t}\pitchfork\Gamma=\bigcup^{n_{\Delta}}_{k=1}c_{\Delta,k}$ is a finite union of disjoint simple loops in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$ satisfying $\frac{\mathcal{H}^{1}(\mathrm{Cyl}_{\Delta,z,t}\cap\Gamma)}{\mathcal{H}^{1}(\partial\Delta_{z,t})}\in(1-\tau,1+\tau)\,.$
Moreover, for any $\Delta\in\mathcal{T}^{(2)}$, we can arrange $\{c_{\Delta,k}\}^{n_{\Delta}}_{k=1}$ in such an order that:
  1. (3)
    $c_{\Delta,1}$ satisfies $c_{\Delta,1}]\neq 0\in H_{1}(\mathrm{Cyl}_{\Delta,z,t};\mathbb{Z}_{2})$ and $\frac{\mathcal{H}^{1}(c_{\Delta,1})}{\mathcal{H}^{1}(\partial\Delta_{z,t})}\in(1-\tau,1+\tau)\,;$ we call it the long loop for $\Delta$.
  2. (4)
    Every $c_{\Delta,k}$ with $k\in\{2,\dots,n_{\Delta}\}$ satisfies $c_{\Delta,k}]=0\in H_{1}(\mathrm{Cyl}_{\Delta,z,t};\mathbb{Z}_{2})$, $c_{\Delta,k}\cap\bigcup_{v\in\mathcal{T}^{(0)}}\mathrm{I}_{v,z,t}=\emptyset$, and
    $\frac{\sum^{n_{\Delta}}_{k=2}\mathcal{H}^{1}(c_{\Delta,k})}{\mathcal{H}^{1}(\partial\Delta_{z,t})}\in(0,\tau)\,;$
    we call them the short loops for $\Delta$.
In the following, at each step, we will choose $\delta$ to be successively smaller, depending only on all the objects defined in Step 0, which were chosen in terms of $M$, $\mathbf{g}$, $\varepsilon$, $g_{1}$ and $\Sigma$.
Step 1.1.
Choose small $\delta$ such that $\Sigma$ and $\Gamma$ are homologous to each other.
By the isoperimetric inequality, Lemma A.4, we choose $\delta$ small enough, depending on $\Sigma$ and $\zeta$, such that $\mathbf{F}(\Gamma,\Sigma)\leq\delta$ and $\Gamma\subset N_{\zeta}(\Sigma)$ imply $\Sigma$ and $\Gamma$ are homologous to each other in $N_{\zeta}(\Sigma)$ with $\mathbb{Z}_{2}$ coefficients.
This implies that for any $v\in\mathcal{T}^{(0)}$, $e\in\mathcal{T}^{(1)}$, $\Delta\in\mathcal{T}^{(2)}$, $z\in\mathbb{D}^{2}_{(1+\tau)\xi}$, $t\in(-T,T)$, and $\Gamma$ with $\mathbf{F}(\Gamma,\Sigma)\leq\delta$, if $\mathrm{I}_{v,z,0}$ intersects $\Gamma$ transversely, then
(5.8) $\mathrm{I}_{v,z,0}\cap\Gamma]=[v_{z,0}]\neq 0\in H_{0}(\mathrm{I}_{v,z,0};\mathbb{Z}_{2})\,;$
if $\mathrm{R}_{e,z,t}$ intersects $\Gamma$ transversely, then
(5.9) $\mathrm{R}_{e,z,t}\cap\Gamma]=[e_{z,t}]\neq 0\in H_{1}(\mathrm{R}_{e,z,t};\mathbb{Z}_{2})\,;$
if $\mathrm{Cyl}_{\Delta,z,t}$ intersects $\Gamma$ transversely, then
(5.10) $\mathrm{Cyl}_{\Delta,z,t}\cap\Gamma]=[\partial\Delta_{z,t}]\neq 0\in H_{1}(\mathrm{Cyl}_{\Delta,z,t};\mathbb{Z}_{2})\,.$
Step 1.2
Choose smaller $\delta$ and $W_{\Gamma}\subset\mathbb{D}^{2}_{(1+\tau)\xi}$ of relatively large measure such that for any $v\in\mathcal{T}^{(0)}$ and $z\in W_{\Gamma}$, $I_{v,z,0}\pitchfork\Gamma$ at exactly a single point.
For any $v\in\mathcal{T}^{(0)}$, by the definition of $\psi$, we can define a $C^{1}$ function $\mathfrak{z}_{v}\mathrel{\mathop{\ordinarycolon}}\mathrm{S}_{v}\to\mathbb{D}^{2}_{(1+\tau)\xi}$ by $\psi(\mathfrak{z}_{v}(x),v)=\pi(x)\,$, and in $S_{v}$, the Jacobian
$J^{M}_{\mathfrak{z}_{v}}\mathrel{\mathop{\ordinarycolon}}=\sqrt{\det[(\nabla^{M}\mathfrak{z}^{i}_{v}\cdot\nabla^{M}\mathfrak{z}^{j}_{v})_{i,j\in\{1,2\}}]}>0\,.$
We also write
$J^{\Gamma}_{\mathfrak{z}_{v}}\mathrel{\mathop{\ordinarycolon}}=\sqrt{\det[(\nabla^{\Gamma}\mathfrak{z}^{i}_{v}\cdot\nabla^{\Gamma}\mathfrak{z}^{j}_{v})_{i,j\in\{1,2\}}]}\quad\textrm{ and }\quad J^{\Sigma}_{\mathfrak{z}_{v}}\mathrel{\mathop{\ordinarycolon}}=\sqrt{\det[(\nabla^{\Sigma}\mathfrak{z}^{i}_{v}\cdot\nabla^{\Sigma}\mathfrak{z}^{j}_{v})_{i,j\in\{1,2\}}]}\,.$
It is clear that for any $x\in\Gamma\cap S_{v}$, $J^{\Gamma}_{\mathfrak{z}_{v}}\leq J^{M}_{\mathfrak{z}_{v}}$ and for any $x\in\Sigma\cap S_{v}$, $J^{\Sigma}_{\mathfrak{z}_{v}}=J^{M}_{\mathfrak{z}_{v}}$.
Therefore, by the coarea formula, for any $v\in\mathcal{T}^{(0)}$,
(5.11) $\int_{S_{v}\cap\Sigma}J^{M}_{\mathfrak{z}_{v}}=\int_{S_{v}\cap\Sigma}J^{\Sigma}_{\mathfrak{z}_{v}}=\int_{\mathbb{D}^{2}_{(1+\tau)\xi}}\mathcal{H}^{0}(\mathrm{I}_{v,z,0}\cap\Sigma)\mathrm{d}z=\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\,$,
and
(5.12) $\int_{S_{v}\cap\Gamma}J^{M}_{\mathfrak{z}_{v}}\geq\int_{S_{v}\cap\Gamma}J^{\Gamma}_{\mathfrak{z}_{v}}=\int_{\mathbb{D}^{2}_{(1+\tau)\xi}}\mathcal{H}^{0}(\mathrm{I}_{v,z,0}\cap\Gamma)\mathrm{d}z\,.$
Note that for any $v\in\mathcal{T}^{(0)}$, $\partial S_{v}\pitchfork\Sigma$, so we have $\lim_{r\to 0}\sup_{v\in\mathcal{T}^{(0)}}|\Sigma|(N_{r}(\partial S_{v}))=0\,.$ By Lemma A.3, we can choose $\delta$ even smaller such that, $\mathbf{F}(\Gamma,\Sigma)\leq\delta$ implies $\sup_{v\in\mathcal{T}^{(0)}}\mathbf{F}_{S_{v}}(|\Gamma|,|\Sigma|)$ is so small that for any $v\in\mathcal{T}^{(0)}$,
$\int_{S_{v}\cap\Gamma}J^{M}_{\mathfrak{z}_{v}}\leq(1+\frac{\tau^{2}}{10})\int_{S_{v}\cap\Sigma}J^{M}_{\mathfrak{z}_{v}}\,.$
This together with (5.11) and (5.12) implies that for any $v\in\mathcal{T}^{(0)}$,
(5.13) $\int_{\mathbb{D}^{2}_{(1+\tau)\xi}}\mathcal{H}^{0}(\mathrm{I}_{v,z,0}\cap\Gamma)\mathrm{d}z\leq(1+\frac{\tau^{2}}{10})\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\,.$
By Sard’s theorem, for any $\Gamma$ with $\mathbf{F}(\Gamma,\Sigma)\leq\delta$, there exists an open subset $U_{\Gamma}\subset\mathbb{D}^{2}_{(1+\tau)\xi}$ with $\mathcal{H}^{2}(U_{\Gamma})=\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})$ such that for any $z\in U_{\Gamma}$, $\mathrm{I}_{v,z,0}$ intersects $\Gamma$ transversely. By (5.8), we have $\mathcal{H}^{0}(\mathrm{I}_{v,z,0}\cap\Gamma)\geq 1\,.$ For any $v\in\mathcal{T}^{(0)}$, let $W_{\Gamma,v}\subset U_{\Gamma}$ be the set of any $z$ such that $\mathcal{H}^{0}(\mathrm{I}_{v,z,0}\cap\Gamma)=1$, which is also an open subset. It follows from (5.13) that
$\mathcal{H}^{2}(W_{\Gamma,v})+2(\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})-\mathcal{H}^{2}(W_{\Gamma,v}))\leq(1+{\tau^{2}}/{10})\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\,$,
and thus $\mathcal{H}^{2}(W_{\Gamma,v})\geq(1-{\tau^{2}}/{10})\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\,.$
Hence, we can define an open set $W_{\Gamma}\mathrel{\mathop{\ordinarycolon}}=\bigcap_{v\in\mathcal{T}^{(0)}}W_{\Gamma,v}\,.$ Then for any $v\in\mathcal{T}^{(0)}$ and $z\in W_{\Gamma}$, $I_{v,z,0}$ intersects $\Gamma$ transversely at exactly a point. Since $\mathcal{T}^{(0)}\leq m$, the set satisfies
$\mathcal{H}^{2}(W_{\Gamma})\geq(1-m{\tau^{2}}/{10})\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})>(1-{\tau}/{100})\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\,$,
Step 1.3
Choose $z_{\Gamma}\in\mathbb{D}^{2}_{\xi}$ and $E_{\Gamma}\subset(-T,T)$ of relatively large measure such that for any $v\in\mathcal{T}^{(0)}$ and $t\in E_{\Gamma}$, $I_{v,z_{\Gamma},t}\pitchfork\Gamma$ at exactly a single point.
Fix $\Gamma$ with $\mathbf{F}(\Gamma,\Sigma)\leq\delta$. Let $W^{\prime}_{\Gamma}\mathrel{\mathop{\ordinarycolon}}=W_{\Gamma}\cap\mathbb{D}^{2}_{\zeta}$, which satisfies
$\mathcal{H}^{2}(W^{\prime}_{\Gamma})\geq\mathcal{H}^{2}(\mathbb{D}^{2}_{\xi})-{\tau}{}\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})/100\geq\left(1-3\tau\right)\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\,.$
For any $v\in\mathcal{T}^{(0)}$ and $z\in W^{\prime}_{\Gamma}$, since $W^{\prime}_{\Gamma}$ is open, the set
$E_{\Gamma,v,z}\mathrel{\mathop{\ordinarycolon}}=\{t\in(-T,T)\mathrel{\mathop{\ordinarycolon}}F_{v,t}(z)\in W^{\prime}_{\Gamma}\}\,$,
is also open, where $F_{v,t}$ is defined in (5.7). For any $v\in\mathcal{T}^{(0)}$, we have
$\int_{W^{\prime}_{\Gamma}}\mathcal{H}^{1}(E_{\Gamma,v,z})\mathrm{d}z$ $=\int_{W^{\prime}_{\Gamma}}\int^{T}_{-T}\mathbf{1}_{W^{\prime}_{\Gamma}}(F_{v,t}(z))\mathrm{d}t\mathrm{d}z=\int^{T}_{-T}\int_{W^{\prime}_{\Gamma}}\mathbf{1}_{W^{\prime}_{\Gamma}}(F_{v,t}(z))\mathrm{d}z\mathrm{d}t$
$=\int^{T}_{-T}\mathcal{H}^{2}\left(W^{\prime}_{\Gamma}\cap F^{-1}_{v,t}(W^{\prime}_{\Gamma})\right)\mathrm{d}t\,.$
Since for any $t\in(-T,T)$ and $v\in\mathcal{T}^{(0)}$, $F_{v,t}\mathrel{\mathop{\ordinarycolon}}\mathbb{D}^{2}_{\xi}\to\mathbb{D}^{2}_{(1+\tau)\xi}$ is a diffeomorphism of bi-Lispchitz constant $1+\frac{\tau^{2}}{100}$, the Hausdorff measure
$\mathcal{H}^{2}(W^{\prime}_{\Gamma}\cap F^{-1}_{v,t}(W^{\prime}_{\Gamma}))$ $\geq(1-{\tau^{2}}/{10})\mathcal{H}^{2}(F_{v,t}(W^{\prime}_{\Gamma})\cap W^{\prime}_{\Gamma})$
$\geq(1-{\tau^{2}}/{10})(\mathcal{H}^{2}(F_{v,t}(W^{\prime}_{\Gamma}))+\mathcal{H}^{2}(W^{\prime}_{\Gamma})-\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi}))$
$\geq(1-{\tau^{2}}/{10})[(1-{\tau^{2}}/{10})\mathcal{H}^{2}(W^{\prime}_{\Gamma})+\mathcal{H}^{2}(W^{\prime}_{\Gamma})-\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})]$
$\geq(1-10\tau)\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\,.$
These imply that for any $v\in\mathcal{T}^{(0)}$, the set
$W^{\prime\prime}_{\Gamma,v}\mathrel{\mathop{\ordinarycolon}}=\left\{z\in W^{\prime}_{\Gamma}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}^{1}(E_{\Gamma,v,z})\geq 2T(1-\sqrt{\tau})\right\}\,$,
satisfies
$\mathcal{H}^{2}(W^{\prime\prime}_{\Gamma,v})+(1-\sqrt{\tau})(\mathcal{H}^{2}(W^{\prime}_{\Gamma})-\mathcal{H}^{2}(W^{\prime\prime}_{\Gamma,v}))$ $\geq\frac{\int_{W^{\prime}_{\Gamma}}\mathcal{H}^{1}(E_{\Gamma,v,z})\mathrm{d}z}{2T}\geq(1-10\tau)\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})$
and thus, $\mathcal{H}^{2}(W^{\prime\prime}_{\Gamma,v})\geq(1-10\sqrt{\tau})\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})\,.$
As in Step 1.2, we define $W^{\prime\prime}_{\Gamma}\mathrel{\mathop{\ordinarycolon}}=\bigcap_{v\in\mathcal{T}^{(0)}}W^{\prime\prime}_{\Gamma,v}\,$, which satisfies
$\mathcal{H}^{2}(W^{\prime\prime}_{\Gamma})\geq(1-m\cdot 10\sqrt{\tau})\mathcal{H}^{2}(\mathbb{D}^{2}_{(1+\tau)\xi})>0\,.$
In the following, we set $z_{\Gamma}\in W^{\prime\prime}_{\Gamma}$, and $E_{\Gamma}\mathrel{\mathop{\ordinarycolon}}=\bigcap_{v\in\mathcal{T}^{(0)}}E_{\Gamma,v,z_{\Gamma}}\,.$ Then $E_{\Gamma}$ has measure $\mathcal{H}^{1}(E_{\Gamma})\geq 2T(1-m\sqrt{\tau})$. Moreover, by definition, for any $v\in\mathcal{T}^{(0)}$ and $t\in E_{\Gamma}$, $I_{v,z_{\Gamma},t}$ intersects $\Gamma$ at exactly a single point.
Step 1.4
Choose smaller $\delta$ and $\widetilde{E}_{\Gamma}\subset E_{\Gamma}$ of relatively large measure such that for any $e\in\mathcal{T}^{(1)}$ and $t\in\widetilde{E}_{\Gamma}$, $\mathcal{H}^{1}(\mathrm{R}_{e,z_{\Gamma},t}\cap\Gamma)/\mathcal{H}^{1}(e_{z_{\Gamma},t})\in\left(1-\frac{\tau^{2}}{10},1+\frac{\tau}{2}\right)$. Furthermore, the path $\gamma_{e,z,t}$ in $\mathrm{R}_{e,z,t}\cap\Gamma$ connecting $\pi^{-1}(\partial e_{z,t})$ is unique.
For any $1$-cell $e\in\mathcal{T}^{(1)}$ and $z\in\mathbb{D}^{2}_{\xi}$, by the definition of $\varphi$ in Step 0, we can define a $C^{1}$ function $\mathfrak{t}_{e,z}\mathrel{\mathop{\ordinarycolon}}\mathrm{Cub}_{e,z}\to(-T,T)$ by $\pi(x)\in e_{z,\mathfrak{t}_{e,z}(x)}\,.$ By the coarea formula, for each $e\in\mathcal{T}^{(1)}$, we have
(5.14) $\int_{\mathrm{Cub}_{e,z}\cap\Sigma}|\nabla^{M}\mathfrak{t}_{e,z}|=\int_{\mathrm{Cub}_{e,z}\cap\Sigma}|\nabla^{\Sigma}\mathfrak{t}_{e,z}|=\int^{T}_{-T}\mathcal{H}^{1}(e_{z,t})dt\,$,
and
(5.15) $\int_{\mathrm{Cub}_{e,z}\cap\Gamma}|\nabla^{M}\mathfrak{t}_{e,z}|\geq\int_{\mathrm{Cub}_{e,z}\cap\Gamma}|\nabla^{\Gamma}\mathfrak{t}_{e,z}|=\int^{T}_{-T}\mathcal{H}^{1}(\mathrm{R}_{e,z,t}\cap\Gamma)dt\,.$
Note that for any $e\in\mathcal{T}^{(1)}$ and $z\in\mathbb{D}^{2}_{\xi}$, since $\partial\mathrm{Cub}_{e,z}$ intersects $\Sigma$ transversely and deforms continuously with respect to $z$, we have
$\lim_{r\to 0}\sup_{e\in\mathcal{T}^{(1)},z\in\mathbb{D}^{2}_{\xi}}|\Sigma|(N_{r}(\partial\mathrm{Cub}_{e,z}))=0\,.$
By Lemma A.3 again, we can choose $\delta$ even smaller such that, $\mathbf{F}(\Gamma,\Sigma)\leq\delta$ implies
$\sup_{e\in\mathcal{T}^{(1)},z\in\mathbb{D}^{2}_{\xi}}\mathbf{F}_{\mathrm{Cub}_{e,z}}(|\Gamma|,|\Sigma|)$
is sufficiently so small that for any $e\in\mathcal{T}^{(1)}$ and $z\in\mathbb{D}^{2}_{\xi}$,
$\int_{\mathrm{Cub}_{e,z}\cap\Gamma}|\nabla^{M}\mathfrak{t}_{e,z}|\leq(1+\frac{\tau^{2}}{10})\int_{\mathrm{Cub}_{e,z}\cap\Sigma}|\nabla^{M}\mathfrak{t}_{e,z}|\,.$
This together with (5.14) and (5.15) leads to
(5.16) $\int^{T}_{-T}\mathcal{H}^{1}(\mathrm{R}_{e,z,t}\cap\Gamma)dt\leq(1+\frac{\tau^{2}}{10})\int^{T}_{-T}\mathcal{H}^{1}(e_{z,t})dt\,$,
for any $z\in\mathbb{D}^{2}_{\xi}$.
By Sard’s theorem, for any $z\in\mathbb{D}^{2}_{\xi}$ and $\Gamma$ with $\mathbf{F}(\Gamma,\Sigma)\leq\delta$, there exists an open set $O_{\Gamma,z}\subset(-T,T)$ with $\mathcal{H}^{2}(O_{\Gamma,z})=2T$, such that for any $t\in O_{\Gamma,z}$ and $e\in\mathcal{T}^{(1)}$, $\mathrm{R}_{e,z,t}$ intersects $\Gamma$ transversely. It follows from (5.9) that, $\mathrm{R}_{e,z,t}\cap\Gamma$ has at least one path $\gamma_{e,z,t}$ connecting the two components $\pi^{-1}(\partial e_{z,t})$. Since $\exp^{\perp}$ is $\left(1+\frac{\tau^{2}}{10}\right)$-biLipschitz,
(5.17) $\mathcal{H}^{1}(\mathrm{R}_{e,z,t}\cap\Gamma)\geq\mathcal{H}^{1}(\gamma_{e,z,t})\geq\frac{\mathcal{H}^{1}(e_{z,t})}{1+\frac{\tau^{2}}{10}}>(1-\frac{\tau^{2}}{10})\mathcal{H}^{1}(e_{z,t})\,.$
For any $e\in\mathcal{T}^{(1)}$, $z\in\mathbb{D}^{2}_{\xi}$, and $\Gamma$ with $\mathbf{F}(\Gamma,\Sigma)\leq\delta$, we define
$\widetilde{E}_{\Gamma,e,z}\mathrel{\mathop{\ordinarycolon}}=\{t\in O_{\Gamma,z}\mathrel{\mathop{\ordinarycolon}}\mathcal{H}^{1}(\mathrm{R}_{e,z,t}\cap\Gamma)\leq(1+{\tau}/{2})\mathcal{H}^{1}(e_{z,t})\}\,.$
The inequality (5.17) implies that for any $t\in\widetilde{E}_{\Gamma,e,z}$, the path $\gamma_{e,z,t}$ in $\mathrm{R}_{e,z,t}\cap\Gamma$ connecting $\pi^{-1}(\partial e_{z,t})$ is unique. By (5.5) and (5.17), we have
$\int_{O_{\Gamma,z}}\mathcal{H}^{1}(\mathrm{R}_{e,z,t}\cap\Gamma)\mathrm{d}t\geq$ $\mathcal{H}^{1}(\widetilde{E}_{\Gamma,e,z})(1-\frac{\tau^{2}}{10})^{2}\mathcal{H}^{1}(e)+(2T-\mathcal{H}^{1}(\widetilde{E}_{\Gamma,e,z}))(1+\frac{\tau}{2})(1-\frac{\tau^{2}}{10})\mathcal{H}^{1}(e)\,$,
and
$\int^{T}_{-T}\mathcal{H}^{1}(e_{z,t})\mathrm{d}t\leq 2T(1+{\tau^{2}}/{10})\mathcal{H}^{1}(e)\,.$
These together with (5.16) imply that the Hausdorff measure of $\widetilde{E}_{\Gamma,e,z}$ satisfies $\mathcal{H}^{1}(\widetilde{E}_{\Gamma,e,z})\geq 2T(1-\tau)\,.$
For any $\Gamma$ with $\mathbf{F}(\Gamma,\Sigma)<\delta$, we define
$\widetilde{E}_{\Gamma}\mathrel{\mathop{\ordinarycolon}}=\bigcap_{e\in\mathcal{T}^{(1)}}\widetilde{E}_{\Gamma,e,z_{\Gamma}}\cap E_{\Gamma}\,$,
which satisfies
$\mathcal{H}^{1}(\widetilde{E}_{\Gamma})\geq 2T(1-m\tau-m\sqrt{\tau})\geq 2T(1-\sqrt[4]{\tau})>0\,.$
Moreover, for any $e\in\mathcal{T}^{(1)}$ and $t\in\widetilde{E}_{\Gamma}$,
$\frac{\mathcal{H}^{1}(\mathrm{R}_{e,z_{\Gamma},t}\cap\Gamma)}{\mathcal{H}^{1}(e_{z_{\Gamma},t})}\in(1-\frac{\tau^{2}}{10},1+\frac{\tau}{2})\,$,
and the path $\gamma_{e,z_{\Gamma},t}$ in $\mathrm{R}_{e,z_{\Gamma},t}\cap\Gamma$ connecting $\pi^{-1}(\partial e_{z_{\Gamma},t})$ is unique.
Step 1.5
Choose smaller $\delta$ and $F_{\Gamma,\{\overline{D_{i}}\}_{i}}\subset(-T,T)$ of relatively large measure such that for any $e\in\mathcal{T}^{(1)}$ and $t\in F_{\Gamma,\{\overline{D_{i}}\}_{i}}$, $\mathrm{R}_{e,z_{\Gamma},t}\cap\bigcup^{n}_{i=1}\overline{D_{i}}=\emptyset$.
By the definition of $\varphi_{z,t}$, we choose $\delta$ smaller such that $\sum^{n}_{i=1}\mathrm{diam}^{M}(\overline{D_{i}})<\delta$ implies that for any $e\in\mathcal{T}^{(1)}$ and $z\in\mathbb{D}^{2}_{\xi}$, there exists $F_{e,z,\{\overline{D_{i}}\}_{i}}\subset(-T,T)$ with
$\mathcal{H}^{1}(F_{e,z,\{\overline{D_{i}}\}_{i}})\geq 2T(1-\sqrt[4]{\tau})$
such that for any $t\in F_{e,z,\{\overline{D_{i}}\}_{i}}$, we have $\mathrm{R}_{e,z,t}\cap\bigcup^{n}_{i=1}\overline{D_{i}}=\emptyset\,.$ Then we define
$F_{\Gamma,\{\overline{D_{i}}\}_{i}}\mathrel{\mathop{\ordinarycolon}}=\bigcap_{e\in\mathcal{T}^{(1)}}F_{e,z_{\Gamma},\{\overline{D_{i}}\}_{i}}\,$,
which has $\mathcal{H}^{1}(F_{\Gamma,\{\overline{D_{i}}\}_{i}})\geq 2T(1-m\sqrt[4]{\tau})$. Clearly, for any $e\in\mathcal{T}^{(1)}$ and $t\in F_{\Gamma,\{\overline{D_{i}}\}_{i}}$, we have $\mathrm{R}_{e,z_{\Gamma},t}\cap\bigcup^{n}_{i=1}\overline{D_{i}}=\emptyset\,.$
Step 1.6
We complete the proof. Fix an arbitrary $\Gamma$ with $\mathbf{F}(\Gamma,\Sigma)<\delta$ and disks $\{\overline{D_{i}}\}^{n}_{i=1}$ in $\Gamma$ with $\sum^{n}_{i=1}\mathrm{diam}^{M}(\overline{D_{i}})<\delta$. We define $G_{\Gamma,\{\overline{D_{i}}\}_{i}}\mathrel{\mathop{\ordinarycolon}}=\widetilde{E}_{\Gamma}\cap F_{\{\overline{D_{i}}\}_{i},z_{\Gamma}}\,$, which satisfies
$\mathcal{H}^{1}(G_{\Gamma,\{\overline{D_{i}}\}_{i}})\geq 2T(1-2m\sqrt[4]{\tau})>0\,.$
Therefore, we can pick $t_{\Gamma,\{\overline{D_{i}}\}}\in G_{\Gamma,\{\overline{D_{i}}\}_{i}}$.
For simplicity, in the following, we write $\bar{z}=z_{\Gamma}$ and $\bar{t}=t_{\Gamma,\{\overline{D_{i}}\}}$. For any $v\in\mathcal{T}^{(0)}$, by Step 1.3 and Step 1.5, $I_{v,\bar{z},\bar{t}}\pitchfork\Gamma$ is a single point in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$. This confirms (1). For any $\Delta\in\mathcal{T}^{(2)}$, assume that $\partial\Delta=e^{1}\cup e^{2}\cup e^{3}$. Then by Step 1.4 and Step 1.5, for any $e^{i}$, $\mathrm{R}_{e^{i},z,t}\pitchfork\Gamma$ lies in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$, and
$\frac{\mathcal{H}^{1}(\mathrm{R}_{e^{i},z,t}\cap\Gamma)}{\mathcal{H}^{1}({e^{i}}_{z,t})}\in(1-\tau/2,1+\tau/2)\,.$
This confirms (2).
Moreover, for any $\Delta\in\mathcal{T}^{(2)}$ with $\partial\Delta=e^{1}\cup e^{2}\cup e^{3}$, by Step 1.4, for each $e^{i}$, there exists a unique connected component in $\mathrm{R}_{e^{i},z,t}\cap\Gamma$ connecting the two components $\pi^{-1}(\partial e_{z,t})$. Together with (1), we can conclude that the other connected component in $\mathrm{R}_{e^{i},z,t}\cap\Gamma$ are loops away from $\pi^{-1}(\partial e_{z,t})$. Therefore, the cylinder $C_{\Delta_{z,t}}\pitchfork\Gamma=\bigcup^{n_{\Delta}}_{k=1}c_{\Delta,k}$ has a unique loop $c_{\Delta,1}$ with $c_{\Delta,1}]\neq 0\in H_{1}(\mathrm{Cyl}_{\Delta,z,t})$ and
$\frac{\mathcal{H}^{1}(c_{\Delta,1})}{\mathcal{H}^{1}(\partial\Delta_{z,t})}\in(1-\tau/2,1+\tau/2)\,.$
Every other loop $c_{\Delta,k}$ with $k\geq 2$, satisfies $c_{\Delta,k}]=0\in H_{1}(\mathrm{Cyl}_{\Delta,z,t})$, $c_{\Delta,k}\cap\bigcup_{v\in\mathcal{T}^{(0)}}\mathrm{I}_{v,z,t}=\emptyset$, and
$\frac{\sum^{n_{\Delta}}_{k=2}\mathcal{H}^{1}(c_{\Delta,k})}{\mathcal{H}^{1}(\partial\Delta_{z,t})}\in(0,\tau)\,.$
These confirm both (3) and  (4). This completes the proof of Step 1.
Step 2. Special case
We prove the proposition with an extra assumption that $\Gamma\subset N_{\zeta}(\Sigma)$.
Let $\delta_{0}$ be as in Step 0 and $\delta_{1}=\delta_{1}(M,\mathbf{g},\varepsilon,g_{1},\Sigma)$ be as in Step 1. Note that for any $\Delta\in\mathcal{T}^{(2)}$, $z\in\mathbb{D}^{2}_{\xi}$, $t\in(-T,T)$, since $\mathrm{Cyl}_{\Delta,z,t}\pitchfork\Sigma$ and deforms continuosly with respect to $z$ and $t$, we have
$\lim_{r\to 0}\sup_{\Delta\in\mathcal{T}^{(2)},z\in\mathbb{D}^{2}_{\xi},t\in(-T,T)}|\Sigma|(N_{r}(\mathrm{Cyl}_{\Delta,z,t}))=0\,.$
By Lemma A.3,  (0.i) and (5.6), there exists positive $\delta=\delta_{2}(M,\mathbf{g},\varepsilon,g_{1},\Sigma)\in(0,\min(\delta_{0}/3,\delta_{1}))$ such that for any $\Delta\in\mathcal{T}^{(2)}$, $z\in\mathbb{D}^{2}_{\xi}$, $t\in(-T,T)$ and $\Gamma\subset N_{\zeta}(\Sigma)$ with $\mathbf{F}(\Gamma,\Sigma)<\delta$, we have
(5.18) $\mathcal{H}^{2}(\mathrm{Pri}_{\Delta,z,t}\cap\Gamma)<\delta_{0}/3\,.$
Fix $\Gamma\subset N_{\zeta}(\Sigma)$ with $g_{0}\leq\mathfrak{g}(\Gamma)\leq g_{1}$ and $\mathbf{F}(\Gamma,\Sigma)<\delta$ and closed disks $\{\overline{D_{i}}\}^{n}_{i=1}$ in $\Gamma$ with $\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D_{i})<\delta$ and $\sum^{n}_{i=1}\operatorname{diam}_{M}(\overline{D_{i}})<\delta$. Let $z_{1}\mathrel{\mathop{\ordinarycolon}}=\bar{z}\in\mathbb{D}^{2}_{\xi}$ and $t_{1}\mathrel{\mathop{\ordinarycolon}}=\bar{t}\in(-T,T)$ be determined by Step 1.6. For any $\Delta\in\mathcal{T}^{(2)}$, by (0.ii),  (5.6) and Step 1 (2), we have
(5.19) $\mathcal{H}^{1}(\mathrm{Cyl}_{\Delta,z_{1},t_{1}}\cap\Gamma)<\delta_{0}/3<\varepsilon/2\,$,
and by (5.18)
(5.20) $\mathcal{H}^{2}(\mathrm{Pri}_{\Delta,z_{1},t_{1}}\cap\Gamma)<{\delta_{0}}/{3}\,.$
For any $\Delta\in\mathcal{T}^{(2)}$, by Step 1 (1) and (3), we know that for any $v\in\Delta^{(0)}$, we have $\mathrm{I}_{v,z_{1},t_{1}}\cap\Gamma=\mathrm{I}_{v,z_{1},t_{1}}\cap c_{\Delta,1}\,.$ Therefore, every short loop $c_{\Delta,k}$ with $k\geq 2$ lies in the interior of $\mathrm{R}_{e,z_{1},t_{1}}$ for some $e\in\Delta^{(1)}$. In particular, every short loop for $\Delta\in\mathcal{T}^{(2)}$ coincides with exactly one short loop in another $\Delta^{\prime}\in\mathcal{T}^{(2)}$.
In the following, we regard $\Gamma$ as an intrinsic surface. First, we cut $\Gamma$ along every loop $\bigcup_{\Delta\in\mathcal{T}^{(2)}}\{c_{\Delta,k}\}^{n_{\Delta}}_{k=1}$. Then for any $\Delta\in\mathcal{T}^{(2)}$, $\Gamma_{\Delta}\mathrel{\mathop{\ordinarycolon}}=\Gamma\cap\mathrm{Pri}_{\Delta,z_{1},t_{1}}$ is a surface with boundary; the boundary consists of smooth short loops $\{c_{\Delta,k}\}^{n_{\Delta}}_{k=2}$ and a piecewise smooth long loop $c_{\Delta,1}$. We glue disks $\{D_{\Delta,k}\}^{n_{\Delta}}_{k=1}$ of arbitrary small areas to its boundary and obtain a smooth closed orientable surface $\Gamma^{\prime}_{\Delta}$. By (5.19),  (5.20) and the choice of $\delta$, we have
$\sum^{n_{\Delta}}_{k=1}\mathcal{H}^{1}(\partial D_{\Delta,k})+\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D_{i})<\delta_{0},\quad\mathcal{H}^{2}(\Gamma^{\prime}_{\Delta})<\delta_{0}\,.$
Hence, for any $\Delta\in\mathcal{T}^{(2)}$, since $\mathfrak{g}(\Gamma^{\prime}_{\Delta})\leq g_{1}$, applying Proposition 5.1 to $\Gamma^{\prime}_{\Delta}$, we obtain $\mathfrak{g}(\Gamma^{\prime}_{\Delta})$ many disjoint loops $\{c^{\prime}_{\Delta,k}\}^{\mathfrak{g}(\Gamma^{\prime}_{\Delta})}_{k=1}$ in $\Gamma^{\prime}_{\Delta}\setminus\left(\bigcup^{n}_{i=1}\overline{D_{i}}\cup\bigcup^{n_{\Delta}}_{k=1}\overline{D_{\Delta,k}}\right)$, each of which has
(5.21) $\mathcal{H}^{1}(c^{\prime}_{\Delta,k})\leq\varepsilon/2\,,\quad c^{\prime}_{\Delta,k}\subset\Gamma_{\Delta}\setminus\bigcup^{n}_{i=1}\overline{D_{i}}\,.$
Figure 6. $\Gamma_{\Delta}$ and $\Gamma^{\prime}_{\Delta}$
Then for any $\Delta\in\mathcal{T}^{(2)}$, we perform surgeries on $\Gamma^{\prime}_{\Delta}$ along $\{c^{\prime}_{\Delta,k}\}^{\mathfrak{g}(\Gamma^{\prime}_{\Delta})}_{k=1}$, and remove $D_{\Delta,1}$ whose boundary is the long loop $c_{\Delta,1}$. By Proposition 5.1, the resulting surface is a genus 0 surface, consisting of multiple spheres and a triangle $\widetilde{\Delta}$ with boundary $c_{\Delta,1}$. By Step 1 (1) and (3), one can glue any triangles $\{\widetilde{\Delta}\}_{\Delta\in\mathcal{T}^{(2)}}$ in the same manner as $\mathcal{T}^{(2)}=\{\Delta\}$ to obtain a smooth closed surface $\widetilde{\Gamma}$, which is homeomorphic to $\Sigma$. So $\mathfrak{g}(\widetilde{\Gamma})=\mathfrak{g}(\Sigma)=g_{0}\,.$
Therefore, we can perform surgeries on $\Gamma$ along the disjoint loops
$\bigcup_{\Delta\in\mathcal{T}}\{c_{\Delta,k}\}^{n_{\Delta}}_{k=2}\cup\{c^{\prime}_{\Delta,k}\}^{\mathfrak{g}(\Gamma^{\prime}_{\Delta})}_{k=1}$
to obtain a smooth genus $g_{0}$ closed surface $\widetilde{\Gamma}$ and multiple spheres. There are exactly $g=\mathfrak{g}(\Gamma)-g_{0}$ disjoint simple loops $\{c_{j}\}^{g}_{j=1}$ from $\bigcup_{\Delta\in\mathcal{T}}\{c_{\Delta,k}\}^{n_{\Delta}}_{k=2}\cup\{c^{\prime}_{\Delta,k}\}^{\mathfrak{g}(\Gamma^{\prime}_{\Delta})}_{k=1}$, each of which lies in $\Gamma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}$ such that
  1. (2.i)
    For each $j\in\{1,2,\dots,g\}$, by (5.19) and (5.21), $\mathcal{H}^{1}(c_{j})\leq\varepsilon$.
  2. (2.ii)
    $\{[c_{j}]\}^{g}_{j=1}$ is linearly independent in $H_{1}(\Gamma;\mathbb{Z}_{2})$.
Step 3. General case
Let $\delta=\delta_{3}(M,\mathbf{g},\Sigma,g_{1},\varepsilon)\mathrel{\mathop{\ordinarycolon}}=\delta(M,\mathbf{g},\Sigma,\min(\delta_{2},\zeta))$ as in Lemma 5.9. Then by Lemma 5.9, for $\Gamma$ and $\{\overline{D_{i}}\}^{n}_{i=1}$ satisfying all the conditions in the proposition, there exists $\eta\in(0,\zeta)$ we can cut $\Gamma\pitchfork\partial N_{\eta}(\Sigma)=\bigcup^{n_{\Gamma}}_{j=1}c^{\prime}_{j}$ and attach disks to obtain two disjoint orientable closed surfaces $\Gamma^{1}$ and $\Gamma^{2}$ satisfying:
  1. (2.i)
    $\Gamma^{1}\subset N_{\zeta}(\Sigma)$.
  2. (2.ii)
    $\mathbf{F}(\Gamma^{1},\Sigma)<\delta_{2}$.
  3. (2.iii)
    $\mathcal{H}^{2}(\Gamma^{2})<\delta_{2}<\delta_{0}$.
  4. (2.iv)
    $\sum^{n_{\Gamma}}_{j=1}\mathcal{H}^{1}(\partial D_{j})<\delta_{2}$ and $\sum^{n_{\Gamma}}_{j=1}\operatorname{diam}_{M}(\overline{D_{j}})<\delta_{2}$, where $\{\overline{D_{j}}\}^{n_{\Gamma}}_{j=1}$ is as defined in Lemma 5.9.
Applying Step 2 to $\Gamma^{1}$ and Proposition 5.1 to $\Gamma^{2}$ with closed disks $\bigcup^{n_{\Gamma}}_{j=1}\{\overline{D_{j}}\}$, we obtain $g^{\prime}=\mathfrak{g}(\Gamma^{1})+\mathfrak{g}(\Gamma^{2})-g_{0}$ disjoint simple loops $\{\widetilde{c}_{j}\}^{g^{\prime}}_{j=1}$ in $\Gamma\setminus\bigcup^{n_{\Gamma}}_{j=1}\overline{D_{j}}$. Moreover, if we view $\Gamma$ as an intrinsic surface and perform surgeries on $\Gamma$ along $\{\widetilde{c}_{j}\}^{g^{\prime}}_{j=1}\cup\{c^{\prime}_{j}\}^{n_{\Gamma}}_{j=1}$, we obtain a genus $g_{0}$ closed surface and multiple spheres as in Step 2. Therefore, as in Step 2, we can conclude the existence of $g=\mathfrak{g}(\Gamma)-g_{0}$ disjoint simple loops from $\{\widetilde{c}_{j}\}^{g^{\prime}}_{j=1}\cup\{c^{\prime}_{j}\}^{n_{\Gamma}}_{j=1}$ satisfying the proposition. ∎
5.3. Small balls in punctate surfaces
In the previous two subsections, we’ve proved how to detect simple short loops in a smooth surface. However, neither Proposition 5.1 or Proposition 5.7 can be applied to every surface in our Simon–Smith family (Definition 2.7), as not every surface is smooth therein.
In fact, for any $L\in\mathbb{R}^{+}$, one can construct a genus $1$ punctate surface $T$ of area $1$ but any nontrivial simple loop in $S$ has length at least $L$; see Appendix C for the sketch of such a construction. In particular, Proposition 5.1 fails for a punctate surface of genus $1$. Fortunately, one can still find a nontrivial simple loop of $T$ that lies in a geodesic ball of uniformly bounded radius.
To generalize this statement to all punctate surfaces of any genus, we first prove the following useful lemma, which asserts that finitely many (not necessarily disjoint) small geodesic balls can be covered by relatively larger and pairwise disjoint geodesic balls in a controlled manner.
Lemma 5.11.
In an orientable closed Riemannian $3$-manifold $(M,\mathbf{g})$, for any $q,n\in\mathbb{N}^{+}$, $\zeta\in\left(0,\frac{\operatorname{injrad}(M,\mathbf{g})}{2n(3n)^{q}}\right)$, and a finite set of geodesic balls $\{B(p_{i},\zeta)\}^{q}_{i=1}$ (not necessarily disjoint), there exist disjoint geodesic balls $\{B(p^{\prime}_{i},\zeta^{\prime})\}^{q^{\prime}}_{i=1}$ such that:
  1. (1)
    $q^{\prime}\leq q$, $\zeta^{\prime}=(3n)^{q-q^{\prime}}\zeta$ and $\{p^{\prime}_{i}\}^{q^{\prime}}_{i=1}\subset\{p_{i}\}^{q}_{i=1}$.
  2. (2)
    $\bigcup^{q}_{i=1}B(p_{i},\zeta)\subset\bigcup^{q^{\prime}}_{i=1}B(p^{\prime}_{i},\zeta^{\prime})$.
  3. (3)
    The closed balls $\{\overline{B(p^{\prime}_{i},n\zeta^{\prime})}\}^{q^{\prime}}_{i=1}$ are disjoint from each other.
Proof.
If the closed balls $\{\overline{B(p_{i},n\zeta)}\}^{q}_{i=1}$ are disjoint from each other, then the lemma holds trivially with $q^{\prime}=q$, $p^{\prime}_{i}=p_{i}$ and $\zeta^{\prime}=\zeta$. Otherwise, there exist two indices $k\neq l\in\{1,2,\dots,q\}$ such that $\overline{B(p_{k},n\zeta)}\cap\overline{B(p_{l},n\zeta)}\neq\emptyset$. Then, we have $B(p_{k},\zeta)\cup B(p_{l},\zeta)\subset B(p_{k},3n\zeta)\,.$ Set $\zeta_{1}\mathrel{\mathop{\ordinarycolon}}=3n\zeta$ and $\{p_{1,i}\}^{q-1}_{i=1}\mathrel{\mathop{\ordinarycolon}}=\{p_{i}\}^{q}_{i=1}\setminus\{p_{k}\}$. Note that we have
$\bigcup^{q}_{i=1}B(p_{i},\zeta)\subset\bigcup^{q-1}_{i=1}B(p_{1,i},\zeta_{1})\,.$
Inductively, we define $\zeta_{l+1}=3n\zeta_{l}$ provided that $\left\{\overline{B(p_{l,i},n\zeta_{l})}\right\}^{q-l}_{i=1}$ are not pairwise disjoint. We also define $\{p_{l+1,i}\}^{q-l-1}_{i=1}$ as a subset of $\{p_{l,i}\}^{q-l}_{i=1}$ by removing the center of a ball that intersects another as above. Since the number of balls strictly decreases in the induction, this process stops at defining $\zeta_{l}$ for some $l\in\{1,\dots,q-1\}$. Hence, we can take $\zeta^{\prime}=\zeta_{l}$, $q^{\prime}=q-l$, $\{p^{\prime}_{i}\}^{q^{\prime}}_{i=1}=\{p_{l,i}\}^{q-l}_{i=1}\,$, and complete the proof. ∎
Given two punctate surface $S_{1},S_{2}\in\mathcal{S}(M)$ with punctate sets $P_{1}$ and $P_{2}$, we say that $S_{1}$ and $S_{2}$ intersect transversally if the following holds: If $S_{1}\cap P_{2}=S_{2}\cap P_{1}=\emptyset$, then $S_{1}\setminus P_{1}$ and $S_{2}\setminus P_{2}$ intersect transversely We denote their intersection by $S_{1}\pitchfork S_{2}$. We now show how to approximate a punctate surface by a smooth surface.
Lemma 5.12.
In an orientable closed Riemannian $3$-manifold $(M,\mathbf{g})$, suppose that $\Sigma$ is either an empty set or an embedded smooth orientable closed surface. For any $q\in\mathbb{N}^{+}$ and $\varepsilon\in\left(0,\frac{\operatorname{injrad}(M,\mathbf{g})}{2}\right)$, there exists a constant $\delta=\delta(M,\mathbf{g},\Sigma,q,\varepsilon)\in(0,\varepsilon)$ with the following property. For any punctate surface $S\in\mathcal{S}(M)$ with a punctate set $P$ such that $P\leq q\,$ and $\mathbf{F}(S,\Sigma)<\delta$, there exist an integer $\widetilde{q}\leq q$, a real number $r\in(0,\varepsilon)$ and geodesic balls $\{B(p_{i},r)\}^{\widetilde{q}}_{i=1}$ in $(M,\mathbf{g})$ such that:
  1. (1)
    The closed balls $\{\overline{B(p_{i},r)}\}^{\widetilde{q}}_{i=1}$ are pairwise disjoint geodesic balls.
  2. (2)
    For each $i\in\{1,2,\dots,\widetilde{q}\}$, $\partial B(p_{i},r)\pitchfork\Sigma$ is a union of finitely many loops.
  3. (3)
    $S\setminus\bigcup^{\widetilde{q}}_{i=1}B(p_{i},r)$ is an embedded smooth surface with boundary and genus at most $\mathfrak{g}(S)$.
  4. (4)
    There exists an embedded smooth orientable closed surface $\Gamma$ such that
    1. (a)
      $\Gamma\setminus\bigcup^{\widetilde{q}}_{i=1}B(p_{i},r)=S\setminus\bigcup^{\widetilde{q}}_{i=1}B(p_{i},r)$.
    2. (b)
      $\mathbf{F}(\Gamma,\Sigma)<\varepsilon$.
    3. (c)
      $\Gamma\cap\bigcup^{\widetilde{q}}_{i=1}\overline{B(p_{i},r)}$ is a union of disjoint closed disks $\{\overline{D_{j}}\}^{n}_{j=1}$ satisfying $\sum^{n}_{j=1}\mathcal{H}^{1}(\partial D_{j})<\varepsilon$ and $\sum^{n}_{j=1}\operatorname{diam}(\overline{D_{j}})<\varepsilon\,.$
Proof.
Since $\Sigma$ is either an empty set or a smooth closed surface, there exists a positive constant $C_{\Sigma}>1$ depending only on $M,\mathbf{g}$ and $\Sigma$, such that for any $p\in M$ and $r\in(0,\operatorname{injrad}(M,\mathbf{g})/2)$,
$\mathcal{H}^{2}(\overline{B(p,r)}\cap\Sigma)\leq C_{\Sigma}r^{2}\,,\quad\mathcal{H}^{3}(\overline{B(p,r)})\leq C_{\Sigma}r^{3}\,.$
By the Jordan-Schoenflies theorem and Lemma A.4, there exists $\nu_{M}>0$, depending only on $(M,\mathbf{g})$, such that for any $p\in M$ and $r\in(0,\operatorname{injrad}(M,\mathbf{g})/2)$, every simple loop $c$ in $\partial B(p,r)$ bounds a disk $D\subset\partial B(p,r)$ with
(5.22) $\mathcal{H}^{2}(D)\leq\nu_{M}\mathcal{H}^{1}(c)^{2}\,.$
Let $\eta_{1}=\delta(M,\mathbf{g},\Sigma,\varepsilon/2)/2$ be defined as in Corollary A.2. We choose $\eta_{2}\in(0,\operatorname{injrad}(M,\mathbf{g})/4)$ such that
$(2\nu_{M}+1)\eta^{2}_{2}<\eta_{1}\,,\quad qC_{\Sigma}(\eta_{2})^{3}<\eta_{1}\,,\quad\text{and }2\eta_{2}<\varepsilon\,.$
Let $C_{q}>1$ be a constant depending only on $q$ and determined later. And we set $\eta_{3}=\frac{\eta_{2}}{8qC_{q}C_{\Sigma}}$.
By the definition of $\mathbf{F}$-metric, there exists $\delta\in(0,\varepsilon/2)$ such that for any $S\in\mathcal{S}(M)$ with $\mathbf{F}(S,\Sigma)\leq\delta$ and any $p\in M$,
$\mathcal{H}^{2}(\overline{B(p,\eta_{3})}\cap S)\leq 2\mathcal{H}^{2}(\overline{B(p,2\eta_{3})}\cap\Sigma)\leq 2C_{\Sigma}(2\eta_{3})^{2}=8C_{\Sigma}(\eta_{3})^{2}\,.$
In the sequel, we fix a punctate surface $S$ with a punctate set $P$ satisfying $\mathbf{F}(S,\Sigma)\leq\delta$ and $P\leq q$. By Sard’s theorem and the coarea formula, we can choose a constant $C_{q}>0$, only depending on $q$, and a radius $\widetilde{r}\in\left(\eta_{3}/3^{q},2\eta_{3}/3^{q}\right)$ such that, for any $p\in P$ and any $l\in\{1,3,3^{2},\dots,3^{q}\}$, $\partial B(p,l\widetilde{r})\pitchfork S$ is a finite union of disjoint simple loops and
$\mathcal{H}^{1}(\partial B(p,l\widetilde{r})\cap S)\leq C_{q}\mathcal{H}^{2}(\overline{B(p,2\eta_{3})}\cap\Sigma)/\eta_{3}<8C_{q}C_{\Sigma}\eta_{3}=\eta_{2}/q\,.$
By Lemma 5.11 with $n=1$, there exist $P^{\prime}\subset P$ and a radius $r\in\{\widetilde{r},3\widetilde{r},3^{2}\widetilde{r},\dots,3^{q}\widetilde{r}\}\subset(0,\eta_{3})$ such that $\breve{\Gamma}=S\setminus\bigcup_{p^{\prime}\in P^{\prime}}B(p^{\prime},r)$ is a smooth surface with boundary satisfying
$\sum_{p^{\prime}\in P^{\prime}}\mathcal{H}^{1}(\partial B(p^{\prime},r)\cap S)<(\#P^{\prime})\eta_{2}/q\leq\eta_{2}\,$,
$P\subset\bigcup_{p^{\prime}\in P^{\prime}}B(p^{\prime},r)$ and $\{\overline{B(p^{\prime},r)}\}_{p^{\prime}\in P^{\prime}}$ are disjoint from each other. Clearly, we also have $\mathfrak{g}(\breve{\Gamma})\leq\mathfrak{g}(S)$. These confirms (1)(2) and (3).
Then by (5.22), for each $p^{\prime}\in P^{\prime}$, every simple loop $c$ in $\partial B(p^{\prime},r)\cap S$ bounds a disk $D\subset\partial B(p^{\prime},r)$ with $\mathcal{H}^{2}(D)\leq\nu_{M}\left(\mathcal{H}^{1}(c)\right)^{2}\,.$ In particular, for each $p^{\prime}\in P^{\prime}$, near $B(p^{\prime},r)$, we can remove $B(p^{\prime},r)\cap S$ and attach finitely many disks $\{D_{p^{\prime},j}\}^{m_{p^{\prime}}}_{j=1}$ near $\partial B(p^{\prime},r)$, yielding a smooth closed surface $\Gamma$ of genus $\leq\mathfrak{g}(S)$, confirming (4)(a).
Furthermore, we have
$\sum_{p^{\prime}\in P^{\prime}}\sum^{m_{p^{\prime}}}_{j=1}\mathcal{H}^{2}(D_{p^{\prime},j})$ $\leq 2\sum_{p^{\prime}\in P^{\prime}}\sum^{m_{p^{\prime}}}_{j=1}\nu_{M}(\mathcal{H}^{1}(\partial D_{p^{\prime},j}))^{2}\leq 2\nu_{M}[\sum_{p^{\prime}\in P^{\prime}}\sum^{m_{p^{\prime}}}_{j=1}\mathcal{H}^{1}(\partial D_{p^{\prime},j})]^{2}\leq 2\nu_{M}(\eta_{2})^{2}\,.$
Therefore, we have
$|\mathcal{H}^{2}(\Gamma)-\mathcal{H}^{2}(S)|$ $\leq\sum_{p^{\prime}\in P^{\prime}}\sum^{m_{p^{\prime}}}_{j=1}\mathcal{H}^{2}(D_{p^{\prime},j})+\sum_{p^{\prime}\in P^{\prime}}\mathcal{H}^{2}(\overline{B(p^{\prime},r)}\cap S)$
$<2\nu_{M}(\eta_{2})^{2}+8qC_{\Sigma}(\eta_{3})^{2}<\left(2\nu_{M}+1\right)(\eta_{2})^{2}<\eta_{1}\,$,
and
$\mathcal{F}([\Gamma],[S])\leq\sum_{p^{\prime}\in P^{\prime}}\mathcal{H}^{3}(B(p^{\prime},r))\leq qC_{\Sigma}r^{3}<qC_{\Sigma}(\eta_{2})^{3}<\eta_{1}\,.$
By the choice of $\eta_{1}$ and Corollary A.2, we have $\mathbf{F}(\Gamma,S)<\varepsilon/2\,$, and thus, $\mathbf{F}(\Gamma,\Sigma)<\varepsilon\,$, confirming (4)(b).
Finally, we have
$\sum_{p^{\prime}\in P^{\prime}}\sum^{m_{p^{\prime}}}_{j=1}\mathcal{H}^{1}(\partial D_{p^{\prime},j})=\sum_{p^{\prime}\in P^{\prime}}\mathcal{H}^{1}(\partial B(p^{\prime},r)\cap S)<\eta_{2}<\varepsilon\,$,
and
$\sum_{p^{\prime}\in P^{\prime}}\sum^{m_{p^{\prime}}}_{j=1}\operatorname{diam}_{M}(\overline{D_{p^{\prime},j}})\leq 2\sum_{p^{\prime}\in P^{\prime}}\sum^{m_{p^{\prime}}}_{j=1}\mathcal{H}^{1}(\partial D_{p^{\prime},j})<2\eta_{2}<\varepsilon\,$,
confirming (4)(c). ∎
Here is a generalization of Proposition 5.1 to punctate surfaces, as an application of the previous approximation lemma.
Lemma 5.13 (Existence of small balls I).
In an orientable closed Riemannian $3$-manifold $(M,\mathbf{g})$, for any integers $q,N\in\mathbb{N}^{+}$ and any real number $\varepsilon\in\left(0,\operatorname{injrad}(M,\mathbf{g})/(2N)\right)$, there exists a constant $\delta=\delta(M,\mathbf{g},q,N,\varepsilon)>0$ with the following property. For any punctate surface $S\in\mathcal{S}(M)$ with a punctate set $P$ such that $\mathfrak{g}(S)+\#P\leq q$ and $\mathcal{H}^{2}(S)<\delta$, there exist an integer $\widetilde{q}\leq q$, a real number $r\in(\frac{\varepsilon}{2(3N)^{q}},\varepsilon)$ and geodesic balls $\{B(p_{i},r)\}^{\widetilde{q}}_{i=1}$ in $(M,\mathbf{g})$ such that:
  1. (1)
    The closed balls $\{\overline{B(p_{i},Nr)}\}^{\widetilde{q}}_{i=1}$ are pairwise disjoint geodesic balls.
  2. (2)
    For each $i\in\{1,2,\dots,\widetilde{q}\}$ and $l\in\{1,2,\dots,N\}$, $\partial B\left(p_{i},lr\right)\pitchfork S$ is a union of finitely many disjoint loops.
  3. (3)
    $S\setminus\bigcup^{\widetilde{q}}_{i=1}B\left(p_{i},r\right)$ is a smooth genus 0 surface with boundary.
Proof.
We set $\eta\mathrel{\mathop{\ordinarycolon}}=\frac{\varepsilon}{2(3N)^{q}}\,.$ Let $S$ be a punctate surface with a punctate set $P$ and $P\leq q$. By Sard’s theorem and the coarea formula, there exist a constant $C=C(q)$ and a radius $r\in\left(\frac{\eta}{2\cdot 3^{q}},\frac{\eta}{3^{q}}\right)$ such that for any $p\in P$ and $l\in\{1,3,3^{2},\dots,3^{q}\}$, $\partial B(p,lr)\pitchfork S$ is a finite union of loops and
$\sum_{p\in P}\mathcal{H}^{1}(\partial B(p,lr)\cap S)\leq C{\mathcal{H}^{2}(S)}/{\eta}\,.$
By Lemma 5.11, there exist $P^{\prime}\subset P$ and $r^{\prime}\in\{r,3r,3r^{2},\dots,3^{q}r\}\subset(\frac{\eta}{2\cdot 3^{q}},\eta)$ such that $\Gamma=S\setminus\bigcup_{p^{\prime}\in P^{\prime}}B(p^{\prime},r^{\prime})$ is a smooth surface with boundary satisfying
(5.23) $\sum_{p^{\prime}\in P^{\prime}}\mathcal{H}^{1}(\partial B(p^{\prime},r^{\prime})\cap S)\leq C{\mathcal{H}^{2}(S)}/{\eta}\,,\quad P\subset\bigcup_{p^{\prime}\in P^{\prime}}B(p^{\prime},r^{\prime})\,$,
and $g\mathrel{\mathop{\ordinarycolon}}=\mathfrak{g}(\Gamma)\leq\mathfrak{g}(S)$.
In the sequel, we regard $\Gamma$ as an intrinsic surface, rather than a surface embedded in $(M,\mathbf{g})$. We also let $\widetilde{\delta}=\delta(q,\eta)$ be the constant as determined in Proposition 5.1. By (5.23), taking $\delta\in(0,\widetilde{\delta}/2)$ sufficiently small, if we further assume that $\mathcal{H}^{2}(S)<\delta$, then we can attach disks $\{D_{i}\}^{n}_{i=1}$ with arbitrarily small area to the boundary of $S\setminus\bigcup_{p^{\prime}\in P^{\prime}}B(p^{\prime},r^{\prime})$ to obtain a smooth genus $g$ closed surface $\Sigma$, and
(5.24) $\sum^{n}_{i=1}\mathcal{H}^{1}(\partial D_{i})\leq\widetilde{\delta},\quad\mathcal{H}^{2}(\Gamma)\leq 2\mathcal{H}^{2}(S)<2\delta<\widetilde{\delta}\,$.
Moreover, it follows from Proposition 5.1 with (5.24), there exists $g$ disjoint simple loops $\{c_{j}\}^{g}_{j=1}$ in $\Sigma\setminus\bigcup^{n}_{i=1}\overline{D_{i}}\subset S$ such that
  1. (i)
    For each $j\in\{1,\dots,g\}$, each $c_{j}$ satisfies $\mathcal{H}^{1}(c_{j})\leq\eta$.
  2. (ii)
    $\{[c_{j}]\}^{g}_{j=1}$ is linearly independent in $H_{1}(\Sigma)$.
In particular, each $c_{j}$ is contained in some ball $B(p_{j},\eta)$. Applying Sard’s theorem again, there exists $\eta^{\prime}\in(\eta,2\eta)$ such that for any integer $l\in\{1,2,3,\dots,N(3N)^{q}\}$, each $B(p_{j},l\eta^{\prime})$ with $j\in\{1,2,\dots,g\}$ intersects $S$ transversally, and so does each $B(p^{\prime},l\eta^{\prime})$ with $p^{\prime}\in P^{\prime}$.
In the sequel, we denote $\{B(p,\eta^{\prime})\}_{p\in P}\cup\{B(p_{j},\eta^{\prime})\}^{g}_{j=1}$ by $\{B(p_{j},\eta^{\prime})\}^{q^{\prime}}_{j=1}$, where $q^{\prime}=g+\#P^{\prime}\leq q$. Applying Lemma 5.11 to $\{B(p_{j},\eta^{\prime})\}^{q^{\prime}}_{j=1}$ with $n=N$, we obtain $\widetilde{q}\leq q^{\prime}$, $\{p^{\prime}_{j}\}^{\widetilde{q}}_{j=1}\subset\{p_{j}\}^{q^{\prime}}_{j=1}$, and a radius $\widetilde{r}=(3N)^{q^{\prime}-\widetilde{q}}\eta^{\prime}\in(\frac{\varepsilon}{2(3N)^{q}},\varepsilon)$, such that the closed balls $\{\overline{B(p^{\prime}_{j},N\widetilde{r})}\}^{\widetilde{q}}_{j=1}$ are disjoint from each other. This confirms (1). Finally, the transversality result  (2) follows from the choice of $\eta^{\prime}$, and the genus result  (3) follows from the choice of the balls. ∎
Similarly, we can extend Proposition 5.7 to punctate surfaces.
Lemma 5.14 (Existence of small balls II).
In a $3$-dimensional orientable closed Riemannian manifold $(M,\mathbf{g})$, for any $q,N\in\mathbb{N}^{+}$, $g_{0}\in\mathbb{N}$, $\mathscr{S}_{g_{0}}$ a compact subset of embedded smooth orientable genus $g_{0}$ closed surfaces into $(M,\mathbf{g})$ (endowed with smooth topology), $\varepsilon\in(0,\operatorname{injrad}(M,\mathbf{g})/(2N))$, there exists $\delta=\delta(M,\mathbf{g},\mathscr{S}_{g_{0}},q,N,\varepsilon)>0$ with the following property. For any punctate surface $S\in\mathcal{S}(M)$ with punctate set $P$ that satisfies $\mathfrak{g}(\Sigma)-g_{0}+\#P\leq q$ and any smooth surface $\Sigma\in\mathscr{S}_{\mathfrak{g}_{0}}$ with $\mathbf{F}(S,\Sigma)\leq\delta$, there exist an integer $\widetilde{q}\leq q$, a real number $r\in\left(\frac{\varepsilon}{2(3N)^{q}},\varepsilon\right)$ and geodesic balls $\{B(p_{i},r)\}^{\widetilde{q}}_{i=1}$ in $(M,\mathbf{g})$ such that:
  1. (1)
    The closed balls $\{\overline{B(p_{i},Nr)}\}^{\widetilde{q}}_{i=1}$ are pairwise disjoint geodesic balls.
  2. (2)
    For each $i\in\{1,2,\dots,\widetilde{q}\}$ and $l\in\{1,2,\dots,N\}$, $\partial B(p_{i},lr)\pitchfork S$ is a union of finitely many loops.
  3. (3)
    $S\setminus\bigcup^{\widetilde{q}}_{i=1}B(p_{i},r)$ is a smooth genus $g_{0}$ surface with boundary.
Proof.
Step 1. We prove the lemma when $\mathscr{S}_{g_{0}}=\{\Sigma\}$ is a singleton set. Let $\widetilde{\delta}_{1}(\eta)=\delta(M,\mathbf{g},\Sigma,q,\eta)\in(0,\eta)$ be as determined in Proposition 5.7. We set $\eta_{1}\mathrel{\mathop{\ordinarycolon}}=\frac{\varepsilon}{2(3N)^{q}}$, and $\eta_{2}\mathrel{\mathop{\ordinarycolon}}=\widetilde{\delta}_{1}(\eta_{1})/2\,.$ Let $\widetilde{\delta}_{2}=\delta(M,\mathbf{g},\Sigma,q,\eta_{2})$ be as in Lemma 5.12. If $\mathbf{F}(S,\Sigma)<\delta$, then there exist a finite set $P^{\prime}\subset P$, a radius $r^{\prime}\in(0,\eta_{2})$ and an embedded smooth orientable closed surface $\Gamma$ and a finite set of closed disks $\{\overline{D}_{j}\}^{n}_{j=1}$ in $\Gamma$ such that:
  1. (1)
    $\Gamma\setminus\bigcup^{n}_{j=1}\overline{D_{j}}=S\setminus\bigcup_{p^{\prime}\in P^{\prime}}\overline{B(p^{\prime},r^{\prime})}$ and thus, $g_{1}\mathrel{\mathop{\ordinarycolon}}=\mathfrak{g}(\Gamma)\leq\mathfrak{g}(S)$.
  2. (2)
    $\mathbf{F}(\Gamma,S)\leq 2\eta_{2}$.
  3. (3)
    $\sum^{n}_{j=1}\operatorname{diam}_{M}(\overline{D_{j}})<\eta_{2}$.
  4. (4)
    $\sum^{n}_{j=1}\mathcal{H}^{1}(\partial\overline{D_{j}})<\eta_{2}$.
It follows from Proposition 5.7 that there exist $g=g_{1}-\mathfrak{g}(\Sigma)$ disjoint simple loops $\{c_{j}\}^{g}_{k=1}$ in $\Gamma^{\prime}\setminus\bigcup^{n}_{j=1}\overline{D_{j}}\subset S$, which are homologically independent in $H_{1}(\Gamma^{\prime})$. Moreover, each $c_{k}$ satisfies $\mathcal{H}^{1}(c_{k})\leq\eta_{1}\,$, so there exists $\widetilde{p}_{k}\in M$ and $c_{k}\subset B(\widetilde{p}_{k},\eta_{1})$.
We set $P^{\prime\prime}\mathrel{\mathop{\ordinarycolon}}=P^{\prime}\cup\{\widetilde{p}_{k}\}^{g}_{k=1}$. By Sard’s Lemma, there exists $\widetilde{r}\in(\eta_{1},2\eta_{1})=\left(\frac{\varepsilon}{2(3N)^{q}},\frac{\varepsilon}{(3N)^{q}}\right)$ such that for every $p\in P^{\prime\prime}$, $l\in\{1,2,\dots,N(3N)^{q}\}$, $\partial B(p,l\widetilde{r})\pitchfork\Sigma$ is a union of finitely many loops. Then applying Lemma 5.11 again to $P^{\prime\prime}$ balls $\{B(p,\widetilde{r})\}_{p\in P^{\prime\prime}}$ with $n=N$, we obtain $\widetilde{q}\leq\#P^{\prime\prime}\leq q$, $r\in(\frac{\varepsilon}{2(3N)^{q}},\varepsilon)$ and disjoint geodesic balls $\{B(p_{i},r)\}^{\widetilde{q}}_{i=1}$ satisfying  (1)(2)(3) of the lemma. In particular, we have defined $\delta(M,\mathbf{g},\varepsilon,q,N,\{\Sigma\})\mathrel{\mathop{\ordinarycolon}}=\widetilde{\delta}_{2}$ for any embedded smooth orientable closed genus $g_{0}$ surface $\Sigma$.
Step 2
Let $\mathscr{S}_{g_{0}}$ be a compact set with respect to the smooth topology. Since the smooth topology is finer than the topology induced by the $\mathbf{F}$-distance defined in Definition 2.4, $\mathscr{S}_{g_{0}}\subset\mathcal{S}$ is a compact set with respect to the $\mathbf{F}$-distance. For each $\Sigma\in\mathscr{S}_{g_{0}}$, we define
$B_{\Sigma}\mathrel{\mathop{\ordinarycolon}}=\left\{\Gamma\in\mathscr{S}_{g_{0}}\mathrel{\mathop{\ordinarycolon}}\mathbf{F}(\Sigma,\Gamma)<\frac{\delta(M,\mathbf{g},\{\Sigma\},q,N\varepsilon)}{4}\right\}\,$,
where $\delta(M,\mathbf{g},\{\Sigma\},q,N,\varepsilon)$ is already defined in Step 1. Then the open balls $\{B_{\Sigma}\}_{\Sigma\in\mathscr{S}_{g_{0}}}$ is an open covering of $\mathscr{S}_{g_{0}}$. By compactness, it has a finite subcovering $\{B_{\Sigma_{l}}\}^{m}_{l=1}.$ Therefore, we can define
$\delta(M,\mathbf{g},\mathscr{S}_{g_{0}},q,N,\varepsilon)=\min_{l\in\{1,2,\dots,m\}}\frac{\delta(M,\mathbf{g},\{\Sigma_{l}\},q,N,\varepsilon)}{4}\,$,
and then  (1)(2)(3) of the lemma hold. ∎
6. Pinch-off processes and interpolations
Let $g_{0}\leq g_{1}$ be two natural numbers. Intuitively, for a Simon–Smith family of genus $\leq g_{1}$ sufficiently close to a smooth genus $g_{0}$ closed surface $\Sigma$, we show that one can pinch all the small necks and shrink all small connected components in a continuous way to obtain a Simon–Smith family of genus $g_{0}$. Note that by Lemma 2.6, each surface in the Simon–Smith family has genus at least $g_{0}$. With ingredients in this section, we will be able to prove Theorem 2.21, and Proposition 2.22 and 2.23.
Proposition 6.1 (Pinch-off proposition I).
Let $(M,\mathbf{g})$ be an orientable closed Riemannian $3$-manifold, $X$ be a cubical subcomplex of $I(m,k)$, $g_{0}\in\mathbb{N}$, $\mathscr{S}_{g_{0}}$ be a compact subset of embeddings of a genus $g_{0}$ surface into $M$ (endowed with smooth topology), $q\in\mathbb{N}$ and $\varepsilon\in(0,1)$. Then there exists $\delta=\delta(M,\mathbf{g},\mathscr{S}_{g_{0}},m,q,\varepsilon)\in(0,\varepsilon)$ with the following property. If $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ is a Simon–Smith family of genus $\leq g_{1}$ with $N(P_{\Phi})+g_{1}-g_{0}\leq q\,$, and $\Sigma\mathrel{\mathop{\ordinarycolon}}X\to\mathscr{S}_{g_{0}}$ is a map such that for every $x\in X$, $\mathbf{F}(\Phi(x),\Sigma(x))\leq\delta\,$, then there exists a Simon–Smith family $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}(M)$ of genus $\leq g_{1}$ such that for any $x\in X$:
  1. (1)
    $H(0,x)=\Phi(x)$, $\mathfrak{g}(H(1,x))=g_{0}$.
  2. (2)
    For any $0\leq t\leq t^{\prime}\leq 1$,
    $\mathcal{H}^{2}(\Phi(x))-\varepsilon\leq\mathcal{H}^{2}(H(t,x))\leq\mathcal{H}^{2}(\Phi(x))+\varepsilon\,,\quad\mathcal{F}([H(t,x)],[\Phi(x)])\leq\varepsilon\,$,
    and $\mathfrak{g}(H(t^{\prime},x))\leq\mathfrak{g}(H(t,x))\,.$
  3. (3)
    The Simon–Smith family $\{H(t,x)\}_{t\in[0,1]}$ is a pinch-off process.
Moreover, if $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ consists of punctate surfaces of small area, we can perform pinch-off processes to obtain a Simon–Smith family of genus 0. This can be viewed as a special case of the previous proposition by taking $g_{0}=0$ and $\mathscr{S}_{0}=\{\emptyset\}$.
Proposition 6.2 (Pinch-off proposition II).
Let $(M,\mathbf{g})$ be an orientable closed Riemannian $3$-manifold, $X$ be a cubical subcomplex of $I(m,k)$, $q\in\mathbb{N}$ and $\varepsilon\in(0,1)$. Then there exists $\delta=\delta(M,\mathbf{g},q,\varepsilon)\in(0,\varepsilon)$ with the following property. If $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ is a Simon–Smith family of genus $\leq g_{1}$ with $N(P_{\Phi})+g_{1}\leq q\,$, and $\sup_{x\in X}\mathcal{H}^{2}(\Phi(x))\leq\delta\,$, then there exists a Simon–Smith family $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}(M)$ of genus $\leq g_{1}$ such that for any $x\in X$:
  1. (1)
    $H(0,x)=\Phi(x)$, $\mathfrak{g}(H(1,x))=0$.
  2. (2)
    For any $0\leq t\leq t^{\prime}\leq 1$, $\mathcal{H}^{2}(H(t,x))\leq\mathcal{H}^{2}(\Phi(x))+\varepsilon$ and $\mathfrak{g}(H(t^{\prime},x))\leq\mathfrak{g}(H(t,x))\,.$
  3. (3)
    The Simon–Smith family $\{H(t,x)\}_{t\in[0,1]}$ is a pinch-off process.
6.1. Local deformations
Recall the following local pinch-off processes, Pinching and Shrinking, introduced in [chuLi2024fiveTori].
Lemma 6.3 (Local Deformation A: Pinching).
In an orientable closed Riemannian $3$-manifold $(M,\mathbf{g})$, there exists a constant $C_{\mathbf{g}}$ with the following property. Let $\Lambda>0$, $r\in\left(0,\mathrm{injrad}(M,\mathbf{g})/4\right)$ and $p\in M$. Suppose that $\Phi\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ is a Simon–Smith family and $x_{0}\in X$ so that
$\mathcal{H}^{2}(\Phi(x_{0})\cap A(p;r,2r))\leq\Lambda r^{2}\,.$
Then for any $\eta>0$, there exist $s\in(r,2r)$, $\zeta\in(0,\min\{\Lambda r,s-r,2r-s\}/5)$, an integer $K\geq 0$, a neighborhood $O_{x_{0}}$ of $x_{0}$ in $X$, and a Simon–Smith family $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times O_{x_{0}}\to\mathcal{S}(M)\,$, with the following properties.
  1. (1)
    For any $y\in O_{x_{0}}$, $\Sigma_{y}\mathrel{\mathop{\ordinarycolon}}=\Phi(y)\cap A(p;s-5\zeta,s+5\zeta)$ is a smooth surface, varying smoothly in $y$.
  2. (2)
    For any $s^{\prime}\in[s-4\zeta,s+4\zeta]$ and $y\in O_{x_{0}}$, $\Phi(y)$ intersects $\partial B(p,s^{\prime})$ transversally, and
    $\Phi(y)\cap\partial B(p,s^{\prime})=\mathrel{\mathop{\ordinarycolon}}\bigsqcup_{i=1}^{K}\gamma_{i}(s^{\prime},y)\,,\quad\sum_{i=1}^{K}\mathcal{H}^{1}(\gamma_{i}(s^{\prime},y))\leq 2\Lambda r\,$,
    where each $\gamma_{i}(s^{\prime},y)$ is a simple loop.
If $K=0$, then $H(t,y)\equiv\Phi(y)$ holds true for any $y\in O_{x_{0}}$ and $t\in[0,1]$; and if $K\geq 1$, then:
  1. (3)
    For any $y\in O_{x_{0}}$, $H(0,y)=\Phi(y)$.
  2. (4)
    There exist $0=t_{0}<t_{1}<t_{2}<\dots<t_{K}<t_{K+1}=1$ such that for any $0\leq i\leq K$, $t\in(t_{i},t_{i+1})$ and $y\in O_{x_{0}}$,
    $\Sigma_{t,y}\mathrel{\mathop{\ordinarycolon}}=H(t,y)\cap A(p;s-5\zeta,s+5\zeta)$
    is a smooth surface with boundary, which is diffeomorphic to $\Sigma_{y}\setminus\bigsqcup_{1\leq j\leq i}\gamma_{j}(s,y)$ attached with $2i$ disks. Moreover, $H(1,y)\cap A(p;s-5\zeta,s+5\zeta)$ is also a smooth surface and
    $H(1,y)\cap A(p;s-\zeta,s+\zeta)=\emptyset\,.$
  3. (5)
    For any $i\in\{1,2,\dots,K\}$ and $y\in O_{x_{0}}$, near $t_{i}$, $H(t,y)$ is a surgery process via pinching the cylinder $\bigcup_{s^{\prime}\in(s-2\zeta,s+2\zeta)}\gamma_{i}(s^{\prime},y)$.
  4. (6)
    For any $y\in O_{x_{0}}$ and $t\in[0,1]$, we have
    $H(t,y)\setminus A(p;s-3\zeta,s+3\zeta)$ $=\Phi(y)\setminus A(p;s-3\zeta,s+3\zeta)\,$,
    $\mathcal{H}^{2}(\Phi(y)\cap A(p;s-3\zeta,s+3\zeta))-\eta$ $\leq\mathcal{H}^{2}(H(t,y)\cap A(p;s-3\zeta,s+3\zeta))$
    $\leq\mathcal{H}^{2}(\Phi(y)\cap A(p;s-3\zeta,s+3\zeta))+C_{\mathbf{g}}\Lambda^{2}r^{2}\,$,
    and $\mathcal{F}([\Phi(y))],[H(t,y)])\leq\eta\,.$
Remark 6.4.
For every $y\in O_{x_{0}}$, by (5), $\{H(t,y)\}_{t\in[0,1]}$ is a pinch-off process.
Sketch of proof.
The only difference between [chuLi2024fiveTori, Lemma 8.3] and this lemma is the inequalities in  (6), where we additionally require a lower bound on $\mathcal{H}^{2}(H(t,y))$ and that $\Phi(y)]$ and $[H(t,y)]$ are close in the flat norm. The same proof applies, and one observes that by choosing $\zeta$ sufficiently small, both inequalities are automatically satisfied, since $H(t,y)$ is obtained by pinching cylinders of height at most $\zeta$. ∎
Lemma 6.5 (Local Deformation B: Shrinking, [chuLi2024fiveTori, Lemma 8.4]).
In an orientable closed Riemannian $3$-manifold $(M,\mathbf{g})$, suppose that $p\in M$ and $0<r^{-}<r^{+}<\mathrm{injrad}(M,\mathbf{g})/4$. Then there exists a smooth one-parameter family of maps $\{\mathcal{R}_{t}\mathrel{\mathop{\ordinarycolon}}M\to M\}_{t\in[0,1]}$ such that:
  1. (1)
    $\mathcal{R}_{t}=\mathrm{id}$ for any $t\in[0,1/4]$; $\mathcal{R}_{t}(B(p,r^{-}))=\{p\}$ for any $t\in[1/2,1]$.
  2. (2)
    $\mathcal{R}_{t}$ is a diffeomorphism for any $t\in[0,1/2)$.
  3. (3)
    $\mathcal{R}_{t}|_{M\setminus B(p,r^{+})}=\mathrm{id}$ and $\mathcal{R}_{t}(B(p,r^{\pm}))\subset B(p,r^{\pm})$ for any $t\in[0,1]$.
  4. (4)
    $\mathcal{R}_{t}|_{B(p,r^{-})}$ is $1$-Lipschitz for any $t\in[0,1]$.
Lemma 6.6 (Local Deformation C: Successive shrinkings, [chuLi2024fiveTori, Lemma 8.5]).
[…]
  1. (1)
    There exists a permutation $\sigma\in\mathfrak{S}_{N}$ such that for any $1\leq i<j\leq N$, either $B_{\sigma(i)}^{+}\cap B_{\sigma(j)}^{+}=\emptyset$, or $B_{\sigma(i)}^{+}\subset B_{\sigma(j)}^{-}$. In this case, we call such a $\sigma$ admissible.
  2. (2)
    For any admissible $\sigma$, the map
    $\mathcal{R}^{\sigma(N)}_{t}\circ\mathcal{R}^{\sigma(N-1)}_{t}\circ\cdots\circ\mathcal{R}^{\sigma(2)}_{t}\circ\mathcal{R}^{\sigma(1)}_{t}\mathrel{\mathop{\ordinarycolon}}M\to M$
    is $1$-Lipschitz on every connected component of $M\setminus\bigcup_{j=1}^{N}A_{j}$; and it is independent of the choice of admissible $\sigma$.
6.2. Combinatorial arguments
Recall the following combinatorial arguments in [chuLi2024fiveTori], inspired by [Pit81, Lemma 4.8, Proposition 4.9] and [De_Lellis_Jusuf_2018_min_max, Lemma 5.7].
Lemma 6.7 (Combinatorial Argument I, [chuLi2024fiveTori, §8.3]).
For any positive integers $m\in\mathbb{N}^{+}$ and $q\in\mathbb{N}^{+}$, there exists $N=N(m,q)\in\mathbb{N}^{+}$ with the following property.
Given any $k\in\mathbb{N}^{+}$ and a cubical complex $I(m,k)$, suppose that $\{\mathcal{F}_{\sigma}\}_{\sigma\in I(m,k)}$ is a family of collections of open sets in a closed manifold $M$ assigned to each cell $\sigma$ of $I(m,k)$, where each collection $\mathcal{F}_{\sigma}$ has the form
$\mathcal{F}_{\sigma}=(\mathcal{O}_{\sigma,1},\mathcal{O}_{\sigma,2},\dots,\mathcal{O}_{\sigma,q})$
and each $\mathcal{O}_{\sigma,i}=\{U^{1}_{\sigma,i},U^{2}_{\sigma,i},\dots,U^{N}_{\sigma,i}\}$ satisfies
(6.1) $\operatorname{dist}(U^{r}_{\sigma,i},U^{s}_{\sigma,i})\geq 2\min\{\mathrm{diam}(U^{r}_{\sigma,i}),\mathrm{diam}(U^{s}_{\sigma,i})\}\,$,
for all $\sigma\in I(m,k)$, $i\in\{1,2,\dots,q\}$ and $r,s\in\{1,2,\dots,N\}$ with $r\neq s$. Then we can extract a family of open sets
$\{(U_{\sigma,1},U_{\sigma,2},\dots,U_{\sigma,q})\}_{\sigma\in I(m,k)}$
where $U_{\sigma,i}\in\mathcal{O}_{\sigma,i}$ such that $\operatorname{dist}(U_{\sigma,i}\cap U_{\tau,j})>0$ whenever $\sigma,\tau\in I(m,k)$, $(\sigma,i)\neq(\tau,j)$ and $\sigma,\tau$ are faces of a common cell $\gamma$ of $I(m,k)$.
Moreover, this is also true if the number of collections in each $\mathcal{F}_{\sigma}$ is no greater than $q$.
Lemma 6.8 (Combinatorial argument II, [chuLi2024fiveTori, §8.3]).
In a closed Riemannian manifold $(M,\mathbf{g})$, given any integers $q,N\in\mathbb{N}^{+}$ and any $R_{0}\in(0,\mathrm{injrad}(M,\mathbf{g})/4)$, if $P$ is a finite set of at most $q$ points, then there exists a radius $R\in(5^{-2Nq^{2}}R_{0},5^{-2N}R_{0})$ such that for every $p\in P$, we have $A(p;R,5^{2N}R)\cap P=\emptyset\,.$
6.3. Proof of Proposition 6.1 and Proposition 6.2
The proof will proceed in three steps.
Step 1. Set up
Let $C_{\mathbf{g}}$ be as in Lemma 6.3 and $N\mathrel{\mathop{\ordinarycolon}}=N(m,q)$ as in Lemma 6.7.
In Proposition 6.1, since $\mathscr{S}_{g_{0}}$ is a compact set of embedded smooth closed surfaces, without loss of generality, we assume that $\varepsilon<\min\left\{\mathrm{injrad}(M,\mathbf{g}),1\right\}\,$, such that for any $r\in(0,\varepsilon/1000)$, we have $\sup_{p\in M}\mathcal{H}^{3}(B(p,2r))$ and
(6.2) $\sup_{\Sigma\in\mathscr{S}_{g_{0}},\ p\in M}\left\{\mathcal{H}^{2}\left(\Sigma\cap B(p,2r)\right)\right\}\leq 50r^{2}\,.$
We set
$\varepsilon_{1}\mathrel{\mathop{\ordinarycolon}}=\frac{\varepsilon}{10000Nq\cdot 3^{m}\max(C_{\mathbf{g}},1)}\quad\textrm{ and }\quad\varepsilon_{2}\mathrel{\mathop{\ordinarycolon}}=\frac{\varepsilon_{1}}{2(3N)^{q}q^{2}\cdot 5^{2N}}\,.$
By (6.2), we choose $\delta_{0}=\delta_{0}(\mathscr{S}_{g_{0}})>0$ such that for any $S\in\mathcal{S}(M)$, if $\mathbf{F}(S,\Sigma)<\delta_{0}$ for some $\Sigma\in\mathscr{S}_{g_{0}}$, then for any $r\in(\varepsilon_{2},\varepsilon_{1})$,
(6.3) $\sup_{p\in M}\left\{\mathcal{H}^{2}\left(S\cap B(p,2r)\right)\right\}\leq 100r^{2}\,$,
and in particular,
(6.4) $\sup_{p\in M}\left\{\mathcal{H}^{2}\left(S\cap A(p;r,2r)\right)\right\}\leq 100r^{2}\,.$
Note that these inequalities also hold for Proposition 6.2.
In the following, we choose $\delta_{1}\mathrel{\mathop{\ordinarycolon}}=\delta(M,\mathbf{g},q,N,2(3N)^{q}\varepsilon_{2})$ as defined in Lemma 5.13 and $\delta_{2}\mathrel{\mathop{\ordinarycolon}}=\delta(M,\mathbf{g},\mathscr{S}_{g_{0}},q,N,2(3N)^{q}\varepsilon_{2})$ in Lemma 5.14. Set $\delta\mathrel{\mathop{\ordinarycolon}}=\min\{\delta_{0},\delta_{1},\delta_{2}\}\,.$
For a Simon–Smith family $\Phi$ as in Proposition 6.1 or Proposition 6.2, by Lemma 5.13 or Lemma 5.14 with $N=1$ therein, for any $x\in X$, there exist $r_{x}\in(\varepsilon_{2},2(3N)^{q}\varepsilon_{2})$ and a finite set of balls $\{B(p_{x,i},r_{x})\}^{q_{x}}_{i=1}$ such that
  1. (i)
    $q_{x}\leq q$;
  2. (ii)
    $\Phi(x)\setminus\bigcup^{q_{x}}_{i=1}B(p_{x,i},r_{x})$ is a smooth surface of genus $g_{0}$ with boundary.
Note that when $\Phi$ is a family as in Proposition 6.2, $g_{0}=0$.
We write $\widetilde{P}(x)\mathrel{\mathop{\ordinarycolon}}=\{p_{x,i}\}^{q_{x}}_{i=1}$ and for every $l\in\{1,2,\dots,N\}$, $r_{x,l}\mathrel{\mathop{\ordinarycolon}}=5^{2(l-1)}r_{x}\,.$ By (6.3) and (6.4), for every $1\leq l\leq N$, $x\in X$ and $p\in\widetilde{P}(x)$, since $r_{x,l}\in(\varepsilon_{2},\varepsilon_{1})$, we have
(6.5) $\mathcal{H}^{2}\left(\Phi(x)\cap B(p,2r_{x,l})\right)\leq 100r_{x,l}^{2}\,$,
and
(6.6) $\mathcal{H}^{2}\left(\Phi(x)\cap A(p;r_{x,l},2r_{x,l})\right)\leq 100r_{x,l}^{2}\,.$
Thus, applying Lemma 6.3 with $\eta=\varepsilon/2$ to $\Phi(x)$ in $A(p;r_{x,l},2r_{x,l})$ and $\Lambda=100$, we obtain the following data:
  1. (1)
    $s_{x,l,p}\in(r_{x,l},2r_{x,l})$,
  2. (2)
    $\zeta_{x,l,p}\in(0,\min\{100r_{x,l},s_{x,l}-r_{x,l},2r_{x,l}-s_{x,l}\}/5)$,
  3. (3)
    a neighborhood $O_{x,l,p}\subset X$ of $x$,
  4. (4)
    and a Simon–Smith family $H_{x,l,p}\mathrel{\mathop{\ordinarycolon}}[0,1]\times O_{x,l,p}\to\mathcal{S}(M)\,.$
In addition, for each $x\in X$, we fix a neighborhood of $x$
$O_{x}\subset\bigcap_{l\in\{1,2,\dots,N\},\ p\in\widetilde{P}(x)}O_{x,l,p}\,$,
such that $y\mapsto\Phi(y)\setminus\bigcup_{p\in\widetilde{P}(x)}B(p,r_{x})$ is continuous in the smooth topology for $y\in O_{x}$.
Step 2. Refinement
Since $X$ is a compact cubical subcomplex of $I(m,k)$, its open cover $\{O_{x}\}_{x\in X}$ has a finite subcover $\{O_{x_{j}}\}$.
Then we refine $X$ and assign to each cell $\sigma$ of $X$ $x_{\sigma}\in\{x_{j}\}$ such that: Every cell $\tau$ of $X$ is a subset of $O_{x_{\sigma}}$ provided that $\sigma$ and $\tau$ are faces of some cell $\gamma$ of $X$. Note that for each $\sigma$, the number of such $\tau$ is no more than $5^{m}$. By Lemma 6.7 with
$\mathcal{F}_{\sigma}=(\{A(p;r_{x_{\sigma},i},2r_{x_{\sigma},i})\}^{N}_{i=1})_{p\in\widetilde{P}(x_{\sigma})}\,$,
we obtain, associated with each $\sigma$, a collection of annuli denoted by
$A_{\sigma,p}\mathrel{\mathop{\ordinarycolon}}=A(p,r_{\sigma,p},2r_{\sigma,p})_{p\in\widetilde{P}(x_{\sigma})}$
such that whenever $\sigma$ and $\tau$ are faces of some cell $\gamma$ of $X$ and for any $p\in\widetilde{P}(x_{\sigma})$ and $p^{\prime}\in\widetilde{P}(x_{\tau})$, we have $A_{\sigma,p}\cap A_{\tau,p^{\prime}}=\emptyset\,$, unless $(\sigma,p)=(\tau,p^{\prime})$.
Step 3. Construction of $H$
For each cell $\sigma$ and $p\in\widetilde{P}(x_{\sigma})$, for convenience, we denote the data associated with each $A_{\sigma,p}$, constructed at the end of Step 1, as follows:
  1. (1)
    $s_{\sigma,p}\in(r_{\sigma,p},2r_{\sigma,p})$,
  2. (2)
    $\zeta_{\sigma,p}\in(0,\min\{s_{\sigma,p}-r_{\sigma,p},2r_{\sigma,p}-s_{\sigma,p}\}/5)$,
  3. (3)
    a neighborhood $O_{\sigma}\mathrel{\mathop{\ordinarycolon}}=O_{x_{\sigma}}$,
  4. (4)
    and a Simon–Smith family $H_{\sigma,p}\mathrel{\mathop{\ordinarycolon}}[0,1]\times O_{\sigma}\to\mathcal{S}(M)\,.$
We also denote
$r_{\sigma,p}^{\pm}\mathrel{\mathop{\ordinarycolon}}=s_{\sigma,p}\pm\zeta_{\sigma,p}\,,\quad B_{\sigma,p}^{\pm}\mathrel{\mathop{\ordinarycolon}}=B_{r_{\sigma,p}^{\pm}}(p)\,,\quad\hat{A}_{\sigma,p}\mathrel{\mathop{\ordinarycolon}}=\overline{B_{\sigma,p}^{+}\setminus B_{\sigma,p}^{-}}\subset A_{\sigma,p}\,.$
After the refinement in Step 2, suppose that $X$ is a cubical subcomplex of the unit cube $I(m,k^{\prime})$. For each cell $\sigma$ of $X$ and for each $x\in X$, we define $\mathbf{d}_{\sigma}(x)\mathrel{\mathop{\ordinarycolon}}=\min\left\{{2\|x-c_{\sigma}\|_{\ell_{\infty}}}/{3^{-k^{\prime}}},1\right\}\,$, to be the normalized $l^{\infty}$ distant function to $\sigma$, where $c_{\sigma}$ is the center of $\sigma$.
For each $x\in X$, we define $H(t,x)$, $0\leq t\leq 1/2$, using pinching surgeries as described in Lemma 6.3: For every cell $\sigma$ of $X$ and every $p\in\widetilde{P}(x_{\sigma})$,
$H(t,x)\cap A_{\sigma,p}=H_{\sigma,p}\left(\min\{8t(1-\mathbf{d}_{\sigma}(x)),1\},x\right)\cap A_{\sigma,p}$
and $H(t,x)=\Phi(x)$ outside $\bigcup_{\sigma\in X,p\in\widetilde{P}(x_{\sigma})}A_{\sigma,p}$.
The map has the following properties for every $x\in X$:
  • (i)
    The set $Z_{x}\mathrel{\mathop{\ordinarycolon}}=\{\sigma\mathrel{\mathop{\ordinarycolon}}\mathbf{d}_{\sigma}(x)<1\}$ satisfies that every cell $\sigma$ in $Z_{x}$ is a face of the cell $\gamma$, which is the smallest cell containing $x$. By Step 2, the corresponding annuli $\{A_{\sigma,p}\}_{\sigma\in Z_{x},p\in\widetilde{P}(x_{\sigma})}$ are pairwise disjoint, and the number is no greater than $3^{m}\cdot q$. In particular, the map $H(t,x)$ is well-defined.
  • (ii)
    By Lemma 6.3 (4) and (5), $t\mapsto\mathfrak{g}(H(t,x))$ is non-increasing and $t\mapsto H(t,x)$ is a pinch-off process.
  • (iii)
    By (6.6) and Lemma 6.3 (6), for every $t\in[0,1/2]$,
    $\mathcal{H}^{2}(H(t,x))$ $\leq\mathcal{H}^{2}(\Phi(x))+(3^{m}\cdot q)\cdot C_{\mathbf{g}}100^{2}\varepsilon^{2}_{1}\leq\mathcal{H}^{2}(\Phi(x))+\varepsilon/2\,$,
    $\mathcal{H}^{2}(H(t,x))\geq\mathcal{H}^{2}(\Phi(x))-\varepsilon/2\,$,
    and $\mathcal{F}([H(t,x)],[\Phi(x)])\leq\varepsilon/2\,$; in fact, for every $x\in X$, $\sigma\in Z_{x}$ and $p\in\widetilde{P}(x)$,
    (6.7) $\mathcal{H}^{2}(H(t,x)\cap B^{+}_{\sigma,p})\leq 100(r^{+}_{\sigma,p})^{2}+\varepsilon/(100q\cdot 3^{m})<\varepsilon/(10q\cdot 3^{m})\,.$
  • (iv)
    If $\mathbf{d}_{\sigma}(x)\leq 3/4$, then for each $p\in\widetilde{P}(x_{\sigma})$, in $A_{\sigma,p}$, $H(1/2,x)=H_{\sigma,p}(1,x)$, and thus by Lemma 6.3 (4), we know that $H(1/2,x)\cap\hat{A}_{\sigma,p}=\emptyset\,.$
Finally, for each $t\in[1/2,1]$ and each $x\in X$, we define $H(t,x)$ using the shrinking process as described in Lemma 6.5: For each cell $\sigma$ of $X$ and $p\in\widetilde{P}(x_{\sigma})$, we let $\mathcal{R}_{(\sigma,p),t}\mathrel{\mathop{\ordinarycolon}}S^{3}\to S^{3}$ be the one-parameter family of shrinking deformation from Lemma 6.5 with $r^{-}=r^{-}_{\sigma,p}$ and $r^{+}=r^{+}_{\sigma,p}$. Then for each $x\in X$, we label $\{(\sigma,p)\mid\sigma\in Z_{x},\ p\in P^{\prime}(x_{\sigma})\}$ as $\{(\sigma_{i},p_{i})\}_{1\leq i\leq N^{\prime}}$ such that for every $1\leq i<j\leq N^{\prime}$, either $B_{\sigma_{i},p_{i}}^{+}\cap B_{\sigma_{j},p_{j}}^{+}=\emptyset$, or $B_{\sigma_{i},p_{i}}^{+}\subset B_{\sigma_{j},p_{j}}^{-}$, where $N^{\prime}\leq 3^{m}\cdot q$. The existence of such labeling follows from Lemma 6.6 (1). Then, we denote for simplicity,
$\hat{\mathcal{R}}^{(i)}_{t,x}\mathrel{\mathop{\ordinarycolon}}=\mathcal{R}_{(\sigma_{i},p_{i}),(2t-1)(1-\mathbf{d}_{\sigma_{i}}(x))}\,.$
Note that
  • By Lemma 6.5 (1), if $\mathbf{d}_{\sigma_{i}}(x)\geq 3/4$ or $t\in[1/2,5/8]$, then $\hat{\mathcal{R}}^{i}_{t,x}=\mathrm{id}$; when $\mathbf{d}_{\sigma_{i}}(x)<3/4$, by the construction above, $H(1/2,x)\cap\hat{A}_{\sigma_{i},p_{i}}=\emptyset$.
  • if $\mathbf{d}_{\sigma_{i}}(x)\leq 1/2$, then $\hat{\mathcal{R}}^{(i)}_{1,x}(B_{\sigma_{i},p_{i}}^{-})\subset\{p_{i}\}$.
Now for $t\in[1/2,1]$ and $x\in X$, we define
$H(t,x)\mathrel{\mathop{\ordinarycolon}}=\hat{\mathcal{R}}^{(N^{\prime})}_{t,x}\circ\cdots\circ\hat{\mathcal{R}}^{(2)}_{t,x}\circ\hat{\mathcal{R}}^{(1)}_{t,x}\left(H(1/2,x)\right)\,.$
Intuitively, $\{H(t,x)\}_{t\in[1/2,1]}$ is obtained by shrinking some connected components of $H(1/2,x)$ to points. Therefore, for every $x\in X$, $t\mapsto\mathfrak{g}(H(t,x))$ is non-increasing, and $\{H(t,x)\}_{t\in[1/2,1]}$ is a pinch-off process. This confirms statement (3) of both propositions.
By Lemma 6.5 (3), (4) and Lemma 6.6 (2), we see that $\mathcal{H}^{2}(H(t,x))$ is non-increasing for $t\in[1/2,1]$. In particular, for every $t\in[1/2,1]$,
$\mathcal{H}^{2}(H(t,x))\leq\mathcal{H}^{2}(\Phi(x))+\varepsilon\,.$
Also, by the estimate (6.7), one also have
$\mathcal{H}^{2}(H(t,x))\geq\mathcal{H}^{2}(H(1/2,x))-\varepsilon/(2\cdot 3^{q}\cdot m)\cdot N^{\prime}\geq\mathcal{H}^{2}(\Phi(x))-\varepsilon\,$,
and
$\mathcal{F}([H(t,x)],[\Phi(x)])\leq\mathcal{F}([H(t,x)],[H(1/2,x)])+\mathcal{F}([H(1/2,x)],[\Phi(x)])\leq\varepsilon/2+N^{\prime}(\varepsilon_{1})^{2}\leq\varepsilon\,$,
due to the essence of shrinking $N^{\prime}$ balls. This confirms statement (2) of both propositions.
Furthermore, by the construction above, $\mathfrak{g}(H(1,x))=g_{0}$. This verifies statement (1) of both propositions and completes the proof.
6.4. Proof of Theorem 2.21 and Proposition 2.22
Equipped with the ingredients above, the proof of Theorem 2.21 proceeds similarly to [chuLi2024fiveTori, §8.5]: The details are included in Appendix D. As for the proof of Proposition 2.22, we suppose that $X=\operatorname{dmn}(\Phi)$ is a cubical subcomplex of $I(m,k)$. Let $q=N(P_{\Phi})+g$, and $\delta=\delta(\varepsilon/2,q,m)$ as in Proposition 6.2. Since $\mathbf{L}(\Lambda(\Phi))=0$, there exists $\Phi^{\prime}\in\Lambda(\Phi)$ with $\sup_{x\in X}\mathcal{H}^{2}(\Phi^{\prime}(x))<\delta\,.$ Then, the proposition follows from applying Proposition 6.2 on $\Phi^{\prime}$.
6.5. Proof of Proposition 2.23
Following the proof of Proposition 6.1 and Proposition 6.2, we can similarly prove Proposition 2.23 in three steps. This argument is much simpler, as it does not involve area control, and the family is “continuous in the Simon–Smith sense” rather than merely with respect to the $\mathbf{F}$-metric.
Step 1. Set up
Let $q=N(P_{\Phi})$ and $m$ be the ambient dimension of $X$, and $N\mathrel{\mathop{\ordinarycolon}}=N(m,q)$ as in Lemma 6.7. Fix $\varepsilon_{1}\mathrel{\mathop{\ordinarycolon}}={\min\{\mathrm{injrad}(M,\mathbf{g}),1\}}/{3^{m}}$ and $\varepsilon_{2}\mathrel{\mathop{\ordinarycolon}}={\varepsilon_{1}}/({3^{q}5^{2qN}5^{2N}})$.
For every $x\in K$, let $P_{\Phi}(x)$ denote a punctate set associated with $\Phi(x)$. By Sard’s theorem and Lemma 5.11 with $n=5^{2N}$, there exist $q_{x}\leq q$, $\widetilde{P}(x)\mathrel{\mathop{\ordinarycolon}}=\{p_{x,i}\}^{q_{x}}_{i=1}\subset P_{\Phi}(x)$ and $r_{x}\in(\varepsilon_{2},3^{q}5^{2qN}\varepsilon_{2})$ such that:
  1. (1)
    $\{\overline{B(p_{x,i},5^{2N}r_{x})}\}$ are disjoint from each other;
  2. (2)
    For every $k\in\{1,2,\cdots,5^{2N}\}$, $\Phi(x)\setminus\bigcup^{q_{x}}_{i=1}B(p_{x,i},kr_{x})$ is an embedded smooth surface with boundary and genus at most $g_{0}$.
For every $l\in\{1,2,\cdots,N\}$, we write $r_{x,l}\mathrel{\mathop{\ordinarycolon}}=5^{2(l-1)}r_{x}\,.$
For every $x\in K$, $p\in\widetilde{P}(x)$ and $l\in\{1,2,\cdots,N\}$, applying Lemma 6.3 with $\eta=1$ to $\Phi(x)$ in $A(p;r_{x,l},2r_{x,l})$ with suitable $\Lambda$, we obtain the following data:
  1. (1)
    $s_{x,l,p}\in(r_{x,l},2r_{x,l})$,
  2. (2)
    $\zeta_{x,l,p}\in(0,\min\{100r_{x,l},s_{x,l}-r_{x,l},2r_{x,l}-s_{x,l}\}/5)$,
  3. (3)
    a neighborhood $O_{x,l,p}\subset X$ of $x$,
  4. (4)
    and a Simon–Smith family $H_{x,l,p}\mathrel{\mathop{\ordinarycolon}}[0,1]\times O_{x,l,p}\to\mathcal{S}(M)\,.$
In addition, for each $x\in K$, we fix a neighborhood of $x$
$O_{x}\subset\bigcap_{l\in\{1,2,\dots,N\},\ p\in\widetilde{P}(x)}O_{x,l,p}\,$,
such that $y\mapsto\Phi(y)\setminus\bigcup_{p\in\widetilde{P}(x)}B(p,r_{x})$ is continuous in the smooth topology for $y\in O_{x}$; in particular, each $\Phi(y)\setminus\bigcup_{p\in\widetilde{P}(x)}B(p,r_{x})$ has genus $\leq g_{0}$.
Step 2. Refinement
Since $K$ is compact and $\{O_{x}\}_{x\in K}$ is an open cover, there exists a finite subcover $\{O_{x_{j}}\}_{j\in J}\,.$ Up to refinement of $X$, there exists a cubical subcomplex $Z$ of $X$ such that
$K\subset\operatorname{int}(Z)\subset Z\subset\bigcup_{j\in J}O_{x_{j}}\,.$
For simplicity, we denote $\Phi|_{Z}$ by $\Psi$, $O_{x_{j}}\cap Z$ by $U_{x_{j}}$ for every $j\in J$.
Then we refine $Z$ and assign to each cell $\sigma$ of $Z$ an $x_{\sigma}\in\{x_{j}\}_{j\in J}$ such that: Every cell $\tau$ of $Z$ is a subset of $U_{x_{\sigma}}$ provided that $\sigma$ and $\tau$ are faces of a common cell $\gamma$ of $X$. By Lemma 6.7 with
$\mathcal{F}_{\sigma}=\left(\{A(p;r_{x_{\sigma},l},2r_{x_{\sigma},l}\}^{N}_{l=1}\right)_{p\in\widetilde{P}(x_{\sigma})}\,$,
we obtain, associated with each cell $\sigma$ in $Z$, a collection of annuli denoted by
$\left(A_{\sigma,p}\mathrel{\mathop{\ordinarycolon}}=A(p;r_{\sigma,p},2r_{\sigma,p})\right)_{p\in\widetilde{P}(x_{\sigma})}$
such that whenever $\sigma$ and $\tau$ are faces of a common cell of $Z$ and for any $p\in\widetilde{P}(x_{\sigma})$ and $p^{\prime}\in\widetilde{P}(x_{\tau})$, we have $A_{\sigma,p}\cap A_{\tau,p^{\prime}}=\emptyset,$ unless $(\sigma,p)=(\tau,p^{\prime})$.
Step 3. Construction of $H$
Following the same construction as in Step 3 of the proof of Proposition 6.1 and Proposition 6.2, without keeping track of the area estimates, one obtain $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times Z\to\mathcal{S}(M)$ such that $\Psi^{\prime}\mathrel{\mathop{\ordinarycolon}}=H(1)$ is a Simon–Smith family of genus $\leq g_{0}$, and $H$ is a deformation via pinch-off processes from $\Psi$ to $\Psi^{\prime}$.
Appendix A Some geometric measure theory results
In an $N$-dimensional closed Riemannian manifold $(M,\mathbf{g})$, let $G_{n}(M)$ denote its (unoriented) Grassmanian $n$-plane bundle. Given two $n$-dimensional varifolds $V,W\in\mathcal{V}(M)$, their $\mathbf{F}$-distance is defined by
$\mathbf{F}(V,W)\mathrel{\mathop{\ordinarycolon}}=\sup\{V(f)-W(f)\mathrel{\mathop{\ordinarycolon}}f\in C(G_{n}(M)),|f|\leq 1,\operatorname{Lip}(f)\leq 1\}\,;$
for a Borel set $B\subset M$, we also define $\mathbf{F}_{B}(V,W)\mathrel{\mathop{\ordinarycolon}}=\mathbf{F}(V\llcorner G_{n}(B),W\llcorner G_{n}(B))\,$, where $G_{n}(B)$ denotes the restriction of $G_{n}(M)$ to the base $B$.
Let us first recall a result by Pitts relating the $\mathbf{F}$-norm and the $\mathcal{F}$-norm on the space of modulo $2$ integral currents.
Lemma A.1 ([Pit81, 68].).
In a closed $N$-dimensional Riemannian manifold $(M,\mathbf{g})$, given a sequence of modulo $2$ integral currents $\{S_{i}\}^{\infty}_{i=1}$ in $\mathbf{I}_{n}(M;\mathbb{Z}_{2})$ and $S\in\mathbf{I}_{n}(M;\mathbb{Z}_{2})$, we have $\lim_{i\to\infty}\mathbf{F}(S_{i},S)=0\,$, if and only if
$\lim_{i\to\infty}\mathcal{F}(S_{i},S)=0\quad\textrm{ and }\quad\lim_{i\to\infty}\mathbf{M}(S_{i})=\mathbf{M}(S)\,.$
A direct consequence of this lemma is the following quantitative corollary.
Corollary A.2.
In a closed $N$-dimensional Riemannian manifold $(M,\mathbf{g})$, for any $S\in\mathcal{S}(M)$ and $\varepsilon>0$, there exists $\delta=\delta(M,\mathbf{g},S,\varepsilon)>0$ such that for any $S^{\prime}\in\mathcal{S}(M)$,
  • if $\mathbf{F}(S^{\prime},S)<\delta\,$, then $\mathcal{F}([S^{\prime}],[S])<\varepsilon$ and $|\mathcal{H}^{2}(S^{\prime})-\mathcal{H}^{2}(S)|<\varepsilon$,
  • if $\mathcal{F}([S^{\prime}],[S])<\delta$ and $|\mathcal{H}^{2}(S^{\prime})-\mathcal{H}^{2}(S)|<\delta\,$ then $\mathbf{F}(S^{\prime},S)<\varepsilon\,.$
Proof.
We argue by contradiction. Suppose there exists a sequence $\{S_{i}\}^{\infty}_{i=1}$ in $\mathcal{S}(M)$ such that $\mathbf{F}(S_{i},S)<\frac{1}{i}$, but $\mathcal{F}([S_{i}],[S])\geq\varepsilon$. Then we have $\lim_{i\to\infty}\mathbf{F}(S_{i},S)=0$ and $\liminf_{i\to\infty}\mathcal{F}([S_{i}],[S])>\varepsilon$, contradicting Lemma A.1. The other direction can be proved similarly. ∎
Lemma A.3.
In an $N$-dimensional closed Riemannian manifold $(M,\mathbf{g})$, for any $V\in\mathcal{V}_{n}(M)$, $\varepsilon>0$ and any family of Borel sets $\{B_{\alpha}\}_{\alpha\in I}$ such that
$\lim_{r\to 0}\sup_{\alpha\in I}V(\overline{N_{r}(\partial B_{\alpha})})=0\,$,
where $\overline{N_{r}(\partial B_{\alpha})}=\{x\in M\mathrel{\mathop{\ordinarycolon}}\operatorname{dist}_{M}(x,\partial B_{\alpha})\leq r\}$, there exists a constant $\delta=\delta(M,\mathbf{g},V,\varepsilon,\{B_{\alpha}\}_{\alpha\in I})$ satisfying the following property: For any $W\in\mathcal{V}_{n}(M)$, if $\mathbf{F}(V,W)<\delta$, then $\sup_{\alpha\in I}\mathbf{F}_{B_{\alpha}}(V,W)\leq\varepsilon\,.$
Proof.
Since $\lim_{r\to 0}\sup_{\alpha\in I}V(\overline{N_{r}(\partial B_{\alpha})})=0$, we can pick $r_{0}\in(0,1)$ such that $\sup_{\alpha\in I}V(\overline{N_{r_{0}}(\partial B_{\alpha})})<\varepsilon/10$. We set $\delta\mathrel{\mathop{\ordinarycolon}}=r_{0}\varepsilon/1000\,.$
Firstly, we claim that
$\sup_{\alpha\in I}W(\overline{N_{r_{0}/2}(\partial B_{\alpha})})<\varepsilon/5\,.$
Suppose by contradiction that there exists $\alpha_{0}\in I$ such that $W(\overline{N_{r_{0}/2}(\partial B_{\alpha_{0}})})\geq\varepsilon/5$. We can define a cut-off function $h\mathrel{\mathop{\ordinarycolon}}M\to[0,1]$ with $h\equiv 1$ in $\overline{N_{r_{0}/2}(\partial B_{\alpha_{0}})}$, $g\equiv 0$ outside $\overline{N_{r_{0}}(\partial B_{\alpha_{0}})}$ and $\operatorname{Lip}(g)\leq\frac{3}{r_{0}}$. Then we have
$\frac{\varepsilon}{10}>V(\overline{N_{r_{0}}(\partial B_{\alpha})})\geq V(g)\geq W(g)-\mathbf{F}(V,W)\operatorname{Lip}(g)\geq\frac{\varepsilon}{5}-\frac{\delta r_{0}}{3}\geq\frac{\varepsilon}{10}\,$,
a contradiction.
Then fix an arbitrary $\alpha\in I$. For any Lipschitz function $f\mathrel{\mathop{\ordinarycolon}}M\to\mathbb{R}$ with $\sup|f|\leq 1$ and $\operatorname{Lip}(f)\leq 1$, define
$\widetilde{f}(x)=\begin{cases}f(x),&x\in B_{\alpha}\\ 0,&\operatorname{dist}_{M}(x,\partial B_{\alpha})\geq r_{0}\,.\end{cases}$
We can extend $\widetilde{f}$ to a function defined on $M$ with Lipschitz constant $\frac{1}{r_{0}}$ and $L^{\infty}$-norm $1$, still denoted by $\widetilde{f}$. Then we have
$\quad W\llcorner G_{n}(B_{\alpha})(f)-V\llcorner G_{n}(B_{\alpha})(f)$
[…]
$\leq|W(\widetilde{f})-V(\widetilde{f})|+\|W\|(N_{r_{0}}(\partial B_{\alpha})\setminus B_{\alpha})+\|V\|(N_{r_{0}}(\partial B_{\alpha})\setminus B_{\alpha})$
$\leq\mathbf{F}(W,V)\operatorname{Lip}(\widetilde{f})+\|W\|(N_{r_{0}}(\partial B_{\alpha})\setminus B_{\alpha})+\|V\|(N_{r_{0}}(\partial B_{\alpha})\setminus B_{\alpha})$
$\leq\mathbf{F}(W,V)/{r_{0}}+\varepsilon/5<\varepsilon\,.$
This implies that $\mathbf{F}_{B_{\alpha}}(V,W)\leq\varepsilon$. Since $\alpha\in I$ is arbitrary, we conclude that $\sup_{\alpha\in I}\mathbf{F}_{B_{\alpha}}(V,W)\leq\varepsilon$. ∎
Lemma A.4 (Isoperimetric inequality, [Federer_Fleming_60_currents, 6.2]).
In a closed $N$-dimensional Riemannian manifold $(M,\mathbf{g})$, there exists $\delta=\delta(M,\mathbf{g})>0$ and $\nu>0$ such that, if $T\in\mathcal{Z}_{n}(M;\mathbb{Z}_{2})$ and $\mathcal{F}(T)<\delta$, there exists $S\in\mathbf{I}_{n+1}(M;\mathbb{Z}_{2})$ with $\partial S=T$ and $\mathbf{M}(S)\leq\nu\mathbf{M}(T)^{\frac{n+1}{n}}\,.$
Appendix B A genus inequality
Lemma B.1.
Let $\Sigma$ be a smooth closed oriented surface, $S\subset\Sigma\times[0,1]$ be an embedded, smooth, oriented surface homologous to $\Sigma\times\{0\}$ in $H_{2}(\Sigma\times[0,1];\mathbb{Z}_{2})$. Then $\mathfrak{g}(S)\geq\mathfrak{g}(\Sigma)\,.$
Proof.
Without loss of generality, we assume $\Sigma$ and $S$ are connected and $S\cap(\Sigma\times\{0\})=\emptyset$. Consider the projection map onto the first factor $\varpi\mathrel{\mathop{\ordinarycolon}}\Sigma\times[0,1]\to\Sigma$. Since $S$ is homologous to $\Sigma\times\{0\}$, we claim that the mod $2$ degree of $\pi\mathrel{\mathop{\ordinarycolon}}=\varpi|_{S}$ is nonzero. To see this, let $x\in\Sigma$ be a regular value of $\pi$, $D\subset\Sigma\times[0,1]$ be the domain such that $\partial[D]=[S]-[\Sigma\times\{0\}]$ in $\mathbf{I}_{2}(\Sigma\times[0,1],\mathbb{Z}_{2})$. Then $D\cap\varpi^{-1}(x)$ is a finite union of line segments. Note, exactly one endpoint in $D\cap\varpi^{-1}(x)$ lies in $\Sigma\times\{0\}$, and so there is an odd number of remaining endpoints, which together form the set $\pi^{-1}(x)$.
Now we shall show that $\pi^{*}\mathrel{\mathop{\ordinarycolon}}H^{1}(\Sigma;\mathbb{Z}_{2})\to H^{1}(S;\mathbb{Z}_{2})$ is injective, which immediately implies the genus inequality. Let $\alpha\in\ker\pi^{*}\cap H^{1}(\Sigma;\mathbb{Z}_{2})$ be arbitrary, for every $\beta\in H^{1}(\Sigma;\mathbb{Z}_{2})$,
$\langle\alpha\cup\beta,[\Sigma]\rangle=\langle\alpha\cup\beta,\pi_{*}[S]\rangle=\langle\pi^{*}\alpha\cup\pi^{*}\beta,[S]\rangle=0\,.$
By the nondegeneracy of cup product, it follows that $\alpha=0$. ∎
Proof of Lemma 2.6.
Let $\eta$ be defined as in Corollary 5.10, and $\varepsilon=\delta(M,\mathbf{g})$ as in Lemma A.4. We also let $r=\delta(M,\mathbf{g},\Sigma,\varepsilon)$ as defined in Corollary 5.10.
By Corollary 5.10, there exists a punctate surface $S^{\prime}$ such that:
  1. (i)
    $S^{\prime}\subset N_{\eta}(\Sigma)$.
  2. (ii)
    $\mathcal{F}(S^{\prime}-\Sigma)\leq\mathbf{F}(S^{\prime},\Sigma)<\varepsilon$.
  3. (iii)
    $\mathfrak{g}(S^{\prime})\leq\mathfrak{g}(S)$.
By (i),  (ii) and Lemma A.4, $S^{\prime}$ and $\Sigma$ are homologous in $H_{2}(N_{\eta}(\Sigma),\mathbb{Z}_{2})$. Since $N_{\eta}(\Sigma)=\exp^{\perp}(\Sigma,(-\eta,\eta))$ is diffeomorphic to $\Sigma\times(0,1)$, by  (iii) and Lemma B.1, we can conclude that $\mathfrak{g}(S)\geq\mathfrak{g}(S^{\prime})\geq\mathfrak{g}(\Sigma)\,.$
Appendix C Punctate surfaces with arbitrarily long loops
In this section, we show that for all $L\in\mathbb{R}^{+}$, there exists a punctate surface $S$ of genus $1$ in $\mathbb{R}^{3}$ such that its area is $1$ but any nontrivial loop in $S$ has length at least $L$.
We start with a flat torus whose meridian has length $L$ and longitude $1/L$. Then we remove $9/10$ of an orbit of slope $2L^{2}$ to obtain a “slit torus”, denoted by $T_{1}$, as illustrated in Figure 7. It is easy to see that every nontrivial loop in the slit torus has length more than $L$ while the slit torus still has area $1$. However, note that this is not a punctate surface in $\mathbb{R}^{3}$, as it is not embedded in $\mathbb{R}^{3}$ and the singular set is a slit instead of a finite set of points.
Figure 7. A slit torus
Next, in $\mathbb{R}^{3}$, we can choose a sufficiently small embedded torus with one point removed, denoted by $T_{2}$, and a smooth embedding $\Phi\mathrel{\mathop{\ordinarycolon}}T_{1}\to\mathbb{R}^{3}$ which maps the slit in $T_{1}$ to the point in $T_{2}$. Moreover, we require that $\Phi$ is a strictly short map in the sense that its Lipschitz constant is less than $1$. Finally, one can follow Nash-Kuiper’s spiral-strain construction [Kuiper_1955_C1_isometric] to deform $\Phi$ and, without taking the limit, to obtain a smooth almost isometry $\Phi^{\prime}\mathrel{\mathop{\ordinarycolon}}T_{1}\to\mathbb{R}^{3}$ such that $T^{\prime}_{2}\mathrel{\mathop{\ordinarycolon}}=\Phi^{\prime}(T_{1})$ is still a torus with a point removed. One can also verify that $T^{\prime}_{2}$ is a punctate surface that we need. The details are left to the readers.
Appendix D Proof of Theorem 2.21
In this section, we use Proposition 6.1 and the following local min-max theorem to finish the proof of Theorem 2.21. The proof closely follows that in [chuLi2024fiveTori, § 8.5], with a slight improvement: we no longer modify the parameter space. For completeness, we include the details below.
Let ${\mathbb{B}}^{k}$ be the open unit $k$-ball, and $\overline{\mathbb{B}}^{k}$ the closed unit $k$-ball.
Theorem D.1 (Local min-max theorem [MN21, Theorem 6.1]).
Let $\Sigma$ be a closed, smooth, embedded non-degenerate minimal surface with Morse index $k$ and multiplicity one, in a closed $3$-dimensional manifold $(M,g)$. For every $\beta>0$, there exists $\varepsilon_{0}>0$ and a smooth family $\{F_{v}\}_{v\in\overline{\mathbb{B}}^{k}}\subset\operatorname{Diff}^{\infty}(M)$ such that
  1. (1)
    $F_{0}=\operatorname{Id}$, $F_{v}=F^{-1}_{v}$ for all $v\in\overline{\mathbb{B}}^{k}$.
  2. (2)
    $\|F_{v}-\operatorname{Id}\|_{C^{1}}<\beta$ for all $v\in\overline{\mathbb{B}}^{k}$.
  3. (3)
    The function $A^{\Sigma}\mathrel{\mathop{\ordinarycolon}}\overline{\mathbb{B}}^{k}\to[0,\infty]$ given by $v\mapsto\mathcal{H}^{2}((F_{v})_{\#}\Sigma)$ is strictly concave.
  4. (4)
    For every $T\in\mathcal{Z}_{2}(M;\mathbb{Z}_{2})$ with $\mathcal{F}(T,[\Sigma])<\varepsilon_{0}$, we have
    $\max_{v\in\overline{\mathbb{B}}^{k}}\mathbf{M}((F_{v})_{\#}T)\geq\mathcal{H}^{2}(\Sigma)$
    with equality only if $\Sigma]=(F_{v})_{\#}T$ for some $v\in\overline{\mathbb{B}}^{k}$.
We proceed with the proof in three steps.
Step 1. Set up
Since $\mathcal{W}_{L,\leq g}(M,\mathbf{g})$ consists of a varifold associated with a non-degenerate multiplicity-one minimal surface $\Gamma$, we define $\delta_{0}\mathrel{\mathop{\ordinarycolon}}=\delta(M,\mathbf{g},\Gamma,r/10),$ where the function $\delta(\cdot,\cdot,\cdot,\cdot)$ is as defined in Corollary A.2. Let $\varepsilon_{0}>0$ and $\{F_{v}\}_{v\in\overline{\mathbb{B}}^{k}}$ be the smooth family associated with $\Gamma$ given by Theorem D.1 such that:
  • (i)
    For every $S_{1},S_{2}\in\mathcal{S}(M)$ and $v\in\overline{\mathbb{B}}^{k}$,
    $\mathbf{F}(F_{v}(S_{1}),F_{v}(S_{2}))\leq 2\mathbf{F}(S_{1},S_{2})\quad\textrm{ and }\quad\mathbf{F}(F_{v}(S_{1}),S_{1})\leq\frac{\delta_{0}}{10L}\mathcal{H}^{2}(S_{1})\,.$
We can choose smaller $\varepsilon_{0}>0$ such that:
  • (ii)
    For every $S\in\mathcal{S}(M)$ with $\mathbf{F}(S,\Gamma)<\varepsilon_{0}$, the function $A^{T}\mathrel{\mathop{\ordinarycolon}}\overline{\mathbb{B}}^{k}\to[0,\infty]$ given by $v\mapsto\mathcal{H}^{2}(F_{v}(S))$ is strictly concave with a unique maximum in $\mathbb{B}^{k}_{1/2}$.
  • (iii)
    We have
    $\varepsilon_{1}\mathrel{\mathop{\ordinarycolon}}=\min_{v\in\partial\mathbb{B}^{k}}(\mathcal{H}^{2}(\Gamma)-\mathcal{H}^{2}(F_{v}(\Gamma)))>0\,.$
Let $\mathscr{S}\mathrel{\mathop{\ordinarycolon}}=\{F_{v}(\Gamma)\}_{v\in\overline{\mathbb{B}}^{k}}$, which is a compact subset of embeddings into $(M,g)$. We set
$\delta_{1}\mathrel{\mathop{\ordinarycolon}}=\delta(M,\mathbf{g},\mathscr{S},m+1,N_{P}(\Phi)+g,\min(\delta_{0},\varepsilon_{0},\varepsilon_{1})/10)$
as defined in Proposition 6.1.
Step 2. Initial Simon–Smith family
Let $r_{0}\mathrel{\mathop{\ordinarycolon}}=\delta_{1}/10$. It follows from Theorem 2.17 with $r=r_{0}$ that there exist $\eta\in(0,\delta_{0}/100)$ and $\Phi_{1}\in\Lambda(\Phi)$ such that $\mathcal{H}^{2}(\Phi_{1}(x))\geq L-4\eta$ implies $\Phi_{1}(x)]\in\mathbf{B}^{\mathbf{F}}_{r_{0}}([\Gamma])\,.$ In particular, there exists a Simon–Smith family $H_{1}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}(M)$ such that $H_{1}(0,\cdot)=\Phi$ and $H_{1}(1,\cdot)=\Phi_{1}$. In fact, for every $x\in X$, $H_{1}(\cdot,x)$ is a pinch-off process, because it is induced by a one-parameter group of diffeomorphisms.
We refine the cubical subcomplex $X$ so that, for each cell $\sigma$ of $X$, exactly one of the following conditions holds:
  • There exists a point $x_{0}\in\sigma$ such that $\mathcal{H}^{2}(\Phi_{1}(x_{0}))\geq L-2\eta$; Moreover, in this case, for every cell $\sigma^{\prime}$ with $\sigma\cap\sigma^{\prime}\neq\emptyset$ and $x\in\sigma^{\prime}$, $\mathcal{H}^{2}(\Phi_{1}(x))\geq L-4\eta$.
  • For every $x\in\sigma$, $\mathcal{H}^{2}(\Phi_{1}(x))<L-2\eta$.
We denote the side length of each cell in $X$ by $\ell$. Let $Y$ be the smallest cubical subcomplex of $X$ containing all cells that satisfy the first condition, and let $\widetilde{Y}$ be the cubical subcomplex of $X$ whose underlying set is $\{x\in X\mathrel{\mathop{\ordinarycolon}}d(x,Y)\leq\ell\}\,.$ Let $Z\mathrel{\mathop{\ordinarycolon}}=\overline{\widetilde{Y}\setminus Y}$, which is compact. Clearly, for every $z\in Z$,
$L-4\eta\leq\mathcal{H}^{2}(\Phi_{1}(z))\leq L-2\eta\,.$
Step 3. Deformation
For every $z\in Z$, let $A^{z}\mathrel{\mathop{\ordinarycolon}}\overline{\mathbb{B}}^{k}\to[0,\infty)$ be the function given by $v\mapsto\mathcal{H}^{2}(F_{v}(\Phi_{1}(z))).$ By (ii) of Step 1, $A^{z}$ is strictly concave and has a unique maximum $m(z)\in\mathbb{B}^{k}_{1/2}$. As shown in [chuLi2024fiveTori, §2], the Simon–Smith family $\Phi_{1}$ is continuous in the $\mathbf{F}$-metric, so the function $m\mathrel{\mathop{\ordinarycolon}}Z\to\mathbb{B}^{k}_{1/2}$ is also continuous.
It follows from Theorem D.1 (4) and $\mathcal{H}^{2}(\Phi_{1}(z))<L$, that $m(z)\neq 0$ for every $z\in Z$. Hence, there exists $\alpha>0$ such that $\alpha\leq|m(z)|<1/2\,.$ Consider the one-parameter flow $\{\phi^{z}(\cdot,t)\}_{t\geq 0}\subset\operatorname{Diff}(\overline{\mathbb{B}}^{k})$ generated by $v\mapsto-(1-|v|^{2})\nabla A^{z}(v)\,.$ For every $v\in\overline{\mathbb{B}}^{k}\setminus m(y)$, $t\mapsto A^{z}(\phi^{z}(v,t))$ is decreasing, and the limit $\lim_{t\to\infty}\varphi^{z}(v,t)\in\partial\mathbb{B}^{k}$. In particular, by (iii) of Step 1, we have
$A^{z}(0)-\lim_{t\to\infty}A^{z}(\phi^{z}(v,t))\geq\varepsilon_{1}\,.$
By the compactness of $Z$, we can choose $T_{0}>0$ such that for every $z\in Z$, $t\mapsto A^{z}(\phi^{z}(v,t))$ is decreasing along $0,T_{0}]$ and
$A^{z}(0)-A^{z}(\phi^{z}(v,T_{0}))\geq\varepsilon_{1}/2\,.$
Now, let us define a bump function $\xi\mathrel{\mathop{\ordinarycolon}}[0,\infty)\to[0,1]$ by
$\xi(s)\mathrel{\mathop{\ordinarycolon}}=\begin{cases}3s/\ell&s\in[0,\ell/3)\\ 1&s\in[\ell/3,2\ell/3)\\ 3-3s/\ell&s\in[2\ell/3,\ell)\\ 0&s\in[\ell,\infty)\,.\end{cases}$
Then we define a new Simon–Smith family $\Phi_{2}\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ by
$\Phi_{2}(x)\mathrel{\mathop{\ordinarycolon}}=F_{\phi^{x}(0,\xi(d(x,Y))\cdot T_{0})}(\Phi_{1}(x))\,.$
We also define a deformation $H_{2}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}(M)$ from $\Phi_{1}$ to $\Phi_{2}$ by
$H_{2}(t,x)$ $=F_{\phi^{x}(0,t\cdot\xi(d(x,Y))\cdot T_{0})}(\Phi_{1}(x))\,.$
Note that when $x\notin Z$, we abuse the notation by setting $F_{\phi^{x}(0,0)}=\operatorname{Id}$.
Similarly, for every $x\in X$, $H_{2}(\cdot,x)$ is also induced by an isotopy, and thus, a pinch-off process. In particular, we have
(D.1) $N_{P}(\Phi)=N_{P}(\Phi_{1})=N_{P}(\Phi_{2})\,.$
Furthermore, $\Phi_{2}$ has the following properties: For every $x\in X$,
  • (i)
    if $d(x,Y)>0$, then $\mathcal{H}^{2}(\Phi_{2}(x))\leq L-2\eta\,;$
  • (ii)
    if $d(x,Y)\leq 2\ell/3$, by (i) of Step 1,
    $\mathbf{F}(\Phi_{2}(x),\mathscr{S})\leq 2\mathbf{F}(\Phi_{1}(x),\Gamma)<2r_{0}<\delta_{1}\,$,
    $\mathbf{F}(\Phi_{2}(x),\Gamma)$ $\leq\mathbf{F}(\Phi_{2}(x),\mathscr{S})+\mathbf{F}(\mathscr{S},\Gamma)<\delta_{1}+\delta_{0}/10\leq\delta_{0}/5,$
    and thus, $\mathcal{F}([\Phi_{2}(x)],[\Gamma])\leq\delta_{0}/10\,;$
  • (iii)
    if $d(x,Y)\in[\ell/3,2\ell/3]$, then $\mathcal{H}^{2}(\Phi_{2}(x))\leq L-2\eta-\varepsilon_{1}/2\,.$
Step 4. Interpolation
Let $Y^{\prime}\mathrel{\mathop{\ordinarycolon}}=\{x\in X\mathrel{\mathop{\ordinarycolon}}d(x,Y)\leq 2\ell/3\}$. By Step 3 (ii), we can apply Proposition 6.1 to $\Phi_{2}|_{Y^{\prime}}$ and obtain a deformation $H_{3,Y^{\prime}}\mathrel{\mathop{\ordinarycolon}}[0,1]\times Y^{\prime}\to\mathcal{S}(M)\,.$ Note that by the choice of $\delta$, we can assume $\varepsilon=\min(\delta_{0},\varepsilon_{0},\varepsilon_{1})/10$ in Proposition 6.1.
Consequently, we can define a Simon–Smith family $H_{3}\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}(M)$ as
$H_{3}(t,x)=\begin{cases}H_{3,Y^{\prime}}(t,x)&d(x,Y)\leq\ell/3\\ H_{3,Y^{\prime}}\left((2-3d(x,Y)/\ell)t,x\right)&x\in Y^{\prime}\\ \Phi_{2}(x),&x\notin Y^{\prime}\,.\end{cases}$
We denote $H_{3}(1,\cdot)\mathrel{\mathop{\ordinarycolon}}X\to\mathcal{S}(M)$ by $\Phi_{3}$.
By Proposition 6.1 (1), for every $x\in X$ with $d(x,Y)\leq\ell/3$, we have $\mathfrak{g}(\Phi_{3}(x))=\mathfrak{g}(H_{3}(1,x))=\mathfrak{g}(\Gamma)\,$, and
$\mathcal{H}^{2}(\Phi_{3}(x))=\mathcal{H}^{2}(H_{3,Y^{\prime}}(1,x))\leq\mathcal{H}^{2}(\Phi_{2}(x))+\delta_{0}/10\,.$
By Proposition 6.1 (2), Step 3 (ii) and (iii), for every $x\in X$ with $d(x,Y)\in[\ell/3,2\ell/3]$, we have
$\mathcal{H}^{2}(\Phi_{3}(x))=\mathcal{H}^{2}\left((2-3d(x,Y)/\ell)t,x\right)\leq\mathcal{H}^{2}(\Phi_{2}(x))+\varepsilon_{1}/10\leq L-2\eta\,.$
By Step 3 (i), for every $x\in X$ with $d(x,Y)\geq 2\ell/3$,
$\mathcal{H}^{2}(\Phi_{3}(x))=\mathcal{H}^{2}(\Phi_{2}(x))\leq L-2\eta\,.$
Therefore, for every $x\in X$ with $\mathcal{H}^{2}(\Phi_{3}(x))\geq L-\eta$, we have $d(x,Y)\leq\ell/3$, which implies:
  • (i)
    $\mathfrak{g}(\Phi_{3}(x))=\mathfrak{g}(\Gamma)$;
  • (ii)
    $|\mathcal{H}^{2}(\Phi_{3}(x))-L|\leq\delta_{0}$;
  • (iii)
    $\mathcal{F}([\Phi_{3}(x)],[\Gamma])\leq\mathcal{F}([\Phi_{3}(x)],[\Phi_{2}(x)])+\mathcal{F}([\Phi_{2}(x)],[\Gamma])<\delta_{0}\,.$
In particular, by (ii), (iii) and Corollary A.2, we can conclude that $\mathbf{F}(\Phi_{3}(x),\Gamma)<r\,.$
Moreover, by Proposition 6.1 (3), for every $x\in X$, $\{H_{3}(t,x)\}_{t\in[0,1]}$ is also a pinch-off process. Finally, we can set $\Phi^{\prime}=\Phi_{3}$, which satisfies (1) of the theorem. The deformation map $H\mathrel{\mathop{\ordinarycolon}}[0,1]\times X\to\mathcal{S}(M)\,$, is simply the concatenation of $H_{1}$, $H_{2}$ and $H_{3}$ defined above. Clearly, it is a Simon–Smith family of genus $\leq g$ satisfying (2)(a) and (2)(b) of the theorem. This completes the proof.